A Complete Bibliography of the Journal of Number Theory (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

06 April 2024 Version 1.08

Title word cross-reference

(1^{3} + q^{3})(2^{3} + q^{3}) ⋯ (n^{3} + q^{3}) [2054]. (1^{ℓ} + 1)(2^{ℓ} + 1) ⋯ (n^{ℓ} + 1) [764, 1080]. (1^{k} + 1)(2^{k} + 1) ⋯ (n^{k} + 1) [569]. (1^{μ} + 1)(2^{μ} + 1) ⋯ (n^{μ} + 1) [24]. (2, 3) [2480]. (5, 5) [1319]. (a, b) [2310]. (a^{2}cx^{k} − 1)(b^{2}cy^{k} − 1) = (abcz^{k} − 1)^{2} [1335]. ({a_{n}α})_{n≥1} [2488]. (ax^{k} − 1)(by^{k} − 1) = abz^{k} − 1 [925]. (χ, b) [1520]. (d ≥ 4) [1204]. (ℓ_{1}, ℓ_{2}) [1475]. (k + 1) [1838]. (m^{2} − 1)^{r} + b^{2} = c^{2} [1323]. (m^{2} − 1)^{x} + b^{y} = c^{z} [1323]. (GSp_{4}, GL_{2×_{GL_{1}}} GL_{2}) [2121]. (Sp_{4n}(F), Sp_{2n}(E)) [194]. (nα) [328]. (ϕ, Γ) [2638]. (ρ, q) [1799]. (S, {2}) [1426, 2509]. (s, t) [1798]. (φ, Ĝ) [849]. (φ, N) [1590]. (x + 1)^{k} + (x + 2)^{k} + ⋯ + (2x)^{k} = y^{n} [2131]. −1 [2446]. −2 [1493]. −3 [1260]. 0 [1669, 2375]. 1 [55, 214, 285, 381, 536, 788, 901, 1248, 1281, 1669, 1701, 1891, 2124, 2189, 2438, 2451]. 1 + 2^{a} + x^{b} = y^{n} [1043]. {1, k^{2} + 1, k^{2} + 2k + 2} [230]. 1/π [548, 746, 824, 2254]. 10 [842, 2156]. 11 [995]. 13 [975, 1110, 2007]. 14 [1451]. 16 [1614]. 16p^{2} [419]. 1^{k} + 2^{k} + ⋯ + x^{k} = y^{n} [1571]. 1^{k} + ⋯ + x^{k} = y^{n} [1359, 1627]. 2 [112, 213, 273, 285, 286, 325, 430, 572, 605, 618, 637, 662, 664, 688, 726, 818, 846, 989, 999, 1011, 1030, 1275, 1319, 1322, 1339, 1381, 1402, 1496, 1499, 1523, 1594, 1788, 1789, 1876, 1878, 1900, 1916, 1992, 2001, 2012, 2028, 2051, 2075, 2079, 2088, 2139, 2162, 2181, 2189, 2214, 2358, 2369, 2370, 2374, 2438, 2580, 2588, 2611]. 20 [2254]. 21 [632]. 23 [2214]. 25 [2214]. 27 [1800]. 2^{a} [1044]. 2^{m} + nx^{2} = y^{n} [567]. 2d + 1 [1204]. 2k [2019, 2468]. 2p [515]. 3 [75, 214, 217, 262, 726, 790, 843, 890, 972, 1011, 1030, 1034, 1192, 1242, 1521, 1590, 1778, 1889, 1956, 2189, 2200, 2309, 2438, 2558, 2582]. 3/2 [768, 994, 1471]. 34 [1260]. 36 [381]. 3^{n} [1436]. 3^{τ} · 2 [1978]. 3k [2468]. 3p^{n+1} [653]. 3x + 1 [1226]. 4 [262, 381, 1088]. 5 [973, 1207, 1402, 1788, 2230, 2518]. 6 [1198, 2156, 2189, 2539]. 64 [2170]. 7 [1011, 1321, 1451, 1487, 1586, 2479]. 72 [125, 1533]. 8 [2295, 2480]. 9 [381, 1030, 1189, 1357]. > 3 [650]. [0, 1]^{d} [2272]. _{3}F_{2} [1324]. A!B! = C! [2227]. A [497, 749, 1033, 1048, 1245, 1657, 1765, 1854, 2601]. a(n, k) [355]. A(q) [816]. A + A [2532]. A + A + A [497]. A + B [146]. A_{m} [456]. A_{n} [1639]. AAA [2532]. α [255, 1452, 2124, 2496]. αp^{γ} + β [1941]. B [1671, 1718]. B_{dR}^{+} [134]. S̄pec(Z) [1561]. β [855, 1669, 1979, 2554]. \binom{2k}{k}^{2}\binom{3k}{k} [709]. mod1 [1308, 1840]. mod11 [979]. mod2 [1131]. mod3 [83]. mod4 [2241]. modℓ [321, 1859]. modl [1943]. modP^{⇕} [608]. modp^{m} [57]. modn [867]. modp [466, 1018, 1033, 1165, 1358, 2263, 2449]. modp^{n} [866]. mod℘ [1470]. C [18]. C_{1} [2585]. C_{m}[ltimes]_{ϕ}C_{mn} [2470]. C_{q}[rtimes]_{s}C_{m} [2609]. D [32, 1204, 1279, 1604, 1808, 2091]. D(−1) [230]. D(1) [230]. D(4) [2384]. D(n) [2159]. D_{n} [554]. Δ [1384]. Δ(prime)^{2} [1384]. Δ_{\overline{O(⇕)}_{∞}} [863]. E [706, 899, 1168]. e^{γ} [1536]. ℓ [646, 720, 828, 829]. ϵ(k) [1885]. er = p − 1 [1642]. F [582, 1012, 1713]. f(p + q + n_{0}) = f(p) + f(q) + f(n_{0}) [1641]. F_{4} [2506]. \frac{n+3}{2} [2537]. G [893, 1915]. g^{2} [2124]. Γ [14, 542, 1561, 2466]. Γ^{∗}(k) [359]. Γ_{0}(2) [2612]. Γ_{0}(3) [2250]. Γ_{0}(p) [1027]. Γ_{0}(N) [2296]. Γ_{0}(p^{α}) [245]. Γ_{0}^{+}(2) [1673]. γ_{α}(z) [602]. h [1373]. H(\sqrt{N}) [111]. j [58, 1261, 1280]. K [355, 482, 598, 659, 688, 803, 811, 982, 1247, 1282, 1312, 1338, 1542, 1661, 1718, 1730, 1763, 1992, 2167, 2230, 2259, 2430]. k(t) [1225]. k + l [2468]. k/12 [2348]. K3 [122, 261, 262, 2309]. k = mp^{s} [1282]. K_{2} [1104]. ka^{x} + lb^{y} = c^{z} [2233]. L [40, 65, 84, 168, 182, 219, 245, 249, 268, 292, 296, 301, 312–314, 320, 326, 342, 347, 384, 387, 407, 468, 503, 524, 586, 621, 645, 666, 668, 671, 766, 769, 880, 884, 905, 978, 1017, 1031, 1035, 1040, 1081, 1093, 1114, 1120, 1162, 1167, 1170, 1184, 1217, 1228, 1301, 1345, 1368, 1387, 1416, 1465, 1488, 1498, 1536, 1694, 1705, 1764, 1802, 1858, 1859, 1865, 1909, 1920, 1938, 1973, 2004, 2091, 2096, 2117, 2118, 2140, 2158, 2211, 2212, 2223, 2243, 2258, 2278, 2285, 2288, 2300, 2316, 2322, 2380, 2425, 2437, 2451, 2471, 2508, 2529, 2575, 2583, 2611, 2616, 2621, 2626, 2635]. L [68, 101, 110, 251, 408, 414, 555, 742, 851, 983, 991, 1032, 1299, 1380, 1555, 1616, 1693, 1809, 2021, 2312, 2388, 2426, 2591]. L(1, χ) [1969]. L(1, χ_{s})/Π [2231]. L(1/2, f × χ_{q}) [705]. L(\frac{1}{2}, χ) [581, 775]. L(\frac{1}{2}, χ_{d}) [2464]. L(\frac{1}{2}, Sym^{2}f × g [2649]. L(\frac{1}{2}, Sym^{2}f × g) [2557]. L^{2}(Γ \ Sp(2, R)) [822]. L_{2} [1671]. L_{n}^{(−1−n−r)}(x) [2133].

L_{n}^{( \frac{1}{2} +u)}(x) [1480]. L_{⊙}(β) [2020]. Λ [542, 965, 1174, 1243]. Λ^{2}Λ^{−} [363]. λ_{1}x_{1}^{k_{1}} + ⋯ + λ_{t}x_{t}^{k_{t}} ≡ c mod n [2224]. [ − m, n] [2584]. ⌊n^{c}⌋ [520]. lm [148]. log 3 [1037]. m [456, 477, 544, 580, 867, 1750, 2188]. M_{2} [1674]. M_{2k}(Γ_{0}(N)) [2416]. C [782]. F_{2} [923]. F_{2}[T] [722]. F_{p} [2458, 2528]. F_{p}(t) [1047]. F_{q} [2354]. F_{q}((x^{−1})) [2020]. F_{q}(T) [241, 2234]. F_{q}[[T]] [976]. F_{q}[t] [86, 974, 1213]. F_{q}^{2} [2062]. G_{m} \ μ_{N} [1491]. N^{d} [1174]. N^{r} [2232]. P^{2} [398]. P^{N} [2282, 2400]. Q [275, 720, 1158, 1383]. Q(i\sqrt{2}) [364]. Q(ζ_{p^{n}+1})^{+} [1488]. Q_{p} [606]. R^{d} [2325]. U_{4}(F_{p}) [1942]. Z/3 [1767]. Z/mZ [1053]. Z[i] [121]. Z^{3} [231]. Z_{2} [1281, 1756]. Z_{l} [1104]. Z_{p} [226, 228, 457, 859, 1263, 1837, 2243, 2249]. Z_{p}^{d} [2581]. Z_{p}^{n} [542]. Z_{p_{1}} × ⋯ × Z_{p_{s}} [720]. P^{n}(C)(n ≥ 1) [2297]. Q [800]. Z_{m} [64]. Z_{p} [2236, 2368]. D [672]. F_{∞,∈} [1342]. H × C^{(},∞)} [2391]. I_{⌋}^{(⨿)} [2120]. L [1185]. P [1489]. gl(n) [1147]. p [323, 1938]. v [929]. GL(2 ) [485, 1380, 2035]. GL(3 ) [192, 320, 1220]. GL(3 ) × GL(2 ) [320]. GL_{2} [621]. GL_{m} [1228]. GL_{m}(D) [823]. SL(2 , C) [1132]. SL(2 , Z[i]) [1000]. SL_{2} (Q_{p}) [693, 2489]. SL_{m}(Z) [1587]. SO_{0} (m, 2 ) [555]. SO_{2n} × GL_{n} [147]. Sp(4 ) [1139]. Gal(k((t))/k) [1638]. Gal(K_{2}^{(2)}/K) [1992]. GCF_{ϵ} [1885]. GL(2) [1509, 1621]. GL(2, D) [1279]. GL(2, F_{q}[T]) [1696]. GL(3) × GL(2) [2113]. GL(4) [825, 1131]. GL(4) × GL(2) [2626]. GL(n) [1144, 1559, 2134]. GL(r) [1535]. GL(r) × GL(r + s) × GL(s) [879]. GL_{2} [778, 1897, 1938]. GL_{2}(A) [1027]. GL_{2}(F) [1018]. GL_{n} [1529, 2212]. GL_{n}(F) [1482]. GO_{4} [1518]. GSp(4) [920, 1134]. GSp(4, F) [2426]. GSp_{4} [2072, 2493]. lcm(a, b, c) ≤ 6 [1729]. PGL(2) [623]. PGL(2, F) [633]. PSL_{2} [2371]. PSL_{2}(R)^{n} [751]. PSL_{4}(Z) [394]. Res < 1/2 [875]. SL(2, Z) [2324]. SL(n, Z) [1316]. SL_{2} [1460, 1556]. SL_{2}(Z) [2274]. SL_{n} [2256]. SL_{n}(A) [2338]. SO_{4} [2604]. Sp(4, F) [2487]. Sp_{2m} × Sp_{2n} [229]. Sp_{4} [2604]. U(1) [680]. U(1, 1) [680]. U_{2} [2604]. M \ L [2396]. μ [904]. μ = 0 [1822]. N [353, 355, 482, 531, 589, 604, 807, 1243, 1351, 1502, 1657, 1777, 1914, 2159, 2242, 2425]. n/φ(n) [588]. N = 2, 3, 4, 6, 8 [1491]. n = p^{α}q^{β} [862]. n = x^{2} + Ny^{2} [111]. n^{2} = x^{2} + by^{2} + cz^{2} [649]. nx^{2} + 2^{2m} = y^{n} [327, 1001]. NX^{2} + 2^{L}3^{M} = Y^{N} [1016]. Ω [63, 1316, 2551]. P [2, 14, 55, 65, 91, 243, 285, 298, 309, 331, 349, 372, 374, 457, 478, 492, 512, 525, 542, 593, 635, 690, 693, 702, 707, 733, 754, 757, 788, 819, 829, 837, 866, 869, 877, 881, 885, 892, 905, 912, 924, 957, 990, 993, 1013, 1023, 1081, 1087, 1091, 1183, 1208, 1268, 1279, 1301, 1310, 1412, 1433, 1435, 1474, 1488, 1503, 1556, 1558, 1655, 1668, 1726, 1735, 1738, 1751, 1765, 1775, 1786, 1805, 1811, 1842, 1846, 1874, 1875, 1902, 1927, 1957, 1965, 1966, 1970, 2040, 2072, 2083, 2091, 2236, 2239, 2300, 2323, 2373, 2412, 2415, 2453, 2502, 2526, 2552, 2581, 2614, 2630]. p [154, 251, 539, 1190, 1352]. p + 1|n + 1 [1155]. p + 2^{ℓ} [2606]. p = 2 [949, 2622]. p = a^{2} ± ab ± b^{2} [1668]. p = x^{2} + 3y^{2} [791]. p^{2} [419, 837, 1282]. p^{3} [1282, 1297]. p^{ℓ} [258]. p^{n} [1383]. Φ_{T} (X, Y) [311]. π [358, 906, 967]. π^{−2} [1443, 1882]. π^{4} [1997]. π^{6} [1997]. ±ra^{x} ± sb^{y} = c [293]. px^{2} + q^{2m} = y^{p} [7, 189]. px^{2} + q^{2n} = y^{p} [336]. Q [40, 413, 647, 819, 836, 844, 900, 1111, 1281, 1296, 1386, 1464, 1511, 1743, 1880, 1887, 1904, 1914, 1967, 2066, 2140, 2512, 2534]. q + 1 [2062]. r [22, 1661, 1974, 2269, 2499]. S [47, 133, 963, 1460, 1605, 1894, 2092, 2093, 2436]. s = −1 [1045]. s = 0 [516]. s = 1 [671]. s = 1/2 [1536]. S_{3}(Γ_{0}(24), χ) [2189]. S_{3}(Γ_{2}(2, 4, 8)) [407]. S_{4} [2460]. S_{n}(σ, t) [2516]. [Cyrillicsha] [236, 379]. σ(n)/n [588]. ∑_{ak+bl+cm=n} σ(k)σ(l)σ(m) [1729]. ∑_{i=1}^{11} \frac{1}{x_{i}} = 1 [732]. ∑_{n=1}^{∞} σ_{k}(n)/n!

[275]. ∑_{r=1}^{τ} e(a2^{r}/q) [566]. T [537, 722, 861, 984, 1167, 1343, 1672, 1696, 1830, 2053, 2100, 2223, 2422]. t^{2} = G(x, y, z) [1454]. θ [1801, 2402, 2551]. θ_{10} [1327]. U [1336, 1480]. U(2, 1) [2412]. U(3) × U(2) [1645]. U(n) [1611]. U(n, 1) [1749]. U_{p} [55]. Υ [1203]. v [537, 820]. φ(n) [1506]. φ^{4} [2160]. ∥e^{n}∥ [139]. w [432, 716]. X [1571, 2588, 2589]. x(ax + 1) + y(by + 1) + z(cz + 1) [1814]. x(ax + b) + y(ay + c) + z(az + d) [1814]. x^{2} + ny^{2} [74]. x^{2} − (a^{2} − 1)y^{2} = 1 [2150]. x^{2} − 24y^{2} = 1 [1157]. X_{0}^{∗}(p) [2420]. x^{d} + ax [1498]. x^{n} − 1 [1700, 1774]. x^{q} = a [606]. X_{0}(N) [817, 2579]. X_{0}(p) [55]. X_{0}^{+}(37M) [181]. X_{0}^{+}(N) [2183]. x_{1}^{2} + 2x_{2}^{2} + x_{3}^{2} + x_{4}^{2} + x_{1}x_{3} + x_{1}x_{4} + x_{2}x_{4} [395]. x_{1}^{2} + ⋯ + x_{k}^{2} ≡ λ(modn) [2521]. x_{1}x_{2} ≡ x_{3}x_{4}(modm) [67]. x_{1}x_{2} ≡ x_{3}x_{4} mod q [2041]. X_{2}(2) [2582]. x ≡ 1 mod N [74]. Ξ [1256, 1731]. X ↦ K · (X + X^{−1}) [1038]. X ↦ X + X^{−1} [686]. y^{2} + xy − t^{d}y = x^{3} [1714]. y^{2} − pz^{2} = 1 [1157, 2150]. y^{2} = x^{3} + n [435]. y^{2} = x^{3} − nx [712]. y^{n} = x^{m} + x [1098]. y^{q} = x^{p} + a [1118]. y ≡ 0 mod N [74]. Z [1896, 2164]. z^{2} = f(x)^{2} ± g(y)^{2} [1899]. Z_{p^{2}}^{∗} [196]. ζ(2) [1116]. ζ({{2}^{m}, 1, {2}^{m}, 3}^{n}, {2}^{m})/π^{4n+2m(2n+1)} [1230]. ζ(3) [1116]. ζ(s) [248, 1643, 2173]. ζ′(s) [541, 1643]. |ζ(\frac{1}{2} + it)| [1197].

-additive [1511]. -Adic [2, 14, 55, 65, 101, 213, 243, 251, 309, 323, 349, 478, 492, 512, 525, 537, 542, 593, 635, 646, 690, 702, 707, 722, 754, 788, 819, 820, 829, 869, 881, 885, 905, 912, 929, 957, 984, 990, 1023, 1081, 1087, 1091, 1183, 1190, 1208, 1243, 1268, 1275, 1301, 1310, 1322, 1352, 1433, 1474, 1488, 1503, 1521, 1556, 1558, 1655, 1672, 1735, 1738, 1751, 1786, 1805, 1842, 1846, 1894, 1902, 1938, 1965, 1970, 2040, 2075, 2083, 2091, 2181, 2223, 2239, 2300, 2358, 2453, 2502, 2526, 2552, 2580, 2611, 2614, 2630]. -adique [2, 154, 820]. -adischen [2580]. -algebra [2312]. -analog [844]. -analogue [1887]. -analogues [1111]. -ary [877, 1750, 2019, 2188]. -bases [1373, 1916]. -Bell [1974]. -Bernoulli [1412]. -binomial [900]. -Class [664, 1192, 1339, 1499, 1811, 1876, 1992, 2079, 2093]. -coefficients [1767]. -color [1402, 1594]. -colored [2088, 2167, 2230]. -complete [1604]. -continued [255, 807]. -convergence [2120]. -coordinates [2588, 2589]. -core [861, 972, 1011, 2100]. -cores [1242]. -covering [544]. -covers [2243]. -crystals [582]. -decompositions [1012]. -deformed [819]. -derived [1730]. -diamond [843, 890, 995, 1011, 1788]. -differences [2499]. -digit [2181]. -dimensional [32, 133, 430, 604, 1487, 1956]. -divisible [967]. -division [790, 2518]. -eigenspaces [1248]. -elliptic [672]. -expansion [1669, 1979]. -expansions [749, 1245, 1801, 2124, 2402, 2496, 2554]. -exponents [836]. -extension [230, 720]. -extensions [91, 226, 457, 859, 924, 1104, 1263, 1837, 1942, 2236, 2249, 2369]. -factors [1520, 2426, 2466]. -family [1048]. -fields [2585]. -fold [982, 1357, 2582]. -folds [973]. -forms [1489]. -formula [2348]. -free [1502]. -friable [1718]. -Frobenius [1464, 1743]. -function [251, 296, 342, 387, 432, 766, 893, 905, 983, 991, 1217, 1256, 1280, 1731, 1896, 1915, 2164]. -Functions [40, 65, 68, 110, 182, 219, 245, 249, 268, 292, 301, 312, 320, 326, 347, 384, 407, 414, 468, 503, 555, 586, 621, 645, 668, 671, 742, 769, 851, 884, 978, 1017, 1031, 1081, 1093, 1114, 1120, 1162, 1170, 1184, 1228, 1301, 1345, 1368, 1380, 1387, 1416, 1488, 1555, 1616, 1693, 1694, 1705, 1764, 1802, 1809, 1858, 1865, 1909, 1920, 1938, 1973, 2004, 2021, 2096, 2117, 2118, 2140, 2158, 2211, 2212, 2223, 2243, 2258, 2278, 2285, 2288, 2300, 2316, 2322, 2380, 2388, 2425, 2437, 2451, 2471, 2508, 2529, 2575, 2583, 2591, 2616, 2621, 2626, 2635]. -Gamma [1386]. -gcd [1718]. -group [2323]. -groups [811, 866, 1966]. -hypergeometric [1033, 1904]. -integer [993]. -integers [1460]. -integral [2092]. -invariant [58, 314, 904, 1261]. -invariants [542, 965, 1174, 1185]. -isogenous [2309]. -isogeny [1319]. -Isomorphism [1312]. -Iwasawa [1426]. -lattices [554]. -level [2425]. -linear [275, 2234]. -Lipschitz [788, 1281]. -Littlewood [1475]. -loop [2438]. -Lüroth [1452]. -Mahler [1281]. -manifolds [2558]. -map [2601]. -modular [829]. -modules [849, 1167, 1343, 1590, 2638]. -motives [536, 537, 2053]. -multisets [1889]. -MZSV [413]. -Narayana [1967]. -nomial [18]. -numbers [1336, 1605]. -orderings [1874]. -ordinary [2236]. -parallelogram [2551]. -part [214, 720]. -power [688, 1013]. -powers [688]. -prime [22]. -primitive [2373]. -quintuples [2384]. -Raabe [647, 1880]. -ramified [1900, 2012]. -rank [258, 285, 1435, 1778]. -ranks [1674, 2422, 2581]. -refined [2509]. -regular [1030, 1044, 1299, 1798]. -results [63, 1316]. -rings [1561]. -sequences [328, 706]. -series [666, 1035, 1465, 1536, 2066, 2534]. -sets [1713, 2159]. -sheaves [2458]. -stage [618]. -th [867, 1247, 1338, 1808, 1838, 1914, 1927, 2091, 2242, 2259, 2269]. -theory [659]. -torsion [372, 1383, 1878]. -towers [1499, 2581]. -transform [14]. -transforms [542]. -triangle [2551]. -triple [230]. -tuple [2430]. -tuples [531, 1661]. -type [1518, 1897]. -unit [47]. -units [963, 2436]. -values [84, 168, 313, 880, 1032, 1040, 1167, 1854, 1859, 2091, 2611]. -Volkenborn [1799]. -wise [1661]. -zeta [2512].

107 [107, 1271]. 109 [1128]. 125 [93]. 128 [189]. 129 [94, 106, 1304]. 130 [129, 302, 2140]. 131 [384, 517, 1001]. 132 [550, 655]. 133 [738, 739, 808, 939, 957, 1471]. 137 [2080]. 148 [1291, 1303]. 155 [1627]. 156 [1628]. 158 [1882]. 159 [1743]. 167 [1770, 2642]. 171 [1848]. 173 [2079]. 174 [2012]. 178 [2139]. 179 [2141]. 186 [2644]. 189 [2648]. 195 [2646]. 196 [2643]. 197 [2548].

200 [2647]. 202 [2645, 2649]. 297 [2645].

56 [909].

A. [39]. A1 [1993]. A2 [1993]. ABC [719, 974, 1332, 2110]. Abelian [10, 72, 90, 128, 184, 217, 226, 270, 356, 379, 559, 637, 653, 694, 728, 785, 800, 805, 896, 912, 918, 924, 1013, 1029, 1074, 1161, 1167, 1179, 1194, 1224, 1297, 1304, 1398, 1468, 1470, 1513, 1566, 1584, 1723, 1842, 1966, 2132, 2190, 2209, 2318, 2320, 2323, 2369, 2414, 2429, 2449, 2478, 2514, 2571, 2617, 2618, 2646]. abéliennes [1161, 1297]. Abhyankar [1495]. absence [1390]. Absolute [971, 1068, 1092, 1561, 1605, 1780]. Abu [189]. abundant [2546]. according [1183]. account [1969]. accurate [1626]. accurately [1839]. acting [55]. action [711, 939, 953, 1132, 1345]. actions [49, 2023, 2141]. Adamchik [1089, 1591]. Addendum [93, 129, 384, 738, 739, 2139, 2140, 2642]. Adding [648]. Addison [166, 1397]. Addison-type [166, 1397]. Addition [282, 284, 411, 1351, 1777, 2466]. Additive [265, 335, 499, 609, 616, 891, 922, 1008, 1373, 1401, 1511, 1916, 1946, 1986, 2009, 2055, 2107, 2144, 2353, 2583]. adèle [235]. Adelic [66, 498, 508]. Adic [2, 14, 55, 65, 101, 213, 243, 251, 309, 323, 349, 478, 492, 512, 525, 537, 542, 593, 635, 646, 690, 702, 707, 722, 754, 788, 819, 820, 829, 869, 881, 885, 905, 912, 929, 957, 984, 990, 1023, 1081, 1087, 1091, 1183, 1190, 1208, 1243, 1268, 1275, 1301, 1310, 1322, 1352, 1433, 1474, 1488, 1503, 1521, 1556, 1558, 1655, 1672, 1735, 1738, 1751, 1786, 1805, 1842, 1846, 1894, 1902, 1938, 1965, 1970, 2040, 2075, 2083, 2091, 2181, 2223, 2239, 2300, 2358, 2453, 2502, 2526, 2552, 2580, 2611, 2614, 2630]. adique [2, 154, 820]. adischen [2580]. adjacent [716]. adjoint [314, 1301, 1990, 2541]. admit [1178, 1856, 2423]. Advanced [1836]. affine [193, 587, 745, 1577, 1704, 2517]. affine-like [587]. after [2553]. AGM [782]. Akashi [2367]. algebra [812, 1268, 1279, 1561, 2312, 2431, 2513, 2539]. Algebraic [2, 3, 22, 30, 82, 98, 107, 123, 156, 172, 195, 204, 340, 404, 559, 560, 603, 671, 820, 839, 854, 956, 1069, 1126, 1194, 1211, 1236, 1287, 1347, 1355, 1390, 1441, 1557, 1656, 1662, 1699, 1704, 1780, 1782, 1820, 1832, 1875, 1934, 1976, 2093, 2109, 2161, 2163, 2431, 2552, 2593, 2636, 2642]. algebras [174, 244, 462, 476, 491, 557, 882, 1206, 1284, 1739, 1950, 2090, 2127, 2186, 2286, 2639]. algèbre [1279]. algébriques [2, 154, 820]. algorithm [618, 631, 705, 841, 1474]. algorithms [283, 744, 846, 1741]. allied [1203]. Almost [360, 447, 480, 936, 1015, 1330, 1364, 1381, 1760, 1793, 1823, 1954, 2136, 2147, 2411, 2606]. Almost-prime [1330, 1793, 2147]. almost-primes [2136]. alternate [2627]. alternating [1340, 1382, 1528, 1839, 1967]. alternative [548, 965, 1614, 1757]. ambiguities [350]. among [236, 1615, 2627]. ample [1929]. amplitudes [1078]. analog [844, 1179, 1287]. Analogies [144]. analogous [1490]. Analogue [338, 667, 725, 975, 1183, 1614, 1732, 1887, 2153, 2348, 2366]. Analogues [155, 724, 1105, 1111, 1156, 1492, 1582, 2456, 2502]. Analysis [348, 716, 843, 1448, 2226, 2349]. Analytic [338, 445, 670, 1227, 1742, 1805, 2091, 2152]. Analytically [1370]. Anderson [2053]. Andrews [1622, 1637, 1869, 2100]. angle [1705]. angled [630]. Anisotropy [926]. Ankeny [2345]. Annihilation [1090]. annihilators [754]. annulation [408]. Anomalous [2531]. anticyclotomic [1859, 2300]. Aomoto [331]. aperiodic [1357]. Apéry [509, 571, 1056, 1189]. Apollonian [389]. Apostol [812, 1544, 1593, 2089]. Apostol-type [2089]. appendix [262, 485]. Application [47, 72, 409, 680, 897, 971, 1106, 1715, 2075, 2403, 2495, 2508]. Applications [196, 274, 378, 621, 711, 765, 916, 939, 976, 1082, 1156, 1190, 1278, 1313, 1437, 1599, 1748, 1802, 1933, 1984, 2002, 2108, 2219, 2416, 2582, 2601]. applying [2557, 2649]. approach [147, 965, 1474, 1548, 1742, 2047, 2053, 2195]. approaches [1600, 1681]. approximable [745, 1605]. approximate [2320]. Approximating [548, 763]. Approximation [312, 364, 389, 820, 824, 839, 845, 935, 992, 1091, 1154, 1158, 1160, 1253, 1309, 1335, 1626, 1820, 1913, 2032, 2046, 2157, 2259, 2535, 2554]. Approximations [81, 1089, 1116, 1591, 1600, 1925, 1968, 2029, 2329]. April [70, 263, 437, 676, 940, 1234, 1849, 2168, 2439]. Arabia [189]. Arakawa [292, 1052, 1266, 1352, 1970]. Arakelov [237, 1689]. Arbitrarily [622, 2209]. arbitrary [3, 107, 177, 787, 802, 1664, 2051, 2201, 2268, 2452]. Archimedean [1749, 2039, 2212, 2238, 2331]. area [1613, 2044, 2551]. areal [1287]. argument [358, 886, 2284, 2507]. arguments [773, 1233, 1291, 1988, 2073, 2147, 2261, 2336]. arising [72, 113, 114, 538, 585, 655, 661, 849, 1342, 1452, 1881, 2160, 2309, 2460]. Arithmetic [41, 69, 85, 131, 222, 228, 237, 267, 360, 370, 394, 398, 416, 426, 431, 461, 497, 506, 507, 719, 730, 758, 780, 841, 1023, 1092, 1109, 1123, 1166, 1242, 1255, 1298, 1307, 1314, 1456, 1478, 1576, 1635, 1696, 1758, 1798, 1815, 1853, 1893, 1950, 2048, 2123, 2130, 2137, 2240, 2307, 2434, 2473, 2484, 2543, 2547, 2579, 2602, 2623]. Arithmetical [1320, 1548, 1890, 2120, 2171]. Arithmeticity [1289]. arrangements [2062]. Arthur [262, 1327]. Artin [296, 521, 547, 694, 914, 1114, 1221, 1344, 1694, 1781, 1858, 1872, 1978, 2345, 2379, 2629]. ary [877, 1750, 2019, 2188]. aspect [320, 1040, 2575]. associated [29, 291, 348, 370, 380, 495, 539, 686, 730, 863, 914, 930, 1093, 1174, 1200, 1246, 1275, 1384, 1484, 1517, 1739, 1786, 1842, 1854, 1893, 1957, 2088, 2288, 2322, 2582]. associates [961]. associée [380, 863]. assumption [1822]. Asymptotic [54, 99, 257, 335, 526, 774, 872, 893, 917, 998, 1077, 1112, 1252, 1380, 1494, 1564, 1582, 1601, 1644, 1674, 1753, 1791, 1834, 1915, 2052, 2287, 2320]. Asymptotically [2148]. Asymptotics [200, 458, 559, 1621, 1622, 2490, 2520, 2647]. atom [1863]. attached [14, 41, 57, 101, 270, 423, 443, 694, 1185, 1358, 1415, 1938, 1948, 2059, 2237, 2278, 2338, 2555]. attempt [271]. August [137, 315, 501, 761, 1002, 1305, 1629, 1962, 2270, 2523]. automata [2231]. automatic [520, 2242]. automorphes [408]. Automorphic [68, 216, 219, 370, 408, 423, 555, 599, 621, 825, 826, 851, 978, 1099, 1140, 1184, 1228, 1380, 1509, 1520, 1675, 1809, 1865, 2338, 2456, 2471, 2506, 2576, 2634]. Automorphism [1768, 2476, 2618]. Automorphisms [827, 1533, 2448]. automorphy [1518]. aux [81, 380]. Average [187, 326, 361, 384, 477, 555, 662, 850, 892, 908, 1394, 1657, 1763, 1765, 2074, 2179, 2190, 2211, 2217, 2315, 2382, 2574, 2644]. Averages [258, 347, 362, 769, 1870, 1959, 2050, 2113, 2346, 2497]. Averaging [1097, 1226]. avoids [50]. away [112]. Ax [439].

backward [308]. bad [800]. Badly [745, 1605]. Baker [2112]. balanced [927, 2196]. ball [1791]. ballot [102]. Banach [146, 1557, 1695, 2408, 2431].

Baragar [262]. Barban [529]. Barnes [61, 893, 1915, 2502]. Baruah [1198]. base [40, 114, 855, 876, 983, 1671, 2140, 2161, 2322, 2559]. based [774, 812, 841, 846, 1112, 1248]. Bases [54, 548, 697, 716, 878, 1178, 1180, 1373, 1614, 1648, 1856, 1857, 1916, 2410, 2423, 2628]. basic [553, 1462, 1881]. Basis [64, 232, 245, 404, 494, 543, 635, 692, 722, 1281, 2612]. be [2179]. Beatty [544, 1385, 1754, 2319, 2390]. Becker [1912]. behavior [257, 361, 370, 850, 1939]. behaviors [1380]. behaviour [65]. Bell [708, 774, 1423, 1974]. Belyi [789, 898]. Bendersky [1089, 1591]. Berggren [2180]. Berkovich [2408]. Bernoulli [51, 180, 292, 414, 440, 954, 1052, 1200, 1288, 1296, 1374, 1412, 1453, 1593, 1944, 2145, 2368]. Bernstein [829]. Bernstein-blocks [829]. Berthelot [1663]. Bessel [920, 1136, 1783, 2198]. beta [1404, 1589, 2308]. beta-expansion [1589]. beta-expansions [2308]. between [2, 18, 57, 150, 227, 283, 374, 377, 488, 678, 696, 791, 838, 970, 1110, 1288, 1543, 1643, 1681, 1702, 1821, 1919, 2057, 2091, 2234, 2273, 2438, 2564, 2583]. Beurling [63, 225, 271, 549, 1619, 2253]. Beyond [342, 983, 1660, 2134, 2206]. BGG [338]. bi [64, 2513]. bi-basis [64]. bi-brackets [2513]. bias [357, 2248]. biased [1844]. Biases [88, 2175]. bicyclic [1982]. bidirectional [102]. Bielliptic [1996, 2183]. biextensions [536]. Bighash [2560]. Bilinear [1734, 2257]. billiard [474]. Binary [25, 62, 143, 339, 433, 852, 901, 1219, 1384, 1436, 1986, 2016, 2219, 2336, 2373]. binomial [23, 200, 208, 213, 242, 382, 425, 614, 804, 900, 1159, 1266, 1724, 1928, 1999, 2064, 2135, 2339, 2344]. binomials [830, 1151, 1341, 2083, 2257]. bipartition [1798]. bipartitions [972, 1242]. Biquadratic [962, 1061, 1873, 1982, 2599]. Birch [187, 214, 270, 2232]. bisectional [1405]. Bits [610, 1436]. Bivariate [1739]. Bloch [331, 766]. blocks [829, 1635]. Blueprints [1092]. Board [12, 27, 44, 71, 96, 109, 120, 138, 153, 171, 191, 206, 221, 234, 247, 264, 277, 289, 304, 316, 333, 346, 366, 386, 403, 422, 429, 438, 460, 473, 487, 502, 519, 533, 552, 574, 601, 626, 656, 677, 701, 718, 741, 762, 784, 810, 833, 857, 871, 889, 911, 941, 959, 969, 981, 1003, 1025, 1042, 1071, 1095, 1129, 1149, 1202, 1235, 1258, 1273, 1293, 1306, 1329, 1350, 1366, 1392, 1421, 1446, 1473, 1515, 1539, 1568, 1603, 1630, 1651, 1683, 1712, 1745, 1772, 1797, 1825, 1850]. Board [1884, 1918, 1936, 1963, 1994, 2014, 2034, 2082, 2095, 2115, 2143, 2169, 2194, 2221, 2247, 2271, 2290, 2314, 2333, 2356, 2377, 2399, 2418, 2440, 2463, 2482, 2505, 2524, 2550, 2573, 2595, 2625]. bodies [1407, 2637]. Bombieri [99, 1929]. bonnes [210]. Borcea [1300]. Borcherds [826, 1214]. Borne [380]. both [583, 1953]. bound [322, 380, 576, 689, 742, 886, 1197, 1338, 1373, 1759, 1916, 2119, 2285, 2379, 2400, 2628, 2630, 2635]. Bounded [145, 956, 1119, 1256, 1330, 1782, 1821, 2027, 2488]. Boundedness [528]. Bounding [248, 789, 848, 2340]. Bounds [63, 105, 117, 307, 468, 608, 673, 802, 816, 903, 923, 1039, 1062, 1255, 1406, 1463, 1579, 1698, 1702, 1920, 1931, 1934, 1969, 2085, 2201, 2250, 2258, 2266, 2282, 2288, 2326, 2330, 2362, 2385, 2529, 2566, 2574, 2620, 2640]. Bourgain [2357]. Bowman [527]. boxing [2610]. bracelet [1282]. brackets [2513]. Bradley [527]. branch [1975]. branches [2154]. Brauer [1179, 1577, 2109, 2366, 2593]. breaks [1026]. Bressoud [1716]. Breuil [553, 949, 1403, 1642]. bricks [796]. Briggs [1662]. Brocard [683]. broken [843, 890, 995, 1011, 1788]. Brumer [427, 663, 1029]. Büchi [1255]. Bugeaud [154]. buildings [19]. bundle [1929]. bundles [453]. Burgess [1579]. Burnside [935, 1112]. butterfly [2365].

Calabi [1300, 2160]. Calcul [2214]. calculation [1465]. calculations [2483]. Calculus [298, 812, 874, 1171, 1200, 1325]. Caliber [201]. can [2179]. Cancellation [163, 408]. Canonical [335, 578, 697, 885, 949, 1261, 1410, 2298, 2538, 2539, 2612, 2640]. canonique [1410]. canoniques [1318]. capacity [1431, 1808]. capitulation [2568]. Capparelli [2530]. caractéristique [1147]. cardinal [1476]. Carlitz [250, 440, 713, 739, 1049, 1105, 1581, 1585, 1903]. Carmichael [1155, 1632]. carré [820]. Cartan [218, 2405]. case [134, 199, 254, 399, 1163, 1642, 1645, 2300, 2464, 2576, 2617]. cases [368, 692, 1396, 1984]. Catalan [23, 151, 490, 1755, 2022]. Cauchy [1084, 1581]. center [643, 1417]. Centers [1206]. Central [182, 208, 320, 382, 468, 555, 557, 614, 684, 769, 1220, 1368, 1380, 1858, 1859, 2091, 2316, 2322, 2591, 2611, 2634]. centre [643]. Certain [52, 91, 169, 244, 254, 256, 261, 262, 268, 274, 348, 370, 423, 492, 576, 665, 725, 873, 901, 915, 917, 973, 1004, 1019, 1038, 1072, 1105, 1159, 1173, 1320, 1354, 1355, 1383, 1387, 1428, 1504, 1506, 1518, 1552, 1617, 1646, 1673, 1723, 1738, 1742, 1846, 1858, 1875, 1898, 1899, 1923, 1957, 1972, 2006, 2019, 2025, 2056, 2065, 2075, 2125, 2260, 2304, 2330, 2414, 2457, 2459, 2468, 2506, 2587, 2623, 2646]. Cesàro [2067, 2179]. chains [375, 517]. champion [2430]. Change [40, 887, 983, 1885, 2140, 2301, 2322, 2510, 2559]. changes [836, 1108, 1828, 1973, 2590]. Chaotic [1905]. Character [67, 69, 104, 580, 599, 987, 1326, 1374, 1501, 1509, 1579, 1702, 1714, 2000, 2017, 2328, 2472]. Characteristic [217, 283, 298, 331, 374, 445, 498, 528, 535, 539, 615, 650, 665, 686, 698, 710, 733, 837, 839, 1035, 1038, 1147, 1163, 1250, 1388, 1419, 1726, 1875, 1957, 2042, 2126, 2161, 2211, 2367, 2374, 2412, 2415, 2429]. Characterization [506, 722, 731, 1015, 1019, 1281, 1344, 1381, 1868, 1900, 2012, 2234, 2244]. Characterizations [1725]. characterized [376, 1399]. Characterizing [146]. characters [53, 76, 158, 661, 1206, 2352, 2353, 2583]. Chebotarev [662, 2432, 2522]. Chebyshev [276, 357, 549, 760, 1294, 1476, 1756, 2587]. Chen [31, 1760]. Chevalley [806, 1046]. Chief [1995]. Ching [1514]. Choi [1077]. choice [1885]. chosen [2062]. Chow [973]. Chowla [343, 736, 2345]. Christol [1818]. circle [2515]. circles [1852]. Circular [100, 274, 351, 475, 805, 878, 1584, 1607]. Class [1, 6, 21, 25, 33, 62, 73, 80, 92, 186, 218, 257, 296, 350, 390, 427, 436, 489, 507, 515, 559, 576, 664, 669, 675, 687, 725, 730, 736, 830, 831, 896, 902, 909, 932, 947, 1069, 1117, 1161, 1192, 1224, 1248, 1263, 1311, 1339, 1347, 1348, 1354, 1383, 1387, 1439, 1447, 1497, 1499, 1503, 1525, 1532, 1546, 1563, 1592, 1649, 1658, 1666, 1686, 1687, 1778, 1811, 1832, 1862, 1863, 1876, 1886, 1891, 1919, 1952, 1953, 1983, 1992, 2001, 2002, 2079, 2093, 2122, 2139, 2149, 2184, 2203, 2241, 2347, 2385, 2442, 2509, 2544, 2567, 2581, 2599, 2622]. Classe [6, 93, 1161]. Classes [98, 148, 215, 235, 291, 323, 355, 373, 377, 484, 516, 578, 643, 644, 685, 754, 877, 942, 1297, 1312, 1413, 1511, 1980, 2187, 2214, 2321, 2593]. Classical [238, 269, 641, 847, 1014, 1141]. Classification [127, 160, 970, 1347, 1639, 1889, 2036]. Clausen [1464, 1743]. cliques [1563]. close [2446]. closed [1150, 1233, 1291, 1453, 2108]. closed-form [1233, 1291]. closing [2206]. closure [42, 62, 260, 926, 2047]. closures [1919]. Clustering [2071]. CM [80, 90, 156, 195, 285, 515, 613, 721, 1090, 1300, 1413, 1938, 2262, 2371, 2420, 2582, 2622]. CM-extensions [1090]. CM-fields [195, 515, 613]. CM-type [90]. CM-types [195]. Coates [733]. Cocks [280]. cocycles [2616]. code [157, 1237]. codes [877]. coding [2311]. Coefficients [3, 18, 107, 178, 208, 257, 324, 355, 361, 381, 382, 415, 483, 485, 499, 599, 609, 623, 682, 850, 900, 901, 937, 950, 971, 1005, 1096, 1108, 1140, 1142, 1195, 1217, 1228, 1229, 1314, 1326, 1360, 1443, 1467, 1535, 1580, 1624, 1690, 1693, 1700, 1703, 1724, 1738, 1767, 1789, 1802, 1866, 1870, 1882, 1948, 1973, 2000, 2051, 2064, 2068, 2111, 2135, 2138, 2189, 2229, 2250, 2301, 2330, 2338, 2344, 2490, 2540, 2564, 2583, 2619]. cohomological [1895]. Cohomology [349, 690, 873, 1103, 2048, 2238]. Coincidences [737, 808]. Coleman [183, 2091]. Collatz [565]. Collatz-type [565]. Colmez [504, 2371]. color [1402, 1594]. colored [2088, 2167, 2230]. colors [1198]. combinations [1256, 2071, 2468]. Combinatorial [151, 258, 538, 1479, 1792, 1889]. combinatoric [1956]. common [1076, 1334, 1613, 1763, 1931, 1988, 2044, 2551, 2600]. commutative [91, 970, 1843, 2495, 2636]. commutativity [457]. commute [2458]. compact [383]. compactness [1894]. companion [883]. Comparison [2144]. complement [922]. Complementary [1141]. complements [1008, 2037, 2055]. Complete [160, 282, 330, 786, 1084, 1157, 1425, 1604, 1889, 1940, 2026]. Completely [1511, 2373, 2639]. completion [912]. completions [1158]. complex [141, 358, 662, 782, 1122, 1222, 1410, 1612, 1739, 1841, 1845, 2291, 2293, 2297]. complexes [1410, 1529]. complexity [2242]. components [158, 1361, 1469, 1470, 1661, 2293, 2449, 2580]. composite [143, 280, 834, 2637]. composition [266, 679]. Compositions [589]. compositum [351, 2637]. compound [2037]. Computation [136, 147, 158, 281, 674, 1369, 1404, 1625, 1648, 2020, 2152, 2198, 2204, 2214, 2521]. Computations [853, 965]. compute [631, 705, 803]. Computing [30, 279, 283, 286, 481, 582, 650, 987, 1851, 2382, 2570, 2571]. concavity [1285, 1404]. concerning [132, 641, 727, 798, 1014, 1021, 1337, 1359, 1379, 1627]. condition [1455, 2345]. Conditional [541, 1177, 1643, 2201]. conditions [74, 589, 1506, 1934, 2101]. Condolence [1995]. conductor [241, 653, 905, 1015, 1337, 1381, 2190, 2465]. cones [504, 1122]. configurations [1543, 1647]. conformal [1933]. Congruence [67, 115, 412, 523, 844, 979, 1000, 1073, 1331, 1455, 1506, 1517, 1658, 2041, 2251, 2265, 2359, 2469]. Congruences [57, 79, 132, 164, 168, 208, 299, 325, 368, 382, 418, 434, 451, 509, 571, 614, 678, 696, 708–710, 726, 727, 861, 880, 881, 989, 1011, 1030, 1059, 1207, 1229, 1242, 1282, 1299, 1402, 1530, 1594, 1622, 1631, 1637, 1652, 1677, 1750, 1788, 1800, 1806, 1869, 1877, 1904, 1922, 2008, 2072, 2091, 2100, 2103, 2188, 2230, 2274, 2421, 2493, 2500]. Congruent [622, 678, 2225, 2364, 2433]. conic [2539]. conjectural [663, 2022, 2234, 2267]. conjecture [115, 118, 179, 187, 207, 211, 214, 270, 294, 295, 301, 311, 349, 400, 427, 446, 448, 458, 489, 562, 565, 629, 640, 641, 715, 719, 733, 738, 753, 756, 766, 794, 795, 798, 944, 993, 1014, 1021, 1029, 1075, 1087, 1134, 1136, 1142, 1181, 1198, 1231, 1264, 1309, 1336, 1337, 1339, 1346, 1359, 1361, 1377– 1379, 1393, 1400, 1403, 1423, 1436, 1487, 1493, 1495, 1507, 1540, 1627, 1628, 1645, 1663, 1715, 1722, 1757, 1781, 1807, 1822, 1837, 1912, 1978, 2011, 2080, 2106, 2152, 2190, 2282, 2345, 2357, 2371, 2379, 2560, 2629, 2631]. Conjectures [25, 86, 295, 493, 509, 548, 1031, 1327, 1355, 1677, 2235, 2604]. conjointes [2252]. conjugacy [373, 2187]. conjugate [651, 985]. Connected [1361, 1690, 2018, 2293]. connecting [1384]. connection [202, 1288, 1548, 1707, 1722, 2501]. Connections [791]. Connes [39, 1707]. Consani [1707]. Consecutive [227, 354, 575, 598, 695, 1166, 1367, 1389, 1463, 1481, 1669, 1914, 2319, 2459, 2499]. consequence [1837]. consequences [960]. Constancy [314]. constant [209, 255, 356, 371, 602, 652, 774, 785, 928, 1154, 1233, 1291, 1315, 1353, 1566, 1582, 1601, 1644, 1654, 1834, 1843, 1925, 2052, 2208]. constants [166, 211, 673, 695, 763, 1077, 1089, 1113, 1395, 1397, 1443, 1475, 1591, 1662, 1882, 2181, 2282, 2603]. constructible [2458]. Constructing [102, 287, 728, 2222]. Construction [274, 285, 464, 929, 2552]. constructions [103, 1679, 2506]. containing [1330, 2444]. continuation [613, 2390]. Continued [77, 255, 401, 495, 511, 590, 652, 743, 807, 930, 975, 1117, 1124, 1152, 1154, 1210, 1250, 1276, 1278, 1342, 1353, 1540, 1589, 1591, 1595, 1617, 1679, 1691, 1732, 1905, 1925, 1933, 1956, 1977, 2065, 2066, 2157, 2176, 2239, 2264, 2277, 2450, 2518, 2648]. Continuous [2266]. convergence [1548, 2120, 2402]. convergent [928]. converse [1748]. convex [19, 418, 1055, 1803]. convolution [530, 766, 813, 872, 879, 954, 1479, 1729, 1890, 2113, 2171, 2324]. convolutions [714, 1537, 1776, 1895, 2497]. Cooper [1423]. coordinates [2588, 2589]. coprimality [589]. core [651, 861, 972, 985, 1011, 2100]. cores [1242]. corps [1279, 1407, 2637]. correction [1253]. correlated [1153]. correlation [1283, 1430, 1757, 2197, 2556]. correlations [520, 1898, 2104, 2138, 2477]. correspondence [203, 216, 334, 377, 535, 633, 799, 1165, 1611, 2511]. correspondences [243, 1054, 1145, 1422, 2045]. corresponding [239, 2349]. Corrigendum [93, 94, 106, 107, 302, 517, 550, 655, 808, 909, 939, 957, 1128, 1271, 1303, 1304, 1471, 1627, 1628, 1743, 1770, 1848, 1882, 2012, 2079, 2080, 2140, 2141, 2644–2649]. cotangent [1929]. count [2026]. countable [1429]. countably [1178, 2423]. counterexample [629, 738, 1069]. Counting [145, 184, 510, 933, 1553, 1661, 1782, 1867, 1892, 1942, 2097, 2224, 2279, 2354]. courbes [380]. course [278]. cousins [61]. covering [544, 827, 1835, 1975]. coverings [45, 2473]. covers [510, 1194, 1873, 2243, 2283, 2405]. crank [1215, 2167, 2294, 2302]. critère [6, 93]. Criteria [635, 1492, 1912]. criterion [6, 73, 92, 321, 1387, 1427, 1619, 1894, 2253, 2531, 2578]. Critical [10, 248, 306, 442, 513, 886, 905, 964, 991, 1232, 1304, 1616, 1640, 1660, 1676, 1684, 1770, 1895, 2258, 2516]. Cross [798, 918, 1398]. crucial [2039]. cryptography [278]. crystalline [2492]. crystals [582]. cubed [1389]. Cubes [454, 480, 495, 1501, 2030, 2547]. Cubic [4, 34, 76, 82, 149, 254, 302, 377, 405, 496, 504, 568, 607, 724, 753, 897, 913, 1205, 1271, 1280, 1371, 1508, 1658, 1678, 1684, 1692, 1706, 1717, 1725, 1751, 1770, 1800, 1817, 1848, 1907, 1952, 2073, 2129, 2157, 2279, 2302, 2305, 2425, 2517]. Cullen [583, 2569]. Cunningham [375, 517]. cup [1767]. curious [79, 1073, 1331]. Curvature [1709, 1710]. curve [236, 278, 279, 301, 372, 435, 465, 650, 685, 712, 827, 921, 1098, 1163, 1295, 1344, 1368, 1592, 1924, 2370, 2424]. curves [21, 40, 75, 105, 122, 160, 175, 198, 214, 218, 224, 237, 259, 280, 282– 284, 286, 287, 380, 405, 453, 484, 507, 547, 595, 612, 627, 636, 667, 699, 735, 770, 782, 801, 821, 846, 907, 908, 913, 955, 977, 990, 996, 997, 1110, 1118, 1169, 1212, 1319, 1345, 1363, 1390, 1394, 1395, 1413, 1415, 1469, 1505, 1512, 1521– 1524, 1609, 1618, 1623, 1686, 1737, 1747, 1752, 1845, 1897, 1910, 1991, 2087, 2096, 2137, 2140, 2183, 2190, 2200, 2238, 2262, 2283, 2300, 2309, 2366, 2369, 2375, 2381, 2387, 2401, 2405, 2460, 2511, 2528, 2562, 2565, 2611]. curves [169, 241, 244, 285, 390, 436, 449, 493, 790, 802, 887, 892, 951, 1054, 1194, 1241, 1277, 1301, 1383, 1483, 1517, 1819, 1878, 2129, 2210, 2279, 2321, 2587]. cusp [168, 182, 197, 320, 415, 417, 485, 525, 555, 850, 971, 989, 1019, 1040, 1142, 1227, 1362, 1467, 1535, 1587, 1624, 1690, 1735, 1828, 1851, 1866, 1870, 1948, 2138, 2248, 2250, 2301, 2388, 2391]. cuspforms [1289]. cuspidal [124, 825, 931, 1245, 1520, 1621, 2634]. cuspidality [229]. cusps [2255]. Cycle [1009]. cycles [1996, 2582]. Cyclic [165, 256, 413, 427, 510, 615, 648, 664, 862, 877, 878, 1205, 1470, 1706, 1717, 1751, 1876, 1924, 1985, 1991, 1992, 2059, 2079, 2129, 2449, 2465, 2544]. Cyclotomic [117, 150, 160, 178, 355, 433, 457, 545, 564, 638, 695, 760, 885, 919, 1004, 1049, 1105, 1263, 1274, 1438, 1441, 1580, 1724, 1768, 1878, 2249, 2369, 2525, 2566]. Cyclotomy [5, 1128, 1341].

Dagger [1557]. Daniels [1526]. dans [380, 2637]. Darboux [1911]. Darmon [203, 1087]. Dasgupta [203]. data [2415]. Davenport [529, 1315, 1826, 1843, 2058, 2112]. David [657, 1995, 2015]. death [1995]. December [205, 385, 573, 856, 1094, 1391, 1744, 2081, 2355, 2624]. decimal [1150]. decimals [687]. Decomposability [1244]. Decomposing [1390]. Decomposition [765, 772, 854, 1082, 1123, 1259, 1703, 1863, 1932, 2455, 2548]. decompositions [885, 1012]. Dedekind [144, 361, 426, 455, 503, 514, 543, 725, 919, 945, 952, 1045, 1086, 1199, 1374, 1544, 1606, 1722, 1805, 1826, 1840, 1864, 1921, 1998, 2266, 2447, 2601]. Dedekind-type [1722]. defined [770, 1370, 1413]. definite [1185, 2186]. definition [35, 1684, 1770, 2387]. definition/field [2400]. Deformation [310, 646, 1369, 2445, 2492, 2580]. Deformations [331, 2084, 2406]. Deformationsräumen [2580]. deformed [819]. Degenerate [1139, 1675]. degré [148, 1297]. Degree [145, 148, 325, 417, 443, 469, 478, 515, 572, 591, 632, 685, 789, 802, 823, 837, 956, 1036, 1204, 1243, 1250, 1260, 1297, 1321, 1552, 1698, 1789, 1971, 1978, 1996, 2051, 2479, 2480, 2583, 2591, 2593, 2597]. degrees [637]. del [1552, 2593]. Delannoy [396, 1721, 2123]. delicate [1728]. Deligne [1842, 2187]. Demailly [2282]. Demazure [1133]. Denominators [1236, 2343, 2447, 2519, 2641]. Dense [54, 1034, 1370, 2640]. denseness [2453]. densities [1911, 2425]. Density [116, 146, 267, 397, 454, 455, 560, 662, 851, 872, 1017, 1067, 1449, 1733, 1924, 1946, 2244, 2267, 2383, 2388, 2471, 2522, 2578]. Department [189]. depth [1143, 2162]. depth-zero [1143]. derangement [708]. derangements [1576]. derivations [1950]. Derivative [14, 1338, 1990, 2158]. derivatives [362, 722, 875, 1404, 1536, 2117, 2118, 2164]. derived [1283, 1430, 1730, 2023, 2141]. Descending [2124, 2493]. Descent [59, 1146, 1319, 2514]. descents [1491]. Description [1269]. designated [792, 1456, 2155]. designs [1020]. dessous [1561]. detection [2172]. determinant [1171, 1975, 2145, 2214]. Determinants [660, 1868, 2297]. Determination [320, 323, 330, 485, 1040, 1151, 1205, 1220, 1301]. determined [1694]. Determining [182]. Deterministically [1512]. Deutsch [1072]. development [930]. deviation [1382]. d’extensions [148, 643, 1297]. Diagonal [149, 915, 1097, 1158, 1577, 1978, 2358]. diagrams [2438]. Diamond [843, 869, 890, 957, 995, 1011, 1788]. Dickman [2018]. Dickson [1838, 2228, 2260, 2424]. difference [64, 131, 399, 526, 867, 2365, 2430, 2628]. Differences [267, 376, 470, 848, 999, 1153, 1172, 1322, 1399, 1586, 2038, 2057, 2199, 2499]. different [482, 1051, 1513, 1542, 1790, 2062]. Differential [228, 347, 463, 466, 585, 655, 907, 938, 1193, 1206, 1254, 1358, 1932, 2024, 2089]. differentiation [2380]. Digit [193, 425, 722, 1818, 2181]. digital [903, 923]. Digitally [1728]. digits [121, 240, 578, 1150, 1196, 1669, 2239, 2276, 2423]. dihedral [168, 296, 612, 1169]. dilated [1548]. dilating [1399]. dilogarithm [620, 1223]. dilogarithmic [2473]. dilogarithms [331, 2596]. Dimension [417, 689, 1533, 1590, 1691, 1872, 1885, 1933, 1945, 1975, 2076, 2214, 2429, 2618]. dimensional [32, 49, 133, 383, 426, 430, 604, 855, 923, 1484, 1487, 1579, 1593, 1609, 1956, 2086, 2304]. dimensions [217, 989, 1028, 2192, 2214, 2248]. Dinakar [485]. Diophantine [7, 47, 78, 189, 327, 336, 337, 341, 354, 364, 567, 594, 649, 697, 759, 814, 839, 845, 915, 925, 963, 1001, 1016, 1043, 1116, 1127, 1246, 1323, 1325, 1335, 1367, 1406, 1414, 1417, 1454, 1666, 1790, 1793, 1820, 1899, 1945, 2046, 2075, 2119, 2122, 2131, 2185, 2259, 2535, 2545, 2554, 2561, 2641]. Diophantus [534]. Dipendra [485]. Direct [777, 1218, 1303, 2435, 2457]. directed [1933]. Dirichlet [15, 35, 53, 63, 312, 489, 623, 872, 987, 1096, 1162, 1195, 1270, 1285, 1404, 1537, 1541, 1764, 1909, 1990, 2056, 2060, 2158, 2285, 2390, 2425, 2451, 2495, 2529, 2555, 2578, 2635]. Discrepancy [328, 561, 747, 903, 923, 1607, 1671]. Discrete [9, 106, 528, 628, 822, 823, 1058, 1333]. discriminant [254, 340, 539, 1384, 1975, 2010, 2016, 2480, 2539]. Discriminants [419, 1006, 1241, 1553, 1717, 2202, 2342]. discriminating [1507]. disjoint [654, 1635]. d’isogénie [1161]. d’isomorphisme [2214]. dispersed [2272]. distance [2, 154, 556, 793]. distances [1428, 1643, 1879, 2181]. Distinct [118, 732, 1222, 1481, 1879, 1892, 1926, 1937, 2207, 2365, 2499]. Distinction [1143, 2560]. distinguished [1005, 1409, 2559]. Distinguishing [2005, 2409]. distributed [835, 2244]. Distribution [236, 306, 388, 397, 465, 521, 541, 578, 588, 640, 787, 839, 867, 917, 927, 942, 996, 1044, 1057, 1194, 1221, 1248, 1308, 1362, 1368, 1417, 1457, 1459, 1511, 1558, 1616, 1701, 1774, 1854, 1896, 1907, 1913, 1924, 1941, 1991, 2027, 2043, 2116, 2451, 2471, 2507]. Distributions [1761, 1938, 2262, 2291, 2308]. divergent [1176]. Divisibility [126, 467, 811, 900, 937, 999, 1056, 1159, 1324, 1377, 1569, 1868, 1897, 1974, 2064, 2135, 2149, 2166, 2184, 2344, 2636]. divisible [905, 967, 1953]. division [476, 790, 1268, 1279, 1474, 2090, 2518]. divisor [63, 174, 344, 1334, 1727, 1986, 2003, 2153, 2293, 2346, 2347, 2372, 2477, 2497]. divisors [204, 465, 469, 794, 944, 1034, 1050, 1240, 1267, 1364, 1461, 1479, 1656, 1689, 1700, 1774, 1776, 1783, 1847, 2066, 2074, 2080, 2303, 2379, 2525]. DLW [2235]. Does [1193]. Dolbeault [2238]. domain [35, 1244, 2036]. domaine [35]. domains [952, 1289, 1606, 1720, 1921, 2325, 2601]. dots [1282]. Double [29, 45, 605, 642, 748, 1032, 1213, 1665, 1702, 1787, 1812, 1975, 2292, 2324, 2386, 2405, 2500, 2557, 2574, 2649]. double-sum [2292]. doubles [1588]. doubling [1401]. Dress [1434]. Drinfel’d [38, 105, 260, 311, 324, 498, 539, 660–662, 702, 723, 749, 953, 986, 1005, 1245, 1393, 1410, 1495, 1631, 1830, 2010, 2028, 2268, 2280, 2348]. dual [799, 1145, 1872, 2487]. Duality [36, 505, 570, 1137, 1360, 1424, 2238, 2403, 2617]. Duality/sum [2403]. d’un [820, 1147]. d’une [35]. DVR [1833]. Dwork [2026]. Dyer [187, 214, 270]. Dynamical [276, 587, 635, 794, 1179, 1378, 1628, 1756, 2400, 2552]. Dynamics [398, 432, 1225, 2039, 2601]. dynatomic [2245].

each [456]. edge [1616, 2258]. Editor [1995]. Editor-in-Chief [1995]. Editorial [12, 27, 44, 71, 96, 109, 120, 138, 153, 171, 191, 206, 221, 234, 247, 264, 277, 289, 304, 316, 333, 346, 366, 386, 403, 422, 429, 438, 460, 473, 487, 502, 519, 533, 552, 574, 601, 626, 656, 677, 701, 718, 741, 762, 784, 810, 833, 857, 871, 889, 911, 941, 959, 969, 981, 1003, 1025, 1042, 1071, 1095, 1129, 1149, 1202, 1235, 1258, 1273, 1293, 1306, 1329, 1350, 1366, 1392, 1421, 1446, 1473, 1515, 1539, 1568, 1603, 1630, 1651, 1683, 1712, 1745, 1772, 1797, 1825, 1850]. Editorial [1884, 1918, 1936, 1963, 1994, 2014, 2034, 2082, 2095, 2115, 2143, 2169, 2194, 2221, 2247, 2271, 2290, 2314, 2333, 2356, 2377, 2399, 2418, 2440, 2463, 2482, 2505, 2524, 2550, 2573, 2595, 2625]. Edwards [282, 484]. effect [236, 2380]. Effective [2, 154, 224, 321, 576, 845, 1040, 2330, 2454, 2610, 2640]. efficient [2610]. Egyptian [339, 452]. EGZ [1566]. EGZ-constant [1566]. Eichler [1527, 1984, 2048, 2391]. eigencurve [883]. Eigenform [950, 2597]. eigenforms [57, 755, 1440, 1789, 2324]. eigenspaces [1248]. eigenvalue [1828, 1851]. eigenvalues [368, 388, 461, 524, 572, 1051, 1761, 2469, 2497, 2510]. eigenvariety [464]. eight [591]. Eisenstein [167, 229, 412, 478, 525, 572, 591, 617, 905, 1064, 1124, 1138, 1139, 1183, 1243, 1451, 1534, 1696, 1703, 2072, 2106, 2255, 2263, 2296, 2360, 2416, 2468, 2483, 2513, 2537, 2608, 2616]. element [1132, 1531]. Elementary [126, 130, 158, 465, 531, 670, 890, 924, 1462, 1468, 1742, 2369, 2484]. elements [28, 29, 115, 522, 728, 1105, 1488, 1502, 2057, 2237, 2256, 2374, 2408, 2612]. eleven [1437]. Elliott [916]. ellipsoid [1417]. Elliptic [21, 40, 105, 160, 214, 224, 236, 278– 280, 283, 284, 301, 372, 380, 405, 435, 449, 465, 467, 507, 595, 612, 624, 627, 636, 672, 712, 724, 729, 735, 782, 790, 801, 821, 880, 887, 906, 908, 913, 942, 951, 977, 990, 997, 1169, 1183, 1212, 1223, 1241, 1301, 1368, 1383, 1394, 1395, 1413, 1415, 1432, 1505, 1521, 1522, 1609, 1618, 1623, 1686, 1737, 1747, 1752, 1812, 1845, 1878, 1897, 1910, 1924, 2087, 2096, 2134, 2137, 2140, 2190, 2262, 2300, 2309, 2315, 2321, 2366, 2369, 2370, 2375, 2381, 2382, 2460, 2528, 2531, 2562, 2565, 2596, 2611, 2623]. elliptic [169, 197, 348, 493, 892, 1295, 1614, 2129, 2202]. elliptiques [380]. embedding [557, 1107, 1945]. embeddings [776, 2361]. Endomorphism [244, 279, 846, 1013, 2571]. Endomorphisms [1119, 1295]. endoscopy [342, 983, 1134, 2134]. energies [2098]. Engel [1028, 1382, 1617, 2192]. ensemble [581, 775]. ensembles [2215]. entières [2215]. entire [1156, 1388]. entre [2, 154]. entry [590]. enumeration [72, 750, 1971, 2188]. envelope [931]. enveloping [1443, 1882]. epsilon [524]. Epstein [306, 344, 1427]. Equal [447, 535, 930, 948, 1163, 1364, 1389, 1954]. equation [7, 110, 113, 114, 123, 156, 162, 189, 293, 327, 567, 606, 649, 683, 759, 925, 930, 938, 1001, 1016, 1043, 1125, 1193, 1321, 1335, 1359, 1367, 1454, 1571, 1627, 1829, 2131, 2136, 2191, 2216, 2227, 2329, 2393]. equations [4, 16, 47, 78, 253, 336, 337, 341, 354, 377, 495, 585, 632, 655, 780, 818, 864, 915, 961, 1157, 1246, 1271, 1294, 1323, 1414, 1496, 1635, 1666, 1790, 1899, 2042, 2060, 2075, 2089, 2119, 2122, 2150, 2161, 2545, 2567, 2588, 2589]. Equidistribution [224, 945, 1784, 1840, 2039, 2104, 2248]. equivalence [85, 441, 1681, 1889]. equivalent [1609, 1764, 2345]. Equivariant [347, 1822, 1950]. Erdos [211, 295, 356, 477, 556, 793, 982, 1264, 1393, 1436, 2208, 2232, 2349, 2395, 2470, 2533]. Ergodicity [722, 788, 807, 1281]. Erratum [1291, 2548, 2643]. error [387, 541, 1375, 1510, 1865, 2097, 2381]. Escott [1906]. Estes [1493]. estimate [139, 630, 1487, 2388]. estimated [1839]. Estimates [149, 541, 549, 1017, 1217, 1343, 1353, 1425, 1555, 1643, 1964, 2116, 2144, 2253, 2484]. Estimating [209]. estimation [399]. eta [901, 1285, 1404, 1439, 2219, 2490]. eta-quotients [901, 1439, 2219, 2490]. étale [827, 1951, 2109, 2638]. étale-Brauer [2109]. Euclid [1633]. Euclidean [430, 618, 779, 946, 1020, 2216, 2599]. Euler [132, 211, 212, 300, 451, 490, 585, 593, 602, 652, 655, 659, 695, 765, 869, 899, 928, 957, 1086, 1087, 1113, 1154, 1175, 1239, 1443, 1464, 1544, 1575, 1601, 1636, 1644, 1662, 1743, 1755, 1853, 1882, 1888, 1925, 1970, 1987, 2052, 2078, 2178, 2205, 2284, 2298, 2367, 2441, 2501, 2563, 2603, 2645]. Evaluation [548, 1045, 1233, 1291, 1355, 1729]. evaluations [975]. even [383, 399, 748, 773, 1091, 1215, 1251, 1664, 1685, 1937, 1975, 2059, 2170, 2211, 2214, 2261, 2447, 2448]. even-indexed [1685]. Every [469, 2320, 2553]. everywhere [790, 1214, 1522].

Exact [359, 1036, 1192, 1965]. exactly [2410]. example [30]. Examples [1074, 1127, 1403, 1485, 1670, 1929, 2585]. exceeding [1955, 2323]. Exceptional [188, 1110, 1145, 1447, 1664, 1863, 1940, 2192, 2286, 2300, 2529]. Existence [778, 1713, 1893, 2175, 2374]. existential [1388]. expanding [2201]. expansion [13, 526, 652, 729, 998, 1589, 1617, 1669, 1816, 1979, 2065, 2076]. Expansions [242, 463, 743, 749, 774, 807, 893, 1028, 1077, 1178, 1245, 1252, 1375, 1382, 1429, 1443, 1474, 1494, 1510, 1564, 1582, 1593, 1595, 1601, 1679, 1753, 1801, 1856, 1867, 1882, 1890, 1915, 2052, 2124, 2171, 2192, 2255, 2308, 2339, 2368, 2402, 2410, 2423, 2496, 2554]. Explicit [103, 243, 271, 455, 511, 673– 675, 703, 723, 921, 975, 1120, 1197, 1403, 1425, 1495, 1555, 1652, 1714, 1737, 1759, 1861, 2116, 2121, 2258, 2285, 2385, 2483, 2511, 2522, 2529, 2620, 2635]. exponent [89, 892, 1862, 1966]. Exponential [78, 126, 163, 196, 409, 415, 547, 703, 772, 934, 945, 984, 997, 1048, 1057, 1079, 1083, 1093, 1216, 1316, 1414, 1478, 1507, 1624, 1672, 1702, 2029, 2098, 2119, 2223, 2257, 2362, 2443, 2500, 2540]. exponents [836, 839, 1276, 2185]. expressed [2198]. expressing [629, 738]. expression [1668, 1680]. expressions [2145, 2219]. extended [819, 2049, 2442]. extendibility [1127]. Extension [35, 230, 720, 887, 1441, 2204, 2452]. Extensions [33, 34, 45, 91, 98, 117, 148, 183, 226, 296, 302, 323, 357, 457, 507, 521, 536, 559, 570, 643, 644, 654, 746, 837, 847, 859, 877, 914, 924, 951, 1029, 1090, 1104, 1174, 1221, 1263, 1265, 1297, 1312, 1438, 1513, 1670, 1692, 1723, 1833, 1837, 1858, 1902, 1942, 2006, 2087, 2124, 2146, 2203, 2236, 2237, 2249, 2347, 2369, 2458, 2489, 2592]. extensive [843]. Exterior [967, 2212]. extra [2165]. extraordinary [734]. Extremal [420, 523, 566, 1533, 1618, 1845, 2170, 2609]. Extreme [2038, 2516].

F. [666]. facteurs [2252]. factor [249, 1010, 1302, 1463]. factorial [56, 1196, 2205]. factorization [798, 1720]. factorizations [917]. Factors [23, 524, 622, 685, 770, 1141, 1267, 1506, 1520, 1790, 1947, 1967, 2030, 2116, 2207, 2252, 2426, 2466, 2546, 2615]. Fadwa [189]. faithful [2639]. Falconer [556, 793]. false [2292]. Faltings [1598]. Families [87, 103, 224, 544, 612, 622, 645, 812, 961, 1093, 1169, 1242, 1277, 1345, 1611, 1706, 1741, 1779, 1788, 1795, 2162, 2433, 2443, 2494]. family [4, 46, 77, 200, 297, 377, 449, 576, 743, 996, 1004, 1048, 1195, 1271, 1363, 1376, 1531, 1747, 1953, 2018, 2089, 2091, 2125, 2245, 2421, 2460]. Fano [753]. Farey [1249]. Fast [136, 283, 705, 1404, 1625, 2521]. Faster [281, 1253]. February [26, 233, 421, 625, 888, 1148, 1445, 1796, 2114, 2398]. Fekete [176]. Fekete-like [176]. Fermat [159, 685, 1173, 1238]. fermionic [819]. Ferrero [965]. fewnomials [307]. Feynman [2407, 2438]. Fibonacci [157, 252, 506, 737, 803, 808, 1237, 1302, 1367, 1574, 2075, 2108, 2275, 2310, 2589]. Fibonacci-type [2108]. field [34, 42, 134, 199, 279, 291, 302, 323, 364, 404, 445, 465, 496, 559, 586, 587, 653, 668, 675, 690, 699, 728, 805, 895, 905, 912, 919, 1047, 1103, 1107, 1192, 1194, 1236, 1265, 1268, 1274, 1279, 1312, 1339, 1413, 1425, 1441, 1503, 1531, 1546, 1584, 1649, 1658, 1684, 1689, 1696, 1734, 1770, 1781, 1876, 1902, 1907, 1924, 1992, 2036, 2057, 2079, 2087, 2161, 2185, 2226, 2347, 2387, 2400, 2464, 2476, 2571]. field-of-definition [2400]. field-of-definition/field-of-moduli [2400]. field-of-moduli [2400]. Fields [33, 47, 48, 80, 85, 88, 91, 98, 128, 150, 198, 217, 265, 351, 357, 412, 484, 504, 515, 516, 535, 577, 608, 618, 650, 658, 659, 663– 665, 695, 733, 739, 811, 837, 865, 907, 933, 967, 1013, 1031, 1055, 1061, 1122, 1123, 1222, 1336, 1339, 1354, 1390, 1400, 1505, 1540, 1546, 1698, 1706, 1752, 1838, 1857, 1862, 1864, 1874, 1876, 1879, 1891, 1938, 1953, 1964, 2079, 2117, 2184, 2190, 2216, 2228, 2335, 2350, 2367, 2369, 2387, 2419, 2427, 2429, 2452, 2460, 2480, 2484, 2485, 2566, 2570, 2576, 2585, 2592, 2599, 2616, 2622, 2643]. fields [3, 21, 28, 40, 45, 107, 112, 144, 156, 160, 195, 201, 203, 226, 296, 305, 310, 318, 377, 390, 405, 446, 494, 522, 556, 564, 570, 576, 607, 612, 638, 669, 674, 685, 725, 750, 828, 830, 885, 896, 951, 996, 1045, 1060, 1069, 1098, 1105, 1151, 1169, 1224, 1347, 1348, 1376, 1383, 1435, 1441, 1462, 1497, 1512, 1532, 1534, 1609, 1667, 1670, 1678, 1681, 1687, 1694, 1723, 1740, 1751, 1768, 1778, 1813, 1818, 1873, 1878, 1919, 1960, 1980, 1982, 1989, 1991, 2001, 2067, 2132, 2139, 2140, 2149, 2201, 2237, 2276, 2371, 2374, 2375, 2414, 2442, 2448, 2465, 2508, 2618, 2646]. fields [177, 186, 431, 448, 476, 613, 675, 686, 707, 790, 793, 874, 878, 947, 962, 1038, 1295, 1499, 1500, 1503, 1892, 1971, 2058, 2129, 2305, 2466, 2486, 2544]. fifth [360]. figurate [948]. Filling [2479]. Filtered [239, 1590]. filtrations [1723]. Finding [280, 1433]. Fine [904, 1949, 2225]. Finite [10, 72, 165, 184, 198, 217, 265, 279, 347, 356, 441, 484, 494, 496, 528, 535, 556, 614, 685, 686, 699, 750, 793, 830, 839, 874, 918, 922, 931, 933, 952, 996, 1013, 1038, 1060, 1098, 1104, 1151, 1194, 1210, 1304, 1373, 1376, 1390, 1398, 1425, 1431, 1512, 1566, 1606, 1670, 1768, 1818, 1833, 1838, 1854, 1873, 1879, 1890, 1892, 1916, 1921, 1951, 1960, 1971, 1985, 1991, 2038, 2057, 2058, 2087, 2088, 2146, 2171, 2185, 2204, 2228, 2276, 2320, 2323, 2325, 2341, 2374, 2429, 2443, 2484, 2514, 2534, 2536, 2571, 2601, 2618]. First [124, 245, 547, 752, 859, 1208, 1233, 1291, 1535, 1965, 2053, 2158, 2228, 2301, 2510, 2557, 2649]. Fishburn [1530]. Fitting [896]. five [686, 1953, 2547]. fixed [145, 1466, 1571, 1685]. flagged [2492]. flat [104, 310, 528, 535]. Fleck [844]. Fluctuations [198]. fold [982, 1357, 2582]. folds [973]. Fomenko [715]. fonction [81, 863, 1147, 2252]. fonctions [408, 1407]. Ford [1852]. Form [47, 74, 111, 173, 961, 971, 1009, 1204, 1233, 1291, 1608, 1727, 1815, 1899, 1948, 2092, 2274, 2442, 2518, 2606]. formal [49, 823, 1443, 1882]. formed [1703]. forms [25, 52, 168, 245, 324, 381, 419, 466, 540, 609, 616, 660, 692, 696, 704, 716, 721, 745, 768, 853, 880, 882, 883, 901, 902, 907, 929, 985, 994, 1005, 1022, 1051, 1072, 1085, 1097, 1140, 1152, 1185, 1219, 1227, 1254, 1284, 1316, 1358, 1416, 1423, 1442, 1467, 1471, 1484, 1489, 1527, 1535, 1586, 1587, 1624, 1664, 1673, 1734, 1735, 1738, 1748, 1758, 1792, 1859, 1866, 1870, 1927, 1961, 1978, 2016, 2024, 2036, 2073, 2163, 2180, 2182, 2189, 2196, 2250, 2255, 2273, 2278, 2295, 2301, 2330, 2336, 2348, 2359, 2434, 2498, 2506, 2526, 2527, 2555, 2564, 2590, 2597, 2612, 2614]. forms [3, 8, 55, 94, 107, 124, 136, 182, 183, 197, 251, 270, 320, 325, 370, 397, 404, 410, 415, 417, 443, 483, 485, 513, 523, 525, 550, 555, 599, 678, 682, 711, 749, 845, 850, 852, 885, 916, 932, 939, 953, 966, 977, 983, 989, 1019, 1032, 1040, 1054, 1081, 1099, 1100, 1108, 1142, 1220, 1245, 1360, 1384, 1415, 1453, 1478, 1509, 1532, 1554, 1583, 1652, 1690, 1697, 1736, 1784, 1828, 1830, 1851, 1860, 2019, 2051, 2069, 2099, 2111, 2125, 2138, 2144, 2160, 2236, 2248, 2257, 2268, 2288, 2322, 2358, 2363, 2373, 2383, 2388, 2391, 2455, 2548, 2608]. formula [39, 66, 271, 475, 476, 647, 663, 680, 688, 711, 751, 844, 897, 935, 939, 992, 1112, 1132, 1172, 1227, 1261, 1374, 1375, 1427, 1510, 1544, 1558, 1564, 1680, 1785, 1791, 1811, 1861, 1880, 2010, 2026, 2067, 2093, 2099, 2108, 2121, 2287, 2348, 2386, 2452, 2466, 2485, 2542, 2557, 2579, 2626, 2649]. Formulae [906, 1559, 1794, 1864, 2078, 2300]. formulas [135, 392, 417, 548, 703, 723, 736, 746, 822, 824, 1082, 1156, 1191, 1424, 1593, 1674, 1804, 1988, 2251, 2261, 2341, 2351, 2403, 2420, 2518, 2632]. four [165, 454, 636, 759, 862, 1182, 1246, 1552, 1930, 2593]. four-square [1930]. Fourier [3, 72, 107, 415, 463, 483, 499, 599, 609, 682, 729, 850, 901, 1108, 1138, 1140, 1142, 1229, 1249, 1360, 1467, 1535, 1593, 1624, 1690, 1693, 1703, 1789, 1870, 1948, 2051, 2111, 2138, 2255, 2301, 2330, 2334, 2338, 2490, 2540, 2619]. fourth [25, 238, 317, 647, 684, 703, 1490, 1762, 1865, 1908, 2106]. fractal [1448]. fractals [745]. fraction [13, 590, 652, 743, 807, 930, 975, 1124, 1154, 1278, 1353, 1589, 1591, 1595, 1617, 1679, 1732, 1925, 2065, 2518]. fractional [1017, 1088, 1283, 1430]. fractions [77, 255, 339, 401, 452, 495, 511, 747, 1117, 1152, 1210, 1249, 1250, 1276, 1342, 1540, 1691, 1905, 1933, 1956, 1977, 2066, 2157, 2176, 2239, 2264, 2277, 2450, 2648]. framed [1369, 2492]. Frames [579]. framework [2267]. Franel [791]. free [173, 631, 895, 920, 1120, 1286, 1411, 1502, 1775, 1815, 2116, 2190, 2208, 2404, 2609]. Freeness [2412]. Freiman [2396]. French [2, 6, 35, 81, 148, 210, 380, 408, 643, 778, 820, 863, 1106, 1147, 1161, 1279, 1297, 1407, 1410, 1638, 2022, 2028, 2214, 2215, 2252, 2637]. friable [1718]. friendly [287]. Frobenius [391, 436, 512, 585, 655, 756, 990, 1198, 1356, 1464, 1469, 1551, 1609, 1653, 1743, 1794, 1945, 2187, 2218, 2256, 2262]. Frobenius-modules [1653]. Fubini [1877]. Fuchs [780, 982]. Fuchsian [124]. Full [458, 464, 987, 1719, 1878, 2304, 2308]. fully [1161]. Function [9, 28, 46–48, 63, 85, 88, 106, 185, 202, 212, 227, 291, 296, 334, 342, 357, 359, 362, 412, 432, 442, 458, 465, 488, 584, 590, 631, 640, 650, 658, 668, 669, 739, 747, 766, 767, 816, 855, 863, 893, 905, 944, 990, 1031, 1047, 1058, 1156, 1164, 1209, 1217, 1252, 1256, 1285, 1338, 1348, 1400, 1494, 1546, 1558, 1564, 1601, 1626, 1640, 1684, 1687, 1731, 1734, 1755, 1770, 1783, 1816, 1854, 1862, 1864, 1896, 1901, 1924, 1964, 1979, 2010, 2018, 2036, 2080, 2117, 2197, 2204, 2267, 2274, 2302, 2367, 2395, 2419, 2461, 2464, 2507, 2563, 2566, 2583, 2643]. function [37, 45, 81, 144, 174, 186, 199, 251, 305, 306, 318, 344, 361, 369, 387, 476, 490, 512, 531, 559, 564, 578, 591, 602, 613, 638, 647, 675, 828, 875, 886, 895, 919, 951, 983, 991, 992, 998, 1069, 1105, 1112, 1147, 1193, 1229, 1240, 1267, 1278, 1280, 1308, 1324, 1347, 1411, 1462, 1476, 1500, 1534, 1548, 1573, 1600, 1625, 1649, 1696, 1701, 1727, 1753, 1768, 1779, 1861, 1885, 1915, 2029, 2052, 2078, 2164, 2252, 2269, 2303, 2347, 2375, 2414, 2471, 2476, 2554, 2607, 2645, 2646]. Functional [110, 113, 114, 156, 1035, 1285, 1764, 1829, 2060, 2161]. Functions [69, 100, 245, 292, 301, 312, 326, 347, 384, 492, 495, 503, 516, 645, 665, 668, 724, 729, 826, 861, 876, 1031, 1044, 1091, 1093, 1114, 1115, 1120, 1133, 1135, 1156, 1164, 1170, 1172, 1183, 1232, 1266, 1290, 1345, 1355, 1368, 1386– 1388, 1394, 1396, 1397, 1404, 1416, 1427, 1433, 1437, 1488, 1511, 1525, 1544, 1641, 1705, 1726, 1764, 1844, 1865, 1880, 1890, 1893, 1898, 1904, 1920, 1938, 1970, 1972, 1973, 2004, 2117, 2118, 2120, 2121, 2158, 2161, 2171, 2211, 2240, 2278, 2285, 2293, 2334, 2380, 2425, 2437, 2500, 2508, 2528, 2529, 2575, 2583, 2616, 2621, 2626, 2635]. functions [40, 65, 68, 110, 135, 182, 204, 216, 219, 222, 249, 268, 309, 320, 335, 348, 358, 370, 378, 392, 393, 407, 408, 414, 445, 455, 468, 555, 584, 586, 593, 621, 671, 684, 730, 736, 742, 769, 781, 788, 836, 869, 884, 906, 957, 978, 1017, 1045, 1057, 1081, 1086, 1126, 1162, 1184, 1228, 1251, 1281, 1301, 1333, 1380, 1407, 1477, 1498, 1573, 1693, 1694, 1702, 1742, 1766, 1783, 1802, 1858, 1909, 1922, 1957, 1959, 2066, 2071, 2096, 2140, 2201, 2212, 2223, 2243, 2258, 2266, 2288, 2300, 2316, 2322, 2331, 2346, 2351, 2386, 2388, 2390, 2446, 2451, 2471, 2477, 2502]. functions [475, 588, 597, 851, 858, 1052, 1352, 1548, 1555, 1614, 1616, 1636, 1798, 1809, 2002, 2021, 2038, 2061, 2202, 2342, 2428, 2540, 2582, 2591]. functor [1545]. Functorial [243, 2413]. functoriality [1621]. functors [1369]. Fundamental [300, 504, 644, 947, 1491, 1658, 2111]. Fundamentaltheorem [1485]. Fürstenberg [62]. Further [1485].

G.C.D. [2327]. gain [2146]. Gál [1660]. Gál-type [1660]. Gallagher [2206]. Galois [34, 49, 51, 57, 98, 134, 177, 260, 302, 310, 321, 357, 380, 456, 457, 564, 643, 654, 661, 704, 707, 734, 821, 837, 877, 894, 1006, 1029, 1047, 1103, 1131, 1165, 1190, 1205, 1225, 1337, 1345, 1491, 1505, 1513, 1518, 1585, 1609, 1810, 1846, 1919, 1942, 1943, 2005, 2006, 2047, 2084, 2101, 2133, 2146, 2204, 2225, 2245, 2256, 2283, 2406, 2445, 2492, 2586, 2592, 2617, 2633]. galoisiennes [643]. Gamburd [2357]. games [1605, 1695]. Gamma [9, 29, 81, 106, 135, 249, 309, 512, 590, 736, 858, 869, 957, 990, 992, 998, 1112, 1141, 1209, 1252, 1386, 1494, 1558, 1564, 1600, 1626, 1753, 1766, 1913, 2386, 2502]. Gan [1645]. gap [545, 2166, 2206, 2479]. gaps [227, 488, 1286, 1543, 1580, 1702, 1821]. Garcia [1508]. Gauss [592, 611, 743, 897, 1203, 1427, 1801, 2021, 2058, 2402]. Gaussian [392, 494, 801, 844, 1125, 1724, 1821, 2450]. gcd [598, 1660, 1718, 1959]. GCD-s [598]. gcd-sum [1959]. GCF [13, 2076]. Gelfand [194]. Gelfond [89]. genera [564, 932, 1980]. General [399, 829, 930, 1087, 1112, 1170, 1333, 1389, 1490, 1762, 1791, 1811, 2023, 2141, 2177, 2284, 2475, 2497, 2632]. generalisation [85]. generalisations [2372]. Generalised [63, 1189, 2413]. Generalization [31, 36, 39, 135, 223, 232, 401, 444, 562, 563, 841, 844, 1065, 1087, 1238, 1239, 1408, 1427, 1433, 1560, 1619, 1689, 1697, 1781, 2112, 2349, 2359, 2386, 2392, 2603]. Generalizations [641, 715, 765, 1014, 1562, 2364]. Generalized [225, 253, 257, 268, 292, 293, 341, 414, 470, 475, 549, 556, 558, 713, 729, 736, 737, 793, 803, 808, 812, 836, 914, 945, 974, 1006, 1063, 1086, 1113, 1115, 1198, 1233, 1276, 1291, 1294, 1302, 1367, 1418, 1423, 1443, 1461, 1480, 1486, 1495, 1524, 1572, 1620, 1662, 1671, 1691, 1826, 1837, 1840, 1864, 1868, 1882, 1908, 1944, 2050, 2061, 2066, 2075, 2133, 2167, 2199, 2200, 2306, 2428, 2501, 2542, 2569]. Generalizing [1012]. generate [631, 1173]. generated [687, 935, 992, 1205, 1222, 1752, 1929, 2335, 2393, 2537]. Generating [543, 1240, 1512, 2059, 2066, 2103, 2528]. Generation [1687]. Generators [17, 522, 568, 648, 712, 748, 966, 1354, 1886]. Generic [547, 662, 772, 984, 1048, 1409, 1645, 1672, 2072, 2318, 2467]. Genocchi [2629]. Genus [55, 75, 203, 285, 286, 410, 550, 559, 770, 846, 1319, 1523, 1592, 1598, 2045, 2200, 2210, 2370, 2375]. Genus-correspondences [2045]. genus-two [1598]. geodesic [1977, 2474, 2558]. Geometric [1286, 1428, 1929, 2406]. geometrically [973]. Geometry [508, 1557, 1894, 1956, 2172, 2450]. German [2580]. Gilbreath [400]. Ginzburg [356, 2208, 2467, 2470]. Girard [1156]. given [933, 1039, 1050, 1578, 1692, 1773, 1955, 2323]. Glaisher [763, 774, 1077, 1089, 1582, 1591]. global [47, 357, 476, 570, 618, 644, 675, 733, 811, 1135, 1407, 1489, 1500, 1546, 1639, 1734, 1897, 2036, 2205, 2367, 2506]. globally [1929]. Godement [2331]. Goldbach [20, 41, 188, 273, 317, 1102, 1309, 1547, 1733, 1958, 2174]. Goldring [432]. Göllnitz [495]. good [210, 790, 1522, 2085, 2562]. Gordon [495]. Goss [85, 671, 1995, 2015]. grade [367]. graded [966]. Graham [118]. graph [1223, 1633, 1933, 2273, 2334, 2407]. Graphs [686, 1764, 2491]. greatest [1334, 1463]. Green [689, 826, 2582]. Greenberg [1837]. GRH [2522]. Gross [674, 1327, 1645, 2420]. group [25, 389, 435, 456, 464, 496, 528, 535, 627, 643, 690, 728, 805, 829, 862, 878, 931, 960, 1074, 1106, 1206, 1248, 1339, 1484, 1577, 1696, 1704, 1807, 1842, 1862, 1876, 1921, 1992, 2048, 2079, 2090, 2209, 2320, 2323, 2469]. group-invariant [1206, 1807]. groupe [643, 1106]. groupes [1638]. groupoid [1491]. Groups [1, 10, 21, 59, 72, 111, 124, 156, 165, 167, 169, 184, 197, 216, 243, 258, 296, 313, 331, 352, 356, 373, 390, 394, 427, 457, 505, 546, 564, 592, 615, 648, 664, 796, 806, 811, 815, 835, 866, 896, 904, 909, 918, 970, 973, 1006, 1046, 1090, 1099, 1140, 1141, 1146, 1173, 1248, 1304, 1398, 1409, 1419, 1422, 1468, 1504, 1512, 1519, 1520, 1531, 1566, 1639, 1659, 1740, 1768, 1778, 1811, 1924, 1966, 1982, 1983, 1985, 2023, 2042, 2048, 2087, 2102, 2133, 2141, 2165, 2187, 2198, 2225, 2236, 2245, 2367, 2442, 2457, 2486, 2559, 2576, 2592, 2615, 2618, 2622, 2623, 2636]. Growth [87, 465, 779, 938, 1117, 1176, 1595, 1878, 2122, 2509, 2558]. GSp [1161]. Guo [1136]. Guralnick [1493, 2424].

H [1007]. Hadamard [367, 1126]. Halberstam [529]. Half [3, 107, 610, 678, 711, 939, 1527, 1588, 1738, 2278, 2363, 2614]. Half-integral [678, 711, 939, 1527, 1738, 2278, 2363, 2614]. Hamahata [1970]. Hamme [1677]. Hammersley [1671]. Hamming [877]. Hankel [66, 660, 1171]. Hankel-type [660]. Hardy [1270, 1338, 1646, 1715, 1896, 2164, 2542]. harmonic [36, 1009, 1331, 1340, 1416, 1448, 1655, 1758, 1923, 2063, 2519, 2534]. Hasse [1033, 1552, 1634, 1734, 2058]. Hauptmodul [2623]. Hausdorff [1028, 1933, 2192, 2206]. Hauteur [1410, 2028, 2637]. Hauteurs [1318]. having [930, 1068, 1876, 2079, 2599]. Hayes [657]. Hecke [8, 84, 94, 111, 158, 216, 268, 324, 368, 388, 392, 453, 461, 466, 524, 572, 611, 682, 711, 755, 882, 905, 939, 953, 1000, 1017, 1051, 1120, 1137, 1145, 1220, 1415, 1555, 1761, 1816, 1828, 2163, 2214, 2278, 2324, 2412, 2452, 2469, 2491, 2497, 2508, 2510]. Heegner [199, 434, 507, 595]. Hegyvári [786]. Height [145, 172, 1119, 1410, 1782, 1934, 2027, 2028, 2637]. heights [117, 481, 1020, 1521, 1807, 2538, 2640]. Heilbronn [2195]. Heinrich [500]. Heinz [1836]. Heisenberg [796, 2473]. Hendy [2001, 2139]. Hensel [1433]. hereditarily [1740]. Hermite [136, 397, 1608, 1620, 1960, 2329]. Hermitian [325, 525, 1284, 1289, 1317, 1736, 2182, 2383, 2475, 2476]. Heron [2044]. Hesselholt [349]. Hessenberg [319]. heuristic [1248, 2152]. Hickerson [1785]. Hida [2445]. High [125, 356, 436, 1910, 2461]. Higher [368, 387, 426, 670, 692, 714, 715, 742, 797, 896, 1055, 1066, 1180, 1251, 1396, 1547, 1631, 1785, 1809, 1867, 1984, 2010, 2210, 2348, 2429, 2501]. higher-order [2501]. highest [1170]. Hilbert [3, 101, 107, 251, 270, 286, 443, 613, 882, 1100, 1108, 1172, 1185, 1362, 1499, 1501, 1523, 1583, 1649, 1859, 1876, 2079, 2096, 2322, 2517]. Hirst [1691]. Historical [1187]. Hodge [951, 1190, 2084]. holomorphic [347, 555, 692, 1289, 1360, 1673, 1690, 1784, 1948, 2102, 2163, 2330, 2434, 2612]. holonomic [780]. hom [1950]. hom-Lie [1950]. homogeneous [2148]. Homological [536]. homology [394, 1131, 1460, 1503]. homomorphism [1212]. homomorphisms [390]. homothéties [380]. homothies [380]. homotopy [2137]. honor [1514]. Hooley [1560]. houses [1976]. hull [418, 1803]. Humbert [2370]. Hurwitz [37, 51, 393, 584, 593, 1058, 1072, 1115, 1333, 1361, 1563, 1701, 1811, 2031, 2145, 2312, 2351, 2428]. Hurwitz-type [593]. Hutchinson [1852]. Hybrid [378, 751, 1199, 1920, 2288]. hyperbolas [1803]. hyperbolic [383, 604, 1623, 2558]. hyperbolicity [560]. hypercomplex [1227]. hyperderivatives [1245]. Hyperelliptic [75, 287, 390, 436, 481, 581, 667, 775, 906, 955, 1275, 1483, 1512, 2316, 2401, 2587]. Hypergeometric [180, 268, 358, 392, 801, 954, 1033, 1310, 1324, 1649, 1881, 1904, 2026, 2200, 2343]. Hyperharmonic [37, 1175, 1746, 1823]. hypermonogenic [729]. hyperplanes [1596, 2126]. Hyperquadratic [1250]. hyperring [235]. hyperrings [1893]. hyperspecial [1147]. Hyperstructures [1704]. hypersurface [2213]. hypersurfaces [1376, 1867, 1971, 2026]. hypothèse [6, 93]. hypothesis [6, 53, 73, 92, 341, 1007, 1529, 1619, 1757, 1764, 2049, 2253].

ideal [215, 266, 296, 390, 516, 679, 754, 896, 1244, 1274, 1534, 1778, 1862, 2427, 2465, 2599]. ideals [779, 896, 1419, 1708, 1886, 1969, 2639]. idempotent [2404, 2627]. idempotent-product [2404]. identical [240]. Identities [132, 155, 258, 334, 424, 585, 591, 621, 655, 852, 954, 1035, 1124, 1242, 1278, 1310, 1440, 1442, 1528, 1615, 1646, 1716, 1731, 1877, 1949, 2017, 2023, 2031, 2141, 2145, 2273, 2353, 2530, 2597]. identity [300, 563, 591, 952, 1384, 1606, 1769, 1921, 1952]. If [2532]. Igusa [175]. Ihara [702, 816, 1611, 2473, 2539]. II [56, 92, 161, 344, 588, 679, 696, 727, 886, 1014, 1046, 1080, 1127, 1246, 1298, 1315, 1330, 1414, 1430, 1457, 1485, 1510, 1631, 1710, 1860, 2000, 2032, 2042, 2117, 2128, 2171, 2198, 2267, 2269, 2370, 2473, 2519, 2526, 2543, 2617, 2635]. III [1008, 1134, 1620, 2348, 2478]. Ikeda [880, 1137]. image [380, 1585, 1737, 2256, 2453]. Images [1275]. Imaginary [48, 226, 296, 446, 448, 653, 716, 905, 1339, 1499, 1778, 1862, 1891, 1953, 1980, 1982, 1992, 2001, 2139, 2149, 2184, 2197, 2203, 2216, 2237, 2544, 2616]. imaginary-quadratic [1980]. imperfect [134]. implies [2110]. imply [295]. implying [2530]. Impossible [1747]. imprimitive [1484]. Improved [886, 1197, 1916, 2108, 2628]. improvement [322, 439]. Incarnations [1663]. incomplete [133, 282, 867, 1766, 2574]. incompressibility [745]. increase [2304]. increasing [1101]. Indecomposable [1500, 1667]. Independence [199, 204, 222, 275, 671, 1662, 1832, 1875, 2161, 2564]. independent [1972]. index [265, 266, 679, 1013, 1656, 1722, 2210, 2585, 2608]. indexed [272, 520, 798, 1685]. Indices [924, 1026, 1265]. Indivisibility [186, 2203, 2347]. induced [87]. induction [1786]. Inequalities [838, 998, 1000, 1084, 1203, 1252, 1285, 1494, 1564, 1601, 1644, 1791, 1793, 1834, 1836, 1925, 2148]. inequality [916, 1270, 1277, 1475, 1575, 2419, 2561, 2643]. inert [1819]. Infinite [33, 103, 272, 350, 612, 669, 923, 1209, 1210, 1242, 1320, 1363, 1565, 1713, 1788, 1832, 1953, 2267]. infinite-dimensional [923]. infinitely [97]. infiniteness [1177]. information [30]. Inhomogeneous [1381, 1688]. initial [210]. initiales [210]. Injectivity [1212]. Inose [2309]. inradius [1773]. inseparability [924, 1026, 1265]. instances [2357]. instantons [1138]. Integer [136, 308, 329, 397, 419, 558, 619, 773, 819, 867, 868, 993, 1125, 1211, 1449, 1454, 1466, 1479, 1480, 1656, 1933, 1947, 1988, 2077, 2078, 2215, 2311, 2365, 2450, 2542, 2602]. integer-valued [558, 868]. Integers [5, 22, 113, 114, 116, 193, 252, 349, 446, 531, 603, 831, 834, 866, 885, 917, 956, 1034, 1050, 1065, 1076, 1101, 1128, 1150, 1287, 1372, 1388, 1389, 1457, 1460, 1463, 1466, 1604, 1661, 1667, 1699, 1709, 1710, 1718, 1734, 1763, 1780, 1823, 1907, 1976, 2019, 2217, 2411, 2435, 2459, 2527, 2600, 2642, 2644]. Integral [3, 107, 266, 267, 394, 406, 430, 630, 647, 650, 678, 679, 711, 752, 777, 873, 885, 926, 932, 939, 956, 962, 986, 1031, 1084, 1180, 1190, 1374, 1397, 1417, 1527, 1648, 1731, 1738, 1809, 1810, 1857, 1860, 1989, 2036, 2043, 2044, 2069, 2092, 2126, 2278, 2363, 2452, 2517, 2527, 2614]. Integral-valued [956]. intégrales [1147]. Integrality [1099, 1360, 1964, 2237, 2268]. integrals [31, 147, 216, 423, 819, 906, 1009, 1135, 1147, 1527, 1636, 1639, 1749, 1880, 2018, 2134, 2198, 2212, 2331, 2403, 2426]. integration [1799]. interlacing [2125]. Intermediate [2458]. interpolated [2512, 2534]. Interpolation [868, 1887]. interpret [271]. interpretation [536, 1388, 1817, 1848]. interpretations [1479]. intersections [1425, 1747]. Intersective [140]. interval [1870, 1972]. intervals [867, 964, 1216, 1290, 1330, 1450, 1578, 1676, 1708, 1809, 2173, 2262, 2444, 2488, 2500, 2607]. intractable [30]. introduced [2395]. invariance [352]. Invariant [58, 147, 314, 653, 744, 778, 904, 1054, 1205, 1206, 1261, 1807, 2068, 2154, 2215, 2370]. invariantes [778]. invariants [17, 286, 441, 542, 725, 918, 965, 1033, 1078, 1174, 1185, 1523, 1598, 1837, 1981, 2203, 2215, 2249, 2473]. Inverse [545, 982, 1004, 1047, 1218, 1266, 1303, 1513, 1665, 1985, 2435]. inversion [1593]. involution [2127, 2476]. involving [37, 258, 317, 334, 415, 585, 590, 655, 709, 814, 1086, 1331, 1416, 1478, 1542, 1624, 1646, 1721, 1791, 1793, 2022, 2064, 2092, 2476, 2497, 2563, 2632, 2645]. irrational [544, 1968]. irrationalities [2222]. irrationality [56, 511, 1037, 1276, 1997]. irrationals [514, 930]. Irreducibility [242, 815, 1480, 1620, 2133, 2339]. irreducible [751, 806, 1046, 1295, 1409, 1740, 2068, 2553, 2580]. Irreduzible [2580]. Irregular [2629]. isogenies [283, 1686]. isogenous [694, 2309]. isogeny [305, 484, 942, 1161, 1319, 2321]. isolated [2396]. isomorphic [977, 2087]. isomorphism [323, 377, 582, 942, 1110, 1312, 1413, 1984, 2048, 2214]. isomorphy [321]. isosceles [2044]. isotrivial [2315]. iterated [31, 2403]. Iterating [2109]. Iteration [1495]. iterations [1038, 1461]. IV [47]. Iwahori [471]. Iwasawa [91, 493, 542, 653, 896, 965, 1426, 1822, 1837, 1981, 2132, 2203, 2249, 2422, 2622, 2639].

J [93, 94, 106, 107, 129, 189, 302, 384, 517, 550, 655, 738, 739, 808, 909, 939, 957, 1001, 1128, 1271, 1291, 1303, 1304, 1471, 1627, 1628, 1743, 1770, 1848, 1882, 2012, 2079, 2080, 2139–2141, 2548, 2642–2649]. Jack [1526]. Jacobi [8, 94, 158, 647, 899, 938, 1019, 1206, 1254, 1288, 1416, 1465, 1741, 1748, 1880, 2024, 2099, 2196, 2202, 2363, 2391, 2455, 2548, 2590, 2608]. Jacobian [685, 724, 828, 2146, 2299, 2401]. Jacobians [244, 481, 672, 846, 977, 1118, 1275, 1319, 1390]. Jacquet [305, 1136, 2331]. Jager [848]. January [11, 220, 402, 600, 870, 1420, 1771, 2094, 2376]. Jeśmanowicz [641, 1014, 1021, 1379]. Joint [388, 393, 584, 1333, 1511, 1572, 1761, 2252]. Jordan [2286]. Josephus [72]. July [119, 303, 486, 740, 980, 1292, 1602, 1935, 2246, 2504]. jumps [1580]. June [108, 288, 472, 717, 968, 1272, 1567, 1917, 2220, 2481]. just [629, 738].

K. [416, 2394]. Kaneko [292, 1052, 1266, 1352, 1970]. Kaprekar [2181]. Kashio [1558]. Kato [766]. Katz [439, 2017]. Kawazumi [1078, 1598]. kernel [202, 1132, 1435, 2263, 2568]. kernels [811, 1813, 2305, 2544]. Kida [965]. Kim [819]. kind [124, 564, 752, 999, 1084, 1208, 1322, 1838, 1855, 1965, 2392]. Kinkelin [763, 774, 1077, 1089, 1582, 1591]. Kirillov [1132]. Kisin [949]. Klein [1280]. Kloosterman [133, 238, 269, 710, 1186, 1199, 1486, 1490, 1610, 1646, 1762, 1908, 2105, 2565, 2574]. knots [1528]. Koblitz [2190]. Kohnen [1583, 2151]. Komponenten [2580]. Koninck [944, 1346, 2080]. Korselt [2531]. Koszul [1103]. Kronecker [135, 695, 736, 903, 905, 1427, 1631, 2352]. Kronecker-sequences [903]. Kummer [17, 59, 658, 739, 1508, 2347, 2386]. Kurihara [2539]. Kurokawa [2326]. Kuzmin [743, 1801, 2402]. Kuzmin-type [743].

Lagrange [1550, 1930, 2136]. Laguerre [456, 1006, 1480, 1620, 2133]. Lam [1423]. lambda [17, 653, 2204]. Lang [1378, 1628, 1929, 2318]. Langlands [305, 633, 1134, 1165, 1611, 2006]. Laplace [1851]. Laplacians [2297].

Laplacien [863]. Larcombe [2022]. Large [227, 399, 449, 627, 757, 767, 916, 1000, 1382, 1705, 1966, 2101, 2116, 2209, 2419, 2472, 2486, 2643]. largest [1010, 1302, 1595, 2404]. Last [930, 1238]. Lattès [936, 1060]. Lattice [561, 579, 751, 1417, 2049, 2053, 2097, 2170, 2309, 2325, 2608]. lattices [61, 125, 430, 554, 1020, 1023, 1085, 1246, 1500, 1533, 1688, 1980, 2291, 2427, 2445, 2448]. Laurent [154, 1028, 1861, 2185]. Lauricella [906]. law [224, 284, 2381]. laws [282, 1188, 2252]. layers [859, 1104]. Least [477, 712, 840, 1076, 1247, 1481, 1754, 1763, 1775, 1778, 1931, 2016, 2241, 2307, 2600, 2605]. Lebesgue [253]. Lee [2328]. Leech [579]. Lefschetz [685]. Legendre [141, 255, 727, 906, 921, 1063, 1125, 1232, 1524, 2016, 2200, 2216, 2251, 2366]. Lehmer [210, 211, 446, 854, 1113, 1807, 2164, 2280, 2577]. lemma [1433, 1611, 2329]. lemniscate [894]. length [118, 165, 383, 862, 1068, 1088, 1192, 1330, 1955, 1979, 2323, 2554]. lengths [1741]. Lenstra [1563]. Leopoldt [500]. Lerch [1352, 1731, 1827]. Level [128, 243, 320, 368, 408, 523, 582, 605, 692, 704, 920, 975, 1040, 1110, 1163, 1396, 1439, 1451, 1614, 1696, 1719, 1732, 1830, 1851, 1984, 2051, 2072, 2076, 2099, 2254, 2295, 2359, 2425, 2475, 2575]. Levy [1801, 2001, 2139]. Li [6, 73, 92, 93, 257, 1387, 1427, 1514]. Lie [1723, 1739, 1950]. life [657]. lifelong [1514]. lift [1598, 1758, 2475]. Lifting [58, 177, 197, 680, 734, 815, 1261, 1943, 2091]. lifts [499, 604, 766, 826, 880, 1137, 1317, 1554, 2182, 2298, 2326, 2493, 2541]. like [176, 587, 726, 1311, 2162]. Limit [135, 736, 1249, 1427, 1574, 1589, 2137, 2349, 2351]. Limit-Fourier [1249]. Limiting [1066]. limits [363, 1681, 1735]. Lindelöf [2049]. line [442, 453, 886, 991, 1640, 1660, 1873, 2243, 2516]. Linear [143, 145, 243, 275, 404, 585, 655, 689, 829, 931, 960, 1022, 1032, 1067, 1174, 1290, 1662, 1779, 1806, 2023, 2042, 2043, 2071, 2090, 2092, 2141, 2144, 2148, 2234, 2242, 2265, 2336, 2454, 2569, 2615]. linearly [654]. Lines [159, 1173, 2062, 2596]. linkage [2498]. Linnik [68, 916, 1309, 1958]. Linnik-type [68]. Liouville [619, 640]. Lipschitz [788, 1281, 1433]. Littlewood [1270, 1475, 1540, 1715, 2011]. Local [328, 445, 524, 535, 587, 644, 646, 799, 837, 960, 1055, 1059, 1134, 1136, 1146, 1165, 1279, 1312, 1407, 1611, 1645, 1681, 1723, 1861, 1897, 1943, 1981, 2252, 2426, 2466, 2467, 2492, 2617]. locales [2252]. Locally [338, 445, 1055, 1609]. locaux [1407]. locus [1975]. Log [9, 106, 309, 858, 869, 957, 1558]. log-gamma [1558]. logarithm [2327, 2386]. logarithmic [938, 2118]. logarithms [782]. lois [2252]. Long [157, 497, 841]. Longest [2130]. look [1240]. loop [2438]. lost [1210]. Low [523, 561, 586]. low-level [523]. Low-lying [586]. Lower [63, 117, 139, 576, 645, 742, 816, 903, 1007, 1463, 1698, 1911, 1916, 1931, 2266, 2285, 2379, 2385, 2628, 2635, 2640]. lower-order [1007]. lowest [1170]. Lubin [183, 1163]. Lucas [210, 598, 1569, 1790, 1855, 1927, 2327, 2543]. Lüroth [1452]. Lutz [667]. lying [586].

Maass [3, 52, 107, 110, 682, 721, 768, 916, 1009, 1220, 1316, 1317, 1416, 1422, 1471, 1478, 1586, 1587, 1690, 1748, 1758, 1851, 2475, 2503, 2621]. magic [269]. magnitude [1125]. magnitudes [1542]. Mahler [297, 420, 526, 1032, 1066, 1091, 1281, 1287, 1336, 1597, 1894, 1912, 2161, 2368, 2394]. Main [493, 1822]. Malle [1361]. Malmstén [2386]. manifolds [1300, 2558]. Manin [1179]. many [97, 479, 622, 961, 1178, 1773, 1921, 2030, 2423, 2546]. map [512, 659, 686, 1100, 1833, 1990, 2053, 2391, 2568, 2601]. mappings [970, 1739]. maps [183, 398, 722, 1027, 1038, 1060, 1067, 2085, 2154, 2368, 2640]. March [43, 246, 428, 910, 1201, 1472, 1824, 2142, 2417]. Markoff [322, 756, 1294]. Markov [1933, 2396]. Mascheroni [1601, 1644, 1746, 2052]. Mass [476, 1784]. Massey [1942]. massive [2407]. Matching [1145]. mathematical [1397]. Mathematics [189]. Mathieu [1229]. Matrices [15, 136, 319, 397, 446, 777, 1114, 1244, 1725, 1971, 2602]. matrix [592, 835, 1248, 1665]. Max [78, 1414]. Maximal [80, 457, 577, 698, 798, 821, 846, 932, 1438, 1670, 2244, 2587]. Maximum [545, 1338]. May [95, 459, 700, 958, 1257, 1538, 1883, 2193, 2462]. Mazur [58, 1695]. McMullen [1695]. Mean [67, 76, 192, 238, 266, 326, 339, 384, 387, 452, 468, 503, 530, 581, 679, 775, 813, 934, 937, 1083, 1162, 1199, 1394, 1490, 1762, 1908, 1986, 2021, 2117, 2446, 2474, 2610]. mean-value [1394, 2446]. means [1380]. measure [420, 539, 603, 635, 1037, 1066, 1287, 2206]. measure-preserving [635]. measures [14, 222, 297, 511, 526, 542, 1032, 1531, 1946, 2039, 2154]. median [1010, 1913]. medians [735]. Mellin [423, 1232, 2145]. Menon [563, 952, 1606, 1921, 2023, 2141, 2353]. Menon-type [1606, 1921, 2023, 2141, 2353]. Merca [1181]. meromorphic [613, 1156, 1554, 2390]. Mersenne [538, 1105]. message [1995]. metacyclic [148]. métacycliques [148]. Metaplectic [373, 1135]. method [280, 803, 2613]. methods [126, 670]. metric [13, 297]. Metrical [903]. Mézard [1403]. middle [1847]. Midwest [61]. mild [1144]. Milnor [2473]. min [78, 1414]. minima [946, 1688]. Minimal [54, 165, 731, 862, 1047, 1398, 1438, 1945, 1985, 2222, 2584]. minimization [2, 2637]. minimizing [1874]. Minimum [125, 546, 1468, 2480]. Minkowski [322, 1894, 1956, 1960]. Minoration [2, 154, 1410, 2637]. minus [2166, 2434]. mirabolic [960]. Mircea [1181]. missing [172, 267, 376, 1399]. mixed [528, 535, 703, 1083, 1102, 1442, 1475, 1540, 1579, 1688, 2017]. Miyawaki [880, 2182]. Möbius [631, 747, 1573, 1593, 2456, 2583]. mock [768, 1229, 1442, 1471, 1652, 1922, 1961, 2274, 2352]. Mod [693]. model [175, 1132, 1623, 2325, 2387, 2467, 2539]. model-fields [2387]. models [920, 1136, 2198, 2638]. moderate [1382]. modes [1138]. modification [805]. modified [1387, 2536]. Modular [55, 75, 118, 124, 128, 183, 218, 260, 311, 313, 324, 325, 443, 453, 466, 483, 495, 540, 548, 609, 637, 660, 678, 692, 696, 704, 749, 768, 829, 836, 845, 853, 880, 882, 883, 907, 929, 953, 966, 985, 994, 1005, 1051, 1081, 1108, 1110, 1152, 1164, 1172, 1185, 1245, 1358, 1362, 1415, 1437, 1442, 1471, 1484, 1504, 1525, 1527, 1631, 1649, 1652, 1664, 1673, 1696, 1703, 1738, 1784, 1792, 1817, 1830, 1846, 1848, 1859, 1961, 2160, 2163, 2182, 2183, 2236, 2251, 2255, 2273, 2274, 2278, 2288, 2299, 2322, 2330, 2334, 2348, 2359, 2405, 2434, 2438, 2511, 2564, 2597, 2612]. modular [8, 94, 101, 251, 270, 294, 513, 613, 632, 711, 873, 939, 983, 1032, 1100, 1255, 1321, 1360, 1517, 1554, 1583, 1803, 1819, 2125, 2268, 2445, 2494]. modular-type [1164]. Modularity [128, 392, 800, 1732, 2264, 2586]. module [291, 713, 739, 837, 1022, 1105, 1513, 2059]. Modules [38, 49, 105, 239, 498, 553, 661, 662, 723, 849, 896, 949, 961, 967, 986, 1167, 1343, 1393, 1410, 1585, 1590, 1642, 1653, 2028, 2280, 2412, 2422, 2638, 2639]. moduli [286, 544, 610, 634, 906, 1444, 1495, 2335, 2387, 2400, 2419, 2643]. modulo [79, 353, 477, 560, 580, 726, 757, 834, 842, 892, 938, 989, 1011, 1030, 1057, 1073, 1189, 1207, 1268, 1279, 1282, 1299, 1351, 1402, 1408, 1455, 1586, 1634, 1657, 1701, 1724, 1765, 1775, 1777, 1788, 1800, 1819, 1846, 1914, 1941, 2156, 2230, 2500, 2577]. modulus [987, 1481]. Moebius [2154]. Mollification [408]. mollifier [1901]. moment [245, 458, 703, 742, 1865, 1866, 1909, 2106, 2464, 2557, 2565, 2649]. Moments [151, 202, 219, 306, 387, 645, 684, 714, 838, 978, 1017, 1031, 1535, 1610, 1809, 1999, 2118, 2228, 2266, 2327, 2381, 2451, 2469, 2528]. Monochromatic [1053, 1549]. monodromy [646, 698, 2443]. monogenic [2460]. Monogenity [1441, 1857, 2465]. Monomial [398, 1090, 2640]. monotonicity [1084]. Montgomery [916, 1560]. Mordell [353, 435, 568, 650, 951, 1378, 1628, 1659, 1929, 2309, 2566]. Morita [505]. morphisms [2634]. Morse [329, 1171]. Mortici [1564]. Moser [295]. most [1088]. motivated [881, 1168]. motive [2407]. motives [536, 537, 2053]. Motivic [1491, 1503]. Motzkin [2501]. moving [1596, 2213]. MSTD [102, 103, 184]. Mullin [1633]. Multi [272, 702, 1579, 2086, 2515]. multi-dimensional [1579, 2086]. multi-indexed [272]. Multi-rotations [2515]. multi-zeta [702]. multidimensional [1820]. multifractal [2349]. multigraphs [269]. multilinear [409]. multinomial [937]. multinorm [654]. Multiperfect [240]. Multiple [36, 226, 256, 309, 424, 444, 527, 642, 858, 1076, 1082, 1231, 1253, 1313, 1331, 1424, 1439, 1537, 1615, 1680, 1742, 1763, 1787, 1804, 1831, 1887, 1931, 1987, 1988, 2018, 2031, 2088, 2261, 2284, 2341, 2351, 2403, 2411, 2423, 2495, 2502, 2512, 2513, 2534, 2600, 2630, 2632]. Multiple-correction [1253]. multiples [329, 1191, 1988]. multiplication [113, 114, 141, 662, 1241, 1524, 1612, 1845, 2202]. multiplications [1410]. Multiplicative [35, 59, 730, 1123, 1326, 1394, 1418, 1573, 1641, 1702, 1740, 1789, 1790, 1844, 1863, 1898, 1972, 2000, 2042, 2071, 2431, 2446]. multiplicities [933, 2173]. Multiplicity [383, 822, 2152]. multiquadratic [644, 1781, 1891]. multisets [1889]. multivariable [378, 2346]. multivariate [1036, 1803]. multizeta [86, 374, 1875, 2234]. Mumford [1344, 2238]. Muriefah [189]. MZSV [413].

n [164]. Nagell [253, 254, 667]. Nakada [2124]. Nakagawa [2281]. Narayana [1967]. Nathan [1949]. Nathanson [2533]. Natural [397, 597, 697, 732, 964, 2124]. nature [1536]. near [596, 1801, 2396, 2516]. near-optimal [1801]. near-perfect [596]. nearest [1933]. nearly [734, 988]. negative [777, 1896]. negativity [2294]. Neighboring [60, 129]. Néron [175, 481]. net [1432]. nets [1432]. network [1147]. networks [2214]. never [2411]. newform [2296]. Newforms [1358, 1519, 2151, 2189, 2363, 2510]. Newman [1075, 1400]. Newton [936, 1156, 1180, 1498, 1906, 2177]. Nicolas [2395]. Niebur [604, 2619]. Niebur-type [604]. niveau [408]. No [1664, 1811]. nodal [2172]. Noether [2470]. Noguchi [1929]. nombre [820, 2252]. nombres [2, 154, 820]. nomial [18]. Non [75, 91, 218, 408, 427, 515, 585, 622, 643, 655, 691, 716, 719, 728, 751, 758, 790, 905, 961, 1029, 1074, 1081, 1096, 1247, 1297, 1327, 1517, 1536, 1645, 1713, 1859, 2039, 2056, 2111, 2132, 2175, 2238, 2249, 2294, 2308, 2315, 2383, 2433, 2488, 2495, 2510, 2514, 2599, 2605, 2626]. non-abelian [1029, 1297, 2514]. non-adjacent [716]. non-associates [961]. non-bounded [2488]. non-commutative [91, 2495]. non-congruence [1517]. non-congruent [622, 2433]. non-critical [905]. Non-cyclic [427]. non-cyclotomic [2249]. non-full [2308]. non-generic [1645]. Non-hyperelliptic [75]. Non-isotrivial [2315]. non-linear [585, 655]. non-negativity [2294]. non-normal [515]. non-ordinary [1081, 2132]. non-principal [2599]. non-residue [691, 1247]. non-residues [758, 2605]. non-simple [1074]. non-split [218, 2383]. non-square [1074]. non-tempered [1327]. non-totality [2175]. non-trivial [643, 728, 790, 1713]. non-uniform [751]. Non-vanishing [1096, 1536, 1859, 2056, 2111, 2510, 2626]. non-zero [2315]. nonabelian [546, 877]. nonclassical [1469]. noncommutative [237]. Noncongruence [2503]. nondeficient [2317]. nondensity [276]. Nonexistence [2384]. nonholomorphic [52, 1525]. Nonlinear [639, 1888, 2191]. Nonnegative [1411]. nonresidue [477, 2241]. nonresidues [1188, 2620]. Nonrigid [1929]. nonscalar [2545]. nonsingular [1608]. nonsolvable [1670]. nonunits [1351]. nonvanishing [450, 1184, 1989]. nonzero [1219]. norm [47, 125, 310, 961, 1027, 1284, 1634, 2568, 2630]. Normal [136, 232, 397, 494, 515, 1506, 1608, 2518]. Normality [335, 687, 1679]. normalized [1308, 2005, 2409, 2508]. Normalizers [218]. normes [778]. Norms [420, 778, 1667, 2326]. Note [20, 24, 154, 174, 189, 196, 353, 371, 450, 475, 682, 760, 775, 799, 814, 1001, 1028, 1083, 1379, 1501, 1542, 1596, 1642, 1722, 1735, 1784, 1903, 1908, 1957, 2252, 2428, 2487, 2631]. notebook [1210]. notebooks [590]. Nottingham [1106]. November [190, 365, 551, 832, 1070, 1365, 1711, 2033, 2332, 2594]. Novodvorsky [2426]. Number [10, 33, 80, 88, 93, 94, 98, 106, 112, 129, 189, 198, 218, 302, 307, 351, 384, 489, 515, 517, 550, 608, 655, 659, 699, 728, 732, 738, 739, 792, 808, 820, 847, 865, 909, 912, 913, 939, 957, 1001, 1050, 1122, 1128, 1160, 1236, 1260, 1271, 1291, 1303, 1304, 1339, 1394, 1402, 1413, 1439, 1448, 1471, 1505, 1578, 1627, 1628, 1658, 1685, 1698, 1708, 1743, 1770, 1776, 1848, 1874, 1876, 1882, 1891, 1992, 2012, 2032, 2079, 2080, 2140, 2141, 2216, 2217, 2233, 2315, 2350, 2382, 2452, 2480, 2485, 2509, 2548, 2570, 2642–2649]. number [3, 107, 128, 177, 186, 188, 201, 291, 323, 335, 390, 431, 452, 482, 559, 568, 576, 578, 586, 612, 651, 736, 798, 878, 899, 908, 918, 996, 1007, 1039, 1103, 1107, 1150, 1215, 1224, 1228, 1240, 1267, 1332, 1340, 1347, 1348, 1371, 1376, 1398, 1405, 1406, 1425, 1435, 1462, 1466, 1479, 1499, 1502, 1506, 1532, 1552, 1563, 1592, 1678, 1689, 1694, 1750, 1754, 1759, 1794, 1813, 1860, 1919, 1937, 1955, 1982, 2001, 2016, 2067, 2069, 2074, 2093, 2119, 2129, 2139, 2144, 2149, 2155, 2190, 2207, 2218, 2226, 2252, 2265, 2303, 2323, 2385, 2444, 2448, 2470, 2486, 2521, 2542, 2558, 2615]. number [262, 404, 947, 1069, 1280, 2009, 2544]. numbers [2, 30, 51, 80, 127, 132, 143, 149, 151, 180, 240, 290, 361, 416, 490, 538, 560, 565, 583, 628, 631, 639, 678, 697, 732, 756, 803, 820, 854, 902, 948, 964, 994, 1088, 1136, 1155, 1160, 1175, 1177, 1189, 1208, 1236, 1262, 1263, 1288, 1296, 1298, 1307, 1311, 1336, 1367, 1459, 1464, 1476, 1493, 1497, 1530, 1565, 1581, 1605, 1632, 1668, 1676, 1721, 1743, 1746, 1755, 1823, 1835, 1877, 1894, 1934, 1944, 1953, 1965, 1967, 1968, 1974, 1981, 2022, 2032, 2107, 2116, 2145, 2179, 2184, 2275, 2306, 2317, 2347, 2365, 2392, 2410, 2431, 2433, 2501, 2546, 2569, 2577, 2589, 2629]. numbers [104, 111, 142, 319, 396, 414, 440, 451, 549, 569, 582, 596, 597, 622, 629, 681, 738, 752, 791, 874, 887, 999, 1084, 1302, 1322, 1383, 1412, 1655, 1666, 1774, 1820, 1895, 1923, 1952, 2063, 2089, 2105, 2122, 2123, 2203, 2218, 2364, 2519, 2552, 2581]. numerical [391, 1356, 1588]. numerology [2415]. Nyman [2253].

Obituary [2015]. objects [1428]. obstruction [1179, 2514]. obtain [2008]. occurrence [202]. octagonal [2306]. Octahedral [1109, 1337, 1692]. octic [1188]. October [170, 345, 532, 809, 1041, 1349, 1682, 2013, 2313, 2572]. odd [142, 579, 672, 732, 773, 947, 1015, 1038, 1044, 1091, 1250, 1251, 1280, 1295, 1439, 1519, 1665, 1674, 1685, 2024, 2078, 2099, 2267, 2294, 2496, 2562, 2618]. odd-indexed [1685]. Odlyzko [750]. off [2126]. often [2441]. Ohno [444, 642, 2281]. old [245]. One [33, 38, 49, 218, 259, 383, 489, 515, 560, 562, 770, 800, 891, 932, 1061, 1122, 1195, 1250, 1347, 1348, 1393, 1450, 1496, 1531, 1563, 1588, 1592, 1672, 1941, 1996, 2046, 2092, 2136, 2166, 2259, 2544, 2585, 2609]. One-class [932]. one-dimensional [49]. one-parameter [1195]. ones [629, 738]. only [781, 800, 1061, 1232, 1443, 1882]. open [1737, 2593]. openness [498]. opérateurs [2214]. operator [55, 1206, 2263]. operators [8, 94, 228, 324, 347, 388, 392, 453, 466, 711, 939, 953, 976, 1133, 1145, 1254, 1415, 1816, 1932, 2024, 2214, 2491]. Oppenheim [2485]. Optimal [83, 259, 1255, 1338, 1475, 1659, 1801, 2361]. orbifold [1362]. orbit [2640]. orbital [1147, 2134]. orbitales [1147]. Orbits [261, 262, 308, 474, 599, 1140, 1964, 2286, 2357]. order [54, 142, 280, 352, 645, 797, 803, 862, 874, 923, 1007, 1074, 1124, 1211, 1237, 1251, 1437, 1657, 1765, 1867, 2057, 2059, 2284, 2501]. order- [803]. ordering [1174]. orderings [1874]. Orders [19, 213, 254, 322, 557, 577, 914, 1205, 1246, 1477, 1678, 1717, 2090, 2127, 2598]. ordinarity [772]. Ordinary [87, 279, 280, 283, 638, 831, 883, 1081, 2132, 2236, 2571]. ordinates [1643]. origin [1059]. Orthogonal [406, 645, 854, 885, 988, 1409, 1519, 2198, 2406, 2501]. orthogonality [2583]. Oscillations [212, 1467]. oscillatory [370]. other [394, 477, 1309]. outlier [2172]. overconvergence [2638]. overconvergent [55, 2630]. Overpartition [1586]. Overpartitions [726, 838, 1207, 1215, 1251, 1622, 1637, 1716, 1869, 1926, 2100, 2156].

packings [2598]. Padé [2329]. Pages [11, 26, 43, 70, 95, 108, 119, 137, 152, 170, 190, 205, 220, 233, 246, 263, 288, 303, 315, 332, 345, 365, 385, 402, 421, 428, 437, 459, 472, 486, 501, 518, 532, 551, 573, 600, 625, 676, 700, 717, 740, 761, 783, 809, 832, 856, 870, 888, 910, 940, 958, 968, 980, 1002, 1024, 1041, 1070, 1094, 1148, 1201, 1234, 1257, 1272, 1292, 1305, 1328, 1349, 1365, 1391, 1420, 1445, 1472, 1538, 1567, 1602, 1629, 1650, 1682, 1711, 1744, 1771, 1796, 1824, 1849, 1917, 1935, 1962, 1993, 2081]. Pages [1883, 2013, 2033, 2094, 2114, 2142, 2168, 2193, 2220, 2246, 2270, 2289, 2313, 2332, 2355, 2376, 2398, 2417, 2439, 2462, 2481, 2504, 2523, 2549, 2572, 2594, 2624]. Pair [393, 493, 1145, 1283, 1430, 1757, 2104, 2197, 2374, 2389, 2556]. Pairing [281, 287, 846]. Pairing-based [846]. pairing-friendly [287]. Pairings [1416, 1554, 1686]. Pairs [194, 616, 619, 694, 799, 818, 821, 848, 1127, 1182, 1496, 1613, 1719, 1800, 1910, 1953, 1978, 2044, 2092, 2149, 2164, 2175, 2214, 2286, 2302, 2310, 2551]. paper [154, 351, 416]. Paperfolding [1209]. parabolas [1852]. parabolic [1786]. paradigm [278]. parallel [2596]. parallelogram [1613, 2044, 2551]. parameter [1195, 1296, 1327, 1572, 1885]. Parameterizations [617]. parameterized [602]. parameterized-Euler-constant [602]. parameters [825, 1117, 2076, 2567]. parametric [4, 1271]. parametrized [256, 297]. paramodular [2111]. Parity [83, 87, 843, 890, 995, 1831, 1981, 2070]. Parity-induced [87]. Parseval [1375, 1510]. part [214, 720, 1215, 1251, 2198]. Partial [30, 69, 439, 919, 1117, 1250, 1416, 1595, 1844, 1980, 2089, 2304, 2496]. particular [944, 2080]. Partition [155, 538, 861, 979, 1044, 1280, 1442, 1625, 1800, 1949, 2008, 2148, 2267, 2269, 2302, 2530]. Partitions [399, 482, 597, 651, 792, 842, 843, 890, 985, 995, 1011, 1030, 1039, 1198, 1251, 1282, 1299, 1402, 1437, 1456, 1466, 1542, 1570, 1594, 1674, 1685, 1716, 1750, 1759, 1788, 1922, 1937, 2007, 2100, 2155, 2167, 2188, 2230, 2294, 2365, 2542, 2613]. parts [52, 482, 1039, 1088, 1215, 1251, 1283, 1430, 1542, 1674, 1685, 1926, 1937, 2155, 2197, 2274, 2294, 2365]. pattern [2302]. Patterns [1357, 1565]. Pell [162, 319, 930, 1157, 2150, 2393, 2588, 2589]. Pellarin [666]. Pellian [337]. pentagonal [1405]. Perfect [467, 596, 1262, 2166, 2275, 2441, 2547]. perimeter [1613, 2044, 2551]. Period [84, 216, 623, 1396, 1741, 1816, 2035, 2210]. periodic [665, 930, 1086, 1096, 1544, 1741, 2354, 2627]. Periodicity [730, 1977, 2008, 2040]. periods [9, 106, 537, 624, 723, 782, 960, 2222]. permutation [830, 1151, 1940]. permutations [866]. Perrin [97, 319]. Perron [1741]. Perspectives [1961]. perturbed [1901]. perverse [2458]. Petersson [2326, 2557, 2649]. Pezzo [1552, 2593]. Pfister [2498]. phenomenon [835, 2432]. PHKA [1203]. Piatetski [1385, 2174]. Picard [101, 262, 294, 613, 780, 873, 1173, 1512, 1819]. Pierce [2065]. Pila [1343].

Pillai [293, 2128, 2588]. Pinch [280]. Pisot [628, 1088, 1269, 2020]. place [1122]. Plane [198, 1707, 2387]. Platonic [231]. pleinement [1161]. plus [1583, 1947, 2151]. Poincaré [604, 768, 1471, 2238, 2608, 2619]. point [561, 595, 684, 790, 1458, 1671, 2026, 2316, 2613]. Points [49, 90, 122, 145, 172, 173, 181, 198, 199, 261, 262, 276, 308, 434, 481, 507, 510, 513, 595, 630, 650, 802, 828, 883, 908, 921, 986, 996, 997, 1376, 1383, 1394, 1417, 1425, 1428, 1504, 1684, 1714, 1752, 1770, 1782, 1803, 1856, 1867, 1897, 1971, 1991, 2029, 2049, 2078, 2085, 2087, 2092, 2126, 2132, 2279, 2325, 2354, 2357, 2396, 2401, 2414, 2517, 2518, 2623, 2627, 2640, 2646]. Poitou [570]. polarity [508]. polarized [2618]. Poles [110, 167]. Polický [351]. Poly [292, 414, 1296, 1970]. Poly-Bernoulli [292, 414, 1296]. poly-Euler [1970]. Pólya [607, 640, 1061, 1107, 1982, 2592]. polygamma [1766]. Polygons [474, 1180, 1498, 1906]. polylogarithm [1636]. polylogarithms [614, 1231, 1831, 1951, 2536, 2630]. Polynômes [2215]. Polynomial [42, 162, 311, 335, 354, 355, 375, 418, 517, 545, 561, 575, 630, 787, 866, 970, 1052, 1091, 1290, 1440, 1481, 1487, 1647, 1739, 1861, 1931, 2040, 2058, 2103, 2261, 2453, 2484, 2530, 2553, 2631]. Polynomials [78, 83, 84, 130, 132, 140, 141, 176, 217, 292, 330, 433, 456, 469, 509, 558, 571, 585, 655, 708, 710, 727, 731, 770, 774, 789, 795, 812, 814, 895, 933, 954, 956, 984, 1004, 1006, 1015, 1036, 1049, 1056, 1063, 1068, 1152, 1200, 1241, 1270, 1288, 1294– 1296, 1372, 1374, 1381, 1411, 1412, 1414, 1432, 1438, 1439, 1443, 1449, 1453, 1480, 1498, 1525, 1543, 1574, 1593, 1620, 1631, 1634, 1724, 1740, 1756, 1795, 1810, 1838, 1882, 1939, 1940, 1944, 1970, 1974, 2027, 2068, 2074, 2089, 2133, 2147, 2228, 2229, 2245, 2251, 2260, 2429, 2501, 2525, 2578, 2587, 2615]. polynomials [178, 760, 868, 1062, 1892, 2202, 2215, 2411]. polyquadratic [817]. polytopes [19, 1399]. Pomerance [993, 1393, 1562]. Popular [2199]. Positive [116, 254, 298, 445, 531, 603, 615, 636, 665, 1035, 1076, 1150, 1249, 1287, 1372, 1388, 1449, 1463, 1466, 1479, 1661, 1718, 1733, 1763, 1780, 1896, 1988, 2042, 2126, 2307]. Positivity [348, 1395]. Potential [379]. Potentially [239, 1609, 2084]. pour [6, 93, 380, 778, 1407, 1410]. Power [14, 238, 367, 387, 543, 688, 867, 934, 987, 1013, 1015, 1057, 1156, 1247, 1262, 1314, 1337, 1490, 1570, 1610, 1616, 1693, 1762, 1900, 1908, 1912, 2012, 2161, 2259, 2415, 2419, 2441, 2459, 2465, 2469, 2473, 2495, 2528, 2605, 2613, 2643]. power-full [987]. power-sum [1156]. Powerful [569, 795]. Powers [2, 25, 30, 60, 129, 273, 317, 360, 436, 467, 637, 688, 726, 818, 967, 988, 1011, 1030, 1062, 1073, 1088, 1102, 1160, 1164, 1182, 1367, 1496, 1547, 1808, 1954, 1958, 1967, 2007, 2032, 2207, 2230, 2275, 2344, 2547, 2588]. Prasad [485, 1327, 1645, 2560, 2604]. Preface [1516]. Preimages [1100]. premiers [2252]. preperiodic [276, 2085]. prescribed [355, 610, 1169, 1449, 1502, 2016, 2229, 2241, 2444]. Prescribing [2068, 2276]. prescriptions [2311]. presentation [812]. preserving [635]. primaire [408]. primality [624]. primary [408]. Prime [22, 88, 352, 480, 497, 549, 594, 622, 629, 685, 691, 738, 781, 800, 840, 874, 891, 917, 938, 1007, 1010, 1015, 1057, 1073, 1105, 1123, 1182, 1267, 1290, 1295, 1302, 1330, 1337, 1408, 1450, 1455, 1463, 1496, 1506, 1543, 1656, 1661, 1668, 1708, 1754, 1760, 1793, 1819, 1862, 1947, 1948, 1969, 1972, 2009, 2030, 2046, 2073, 2107, 2116, 2136, 2147, 2161, 2175, 2179, 2191, 2207, 2241, 2252, 2259, 2303, 2336, 2357, 2359, 2379, 2430, 2465, 2474, 2500, 2520, 2525, 2546, 2577, 2618, 2620, 2647]. prime-additive [2107]. prime-independent [1972]. Primefree [1855]. Primes [51, 74, 79, 140, 161, 183, 225, 251, 411, 447, 454, 479, 480, 493, 594, 687, 719, 760, 816, 818, 852, 891, 942, 1081, 1105, 1216, 1330, 1385, 1436, 1450, 1467, 1549, 1578, 1584, 1618, 1624, 1664, 1720, 1728, 1733, 1815, 1821, 1845, 1903, 1954, 1958, 2025, 2046, 2070, 2132, 2136, 2174, 2179, 2259, 2312, 2319, 2444, 2446, 2525, 2531, 2540, 2561, 2562, 2606, 2629]. Primitive [38, 53, 232, 245, 619, 759, 794, 840, 1021, 1177, 1354, 1379, 1454, 1461, 1775, 1828, 1835, 1914, 1948, 2175, 2317, 2373, 2374, 2545, 2546, 2558, 2577, 2605, 2629]. principal [338, 412, 634, 1244, 1545, 2599]. Principe [1407]. principle [318, 654, 1079, 1407, 1552, 1734, 2110, 2166, 2456]. principles [1382, 1983]. pro [457, 828]. pro- [457, 828]. Probabilistic [707, 865, 2392]. Probability [22, 1718]. problem [20, 41, 63, 72, 218, 223, 273, 317, 343, 344, 391, 515, 530, 534, 594, 658, 739, 743, 750, 785, 854, 891, 916, 1047, 1102, 1226, 1255, 1325, 1356, 1427, 1429, 1526, 1547, 1548, 1551, 1563, 1592, 1733, 1801, 1906, 1945, 1985, 1986, 2003, 2047, 2070, 2096, 2097, 2128, 2137, 2153, 2174, 2259, 2280, 2287, 2372, 2402, 2454, 2533, 2588]. problems [68, 556, 793, 1218, 1303, 1309, 1448, 1562, 1958, 2009, 2035, 2457, 2492]. process [400]. Processing [400]. produced [1932]. Producing [2406]. Product [249, 531, 546, 613, 663, 897, 977, 1053, 1082, 1150, 1174, 1313, 1394, 1401, 1463, 1644, 1769, 1804, 1947, 2010, 2025, 2077, 2165, 2309, 2404, 2495, 2518, 2597, 2609]. product-one [2609]. Products [24, 180, 290, 354, 436, 755, 764, 860, 978, 1080, 1126, 1166, 1191, 1209, 1214, 1259, 1320, 1374, 1444, 1565, 1587, 1635, 1718, 1737, 1767, 1832, 1942, 1999, 2054, 2255, 2335, 2459, 2589]. progression [360, 431, 758, 1286, 1298, 2208, 2547]. progression-free [1286, 2208]. progressions [267, 416, 461, 497, 506, 719, 841, 1123, 1166, 1307, 1314, 1478, 1635, 1815, 2130, 2307, 2543]. projection [2102]. projective [931, 1782, 1873, 1975, 2243, 2297]. Proof [118, 300, 311, 509, 611, 898, 979, 993, 1181, 1198, 1231, 1337, 1377, 2022, 2035, 2064, 2438, 2465, 2627]. proofs [890, 1880]. proper [1267, 1885]. Properties [13, 69, 195, 225, 265, 492, 752, 780, 865, 900, 963, 1049, 1099, 1159, 1208, 1320, 1355, 1449, 1456, 1486, 1576, 1638, 1653, 1738, 1742, 1798, 1827, 1835, 1905, 1923, 1974, 2123, 2135, 2163, 2200, 2413, 2434, 2614]. property [276, 383, 430, 1911, 1946, 2002, 2177, 2517]. proposed [2394]. proposition [618]. propriétés [1638]. proprieties [1386]. Prouhet [1906]. proving [624]. pseudo [1647]. pseudo-random [1647]. pseudomoments [2461]. pseudoprimes [97]. pseudorandom [2086]. psi [1601, 2052]. puissances [2, 154]. Pullbacks [229, 1317, 2326, 2458]. pure [1497, 1658, 1857, 1919, 2409]. purely [2119]. Put [635, 788, 2368]. Pythagorean [641, 1014, 1021, 1379, 1773].

Quadratic [25, 40, 45, 48, 160, 173, 201, 209, 259, 296, 351, 371, 381, 419, 485, 514, 516, 522, 576, 636, 663, 664, 697, 716, 725, 758, 759, 818, 852, 901, 902, 905, 932, 977, 994, 1015, 1061, 1072, 1085, 1152, 1219, 1225, 1246, 1269, 1274, 1339, 1340, 1353, 1364, 1372, 1381, 1384, 1423, 1532, 1540, 1654, 1667, 1670, 1690, 1697, 1725, 1778, 1815, 1834, 1860, 1862, 1876, 1878, 1909, 1910, 1927, 1953, 1980, 1989, 1999, 2001, 2016, 2036, 2059, 2069, 2074, 2079, 2092, 2139, 2140, 2149, 2180, 2184, 2189, 2216, 2237, 2295, 2369, 2373, 2427, 2498, 2508, 2527, 2545, 2616]. quadratic [186, 236, 299, 322, 354, 390, 446, 448, 477, 674, 790, 930, 947, 1067, 1499, 1500, 1666, 1720, 1727, 1948, 2019, 2021, 2122, 2153, 2203, 2222, 2241, 2342, 2486, 2526]. quadratics [2001, 2139]. quadrics [1577, 2407]. quadrics-motive [2407]. quadrinomials [2362]. quadruples [619]. Quantitative [388, 1121, 1378, 1628, 1761]. quantities [839]. Quantum [5, 113, 114, 1128, 1527, 1792, 2196]. Quartic [16, 82, 226, 607, 612, 1045, 1169, 1505, 1623, 1634, 1725, 1992, 2460, 2544, 2567]. quartics [1730]. Quasi [147, 496, 540, 564, 660, 1086, 1415, 1544]. quasi-cyclotomic [564]. quasi-group [496]. Quasi-modular [540, 660, 1415]. quasi-periodic [1086, 1544]. quasi-split [147]. quasimodular [523, 2555]. quaternary [381, 1085]. Quaternion [167, 174, 462, 491, 1268, 1284, 1422, 1477, 2090, 2127, 2186, 2291, 2539, 2559, 2598]. quaternionic [197, 1524]. quaternions [1185]. Quelques [1638]. question [1597, 2232, 2318, 2394]. questions [58, 2636]. quick [611]. quicker [928, 1089, 1925]. quinary [2295]. quintic [617, 876, 1497]. quintics [46]. quintuple [1769]. quintuples [1406, 2384]. quotient [1595, 2354, 2453, 2526]. quotients [773, 901, 1117, 1250, 1439, 1595, 1659, 1755, 2219, 2304, 2449, 2490, 2496, 2627].

R [657]. Raabe [647, 1880]. Rabinowitsch [330]. race [88]. races [2303]. Rademacher [426, 725]. radical [1332]. Rallis [1130, 2467]. Ramakrishnan [485]. Ramanujan [155, 269, 495, 590, 632, 683, 726, 746, 824, 975, 979, 988, 1059, 1207, 1210, 1278, 1321, 1375, 1437, 1451, 1510, 1528, 1529, 1614, 1731, 1732, 1890, 2050, 2171, 2264, 2359, 2421, 2518, 2542]. Ramanujan-like [726]. Ramanujan-style [1059]. Ramanujan-type [1207]. Ramification [535, 670, 1026, 1047, 1723, 2110, 2415]. ramifications [2541]. ramified [98, 158, 1584, 1900, 1969, 2012, 2132, 2300]. random [22, 136, 771, 1076, 1114, 1153, 1608, 1647, 1844, 2291, 2310, 2600]. randomised [2226]. randomized [2011]. randomly [2062]. Randomness [580, 2456]. rang [2028]. range [157]. Rank [38, 125, 214, 258, 259, 285, 353, 356, 449, 553, 562, 636, 662, 712, 715, 838, 951, 1339, 1393, 1435, 1586, 1631, 1778, 1910, 2010, 2028, 2146, 2170, 2268, 2348, 2457]. rank-one [259]. Rankin [147, 182, 342, 769, 860, 1141, 1587, 1895, 1920, 2198]. Ranks [842, 1483, 1674, 2132, 2156, 2401, 2422, 2566, 2581]. rate [1117, 1176, 1595, 2402]. ratio [1302, 2043]. Rational [28, 81, 122, 181, 216, 259, 369, 412, 452, 510, 565, 595, 650, 668, 735, 802, 828, 858, 895, 973, 1047, 1054, 1160, 1230, 1233, 1283, 1291, 1376, 1396, 1425, 1428, 1430, 1443, 1444, 1458, 1505, 1572, 1623, 1684, 1770, 1818, 1867, 1882, 1968, 1971, 2029, 2030, 2032, 2061, 2132, 2277, 2279, 2286, 2299, 2347, 2401, 2414, 2464, 2641, 2646, 2648]. Rationality [434, 2576, 2614]. rationals [81, 534, 1074, 2562]. rationnelles [81]. rationnels [81]. Ratios [179, 301, 1031, 2502]. ray [725, 2637]. rayon [2637]. real [40, 91, 127, 201, 322, 504, 516, 576, 663, 674, 725, 754, 790, 805, 878, 896, 907, 912, 947, 1160, 1224, 1438, 1441, 1667, 1670, 1876, 1907, 1989, 2032, 2079, 2140, 2197, 2427, 2485, 2508]. réalisables [148, 1297]. Realizable [1297]. realization [633]. realizations [2160]. reals [1607, 2304]. Recaman [223, 1562]. Reciprocal [150, 1164, 1780, 1976, 2317, 2411]. reciprocals [834, 1657]. réciprocité [1106]. reciprocity [611, 1106, 1188, 1374, 1544, 2006, 2514, 2626]. recurrence [143, 209, 371, 804, 1353, 1452, 1572, 1654, 1779, 1834, 1944, 2379, 2436, 2454, 2569]. recurrences [62, 506, 780, 2043]. recursive [492, 688]. reduced [522, 1689, 2130, 2307, 2614]. reducibility [757, 1939]. reducible [699, 704, 2074, 2553]. reduction [790, 800, 1163, 1268, 1410, 1522, 1846, 2085, 2132, 2222, 2268, 2562]. Reductions [465, 1165, 1211, 1699, 1924, 2642]. reductive [464, 505]. Ree [2283]. Refined [733, 1026, 2248, 2370, 2509]. refinement [1434]. refinements [1485, 1836]. Refining [1930]. Reflection [1983]. reflexes [195]. Reflexive [2639]. regarding [944, 2080]. region [1411]. regions [1120]. Regular [772, 1030, 1044, 1085, 1299, 1590, 1691, 1716, 1736, 1798, 1879, 2240]. regularity [1833, 1912]. Regularized [29, 604, 642, 680, 1554, 2106, 2513]. Regulators [1599, 1698, 1751, 1813, 2479]. related [15, 212, 267, 336, 344, 382, 414, 541, 571, 774, 935, 955, 992, 1063, 1171, 1233, 1285, 1291, 1316, 1367, 1395, 1454, 1464, 1564, 1582, 1589, 1592, 1593, 1635, 1639, 1644, 1665, 1714, 1743, 1774, 1834, 1922, 2253, 2327, 2383, 2459, 2506, 2587]. relation [444, 1569, 1764, 2438, 2520, 2647]. Relations [29, 150, 374, 572, 642, 702, 804, 1437, 1658, 1691, 1779, 1789, 1919, 1937, 1944, 1945, 2051, 2145, 2205, 2234, 2240, 2627]. relative [489, 1595, 1811, 2203, 2452]. relatively [22, 1661, 2161]. relatives [2018]. remainder [212, 2488]. Remark [33, 142, 351, 1147, 1187, 1458]. remarkable [1952]. Remarks [409, 483, 605, 657, 946, 1766, 2455, 2548]. remarque [1147]. Rényi [2349]. repeated [1251, 1674, 2294, 2380]. Representation [166, 269, 419, 452, 597, 615, 688, 931, 994, 1084, 1372, 1598, 1830, 1842, 1947, 2016, 2038, 2059, 2145, 2311, 2360, 2506, 2560]. Representations [51, 57, 101, 177, 239, 243, 321, 381, 395, 471, 599, 619, 693, 694, 697, 704, 734, 752, 806, 821–823, 825, 849, 852, 1018, 1046, 1087, 1101, 1131, 1138, 1143, 1165, 1190, 1268, 1275, 1279, 1337, 1397, 1409, 1482, 1484, 1518, 1520, 1556, 1609, 1621, 1675, 1697, 1815, 1816, 1846, 1860, 1872, 1895, 1943, 1957, 2005, 2019, 2069, 2187, 2189, 2225, 2256, 2295, 2338, 2406, 2409, 2413, 2445, 2467, 2487, 2489, 2492, 2527, 2537, 2559, 2576, 2634]. represented [788, 1219, 2161]. representing [1927]. réseau [1147]. réseaux [2214]. resembling [1573]. Residuacity [410, 550]. residually [952, 1606, 1921, 2601]. Residue [134, 355, 578, 687, 691, 1247, 1447, 1511, 1725, 1863, 1886, 2130, 2241, 2303, 2307, 2473]. residues [757, 758, 1188, 1702, 2605]. resolution [338]. resolvents [2586]. Resonance [1587]. Resonances [1316]. respect [420, 697, 2629]. respecting [2045]. Restricted [485, 1082, 1332, 1615, 1680, 1750, 1806, 2265, 2542]. Restricting [8, 94]. restriction [59, 287]. restrictions [193]. result [777, 796, 1550]. resultant [1119]. Resultants [1241, 1438, 2202]. results [63, 86, 242, 273, 528, 641, 779, 890, 955, 972, 995, 1014, 1072, 1141, 1316, 1454, 1462, 1482, 1765, 1972, 2064, 2339, 2454, 2459]. Reversed [1838, 2228]. revisited [343, 1719, 2063]. rhombus [2044]. Riemann [6, 53, 73, 92, 93, 185, 202, 227, 341, 362, 369, 442, 458, 488, 584, 620, 767, 773, 875, 886, 1156, 1164, 1191, 1193, 1256, 1259, 1308, 1411, 1529, 1548, 1598, 1619, 1640, 1731, 1764, 1779, 1811, 1854, 1901, 2078, 2197, 2253, 2461, 2507, 2607]. Riesel [583, 639, 1262]. Riesz [387, 1492, 2098]. Riesz-type [1492]. Right [306, 630, 1613]. right-angled [630]. rigidity [2397]. Rigorous [2226]. ring [21, 279, 446, 463, 507, 543, 819, 846, 885, 1734, 1886, 2442, 2571]. rings [310, 528, 558, 966, 1013, 1023, 1388, 1447, 1561, 1843, 1863, 1952, 2408, 2484, 2633]. Riyadh [189]. Rodriguez [1111, 1904]. Rodriguez-Villegas [1111]. Rogers [495, 975, 1210, 1437, 1528, 1732, 2421, 2518]. Roman [2063]. Root [840, 887, 913, 933, 1136, 1513, 1775, 1981, 2175, 2269, 2315, 2382, 2605, 2629]. rooted [2088]. roots [38, 307, 1433, 1719, 1759, 1914, 2260, 2337, 2577]. Rotated [554]. rotation [2215]. rotationally [2215]. Rotations [462, 2515]. rounded [2427]. rs [164]. RSA [610]. Rubin [115, 562]. run [1979, 2554]. run-length [1979, 2554]. Ruzsa [689].

S [189, 598]. saddle [2613]. saddle-point [2613]. Saito [2326]. Salajan [1507]. Salem [1260, 1311, 1493]. same [531, 597, 913, 1228, 2057]. Samuel [618]. Sárközy [2533]. Sarmah [1198]. Sarnak [2357]. Satake [825]. satisfy [1193]. satisfying [29, 110, 214, 589, 1934, 2161]. Sato [224, 2381, 2432, 2556]. saturation [1371]. Saudi [189]. scaffolds [837, 2101]. scalars [887]. Schäffer [1359, 1627]. Schanuel [1336]. Scharaschkin [1179]. Scharlau [1434]. scheme [130, 372]. schemes [528, 535, 1361, 1495, 1704, 1950]. Schmidt [1596, 1605, 1695, 2213]. Schneider [1545, 2239]. Schreier [521, 547, 914, 1114, 1221]. Schröder [396, 1721, 2123]. Schur [411]. Schur-type [411]. Schweizer [311]. second [531, 684, 691, 999, 1084, 1322, 1460, 1813, 1855, 1866, 1992, 2228, 2365, 2392]. sections [1362]. sector [1068]. Segal [1561]. Selberg [6, 73, 92, 93, 147, 182, 257, 342, 736, 769, 860, 1141, 1172, 1227, 1278, 1587, 1854, 1895, 1920, 2198, 2385]. Selectivity [491, 577, 2186]. Self [312, 651, 985, 1164, 1357, 1872, 2487]. Self-approximation [312]. Self-conjugate [651, 985]. Self-dual [1872, 2487]. Self-reciprocal [1164]. Self-similar [1357]. Selmer [1, 21, 87, 156, 904, 909, 1090, 1419, 1483, 2225, 2236, 2367]. semi [239, 1179, 2084, 2617]. semi-abelian [1179, 2617]. semi-stable [239, 2084]. semigroup [1588, 2037]. semigroups [391, 615, 1315, 1356, 2404]. semisimple [829]. semistable [1277]. separated [770, 1469]. September [152, 332, 518, 783, 1024, 1328, 1650, 1993, 2289, 2549]. septic [2570]. sequence [50, 143, 210, 506, 928, 963, 1385, 1507, 1576, 1754, 1839, 1847, 2037, 2166, 2365]. Sequences [102, 165, 328, 335, 451, 467, 520, 544, 598, 706, 737, 786, 798, 808, 862, 903, 923, 1101, 1171, 1283, 1286, 1295, 1370, 1398, 1430, 1566, 1569, 1604, 1607, 1790, 1855, 1868, 1926, 1927, 1931, 1966, 1985, 2075, 2077, 2108, 2196, 2218, 2242, 2272, 2310, 2319, 2327, 2379, 2397, 2404, 2436, 2454, 2488, 2543, 2556, 2569, 2584, 2609]. série [35]. Series [14, 15, 29, 35, 37, 56, 166, 167, 229, 256, 272, 338, 367, 412, 478, 525, 572, 591, 604, 617, 623, 666, 768, 788, 801, 822, 823, 881, 905, 1028, 1035, 1096, 1124, 1138, 1139, 1176, 1195, 1243, 1266, 1310, 1384, 1397, 1443, 1451, 1465, 1471, 1523, 1536, 1537, 1545, 1617, 1644, 1696, 1703, 1746, 1769, 1783, 1816, 1829, 1834, 1839, 1881, 1882, 1900, 1912, 1990, 2004, 2012, 2020, 2056, 2060, 2065, 2066, 2078, 2103, 2106, 2185, 2205, 2219, 2254, 2255, 2263, 2292, 2296, 2334, 2343, 2360, 2367, 2390, 2415, 2416, 2468, 2513, 2534, 2537, 2555, 2608, 2619, 2632]. series [858, 1052, 1064, 1548, 1888, 2061, 2483]. serpentine [278]. Serre [448, 1363, 1990]. Serret [2176]. sesqui [1009]. sesqui-harmonic [1009]. set [123, 188, 282, 565, 731, 922, 1065, 1452, 1574, 1808, 2109, 2453]. Sets [32, 50, 102, 103, 116, 131, 184, 267, 376, 470, 546, 561, 757, 787, 794, 841, 956, 1060, 1062, 1076, 1097, 1150, 1153, 1334, 1399, 1401, 1418, 1458, 1647, 1671, 1691, 1713, 1733, 1863, 1885, 2038, 2076, 2077, 2159, 2165, 2172, 2192, 2199, 2215, 2244, 2304, 2526]. setting [1310]. settings [1075]. several [47, 2159, 2600]. sextenary [2189]. sextic [607, 616, 2305, 2465]. sextics [1109]. sextuples [2641]. Shafarevich [169, 313, 627, 728, 1074, 1106, 2209, 2486]. Shalika [960]. Shamir [130]. Shanks [207]. Shapiro [1385, 2174]. Sharp [307, 630, 1487, 1494, 1644, 1698, 1834]. sheaves [646, 672, 2458]. Shift [278, 441, 976]. Shifted [202, 530, 813, 1326, 1334, 1394, 1537, 1746, 1855, 1890, 1990, 2000, 2113, 2171, 2324, 2497, 2603]. shifts [441, 1256, 2287]. Shimura [203, 827, 907, 1100, 1984, 2048, 2539, 2582]. Shintani [1122, 2091, 2121, 2508]. Short [163, 300, 979, 1216, 1290, 1450, 1566, 1578, 1579, 1708, 1809, 1870, 1880, 1926, 1966, 1972, 2173, 2262, 2444, 2472, 2607]. Short-interval [1870]. Shortened [1944]. shortest [2107]. shtuka [250]. Shuffle [29, 642, 1082, 1191, 1213, 1804]. Sidon [1418, 2199]. Siegel [8, 94, 229, 313, 417, 443, 478, 525, 572, 680, 696, 704, 711, 826, 853, 920, 939, 1051, 1243, 1583, 1599, 1784, 1789, 1829, 2051, 2329, 2366, 2579, 2582, 2597]. Sierpiński [583, 639]. sieve [99, 363, 916, 1000, 1632, 2226, 2318, 2419, 2643]. Sifting [363]. Sign [69, 350, 836, 1108, 1828, 1973, 2301, 2302, 2510, 2590]. signature [874, 1688, 2480]. Signed [1122, 1468, 2435]. signs [1228, 2248]. similar [451, 1357, 1550]. similitude [799]. simple [557, 685, 979, 1013, 1074, 1482, 1739, 2615, 2618]. simplest [377, 1045]. simplex [1487]. simplices [630, 744]. simultanée [820]. Simultaneous [450, 820, 852, 1116, 1157, 1160, 1184, 1874, 2032, 2150, 2510, 2605]. sine [202, 1476, 2386]. Single [1231, 1245]. single-cuspidal [1245]. Single-valued [1231]. Singular [634, 855, 863, 906, 1425, 1444, 1622, 1637, 1869, 2100, 2335]. singularities [2134]. singulier [863]. Sinnott [733, 1022]. six [2591]. sixth [238, 934, 2565]. size [89, 825, 1468]. sizes [546, 1332]. skew [1284]. skew-Hermitian [1284]. Skolem [2454]. Skorobogatov [1996]. slope [547]. slopes [2062, 2177, 2243]. Slow [2176]. Small [32, 146, 172, 241, 283, 299, 522, 599, 816, 1069, 1138, 1178, 1286, 1401, 1532, 1543, 1643, 1675, 1698, 1856, 1947, 2016, 2191, 2338, 2423, 2506, 2527, 2532, 2552, 2620]. small-degree [1698]. smallest [691, 1215, 1251, 1976, 2274, 2350, 2479]. smooth [198, 290, 416, 510, 2105, 2279, 2387, 2458, 2517]. smoothing [2134]. sociable [142]. solids [231]. solution [1597, 1801]. Solutions [113, 123, 162, 299, 377, 418, 732, 759, 930, 961, 1033, 1125, 1157, 1332, 1389, 1454, 2119, 2191, 2224, 2233, 2265, 2393, 2521, 2545]. Some [9, 13, 69, 79, 82, 86, 90, 106, 132, 195, 213, 242, 253, 313, 362, 409, 424, 441, 446, 471, 509, 515, 589, 592, 619, 657, 687, 710, 769, 847, 878, 900, 960, 963, 972, 998, 1005, 1049, 1072, 1089, 1111, 1120, 1154, 1190, 1203, 1233, 1291, 1299, 1353, 1386, 1437, 1449, 1462, 1478, 1482, 1517, 1547, 1591, 1638, 1639, 1646, 1665, 1670, 1677, 1691, 1869, 1876, 1877, 1885, 1925, 1971, 1992, 2017, 2020, 2031, 2051, 2064, 2078, 2079, 2108, 2135, 2192, 2219, 2249, 2283, 2292, 2295, 2339, 2357, 2410, 2433, 2443, 2558, 2632]. sometimes [936]. Somos [209, 371, 1353, 1654, 1834]. son [820]. Soundararajan [416]. sous [1318, 1638]. sous-groupes [1638]. sous-variétés [1318]. space [286, 604, 692, 699, 854, 1583, 1786, 1975, 2151, 2286, 2383, 2434, 2612]. spaces [383, 417, 556, 748, 989, 1782, 1943, 2297, 2580]. spacing [767]. spacings [18]. spanned [748]. Sparse [131, 433, 2525]. Special [40, 296, 326, 384, 498, 801, 884, 1036, 1140, 1167, 1310, 1416, 1779, 1990, 2004, 2090, 2140, 2165, 2285, 2582, 2616, 2635]. specialization [1212]. specializations [1795]. specified [2084]. spectra [383, 628, 1269, 2408]. Spectral [192, 271, 468, 555, 621, 1000, 1422, 1703, 2621, 2626]. spectroscopy [538]. spectrum [1933, 2396]. speed [2304]. sphere [2598]. Spherical [1020, 2412]. Spiegelungssatz [739]. spin [2520, 2647]. Spinor [1284, 2045]. Split [19, 147, 218, 505, 2198, 2383]. Splitting [659, 1767, 1934]. sporadic [2615]. spt [1215, 1251, 2294]. spt-crank [2294]. Square [168, 173, 326, 361, 384, 387, 499, 631, 742, 820, 874, 895, 920, 944, 991, 1074, 1217, 1513, 1553, 1775, 1930, 2021, 2190, 2212, 2318, 2474, 2641]. Square-free [173, 631, 895, 920, 1775, 2190]. squared [1389]. Squarefree [1307, 1459, 2099, 2229]. Squarefull [1298, 1457]. Squares [121, 161, 269, 334, 406, 419, 431, 447, 594, 850, 891, 962, 1006, 1182, 1219, 1351, 1450, 1455, 1496, 1547, 1549, 1634, 1777, 1860, 1989, 2046, 2055, 2069, 2166, 2179]. Srivastava [1077]. Stability [1134]. Stable [239, 769, 1163, 2084, 2177, 2445]. stage [618]. standard [313]. Stanley [750]. star [527, 1311, 1887, 2411]. star-like [1311]. Stark [115, 427, 434, 562, 663, 674, 1029]. Statistics [436, 1873, 2556]. Staudt [440, 1464, 1743]. Steinberg [2413, 2560]. Steinitz [98, 643, 877, 1980]. Steve [1130]. Stichtenoth [1508]. Stickelberger [28, 659, 797, 1274, 1488, 2237]. Stickelberger-type [797]. Stieltjes [166, 673, 1233, 1291, 2603]. Stirling [752, 999, 1208, 1288, 1322, 1476, 1965, 2178, 2392]. Stochastic [2538]. Stone [706]. Strange’ [1792]. stretched [2325]. Strictly [1085, 1736]. string [1138]. strip [248, 306, 1616, 2258]. stripping [704]. Strong [241, 389, 1572, 1911, 2583]. stronger [295, 2394]. Strongly [1529, 2240, 2488]. Structure [32, 123, 131, 146, 284, 796, 837, 1226, 1390, 1513, 1810, 2165, 2236, 2404]. structures [553, 1055, 1719, 1756, 2163]. stuffle [1313]. Sturmian [855, 1195, 2157, 2390]. style [1059, 1716]. subconvexity [2575]. subcoverings [827]. subcovers [2370]. subfield [1061, 1692]. subfields [485, 1813, 2476]. subgroup [372, 471, 2393]. subgroups [196, 218, 265, 412, 523, 927, 949, 1000, 1169, 1517, 1638, 2023, 2141, 2299, 2369, 2375, 2494, 2503]. subject [1506]. subpartitions [2421]. Subrings [1013]. Subsequence [1218, 1303]. Subsequences [118, 520, 1955, 2323]. Subset [497, 2320, 2633]. subsets [1647, 2086, 2208]. subshifts [1431]. subspace [1121, 1596, 2213]. substitution [2553]. Substitutions [2311, 2631]. subtractive [744]. subvarieties [1179]. Sufficient [2101]. suite [210]. Sum [118, 121, 158, 163, 165, 204, 256, 300, 413, 419, 452, 566, 578, 688, 695, 725, 777, 850, 862, 897, 938, 944, 1053, 1156, 1191, 1196, 1219, 1267, 1367, 1377, 1398, 1424, 1450, 1490, 1544, 1566, 1575, 1595, 1646, 1680, 1722, 1928, 1955, 1959, 1966, 1985, 1988, 2017, 2066, 2080, 2179, 2207, 2261, 2292, 2317, 2323, 2324, 2341, 2403, 2457, 2512, 2534, 2563, 2584, 2632, 2645]. sum-of-digits [578]. sum-of-divisors [2066]. sum-of-proper-divisors [1267]. summands [792, 1456, 2155]. Summation [66, 100, 475, 1310, 1559, 2485]. summations [1233, 1291]. Sums [23, 36, 67, 126, 161, 180, 193, 269, 275, 290, 334, 406, 409, 425, 426, 447, 470, 479, 503, 514, 547, 575, 611, 619, 640, 703, 710, 747, 834, 945, 962, 971, 987, 988, 997, 1079, 1083, 1086, 1093, 1153, 1159, 1175, 1216, 1217, 1314, 1316, 1326, 1341, 1364, 1374, 1399, 1447, 1455, 1465, 1486, 1501, 1549, 1565, 1587, 1610, 1615, 1624, 1660, 1714, 1727, 1783, 1802, 1805, 1826, 1840, 1844, 1864, 1866, 1870, 1886, 1890, 1954, 1959, 1967, 1973, 1989, 1998, 2000, 2030, 2037, 2050, 2083, 2171, 2195, 2306, 2328, 2335, 2337, 2342, 2350, 2436, 2472, 2484, 2500, 2547, 2565, 2633]. sums [104, 133, 144, 196, 200, 213, 238, 376, 415, 454, 571, 580, 592, 614, 772, 813, 919, 934, 984, 1048, 1056, 1186, 1199, 1218, 1303, 1331, 1340, 1389, 1478, 1579, 1636, 1672, 1702, 1729, 1762, 1808, 1888, 1908, 1987, 2021, 2058, 2064, 2098, 2105, 2113, 2178, 2223, 2253, 2257, 2275, 2284, 2304, 2362, 2411, 2443, 2447, 2459, 2497, 2574]. sumset [32, 50]. sumsets [982, 1065, 1408, 1468, 2435]. Sun [25, 509, 1377]. sunset [1223, 2407]. super [2207]. super-powers [2207]. supercharacter [2195]. Supercharacters [1079, 2565]. supercongruence [1721]. Supercongruences [1063, 1111, 1168, 1612, 1881, 1928, 2022]. Supercuspidal [1482, 1556, 2541]. superelliptic [244, 1118, 1483]. supérieurs [1407]. superquadratic [2532]. Supersingular [183, 251, 493, 1018, 1819, 2489, 2615]. superstring [1078]. support [114, 785, 872]. surface [496, 1996, 2309, 2382, 2571]. surfaces [90, 101, 122, 149, 159, 261, 262, 294, 348, 568, 613, 753, 873, 1067, 1074, 1172, 1173, 1371, 1552, 1598, 1819, 1854, 2209, 2315, 2517, 2558, 2593]. surjections [1659]. Suzuki [2283]. Swinnerton [187, 214, 270]. Swinnerton-Dyer [187, 214, 270]. Sylow [866]. Sylvester [1474, 2037]. symbol [2352]. symbols [899, 1703, 2016, 2473, 2617]. Symmetric [130, 168, 216, 387, 499, 531, 742, 804, 814, 991, 1217, 1289, 1531, 1616, 1693, 1734, 1786, 2002, 2038]. symmetric-square [499, 991]. Symmetries [826, 1052, 1432]. symmetrized [1671]. symmetry [1357]. symplectic [87, 645, 777, 1140, 1244, 2102, 2406]. system [1323]. systems [276, 335, 578, 587, 635, 659, 745, 787, 814, 1033, 1087, 1818, 1933, 2130, 2307, 2400]. Szekeres [841]. Szpiro [1277].

T [1001]. table [2479]. tagged [2155]. tails [1787]. taking [781, 1969]. Takloo [2560]. Takloo-Bighash [2560]. Tamagawa [149]. Tame [1435, 1546, 1813, 2305, 2544]. tamely [98]. Tannakian [1653]. targets [2213]. Tarry [1906]. Tate [58, 169, 183, 224, 281, 294, 313, 481, 570, 627, 728, 753, 1074, 1163, 2084, 2209, 2381, 2432, 2486, 2518, 2556, 2617]. Tchebotarev [1144]. Teitelbaum [1185]. tempered [1327]. tensor [1938]. tenth [2096]. term [212, 387, 934, 1375, 1510, 1865, 2097, 2381, 2547]. terms [541, 598, 620, 635, 645, 803, 930, 1007, 1465, 1507, 1927, 2198, 2350]. Ternary [41, 178, 419, 902, 977, 994, 1004, 1015, 1364, 1372, 1381, 1423, 1532, 1697, 1727, 1733, 1736, 1815, 1860, 1948, 2009, 2069, 2119, 2477]. test [179, 301, 2331]. Testing [115]. tetrahedral [2218]. Tetranomial [864]. th [867, 1247, 1338, 1808, 1838, 1914, 1927, 2091, 2242, 2259, 2269]. Thabit [1356]. Thaine [1871]. their [40, 57, 195, 196, 398, 495, 542, 684, 710, 723, 765, 885, 971, 999, 1301, 1345, 1404, 1484, 1703, 1795, 1813, 1835, 1876, 1919, 2005, 2079, 2140, 2163, 2180, 2219, 2403, 2416, 2437]. theme [477, 1949]. Theor [2647]. theorem [192, 216, 232, 254, 322, 393, 411, 439, 527, 529, 557, 662, 667, 689, 706, 786, 898, 982, 1007, 1012, 1058, 1121, 1144, 1187, 1203, 1233, 1238, 1239, 1291, 1333, 1386, 1405, 1434, 1464, 1541, 1550, 1560, 1596, 1634, 1737, 1743, 1748, 1760, 1818, 1826, 1831, 1871, 1930, 1960, 2058, 2112, 2165, 2176, 2206, 2213, 2281, 2349, 2366, 2378, 2383, 2470, 2474, 2522, 2610, 2617, 2621]. theorems [339, 455, 797, 847, 974, 1424, 1589, 1932, 2435, 2530]. theoretic [2465]. theories [1614]. Theory [59, 91, 93, 94, 106, 107, 129, 134, 147, 189, 237, 269, 271, 302, 310, 384, 410, 493, 517, 550, 655, 659, 670, 675, 707, 738, 739, 808, 812, 847, 876, 894, 909, 939, 951, 957, 1001, 1128, 1190, 1271, 1291, 1303, 1304, 1375, 1426, 1448, 1462, 1471, 1503, 1510, 1546, 1627, 1628, 1722, 1743, 1770, 1822, 1830, 1848, 1882, 2009, 2012, 2079, 2080, 2132, 2139– 2141, 2144, 2160, 2222, 2473, 2508, 2548, 2622, 2642–2646, 2648, 2649]. There [97]. thermodynamic [127]. these [2203]. theta [100, 334, 475, 591, 604, 647, 680, 684, 724, 799, 815, 826, 876, 1135, 1229, 1278, 1384, 1492, 1523, 1554, 1598, 1880, 1922, 2219, 2263, 2292, 2455, 2493, 2548]. theta-function [1278]. thin [50, 1097]. third [1237, 2464]. Thomas [1508]. Three [419, 594, 686, 712, 735, 812, 814, 962, 1455, 1484, 1584, 1778, 1794, 1872, 1945, 2046, 2136, 2154, 2264, 2337, 2457, 2567, 2583]. Three-dimensional [1484]. Threefield [852]. threefold [392]. threefolds [753, 2160]. Thue [4, 16, 123, 329, 377, 864, 1171, 1271, 1485, 1939, 2329, 2567]. Titchmarsh [2153, 2372]. top [1853, 2298]. topological [1905]. topologies [1458]. topology [2473]. topos [2602]. toric [1093]. toriques [1318]. Tornheim [1888]. Torsion [49, 90, 105, 160, 236, 259, 372, 405, 595, 849, 908, 1118, 1169, 1383, 1505, 1642, 1752, 1878, 1897, 2028, 2129, 2277, 2299, 2340, 2369, 2375, 2478, 2648]. torsions [582]. Torus [340, 623, 2093, 2438]. totality [2175]. totally [91, 504, 603, 907, 1222, 1287, 1441, 1780, 1907, 2203, 2485, 2552, 2558]. totient [212, 1575, 2441]. tower [1339]. towers [260, 669, 828, 1192, 1265, 1300, 1499, 1508, 1817, 1848, 2177, 2566, 2581]. Trace [39, 392, 494, 751, 885, 990, 1027, 1172, 1227, 1260, 1493, 1502, 1780, 1833, 2099]. traces [436, 721, 2005, 2409, 2541, 2586]. Transcendence [9, 106, 272, 1113, 1300, 2231, 2612]. transcendent [1352, 1731, 1827]. transducers [2176]. transform [14, 1249]. transformation [1452, 1492, 1697, 1864, 2161]. transformations [1164]. transforms [72, 423, 542, 804, 1232, 2145]. Transition [813]. Translation [352]. trees [1311, 2088, 2180, 2361]. tri [2383]. tri-Hermitian [2383]. triangle [23, 1099, 1613, 2044, 2551]. Triangles [735, 1623, 2389, 2535]. triangular [2218]. Trigonometric [729, 1411, 2004, 2061]. Triple [230, 249, 880, 1377, 1727, 1928, 2138, 2473, 2477]. Triples [641, 1014, 1021, 1379, 1760, 1773, 2159]. Trivial [643, 668, 728, 790, 1713, 2148]. tropical [2238]. Trotter [2318]. Truncated [242, 1324, 1769, 1904, 2339]. truncation [806, 1046]. tuple [2430]. tuples [531, 955, 1661, 1921]. Turán [1264]. twelve [1124]. twice [2557, 2649]. Twin [760, 942, 2070]. twist [1274]. Twisted [84, 219, 294, 721, 984, 1134, 1186, 1220, 1477, 1585, 1726, 1909, 1938, 1950, 2223, 2288, 2322, 2425, 2451, 2540, 2575, 2579]. twists [236, 259, 450, 499, 609, 636, 817, 913, 955, 1483, 1859, 1910, 2427]. Two [41, 54, 273, 417, 443, 452, 478, 526, 553, 587, 746, 850, 855, 891, 923, 934, 974, 976, 977, 1064, 1081, 1163, 1182, 1278, 1302, 1330, 1367, 1450, 1453, 1496, 1547, 1592, 1593, 1598, 1609, 1681, 1856, 1928, 1947, 1968, 2022, 2030, 2043, 2179, 2201, 2259, 2275, 2337, 2342, 2344, 2372, 2379, 2386, 2410, 2494, 2562, 2564, 2589, 2597]. two-dimensional [855, 1593]. two-term [934]. Two-variable [1081, 2201, 2379, 2386]. type [68, 90, 166, 337, 411, 441, 565, 593, 604, 619, 660, 743, 763, 797, 812, 824, 915, 916, 1102, 1151, 1161, 1164, 1172, 1185, 1207, 1270, 1375, 1397, 1431, 1437, 1492, 1508, 1510, 1518, 1528, 1574, 1596, 1606, 1639, 1660, 1722, 1758, 1766, 1888, 1897, 1904, 1921, 1958, 1960, 2023, 2084, 2089, 2108, 2141, 2161, 2232, 2325, 2353, 2420, 2478, 2632]. Types [195, 1019, 1556, 2498]. Typically [2340].

U [680]. Ulam [2397]. Ulas [2631]. umbral [812, 1200]. unbounded [1847, 2343]. uncertainty [318, 1079]. undeformed [646]. unequal [698, 1958]. Unicity [1556]. Unified [812, 1600, 2047]. Uniform [380, 458, 751, 802, 1217, 1388, 2400]. uniforme [380]. uniformity [1225]. uniformly [835, 1370, 2244]. unifying [2089, 2108]. unimodal [2196]. Unimodular [776, 1533, 1980, 2291, 2448]. union [172]. unipotent [1140, 1491]. Unique [230, 798, 1101, 1720, 2389]. Uniqueness [83, 860]. unit [47, 274, 1205, 1222, 1653, 1658, 1791, 1979, 2515]. Unitary [179, 197, 799, 815, 825, 1114, 1146, 1520, 2371, 2559, 2576]. units [82, 203, 351, 462, 504, 663, 674, 805, 878, 912, 947, 963, 1269, 1351, 1447, 1584, 1599, 1777, 1863, 1921, 2350, 2436]. unity [1719, 1914, 2337]. univariate [307]. Universal [157, 1015, 1018, 1237, 1261, 1381, 1815, 1943, 2306]. Universality [185, 378, 393, 584, 1058, 1333, 1572, 1701, 2607, 2621]. University [189]. unramified [33, 112, 147, 457, 1670, 2121, 2198, 2237, 2412]. unreasonable [2397]. unstable [2559]. upon [812]. Upper [146, 673, 886, 1039, 1338, 1373, 1407, 1487, 1911, 1969, 2119, 2385]. Using [102, 280, 287, 322, 582, 629, 722, 738, 1101, 1281, 1687, 1839, 1942, 2008].

valeurs [81, 210, 2215]. Valuation [528, 547, 1655, 2040, 2075]. Valuations [1322, 1432, 1751, 1965, 2083, 2611]. valuative [1431, 1808]. Value [192, 266, 306, 339, 452, 468, 531, 581, 679, 731, 775, 1060, 1062, 1066, 1083, 1199, 1394, 1763, 2021, 2420, 2446, 2471, 2610]. value-distribution [2471]. valued [42, 417, 539, 558, 665, 853, 868, 956, 1231, 1283, 1430, 1810, 2455, 2548]. Values [40, 67, 76, 81, 84, 86, 168, 182, 210, 268, 296, 313, 320, 326, 347, 359, 369, 374, 384, 444, 468, 503, 527, 566, 575, 605, 634, 642, 671, 721, 748, 769, 773, 787, 795, 801, 813, 880, 884, 895, 905, 944, 948, 971, 1035, 1040, 1044, 1082, 1162, 1167, 1191, 1213, 1220, 1259, 1266, 1290, 1310, 1313, 1368, 1416, 1424, 1615, 1665, 1680, 1705, 1755, 1779, 1787, 1804, 1812, 1854, 1859, 1875, 1896, 1988, 1990, 1998, 2004, 2073, 2080, 2088, 2091, 2092, 2117, 2140, 2161, 2173, 2211, 2215, 2261, 2267, 2285, 2322, 2336, 2341, 2403, 2451, 2472, 2507, 2512, 2516, 2525, 2611, 2616, 2635]. values [354, 424, 555, 702, 781, 1032, 1067, 1267, 1280, 1380, 1481, 1616, 1727, 1887, 1987, 2031, 2043, 2078, 2234, 2459, 2574, 2582, 2591]. Vandermonde [1868]. Vandiver [658, 739]. Vanishing [21, 1096, 1536, 1859, 2056, 2111, 2203, 2316, 2350, 2510, 2626]. variable [770, 771, 1081, 1672, 2201, 2379, 2386]. variables [41, 47, 759, 1204, 1246, 1364, 1469, 1794, 1948]. variance [461]. variant [2128]. variants [769, 1742]. Variation [1949]. variational [529]. Variations [157, 477, 683, 955, 1237, 1541]. variétés [1161, 1318]. Varieties [128, 145, 172, 217, 270, 379, 637, 694, 785, 800, 828, 1013, 1161, 1470, 1767, 1879, 1897, 1929, 1960, 2132, 2146, 2299, 2318, 2354, 2401, 2414, 2429, 2437, 2449, 2478, 2582, 2585, 2618, 2646]. variety [599, 685, 728, 1179, 1842]. various [1075, 1817, 1848]. Vaughan [916]. vector [417, 556, 748, 853, 2455, 2548]. vector-valued [417, 853, 2455, 2548]. vectors [349, 406, 920, 1675]. version [187, 388, 1597, 1807, 1894, 2011, 2206, 2394]. versions [1761]. vertical [1256]. vertices [1428]. via [59, 126, 855, 949, 1165, 1295, 1319, 1386, 2426, 2437, 2578, 2610]. vicinity [514]. Vigneras [1545]. Villegas [1111, 1904]. Vinogradov [2610]. violating [1552]. Visibility [215]. Visible [2049]. Visual [1486]. Voisin [1300]. Vojta [794, 795]. Volkenborn [1799]. volume [1791, 1854]. Volumes [1023, 1874, 2291]. Voronoi [1559, 1956, 2485].

W. [509, 1377]. Wach [2638]. Waerden [1841]. Waldschmidt [2282]. Wall [61]. Wallis [1839]. Wang [1001]. Waring [20, 273, 317, 608, 750, 1102, 1547, 2174, 2287]. Washington [878]. Watson [1697]. wave [3, 107]. waveforms [1422, 2503]. Weak [52, 721, 768, 1471, 1758]. weakly [692, 1360, 1673, 2107, 2163, 2330, 2434, 2612]. Weber [1687, 2586]. Wehler [262]. Weierstrass [615, 706, 1504, 1747]. weight [3, 107, 678, 711, 748, 768, 901, 939, 994, 1009, 1064, 1471, 1527, 1664, 1665, 1738, 2021, 2278, 2363, 2468, 2537, 2614]. Weighted [69, 575, 597, 640, 1191, 1742, 1949, 2003, 2341, 2632]. weights [41, 1051, 2024, 2059, 2261]. Weil [59, 241, 271, 287, 353, 435, 568, 680, 690, 951, 1183, 1341, 1659, 1842, 2059, 2083, 2187, 2309, 2360, 2431, 2566, 2579, 2637]. Weintraub [2328]. weird [1177, 2546]. Well [2272, 2427]. Well-rounded [2427]. Wen [1514]. Wen-Ching [1514]. Weyl [193, 1132]. which [583, 1178, 1481, 1856, 2159, 2410, 2423, 2589]. Whittaker [492, 1133, 1142, 1675]. whose [50, 381, 1006, 1195, 1778, 2304]. width [716]. width- [716]. Wieferich [51, 719, 1105, 1436, 1903, 2525]. Wielandt [1386]. Wild [698, 811, 1902]. Wilkie [1343]. Wilson [2378]. Windschitl [992]. Winnie [1514]. wise [1661]. without [841, 1251, 1674, 1822, 1926, 2294]. Witt [349, 1221, 2177]. Witten [884]. Wolfgang [500]. Women [189]. Wong [1337]. words [855, 1195, 2308]. work [657, 1514]. Workable [148]. written [2179]. WZ [300]. WZ-proof [300].

Yau [1300, 2160]. Yoshida [766]. Young [1836]. Yue [1187].

Z. [509, 1377]. Z.-W. [509, 1377]. Zêta [863]. zag [1231]. Zagier [642, 1360, 1758, 2222, 2391, 2420]. Zagier-lift [1758]. Zaharescu [1852]. Zaremba [2152]. Zeckendorf [1012, 2217, 2644]. Zero [118, 165, 243, 455, 851, 862, 1009, 1017, 1120, 1143, 1170, 1398, 1411, 1566, 1574, 1722, 1815, 1955, 1966, 1985, 1996, 2300, 2315, 2323, 2388, 2457, 2584]. Zero-cycles [1996]. Zero-density [1017, 2388]. zero-free [1120, 1411, 1815]. zero-sum [118, 862, 1398, 1566, 1722, 1955, 1966, 1985, 2323, 2457, 2584]. Zeroes [412, 665, 1194, 2529]. Zeros [173, 227, 362, 442, 488, 540, 541, 586, 767, 787, 875, 884, 919, 1064, 1068, 1115, 1232, 1256, 1308, 1555, 1640, 1643, 1673, 1892, 2125, 2158, 2173, 2197, 2296, 2358, 2385, 2428, 2468, 2578, 2612]. Zeta [37, 46, 63, 69, 85, 185, 202, 227, 291, 306, 309, 334, 344, 361, 362, 369, 378, 393, 424, 442, 444, 455, 458, 488, 516, 527, 584, 593, 605, 613, 620, 642, 702, 748, 767, 773, 858, 875, 884, 886, 919, 1045, 1052, 1058, 1082, 1115, 1156, 1164, 1172, 1191, 1193, 1194, 1213, 1259, 1266, 1313, 1333, 1411, 1424, 1427, 1477, 1548, 1615, 1640, 1665, 1680, 1701, 1726, 1742, 1749, 1779, 1787, 1804, 1812, 1854, 1861, 1887, 1901, 1970, 1987, 1988, 2031, 2078, 2088, 2145, 2162, 2197, 2212, 2261, 2266, 2331, 2341, 2351, 2390, 2403, 2426, 2461, 2507, 2512, 2607, 2632]. zeta [863, 1308, 1352, 2428]. zeta-function [185, 227, 344, 362, 488, 584, 767, 886, 1058, 1308, 1411, 1640, 1701, 1854, 1901, 2197, 2507, 2607]. zeta-functions [378, 584, 1333, 2266, 2390]. zeta-like [2162]. zeta-star [527, 1887]. zeta-type [2632]. Zhang [1078, 1187, 1598]. Zieve [2424]. zig [1231]. Ziv [356, 2208, 2470]. Zsigmondy [794, 2110]. Zumkeller [681].

References

Schaefer:1996:CGS

[1] Edward F. Schaefer. Class groups and Selmer groups. Journal of Number Theory, 56(1):79–114, January 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X96900062. See corrigendum [909].

Bugeaud:1996:MED

[2] Y. Bugeaud and M. Laurent. Minoration effective de la distance p-adique entre puissances de nombres algébriques. (French) [Effective minimization of the p-adic distance between powers of algebraic numbers]. Journal of Number Theory, 61(2):311–342, December 1996. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X96901523. See

note [154].

Kojima:2004:FCH

[3] Hisashi Kojima. On the Fourier coefficients of Hilbert–Maass wave forms of half integral weight over arbitrary algebraic number fields. Journal of Number Theory, 107(1):25–62, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000575. See

corrigendum [107].

Togbe:2004:PFC

[4] Alain Togbé. A parametric family of cubic Thue equations. Journal of Number Theory, 107(1):63–79, July 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X04000496. See

corrigendum [1271].

Borisov:2004:QIC

[5] Alexander Borisov, Melvyn B. Nathanson, and Yang Wang. Quantum integers and cyclotomy. Journal of Number Theory, 109(1):120–135, November 2004. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X04001180. See corrigendum [1128].

Omar:2007:CLH

[6] Sami Omar and Kamel Mazhouda. Le critère de Li et l’hypothèse de Riemann pour la classe de Selberg. (French) [Li’s criterion and Riemann’s hypothesis for Selberg’s class]. Journal of Number Theory, 125(1):50–58, July 2007. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X06002757. See corrigendum and ad-

dendum [93].

Muriefah:2008:DE

[7] Fadwa S. Abu Muriefah. On the diophantine equation px^{2} + q^{2m} = y^{p}. Journal of Number Theory, 128(6):1670–1675, June 2008. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X08000541. See note [189].

Walling:2009:RHS

[8] Lynne H. Walling. Restricting Hecke–Siegel operators to Jacobi modular forms. Journal of Number Theory, 129(7):1709–1733, July 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09000845. See corrigendum [94].

Gun:2009:TLG

[9] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Transcendence of the log gamma function and some discrete periods. Journal of Number Theory, 129(9):2154–2165, September 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09000511. See

corrigendum [106].

Freeze:2009:CNF

[10] Michael Freeze, Weidong Gao, and Alfred Geroldinger. The critical number of finite abelian groups. Journal of Number Theory, 129 (11):2766–2777, November 2009. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001735. See corrigendum [1304].

Anonymous:2010:PJa

[11] Anonymous. Pages 1–198 (January 2010). Journal of Number Theory, 130(1):??, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBa

[12] Anonymous. Editorial Board. Journal of Number Theory, 130(1): ifc, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002406.

Shen:2010:SMP

[13] Luming Shen and Yuyuan Zhou. Some metric properties on the GCF fraction expansion. Journal of Number Theory, 130(1):1–9, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002005.

Sun:2010:DPS

[14] Hae-Sang Sun. Derivative of power series attached to Γ-transform of p-adic measures. Journal of Number Theory, 130(1):10–26, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002029.

Cardon:2010:MRD

[15] David A. Cardon. Matrices related to Dirichlet series. Journal of Number Theory, 130(1):27–39, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09001723.

Akhtari:2010:QTE

[16] Shabnam Akhtari and Ryotaro Okazaki. Quartic Thue equations. Journal of Number Theory, 130(1):40–60, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002157.

Hubbard:2010:KGL

[17] David Hubbard and Lawrence C. Washington. Kummer generators and lambda invariants. Journal of Number Theory, 130(1):61–81, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001784.

Luca:2010:SBN

[18] Florian Luca, Diego Marques, and Pantelimon Stănică. On the spacings between C-nomial coefficients. Journal of Number Theory, 130(1): 82–100, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002261.

Shemanske:2010:SOC

[19] Thomas R. Shemanske. Split orders and convex polytopes in buildings. Journal of Number Theory, 130(1):101–115, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002042.

Buttcane:2010:NWG

[20] Jack Buttcane. A note on the Waring–Goldbach problem. Journal of Number Theory, 130(1):116–127, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002145.

Longo:2010:VSG

[21] Matteo Longo and Stefano Vigni. On the vanishing of Selmer groups for elliptic curves over ring class fields. Journal of Number Theory, 130(1):128–163, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002078.

Sittinger:2010:PRA

[22] Brian D. Sittinger. The probability that random algebraic integers are relatively r-prime. Journal of Number Theory, 130(1):164–171, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09001863.

Guo:2010:FBS

[23] Victor J. W. Guo and Jiang Zeng. Factors of binomial sums from the Catalan triangle. Journal of Number Theory, 130(1):172–186, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002108.

Gurel:2010:NP

[24] Erhan Gürel and Ali Ulaş Özgür Kişisel. A note on the products (1^{μ} + 1)(2^{μ} + 1) ⋯ (n^{μ} + 1). Journal of Number Theory, 130(1):187– 191, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900211X.

Fitzgerald:2010:SCF

[25] Robert W. Fitzgerald. Sun’s conjectures on fourth powers in the class group of binary quadratic forms. Journal of Number Theory, 130(1):192–197, January 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002017.

Anonymous:2010:PF

[26] Anonymous. Pages 199–468 (February 2010). Journal of Number Theory, 130(2):??, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBb

[27] Anonymous. Editorial Board. Journal of Number Theory, 130(2):ifc, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002716.

Huang:2010:SER

[28] Po-Yi Huang. Stickelberger elements over rational function fields. Journal of Number Theory, 130(2):199–212, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09001814.

Li:2010:GSA

[29] Zhong hua Li. Gamma series associated to elements satisfying regularized double shuffle relations. Journal of Number Theory, 130(2):213– 231, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002376.

Hirvensalo:2010:CPI

[30] Mika Hirvensalo, Juhani Karhumäki, and Alexander Rabinovich. Computing partial information out of intractable: Powers of algebraic numbers as an example. Journal of Number Theory, 130(2):232–253, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002364.

Joyner:2010:GCI

[31] Sheldon Joyner. On a generalization of Chen’s iterated integrals. Journal of Number Theory, 130(2):254–288, February 2010. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002509.

Stanchescu:2010:SDS

[32] Yonutz V. Stanchescu. The structure of d-dimensional sets with small sumset. Journal of Number Theory, 130(2):289–303, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002315.

Brink:2010:RIU

[33] David Brink. Remark on infinite unramified extensions of number fields with class number one. Journal of Number Theory, 130(2):304–306, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002480.

Haberle:2010:CGF

[34] Lothar Häberle. On cubic Galois field extensions. Journal of Number Theory, 130(2):307–317, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002388. See

corrigendum [302].

Saldana:2010:EDD

[35] Amandine Saldana. Extension du domaine de définition d’une série de Dirichlet multiplicative. (French) [Extension of the definition domain of a Dirichlet multiplicative series]. Journal of Number Theory, 130(2):318–346, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002030.

Kawashima:2010:GDM

[36] Gaku Kawashima. A generalization of the duality for multiple harmonic sums. Journal of Number Theory, 130(2):347–359, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09001097.

Mezo:2010:HSI

[37] István Mező and Ayhan Dil. Hyperharmonic series involving Hurwitz zeta function. Journal of Number Theory, 130(2):360–369, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002327.

Yao:2010:PRR

[38] Wei-Chen Yao and Jing Yu. On primitive roots for rank one Drinfeld modules. Journal of Number Theory, 130(2):370–385, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002066.

Li:2010:GCT

[39] Xian-Jin Li. A generalization of A. Connes’ trace formula. Journal of Number Theory, 130(2):386–430, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002121.

Mok:2010:SVF

[40] Chung Pang Mok. Special values of L-functions of elliptic curves over Q and their base change to real quadratic fields. Journal of Number Theory, 130(2):431–438, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002467. See correction and adden-

dum [2140].

Tolev:2010:TGP

[41] D. I. Tolev. The ternary Goldbach problem with arithmetic weights attached to two of the variables. Journal of Number Theory, 130 (2):439–457, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002054.

Chabert:2010:PCV

[42] Jean-Luc Chabert. On the polynomial closure in a valued field. Journal of Number Theory, 130(2):458–468, February 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002650.

Anonymous:2010:PMa

[43] Anonymous. Pages 469–814 (March 2010). Journal of Number Theory, 130(3):??, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBc

[44] Anonymous. Editorial Board. Journal of Number Theory, 130(3): ifc, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000003X.

Liu:2010:DCQ

[45] Cheng Liu and Linsheng Yin. Double coverings for quadratic extensions and function fields. Journal of Number Theory, 130(3):469– 477, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002546.

Goutet:2010:ZFF

[46] Philippe Goutet. On the zeta function of a family of quintics. Journal of Number Theory, 130(3):478–492, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002698.

Gaal:2010:DEG

[47] István Gaál and Michael Pohst. Diophantine equations over global function fields IV: S-unit equations in several variables with an application to norm form equations. Journal of Number Theory, 130(3):493– 506, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0800173X.

Diaz-Vargas:2010:IQF

[48] Javier Diaz-Vargas and Carlos Pompeyo-Gutiérrez. Imaginary quadratic function fields. Journal of Number Theory, 130(3):507–527, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002637.

Strauch:2010:GAT

[49] Matthias Strauch. Galois actions on torsion points of one-dimensional formal modules. Journal of Number Theory, 130(3):528–533, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002807.

Kapoor:2010:SWS

[50] Vishaal Kapoor. Sets whose sumset avoids a thin sequence. Journal of Number Theory, 130(3):534–538, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002674.

Lozano-Robledo:2010:BHN

[51] Álvaro Lozano-Robledo. Bernoulli–Hurwitz numbers, Wieferich primes and Galois representations. Journal of Number Theory, 130(3):539– 558, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900290X.

Dewar:2010:NPC

[52] Michael Dewar. The nonholomorphic parts of certain weak Maass forms. Journal of Number Theory, 130(3):559–573, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002686.

Banks:2010:PDC

[53] William D. Banks, Ahmet M. Güloğlu, and C. Wesley Nevans. On primitive Dirichlet characters and the Riemann hypothesis. Journal of Number Theory, 130(3):574–579, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09001826.

Janczak:2010:DMA

[54] Miroslawa Jańczak and Tomasz Schoen. Dense minimal asymptotic bases of order two. Journal of Number Theory, 130(3):580–585, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002595.

Kilford:2010:OAA

[55] L. J. P. Kilford. On the U_{p} operator acting on p-adic overconvergent modular forms when X_{0}(p) has genus 1. Journal of Number Theory, 130(3):586–594, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002790.

Hancl:2010:IFS

[56] Jaroslav Hančl and Robert Tijdeman. On the irrationality of factorial series II. Journal of Number Theory, 130(3):595–607, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002819.

Chen:2010:CBE

[57] Imin Chen, Ian Kiming, and Jonas B. Rasmussen. On congruences modp^{m} between eigenforms and their attached Galois representations. Journal of Number Theory, 130(3):608–619, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002613.

Finotti:2010:LIQ

[58] Luís R. A. Finotti. Lifting the j-invariant: Questions of Mazur and Tate. Journal of Number Theory, 130(3):620–638, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002856.

Kida:2010:DKT

[59] Masanari Kida. Descent Kummer theory via Weil restriction of multiplicative groups. Journal of Number Theory, 130(3):639–659, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002571.

Beukers:2010:NP

[60] F. Beukers and C. L. Stewart. Neighboring powers. Journal of Number Theory, 130(3):660–679, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09002534. See addendum [129].

Griess:2010:MCB

[61] Robert L. Griess. Midwest cousins of Barnes–Wall lattices. Journal of Number Theory, 130(3):680–695, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002662.

Broughan:2010:FCC

[62] Kevin A. Broughan and Florian Luca. On the Fürstenberg closure of a class of binary recurrences. Journal of Number Theory, 130(3):696– 706, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002844.

Hilberdink:2010:RBZ

[63] Titus W. Hilberdink. Ω-results for Beurling’s zeta function and lower bounds for the generalised Dirichlet divisor problem. Journal of Number Theory, 130(3):707–715, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X0900256X.

Chen:2010:DBB

[64] Yong-Gao Chen and Tang Sun. The difference basis and bi-basis of Z_{m}. Journal of Number Theory, 130(3):716–726, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X09002510.

Metsankyla:2010:BAF

[65] Tauno Metsänkylä. On the behaviour of p-adic L-functions. Journal of Number Theory, 130(3):727–737, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002273.

Li:2010:AHS

[66] Xian-Jin Li. An adelic Hankel summation formula. Journal of Number Theory, 130(3):738–766, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09002133.

Cochrane:2010:CMV

[67] Todd Cochrane and Sanying Shi. The congruence x_{1}x_{2} ≡ x_{3}x_{4}(modm) and mean values of character sums. Journal of Number Theory, 130(3): 767–785, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002339.

Qu:2010:LTP

[68] Yan Qu. Linnik-type problems for automorphic L-functions. Journal of Number Theory, 130(3):786–802, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002558.

Chapdelaine:2010:SAP

[69] Hugo Chapdelaine. Some arithmetic properties of partial zeta functions weighted by a sign character. Journal of Number Theory, 130(3):803– 814, March 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X0900208X.

Anonymous:2010:PAa

[70] Anonymous. Pages 815–1116 (April 2010). Journal of Number Theory, 130(4):??, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBd

[71] Anonymous. Editorial Board. Journal of Number Theory, 130(4): ifc, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000247.

Wilson:2010:AFT

[72] Gregory L. Wilson and Christopher L. Morgan. An application of Fourier transforms on finite Abelian groups to an enumeration arising from the Josephus problem. Journal of Number Theory, 130(4):815– 827, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000089.

Smajlovic:2010:LCR

[73] Lejla Smajlović. On Li’s criterion for the Riemann hypothesis for the Selberg class. Journal of Number Theory, 130(4):828–851, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002881.

Cho:2010:PFC

[74] Bumkyu Cho. Primes of the form x^{2} + ny^{2} with conditions x ≡ 1 mod N, y ≡ 0 mod N. Journal of Number Theory, 130(4):852– 861, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002091.

Gonzalez-Jimenez:2010:NHM

[75] Enrique González-Jiménez and Roger Oyono. Non-hyperelliptic modular curves of genus 3. Journal of Number Theory, 130(4):862– 878, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000181.

Baier:2010:MVC

[76] Stephan Baier and Matthew P. Young. Mean values with cubic characters. Journal of Number Theory, 130(4):879–903, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000119.

Angell:2010:FCF

[77] David Angell. A family of continued fractions. Journal of Number Theory, 130(4):904–911, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000193.

Tung:2010:MMP

[78] Shih Ping Tung. Max–min of polynomials and exponential diophantine equations. Journal of Number Theory, 130(4):912–929, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10000107.

Zhao:2010:SCC

[79] Li-Lu Zhao and Zhi-Wei Sun. Some curious congruences modulo primes. Journal of Number Theory, 130(4):930–935, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000156.

Daileda:2010:MCN

[80] Ryan C. Daileda, Raju Krishnamoorthy, and Anton Malyshev. Maximal class numbers of CM number fields. Journal of Number Theory, 130(4):936–943, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002625.

Rivoal:2010:ARV

[81] Tanguy Rivoal. Approximations rationnelles des valeurs de la fonction Gamma aux rationnels. (French) [Rational approximations of the values of the Gamma function to the rationals]. Journal of Number Theory, 130(4):944–955, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002303.

Louboutin:2010:SCQ

[82] Stéphane Louboutin. On some cubic or quartic algebraic units. Journal of Number Theory, 130(4):956–960, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002455.

Green:2010:UOP

[83] Frederic Green and Amitabha Roy. Uniqueness of optimal mod3 polynomials for parity. Journal of Number Theory, 130(4):961–975, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002893.

Fukuhara:2010:THV

[84] Shinji Fukuhara and Yifan Yang. Twisted Hecke L-values and period polynomials. Journal of Number Theory, 130(4):976–999, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002583.

Cornelissen:2010:AEF

[85] Gunther Cornelissen, Aristides Kontogeorgis, and Lotte van der Zalm. Arithmetic equivalence for function fields, the Goss zeta function and a generalisation. Journal of Number Theory, 130(4):1000–1012, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002297.

Rodriguez:2010:SCR

[86] José Alejandro Lara Rodríguez. Some conjectures and results about multizeta values for F_{q}[t]. Journal of Number Theory, 130(4):1013– 1023, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002285.

Pottharst:2010:PIS

[87] Jonathan Pottharst. Parity-induced Selmer growth for symplectic, ordinary families. Journal of Number Theory, 130(4):1024–1047, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002340.

Cha:2010:BPN

[88] Byungchul Cha and Seick Kim. Biases in the prime number race of function fields. Journal of Number Theory, 130(4):1048–1055, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X09002649.

Shparlinski:2010:SGE

[89] Igor E. Shparlinski. On the size of the Gelfond exponent. Journal of Number Theory, 130(4):1056–1060, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09002522.

Nakamura:2010:TPS

[90] Tetsuo Nakamura. Torsion points on some abelian surfaces of CMtype. Journal of Number Theory, 130(4):1061–1067, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000144.

Hara:2010:ITT

[91] Takashi Hara. Iwasawa theory of totally real fields for certain noncommutative p-extensions. Journal of Number Theory, 130(4):1068– 1097, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002352.

Omar:2010:LCR

[92] Sami Omar and Kamel Mazhouda. The Li criterion and the Riemann hypothesis for the Selberg class II. Journal of Number Theory, 130(4):1098–1108, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002479.

Omar:2010:CAF

[93] Sami Omar and Kamel Mazhouda. Corrigendum et addendum à ^{≪} Le critère de Li et l’hypothèse de Riemann pour la classe de Selberg ^{≫} [J. Number Theory 125 (1) (2007) 50–58]. Journal of Number Theory, 130(4):1109–1114, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002868. See [6].

Walling:2010:CRH

[94] Lynne H. Walling. Corrigendum to “Restricting Hecke–Siegel operators to Jacobi modular forms” [J. Number Theory 129 (7) (2009) 1709–1733]. Journal of Number Theory, 130(4):1115, April 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X09002765. See

[8].

Anonymous:2010:PMb

[95] Anonymous. Pages 1117–1254 (May 2010). Journal of Number Theory, 130(5):??, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBe

[96] Anonymous. Editorial Board. Journal of Number Theory, 130(5): ifc, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000624.

Grantham:2010:TIM

[97] Jon Grantham. There are infinitely many Perrin pseudoprimes. Journal of Number Theory, 130(5):1117–1128, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000120.

Cobbe:2010:SCT

[98] Alessandro Cobbe. Steinitz classes of tamely ramified Galois extensions of algebraic number fields. Journal of Number Theory, 130(5):1129– 1154, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000020X.

Ramare:2010:BAS

[99] O. Ramaré. On Bombieri’s asymptotic sieve. Journal of Number Theory, 130(5):1155–1189, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000211.

Chan:2010:NCS

[100] Song Heng Chan and Zhi-Guo Liu. On a new circular summation of theta functions. Journal of Number Theory, 130(5):1190– 1196, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002601.

Virdol:2010:ARA

[101] Cristian Virdol. On l-adic representations attached to Hilbert and Picard modular surfaces. Journal of Number Theory, 130(5):1197– 1211, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002820.

Zhao:2010:CMSa

[102] Yufei Zhao. Constructing MSTD sets using bidirectional ballot sequences. Journal of Number Theory, 130(5):1212–1220, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10000090.

Miller:2010:ECI

[103] Steven J. Miller, Brooke Orosz, and Daniel Scheinerman. Explicit constructions of infinite families of MSTD sets. Journal of Number Theory, 130(5):1221–1233, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002492.

Xi:2010:CSF

[104] Ping Xi and Yuan Yi. On character sums over flat numbers. Journal of Number Theory, 130(5):1234–1240, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X0900287X.

Breuer:2010:TBE

[105] Florian Breuer. Torsion bounds for elliptic curves and Drinfeld modules. Journal of Number Theory, 130(5):1241–1250, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000132.

Gun:2010:CTL

[106] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Corrigendum to “Transcendence of the log gamma function and some discrete periods”

[J. Number Theory 129 (9) (2009) 2154–2165]. Journal of Number Theory, 130(5):1251, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X09002777. See [9].

Kojima:2010:CFC

[107] Hisashi Kojima. Corrigendum to “On the Fourier coefficients of Hilbert– Maass wave forms of half integral weight over arbitrary algebraic number fields” [J. Number Theory 107 (1) (2004) 25–62]. Journal of Number Theory, 130(5):1252–1253, May 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10000223. See [3].

Anonymous:2010:PJb

[108] Anonymous. Pages 1255–1432 (June 2010). Journal of Number Theory, 130(6):??, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBf

[109] Anonymous. Editorial Board. Journal of Number Theory, 130(6): ifc, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000922.

Raghunathan:2010:FPS

[110] R. Raghunathan. On L-functions with poles satisfying Maass’s functional equation. Journal of Number Theory, 130(6):1255–1273, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1000017X.

Ozgur:2010:NFH

[111] Nihal Yılmaz Özgür. On numbers of the form n = x^{2} + Ny^{2} and the Hecke groups H(\sqrt{N}). Journal of Number Theory, 130(6):1274– 1281, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000387.

Jones:2010:NFU

[112] John W. Jones. Number fields unramified away from 2. Journal of Number Theory, 130(6):1282–1291, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10000521.

Nguyen:2010:SFE

[113] Lan Nguyen. On the solutions of a functional equation arising from multiplication of quantum integers. Journal of Number Theory, 130(6): 1292–1347, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000429.

Nguyen:2010:SBF

[114] Lan Nguyen. On the support base of a functional equation arising from multiplication of quantum integers. Journal of Number Theory, 130(6):1348–1373, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000466.

Roblot:2010:TCC

[115] X.-F. Roblot and D. Solomon. Testing the Congruence Conjecture for Rubin–Stark elements. Journal of Number Theory, 130(6):1374– 1398, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000442.

Grekos:2010:DSS

[116] Georges Grekos, Ladislav Mišík, and János T. Tóth. Density sets of sets of positive integers. Journal of Number Theory, 130(6):1399– 1407, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000375.

Ishak:2010:LBH

[117] M. I. M. Ishak, Michael J. Mossinghoff, Christopher Pinner, and Benjamin Wiles. Lower bounds for heights in cyclotomic extensions. Journal of Number Theory, 130(6):1408–1424, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000545.

Gao:2010:DLM

[118] Weidong Gao, Yahya Ould Hamidoune, and Guoqing Wang. Distinct length modular zero-sum subsequences: a proof of Graham’s conjecture. Journal of Number Theory, 130(6):1425–1431, June 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000570.

Anonymous:2010:PJc

[119] Anonymous. Pages 1433–1684 (July 2010). Journal of Number Theory, 130(7):??, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBg

[120] Anonymous. Editorial Board. Journal of Number Theory, 130(7): ifc, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001009.

Morgenbesser:2010:SDS

[121] Johannes F. Morgenbesser. The sum of digits of squares in Z[i]. Journal of Number Theory, 130(7):1433–1469, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000818.

Baragar:2010:SRC

[122] Arthur Baragar and David McKinnon. K3 surfaces, rational curves, and rational points. Journal of Number Theory, 130(7):1470–1479, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000867.

Grytczuk:2010:ASS

[123] Aleksander Grytczuk and Marek Wójtowicz. The algebraic structure of the set of solutions to the Thue equation. Journal of Number Theory, 130(7):1480–1487, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000879.

Muic:2010:CMF

[124] Goran Muić. On the cuspidal modular forms for the Fuchsian groups of the first kind. Journal of Number Theory, 130(7):1488– 1511, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000831.

Griess:2010:RHM

[125] Robert L. Griess. Rank 72 high minimum norm lattices. Journal of Number Theory, 130(7):1512–1519, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10000740.

Castro:2010:DES

[126] Francis N. Castro, Hugues Randriam, Ivelisse Rubio, and H. F. Mattson. Divisibility of exponential sums via elementary methods. Journal of Number Theory, 130(7):1520–1536, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1000082X.

Garrity:2010:TCR

[127] Thomas Garrity. A thermodynamic classification of real numbers. Journal of Number Theory, 130(7):1537–1559, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000594.

Gonzalez-Jimenez:2010:MLM

[128] Enrique González-Jiménez and Xavier Guitart. On the modularity level of modular abelian varieties over number fields. Journal of Number Theory, 130(7):1560–1570, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000788.

Beukers:2010:ANP

[129] F. Beukers and C. L. Stewart. Addendum to “Neighboring powers” [J. Number Theory 130 (3) (2010) 660–679]. Journal of Number Theory, 130(7):1571, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000909. See [60].

Schinzel:2010:ESP

[130] Andrzej Schinzel, Stanisław Spież, and Jerzy Urbanowicz. Elementary symmetric polynomials in Shamir’s scheme. Journal of Number Theory, 130(7):1572–1580, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1000079X.

Hamel:2010:ASS

[131] Mariah Hamel, Neil Lyall, Katherine Thompson, and Nathan Walters. Arithmetic structure in sparse difference sets. Journal of Number Theory, 130(7):1581–1589, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000880.

Maiga:2010:SIC

[132] Hamadoun Maïga. Some identities and congruences concerning Euler numbers and polynomials. Journal of Number Theory, 130(7):1590– 1601, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000752.

Wang:2010:DIK

[133] Yunjie Wang and Hongze Li. On s-dimensional incomplete Kloosterman sums. Journal of Number Theory, 130(7):1602–1608, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000855.

Ohkubo:2010:GTI

[134] Shun Ohkubo. Galois theory of B_{dR}^{+} in the imperfect residue field case. Journal of Number Theory, 130(7):1609–1641, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000843.

Katayama:2010:GGF

[135] Koji Katayama. A generalization of gamma functions and Kronecker’s limit formulas. Journal of Number Theory, 130(7):1642– 1674, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000478.

Pernet:2010:FCH

[136] Clément Pernet and William Stein. Fast computation of Hermite normal forms of random integer matrices. Journal of Number Theory, 130(7):1675–1683, July 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000727.

Anonymous:2010:PAb

[137] Anonymous. Pages 1685–1888 (August 2010). Journal of Number Theory, 130(8):??, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBh

[138] Anonymous. Editorial Board. Journal of Number Theory, 130(8): ifc, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001198.

Hata:2010:LE

[139] Masayoshi Hata. A lower estimate for ∥e^{n}∥. Journal of Number Theory, 130(8):1685–1704, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000168.

Le:2010:IPP

[140] Thái Hoàng Lê. Intersective polynomials and the primes. Journal of Number Theory, 130(8):1705–1717, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001071.

Morton:2010:LPC

[141] Patrick Morton. Legendre polynomials and complex multiplication, I. Journal of Number Theory, 130(8):1718–1731, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001095.

Pollack:2010:RSN

[142] Paul Pollack. A remark on sociable numbers of odd order. Journal of Number Theory, 130(8):1732–1736, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001113.

Dubickas:2010:BLR

[143] Artūras Dubickas, Aivaras Novikas, and Jonas Šiurys. A binary linear recurrence sequence of composite numbers. Journal of Number Theory, 130(8):1737–1749, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001150.

Okada:2010:ADS

[144] Shozo Okada. Analogies of Dedekind sums in function fields. Journal of Number Theory, 130(8):1750–1762, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000806.

Widmer:2010:CPF

[145] Martin Widmer. Counting points of fixed degree and bounded height on linear varieties. Journal of Number Theory, 130(8):1763–1784, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000582.

Jin:2010:CSW

[146] Renling Jin. Characterizing the structure of A + B when A + B has small upper Banach density. Journal of Number Theory, 130(8):1785– 1800, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000776.

Kaplan:2010:ITA

[147] Eyal Kaplan. An invariant theory approach for the unramified computation of Rankin–Selberg integrals for quasi-split SO_{2n} × GL_{n}. Journal of Number Theory, 130(8):1801–1817, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1000051X.

Sbeity:2010:CRE

[148] Farah Sbeity and Bouchaïb Sodaïgui. Classes réalisables d’extensions métacycliques de degré lm. (French) [Workable classes of metacyclic extensions of degree lm]. Journal of Number Theory, 130(8):1818– 1834, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X09002832.

Elsenhans:2010:ETN

[149] Andreas-Stephan Elsenhans and Jörg Jahnel. Estimates for Tamagawa numbers of diagonal cubic surfaces. Journal of Number Theory, 130(8):1835–1853, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000417.

Helou:2010:RRB

[150] Charles Helou. Reciprocal relations between cyclotomic fields. Journal of Number Theory, 130(8):1854–1875, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000569.

Miana:2010:MCC

[151] Pedro J. Miana and Natalia Romero. Moments of combinatorial and Catalan numbers. Journal of Number Theory, 130(8):1876–1887, August 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000739.

Anonymous:2010:PS

[152] Anonymous. Pages 1889–2098 (September 2010). Journal of Number Theory, 130(9):??, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBi

[153] Anonymous. Editorial Board. Journal of Number Theory, 130(9):ifc, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001277.

Yamada:2010:NPB

[154] Tomohiro Yamada. A note on the paper by Bugeaud and Laurent “Minoration effective de la distance p-adique entre puissances de nombres algébriques”. Journal of Number Theory, 130(9):1889–1897, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000983. See [2].

Chan:2010:NAR

[155] Heng Huat Chan and Pee Choon Toh. New analogues of Ramanujan’s partition identities. Journal of Number Theory, 130(9):1898–1913, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000892.

Hsieh:2010:AFE

[156] Ming-Lun Hsieh. The algebraic functional equation of Selmer groups for CM fields. Journal of Number Theory, 130(9):1914–1924, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10000508.

Basu:2010:LRV

[157] Manjusri Basu and Bandhu Prasad. Long range variations on the Fibonacci universal code. Journal of Number Theory, 130(9):1925–1931, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000533.

Tsushima:2010:ECR

[158] Takahiro Tsushima. Elementary computation of ramified components of the Jacobi sum Hecke characters. Journal of Number Theory, 130 (9):1932–1938, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000405.

Schutt:2010:LFS

[159] Matthias Schütt, Tetsuji Shioda, and Ronald van Luijk. Lines on Fermat surfaces. Journal of Number Theory, 130(9):1939–1963, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000454.

Najman:2010:CCT

[160] Filip Najman. Complete classification of torsion of elliptic curves over quadratic cyclotomic fields. Journal of Number Theory, 130(9):1964– 1968, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10000600.

Harman:2010:SSP

[161] Glyn Harman and Angel Kumchev. On sums of squares of primes II. Journal of Number Theory, 130(9):1969–2002, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001101.

Yokota:2010:SPP

[162] H. Yokota. Solutions of polynomial Pell’s equation. Journal of Number Theory, 130(9):2003–2010, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000363.

Molteni:2010:CSE

[163] G. Molteni. Cancellation in a short exponential sum. Journal of Number Theory, 130(9):2011–2027, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10000764.

Chen:2010:CR

[164] Shi-Chao Chen. Congruences for rs(n). Journal of Number Theory, 130 (9):2028–2032, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1000048X.

Li:2010:MZS

[165] Yuanlin Li, Chris Plyley, Pingzhi Yuan, and Xiangneng Zeng. Minimal zero sum sequences of length four over finite cyclic groups. Journal of Number Theory, 130(9):2033–2048, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000351.

Coffey:2010:ATS

[166] Mark W. Coffey. Addison-type series representation for the Stieltjes constants. Journal of Number Theory, 130(9):2049–2064, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10000399.

Urtis:2010:PES

[167] Çetin Ürtiş. Poles of Eisenstein series on quaternion groups. Journal of Number Theory, 130(9):2065–2077, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000557.

Dummigan:2010:SSV

[168] Neil Dummigan and Bernhard Heim. Symmetric square L-values and dihedral congruences for cusp forms. Journal of Number Theory, 130 (9):2078–2091, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10000491.

Thorne:2010:TSG

[169] Jack Thorne. On the Tate–Shafarevich groups of certain elliptic curves. Journal of Number Theory, 130(9):2092–2098, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10000430.

Anonymous:2010:PO

[170] Anonymous. Pages 2099–2394 (October 2010). Journal of Number Theory, 130(10):??, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBj

[171] Anonymous. Editorial Board. Journal of Number Theory, 130(10): ifc, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001526.

Fukshansky:2010:APS

[172] Lenny Fukshansky. Algebraic points of small height missing a union of varieties. Journal of Number Theory, 130(10):2099–2118, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10001332.

Baker:2010:ZQF

[173] R. C. Baker. The zeros of a quadratic form at square-free points. Journal of Number Theory, 130(10):2119–2146, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001149.

Lu:2010:NDF

[174] Guangshi Lü and Honggang Xia. Note on divisor function for quaternion algebras. Journal of Number Theory, 130(10):2147–2156, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001083.

Liedtke:2010:NMI

[175] Christian Liedtke and Stefan Schröer. The Néron model over the Igusa curves. Journal of Number Theory, 130(10):2157–2197, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001162.

Hare:2010:FLP

[176] Kevin G. Hare and Soroosh Yazdani. Fekete-like polynomials. Journal of Number Theory, 130(10):2198–2213, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001356.

Tomiyama:2010:LGR

[177] Yoshiyuki Tomiyama. Lifting Galois representations over arbitrary number fields. Journal of Number Theory, 130(10):2214–2222, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001174.

Zhao:2010:CTC

[178] Jia Zhao and Xianke Zhang. Coefficients of ternary cyclotomic polynomials. Journal of Number Theory, 130(10):2223–2237, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10001125.

Goes:2010:UTR

[179] John Goes, Steven Jackson, Steven J. Miller, David Montague, Kesinee Ninsuwan, Ryan Peckner, and Thuy Pham. A unitary test of the Ratios Conjecture. Journal of Number Theory, 130(10):2238–2258, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001368.

Kamano:2010:SPH

[180] Ken Kamano. Sums of products of hypergeometric Bernoulli numbers. Journal of Number Theory, 130(10):2259–2271, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001447.

Arai:2010:RP

[181] Keisuke Arai and Fumiyuki Momose. Rational points on X_{0}^{+}(37M). Journal of Number Theory, 130(10):2272–2282, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001393.

Pi:2010:DCF

[182] Qinghua Pi. Determining cusp forms by central values of Rankin– Selberg L-functions. Journal of Number Theory, 130(10):2283–2292, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001460.

Lei:2010:CMM

[183] Antonio Lei. Coleman maps for modular forms at supersingular primes over Lubin–Tate extensions. Journal of Number Theory, 130(10):2293– 2307, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000140X.

Zhao:2010:CMSb

[184] Yufei Zhao. Counting MSTD sets in finite abelian groups. Journal of Number Theory, 130(10):2308–2322, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001435.

Laurincikas:2010:URZ

[185] Antanas Laurinčikas. Universality of the Riemann zeta-function. Journal of Number Theory, 130(10):2323–2331, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001472.

Wong:2010:CNI

[186] Siman Wong. Class number indivisibility for quadratic function fields. Journal of Number Theory, 130(10):2332–2340, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001423.

Goes:2010:TAV

[187] John Goes and Steven J. Miller. Towards an ‘average’ version of the Birch and Swinnerton-Dyer conjecture. Journal of Number Theory, 130 (10):2341–2358, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001381.

Lu:2010:ESG

[188] Wen Chao Lu. Exceptional set of Goldbach number. Journal of Number Theory, 130(10):2359–2392, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001320.

Goss:2010:NDE

[189] David Goss. Note on “On the diophantine equation px^{2} + q^{2m} = y^{p}” [J. Number Theory 128 (6) (2008) 1670–1675]: Fadwa S. Abu Muriefah Riyadh University for Women, Mathematics Department, Saudi Arabia. Journal of Number Theory, 130(10):2393, October 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X0800214X. See [7].

Anonymous:2010:PN

[190] Anonymous. Pages 2395–2634 (November 2010). Journal of Number Theory, 130(11):??, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBk

[191] Anonymous. Editorial Board. Journal of Number Theory, 130(11):ifc, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001915.

Li:2010:SMV

[192] Xiaoqing Li. A spectral mean value theorem for GL(3 ). Journal of Number Theory, 130(11):2395–2403, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001484.

Drmota:2010:WSI

[193] Michael Drmota and Christian Mauduit. Weyl sums over integers with affine digit restrictions. Journal of Number Theory, 130(11):2404–2427, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001459.

Zhang:2010:GP

[194] Lei Zhang. Gelfand pairs (Sp_{4n}(F), Sp_{2n}(E)). Journal of Number Theory, 130(11):2428–2441, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001137.

Oishi-Tomiyasu:2010:SAP

[195] Ryoko Oishi-Tomiyasu. On some algebraic properties of CM-types of CM-fields and their reflexes. Journal of Number Theory, 130(11):2442– 2466, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001496.

Zhou:2010:NES

[196] Binbin Zhou. A note on exponential sums over subgroups of Z_{p^{2}}^{∗} and their applications. Journal of Number Theory, 130(11):2467–2479, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001502.

Yamana:2010:LEC

[197] Shunsuke Yamana. On the lifting of elliptic cusp forms to cusp forms on quaternionic unitary groups. Journal of Number Theory, 130(11): 2480–2527, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001411.

Bucur:2010:FNP

[198] Alina Bucur, Chantal David, Brooke Feigon, and Matilde Lalín. Fluctuations in the number of points on smooth plane curves over finite fields. Journal of Number Theory, 130(11):2528–2541, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1000168X.

Wei:2010:IHP

[199] Fu-Tsun Wei and Jing Yu. On the independence of Heegner points in the function field case. Journal of Number Theory, 130(11):2542–2560, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001642.

Noble:2010:AFB

[200] Rob Noble. Asymptotics of a family of binomial sums. Journal of Number Theory, 130(11):2561–2585, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10001630.

Jun:2010:CNR

[201] Byungheup Jun and Jungyun Lee. Caliber number of real quadratic fields. Journal of Number Theory, 130(11):2586–2595, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10001691.

Kosters:2010:OSK

[202] H. Kösters. On the occurrence of the sine kernel in connection with the shifted moments of the Riemann zeta function. Journal of Number Theory, 130(11):2596–2609, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10001678.

Park:2010:DDU

[203] Jeehoon Park. The Darmon–Dasgupta units over genus fields and the Shimura correspondence. Journal of Number Theory, 130(11):2610–2627, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001654.

Lustig:2010:AIS

[204] Daniel Lustig. The algebraic independence of the sum of divisors functions. Journal of Number Theory, 130(11):2628–2633, November 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10001733.

Anonymous:2010:PD

[205] Anonymous. Pages 2635–2802 (December 2010). Journal of Number Theory, 130(12):??, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2010:EBl

[206] Anonymous. Editorial Board. Journal of Number Theory, 130(12):ifc, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002027.

Trudgian:2010:CS

[207] T. S. Trudgian. On a conjecture of Shanks. Journal of Number Theory, 130(12):2635–2638, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001721.

Tauraso:2010:MCC

[208] Roberto Tauraso. More congruences for central binomial coefficients. Journal of Number Theory, 130(12):2639–2649, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001770.

Mortici:2010:ESQ

[209] Cristinel Mortici. Estimating the Somos’ quadratic recurrence constant. Journal of Number Theory, 130(12):2650–2657, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001897.

Deschamps:2010:BVI

[210] Bruno Deschamps. Sur les bonnes valeurs initiales de la suite de Lucas–Lehmer. (French) [On good initial values for the Lucas–Lehmer sequence]. Journal of Number Theory, 130(12):2658–2670, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10001800.

Murty:2010:ELC

[211] M. Ram Murty and N. Saradha. Euler–Lehmer constants and a conjecture of Erdős. Journal of Number Theory, 130(12):2671–2682, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001836.

Kaczorowski:2010:ORT

[212] Jerzy Kaczorowski and Kazimierz Wiertelak. Oscillations of the remainder term related to the Euler totient function. Journal of Number Theory, 130(12):2683–2700, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001873.

Pan:2010:AOS

[213] Hao Pan and Zhi-Wei Sun. On 2-adic orders of some binomial sums. Journal of Number Theory, 130(12):2701–2706, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001824.

Byeon:2010:ECR

[214] Dongho Byeon. Elliptic curves of rank 1 satisfying the 3-part of the Birch and Swinnerton-Dyer conjecture. Journal of Number Theory, 130 (12):2707–2714, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001769.

Schoof:2010:VIC

[215] René Schoof and Lawrence C. Washington. Visibility of ideal classes. Journal of Number Theory, 130(12):2715–2731, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1000185X.

Ressler:2010:HCT

[216] Wendell Ressler. A Hecke correspondence theorem for automorphic integrals with symmetric rational period functions on the Hecke groups. Journal of Number Theory, 130(12):2732–2744, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001861.

Haloui:2010:CPA

[217] Safia Haloui. The characteristic polynomials of abelian varieties of dimensions 3 over finite fields. Journal of Number Theory, 130(12):2745– 2752, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001848.

Baran:2010:NNS

[218] Burcu Baran. Normalizers of non-split Cartan subgroups, modular curves, and the class number one problem. Journal of Number Theory, 130(12):2753–2772, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001782.

Lau:2010:TMA

[219] Yuk-Kam Lau, Emmanuel Royer, and Jie Wu. Twisted moments of automorphic L-functions. Journal of Number Theory, 130(12):2773–2802, December 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001666.

Anonymous:2011:PJa

[220] Anonymous. Pages 1–158 (January 2011). Journal of Number Theory, 131(1):??, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBa

[221] Anonymous. Editorial Board. Journal of Number Theory, 131(1): ifc, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002209.

Komatsu:2011:IMA

[222] Takao Komatsu, Vichian Laohakosol, and Pattira Ruengsinsub. Independence measures of arithmetic functions. Journal of Number Theory, 131(1):1–17, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001885.

Hajdu:2011:PRG

[223] L. Hajdu and N. Saradha. On a problem of Recaman and its generalization. Journal of Number Theory, 131(1):18–24, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10001794.

Miller:2011:EES

[224] Steven J. Miller and M. Ram Murty. Effective equidistribution and the Sato–Tate law for families of elliptic curves. Journal of Number Theory, 131(1):25–44, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001988.

Olofsson:2011:PBG

[225] Rikard Olofsson. Properties of the Beurling generalized primes. Journal of Number Theory, 131(1):45–58, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10002003.

Itoh:2011:MEI

[226] Tsuyoshi Itoh. On multiple Z_{p}-extensions of imaginary abelian quartic fields. Journal of Number Theory, 131(1):59–66, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1000199X.

Bui:2011:LGB

[227] H. M. Bui. Large gaps between consecutive zeros of the Riemann zeta-function. Journal of Number Theory, 131(1):67–95, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10002131.

Buium:2011:ADO

[228] Alexandru Buium, Claire C. Ralph, and Santiago R. Simanca. Arithmetic differential operators on Z_{p}. Journal of Number Theory, 131(1): 96–105, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002143.

Brown:2011:CPS

[229] Jim L. Brown. On the cuspidality of pullbacks of Siegel Eisenstein series to Sp_{2m} × Sp_{2n}. Journal of Number Theory, 131(1):106–119, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000212X.

He:2011:TUE

[230] Bo He and Alain Togbé. On the D(−1)-triple {1, k^{2} + 1, k^{2} + 2k + 2} and its unique D(1)-extension. Journal of Number Theory, 131(1):120– 137, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000209X.

Ionascu:2011:PS

[231] Eugen J. Ionascu and Andrei Markov. Platonic solids in Z^{3}. Journal of Number Theory, 131(1):138–145, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10002118.

Hsu:2011:GPN

[232] Chih-Nung Hsu and Ting-Ting Nan. A generalization of the primitive normal basis theorem. Journal of Number Theory, 131(1):146– 157, January 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10001708.

Anonymous:2011:PF

[233] Anonymous. Pages 159–362 (February 2011). Journal of Number Theory, 131(2):??, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBb

[234] Anonymous. Editorial Board. Journal of Number Theory, 131(2):ifc, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002350.

Connes:2011:HAC

[235] Alain Connes and Caterina Consani. The hyperring of adèle classes. Journal of Number Theory, 131(2):159–194, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002180.

Quattrini:2011:ETD

[236] Patricia L. Quattrini. The effect of torsion on the distribution of [Cyrillicsha] among quadratic twists of an elliptic curve. Journal of Number Theory, 131(2):195–211, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10002106.

Borek:2011:ATN

[237] Thomas Borek. Arakelov theory of noncommutative arithmetic curves. Journal of Number Theory, 131(2):212–227, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002325.

Wenpeng:2011:FSP

[238] Zhang Wenpeng. The fourth and sixth power mean of the classical Kloosterman sums. Journal of Number Theory, 131(2):228–238, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002313.

Imai:2011:FMC

[239] Naoki Imai. Filtered modules corresponding to potentially semistable representations. Journal of Number Theory, 131(2):239–259, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000226X.

Luca:2011:MNI

[240] Florian Luca and Paul Pollack. Multiperfect numbers with identical digits. Journal of Number Theory, 131(2):260–284, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002283.

Schweizer:2011:SWC

[241] Andreas Schweizer. Strong Weil curves over F_{q}(T) with small conductor. Journal of Number Theory, 131(2):285–299, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002301.

Khanduja:2011:SIR

[242] Sudesh K. Khanduja, Ramneek Khassa, and Shanta Laishram. Some irreducibility results for truncated binomial expansions. Journal of Number Theory, 131(2):300–308, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10002271.

Bushnell:2011:EFC

[243] Colin J. Bushnell and Guy Henniart. Explicit functorial correspondences for level zero representations of p-adic linear groups. Journal of Number Theory, 131(2):309–331, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10002337.

Xue:2011:EAJ

[244] Jiangwei Xue. Endomorphism algebras of Jacobians of certain superelliptic curves. Journal of Number Theory, 131(2):332–342, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10002179.

Ichihara:2011:FMF

[245] Yumiko Ichihara. The first moment of L-functions of primitive forms on Γ_{0}(p^{α}) and a basis of old forms. Journal of Number Theory, 131(2):343–362, February 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10002167.

Anonymous:2011:PM

[246] Anonymous. Pages 363–604 (March 2011). Journal of Number Theory, 131(3):??, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBc

[247] Anonymous. Editorial Board. Journal of Number Theory, 131(3): ifc, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002623.

Carneiro:2011:BCS

[248] Emanuel Carneiro and Vorrapan Chandee. Bounding ζ(s) in the critical strip. Journal of Number Theory, 131(3):363–384, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002258.

Okubo:2011:GFT

[249] Satoshi Okubo. On the gamma factor of the triple product L-functions. Journal of Number Theory, 131(3):385–409, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002295.

Taelman:2011:CS

[250] Lenny Taelman. The Carlitz shtuka. Journal of Number Theory, 131(3):410–418, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1000243X.

Zhang:2011:AFH

[251] Bei Zhang. On the p-adic L-function of Hilbert modular forms at supersingular primes. Journal of Number Theory, 131(3):419–439, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002441.

Luca:2011:FI

[252] Florian Luca, Carl Pomerance, and Stephan Wagner. Fibonacci integers. Journal of Number Theory, 131(3):440–457, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002507.

Zhu:2011:SGL

[253] Hui Lin Zhu and Mao Hua Le. On some generalized Lebesgue– Nagell equations. Journal of Number Theory, 131(3):458–469, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10002490.

Mulay:2011:PDC

[254] S. B. Mulay and Mark Spindler. The positive discriminant case of Nagell’s theorem for certain cubic orders. Journal of Number Theory, 131(3):470–486, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10002477.

Natsui:2011:LCC

[255] Rie Natsui. On the Legendre constant of α-continued fractions. Journal of Number Theory, 131(3):487–507, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X10002489.

Igarashi:2011:CSC

[256] Masahiro Igarashi. Cyclic sum of certain parametrized multiple series. Journal of Number Theory, 131(3):508–518, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002519.

Odzak:2011:ABG

[257] Almasa Odžak and Lejla Smajlović. On asymptotic behavior of generalized Li coefficients in the Selberg class. Journal of Number Theory, 131(3):519–535, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10002428.

Delaunay:2011:AGI

[258] Christophe Delaunay. Averages of groups involving p^{ℓ}-rank and combinatorial identities. Journal of Number Theory, 131(3):536–551, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002593.

Byeon:2011:RTO

[259] Dongho Byeon and Donggeon Yhee. Rational torsion on optimal curves and rank-one quadratic twists. Journal of Number Theory, 131(3):552– 560, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000257X.

Bassa:2011:GCD

[260] Alp Bassa and Peter Beelen. The Galois closure of Drinfeld modular towers. Journal of Number Theory, 131(3):561–577, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002581.

Baragar:2011:OPC

[261] Arthur Baragar. Orbits of points on certain K3 surfaces. Journal of Number Theory, 131(3):578–599, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X10002520. See appendix [262].

vanLuijk:2011:WSP

[262] Ronald van Luijk. Wehler K3 surfaces with Picard number 3 and 4. Appendix to: “Orbits of points on certain K3 surfaces”, by Arthur Baragar. Journal of Number Theory, 131(3):600–603, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X10002465. See [261].

Anonymous:2011:PAa

[263] Anonymous. Pages 605–780 (April 2011). Journal of Number Theory, 131(4):??, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBd

[264] Anonymous. Editorial Board. Journal of Number Theory, 131(4): ifc, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000138.

Berrizbeitia:2011:APS

[265] Pedro Berrizbeitia and Florian Luca. Additive properties of subgroups of finite index in fields. Journal of Number Theory, 131(4):605– 617, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1000260X.

Zhang:2011:MVI

[266] Deyu Zhang and Wenguang Zhai. On the mean value of the index of composition of an integral ideal. Journal of Number Theory, 131(4):618–633, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10002568.

Pandey:2011:DIS

[267] Ram Krishna Pandey and Amitabha Tripathi. On the density of integral sets with missing differences from sets related to arithmetic progressions. Journal of Number Theory, 131(4):634–647, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X10002532.

Otsubo:2011:CVH

[268] Noriyuki Otsubo. Certain values of Hecke L-functions and generalized hypergeometric functions. Journal of Number Theory, 131(4):648– 660, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002544.

Fleming:2011:CKS

[269] Patrick S. Fleming, Stephan Ramon Garcia, and Gizem Karaali. Classical Kloosterman sums: Representation theory, magic squares, and Ramanujan multigraphs. Journal of Number Theory, 131(4):661– 680, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002684.

Virdol:2011:BSD

[270] Cristian Virdol. On the Birch and Swinnerton-Dyer conjecture for abelian varieties attached to Hilbert modular forms. Journal of Number Theory, 131(4):681–684, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000096.

Kamiya:2011:AIW

[271] Yuichi Kamiya and Masatoshi Suzuki. An attempt to interpret the Weil explicit formula from Beurling’s spectral theory. Journal of Number Theory, 131(4):685–704, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000072.

Weatherby:2011:TMI

[272] Chester Weatherby. Transcendence of multi-indexed infinite series. Journal of Number Theory, 131(4):705–715, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000047.

Liu:2011:TRP

[273] Zhixin Liu and Guangshi Lü. Two results on powers of 2 in Waring–Goldbach problem. Journal of Number Theory, 131(4):716– 736, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000084.

Kim:2011:CCC

[274] Jae Moon Kim and Jado Ryu. Construction of a certain circular unit and its applications. Journal of Number Theory, 131(4):737– 744, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000035.

Deajim:2011:LIS

[275] Abdulaziz Deajim and Samir Siksek. On the Q-linear independence of the sums ∑_{n=1}^{∞} σ_{k}(n)/n!. Journal of Number Theory, 131(4):745– 749, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000102.

Ih:2011:NPP

[276] Su ion Ih. A nondensity property of preperiodic points on Chebyshev dynamical systems. Journal of Number Theory, 131(4):750– 780, April 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002556.

Anonymous:2011:EBe

[277] Anonymous. Editorial Board. Journal of Number Theory, 131(5): ifc, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000217.

Koblitz:2011:ECC

[278] Ann Hibner Koblitz, Neal Koblitz, and Alfred Menezes. Elliptic curve cryptography: the serpentine course of a paradigm shift. Journal of Number Theory, 131(5):781–814, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09000481.

Bisson:2011:CER

[279] Gaetan Bisson and Andrew V. Sutherland. Computing the endomorphism ring of an ordinary elliptic curve over a finite field. Journal of Number Theory, 131(5):815–831, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X09002789.

Boneh:2011:FCO

[280] D. Boneh, K. Rubin, and A. Silverberg. Finding composite order ordinary elliptic curves using the Cocks–Pinch method. Journal of Number Theory, 131(5):832–841, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001344.

Arene:2011:FCT

[281] Christophe Arène, Tanja Lange, Michael Naehrig, and Christophe Ritzenthaler. Faster computation of the Tate pairing. Journal of Number Theory, 131(5):842–857, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001757.

Bernstein:2011:CSA

[282] Daniel J. Bernstein and Tanja Lange. A complete set of addition laws for incomplete Edwards curves. Journal of Number Theory, 131(5):858– 872, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002155.

DeFeo:2011:FAC

[283] Luca De Feo. Fast algorithms for computing isogenies between ordinary elliptic curves in small characteristic. Journal of Number Theory, 131(5):873–893, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001812.

Kohel:2011:ALS

[284] David Kohel. Addition law structure of elliptic curves. Journal of Number Theory, 131(5):894–919, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10002672.

OConnor:2011:CCC

[285] Laura Hitt O’Connor, Gary McGuire, Michael Naehrig, and Marco Streng. A CM construction for curves of genus 2 with p-rank 1. Journal of Number Theory, 131(5):920–935, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1000137X.

Lauter:2011:CGC

[286] Kristin Lauter and Tonghai Yang. Computing genus 2 curves from invariants on the Hilbert moduli space. Journal of Number Theory, 131(5):936–958, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X10001745.

Freeman:2011:CPF

[287] David Mandell Freeman and Takakazu Satoh. Constructing pairingfriendly hyperelliptic curves using Weil restriction. Journal of Number Theory, 131(5):959–983, May 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1000171X.

Anonymous:2011:PJb

[288] Anonymous. Pages 985–1148 (June 2011). Journal of Number Theory, 131(6):??, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBf

[289] Anonymous. Editorial Board. Journal of Number Theory, 131(6): ifc, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000400.

Banks:2011:SPS

[290] William D. Banks and David J. Covert. Sums and products with smooth numbers. Journal of Number Theory, 131(6):985–993, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000023.

Gao:2011:ZFA

[291] Xia Gao. On the zeta function associated with module classes of a number field. Journal of Number Theory, 131(6):994–1019, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000114.

Bayad:2011:AKF

[292] Abdelmejid Bayad and Yoshinori Hamahata. Arakawa–Kaneko Lfunctions and generalized poly-Bernoulli polynomials. Journal of Number Theory, 131(6):1020–1036, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000060.

Scott:2011:GPE

[293] Reese Scott and Robert Styer. The generalized Pillai equation ±ra^{x} ± sb^{y} = c. Journal of Number Theory, 131(6):1037–1047, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000059.

Virdol:2011:TCT

[294] Cristian Virdol. Tate conjecture for twisted Picard modular surfaces. Journal of Number Theory, 131(6):1048–1053, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000369.

Kellner:2011:SCI

[295] B. C. Kellner. On stronger conjectures that imply the Erdős– Moser conjecture. Journal of Number Theory, 131(6):1054–1061, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000345.

Konomi:2011:ICG

[296] Yutaka Konomi. The ideal class groups of dihedral extensions over imaginary quadratic fields and the special values of the Artin L-function. Journal of Number Theory, 131(6):1062–1069, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000357.

Samuels:2011:PFM

[297] Charles L. Samuels. The parametrized family of metric Mahler measures. Journal of Number Theory, 131(6):1070–1088, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000321.

Jeong:2011:CPC

[298] Sangtae Jeong. Calculus in positive characteristic p. Journal of Number Theory, 131(6):1089–1104, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000370.

Shparlinski:2011:SSQ

[299] Igor E. Shparlinski. On small solutions to quadratic congruences. Journal of Number Theory, 131(6):1105–1111, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000291.

Apagodu:2011:SWP

[300] Moa Apagodu. A short WZ-proof of Euler’s fundamental sum identity and more. Journal of Number Theory, 131(6):1112–1116, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000382.

Huynh:2011:ECT

[301] Duc Khiem Huynh, Steven J. Miller, and Ralph Morrison. An elliptic curve test of the L-Functions Ratios Conjecture. Journal of Number Theory, 131(6):1117–1147, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000333.

Haberle:2011:CCG

[302] Lothar Häberle. Corrigendum to “On cubic Galois field extensions” [J. Number Theory 130 (2010) 307–317]. Journal of Number Theory, 131(6):1148, June 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1100028X. See [34].

Anonymous:2011:PJc

[303] Anonymous. Pages 1149–1346 (July 2011). Journal of Number Theory, 131(7):??, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBg

[304] Anonymous. Editorial Board. Journal of Number Theory, 131(7): ifc, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000667.

Papikian:2011:JLI

[305] Mihran Papikian. On Jacquet–Langlands isogeny over function fields. Journal of Number Theory, 131(7):1149–1175, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000308.

Sodergren:2011:VDM

[306] Anders Södergren. On the value distribution and moments of the Epstein zeta function to the right of the critical strip. Journal of Number Theory, 131(7):1176–1208, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1100031X.

Avendano:2011:SBN

[307] Martín Avendaño and Teresa Krick. Sharp bounds for the number of roots of univariate fewnomials. Journal of Number Theory, 131(7):1209– 1228, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000527.

Sookdeo:2011:IPB

[308] Vijay A. Sookdeo. Integer points in backward orbits. Journal of Number Theory, 131(7):1229–1239, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000515.

Tangedal:2011:AMZ

[309] Brett A. Tangedal and Paul T. Young. On p-adic multiple zeta and log gamma functions. Journal of Number Theory, 131(7):1240– 1257, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000588.

Kim:2011:GDT

[310] Wansu Kim. Galois deformation theory for norm fields and flat deformation rings. Journal of Number Theory, 131(7):1258–1275, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000540.

Bassa:2011:PCS

[311] Alp Bassa and Peter Beelen. A proof of a conjecture by Schweizer on the Drinfeld modular polynomial Φ_{T} (X, Y). Journal of Number Theory, 131(7):1276–1285, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000539.

Garunkstis:2011:SAD

[312] Ramūnas Garunkštis. Self-approximation of Dirichlet L-functions. Journal of Number Theory, 131(7):1286–1295, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1100062X.

Dummigan:2011:SSM

[313] Neil Dummigan, Tomoyoshi Ibukiyama, and Hidenori Katsurada. Some Siegel modular standard L-values, and Shafarevich–Tate groups. Journal of Number Theory, 131(7):1296–1330, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000606.

Hida:2011:CAI

[314] Haruzo Hida. Constancy of adjoint L-invariant. Journal of Number Theory, 131(7):1331–1346, July 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000552.

Anonymous:2011:PAb

[315] Anonymous. Pages 1347–1546 (August 2011). Journal of Number Theory, 131(8):??, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBh

[316] Anonymous. Editorial Board. Journal of Number Theory, 131(8): ifc, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000898.

Cai:2011:WGP

[317] Yingchun Cai. On Waring–Goldbach problem involving fourth powers. Journal of Number Theory, 131(8):1347–1362, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000618.

Thorne:2011:UPF

[318] Frank Thorne. An uncertainty principle for function fields. Journal of Number Theory, 131(8):1363–1389, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000564.

Yilmaz:2011:HMP

[319] Fatih Yilmaz and Durmus Bozkurt. Hessenberg matrices and the Pell and Perrin numbers. Journal of Number Theory, 131(8):1390–1396, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000576.

Liu:2011:DCF

[320] Sheng-Chi Liu. Determination of GL(3 ) cusp forms by central values of GL(3 ) × GL(2 ) L-functions, level aspect. Journal of Number Theory, 131(8):1397–1408, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000631.

Takai:2011:EIC

[321] Yuuki Takai. An effective isomorphy criterion for modℓ Galois representations. Journal of Number Theory, 131(8):1409–1419, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000643.

Srinivasan:2011:IMB

[322] Anitha Srinivasan. An improvement of the Minkowski bound for real quadratic orders using the Markoff theorem. Journal of Number Theory, 131(8):1420–1428, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000795.

Monge:2011:DNI

[323] Maurizio Monge. Determination of the number of isomorphism classes of extensions of a p-adic field. Journal of Number Theory, 131(8):1429– 1434, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000758.

Armana:2011:CDM

[324] Cécile Armana. Coefficients of Drinfeld modular forms and Hecke operators. Journal of Number Theory, 131(8):1435–1460, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000825.

Kikuta:2011:CHM

[325] Toshiyuki Kikuta. Congruences for Hermitian modular forms of degree 2. Journal of Number Theory, 131(8):1461–1469, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000837.

Alkan:2011:MSA

[326] Emre Alkan. On the mean square average of special values of L-functions. Journal of Number Theory, 131(8):1470–1485, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11000849. See addendum [384].

Wang:2011:DE

[327] Yongxing Wang and Tingting Wang. On the Diophantine equation nx^{2} + 2^{2m} = y^{n}. Journal of Number Theory, 131(8):1486–1491, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000850. See note [1001].

Rocadas:2011:LDS

[328] Luís Roçadas and Johannes Schoißengeier. On the local discrepancy of (nα)-sequences. Journal of Number Theory, 131(8):1492–1497, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000710.

Morgenbesser:2011:TMM

[329] Johannes F. Morgenbesser, Jeffrey Shallit, and Thomas Stoll. Thue– Morse at multiples of an integer. Journal of Number Theory, 131 (8):1498–1512, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000771.

Byeon:2011:CDR

[330] Dongho Byeon and Jungyun Lee. A complete determination of Rabinowitsch polynomials. Journal of Number Theory, 131(8):1513–1529, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000801.

Unver:2011:DBG

[331] Sinan Ünver. Deformations of Bloch groups and Aomoto dilogarithms in characteristic p. Journal of Number Theory, 131(8):1530– 1546, August 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000746.

Anonymous:2011:PS

[332] Anonymous. Pages 1547–1742 (September 2011). Journal of Number Theory, 131(9):??, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBi

[333] Anonymous. Editorial Board. Journal of Number Theory, 131(9):ifc, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001144.

Dutt:2011:ZFI

[334] Somjit Dutt. On zeta function identities involving sums of squares and the zeta–theta correspondence. Journal of Number Theory, 131 (9):1547–1552, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000874.

Madritsch:2011:ANA

[335] Manfred G. Madritsch and Attila Pethő. Asymptotic normality of additive functions on polynomial sequences in canonical number systems. Journal of Number Theory, 131(9):1553–1574, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000862.

Laradji:2011:RDE

[336] A. Laradji, M. Mignotte, and N. Tzanakis. On px^{2} + q^{2n} = y^{p} and related Diophantine equations. Journal of Number Theory, 131 (9):1575–1596, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11000783.

Halter-Koch:2011:DEP

[337] Franz Halter-Koch. Diophantine equations of Pellian type. Journal of Number Theory, 131(9):1597–1615, September 2011. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1100076X.

Jones:2011:ABR

[338] Owen T. R. Jones. An analogue of the BGG resolution for locally analytic principal series. Journal of Number Theory, 131(9):1616–1640, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000813.

Huang:2011:MVT

[339] Jingjing Huang and Robert C. Vaughan. Mean value theorems for binary Egyptian fractions. Journal of Number Theory, 131(9):1641–1656, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000965.

Bitan:2011:DAT

[340] Rony A. Bitan. The discriminant of an algebraic torus. Journal of Number Theory, 131(9):1657–1671, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000989.

Fodden:2011:DEG

[341] Brandon Fodden. Diophantine equations and the generalized Riemann hypothesis. Journal of Number Theory, 131(9):1672–1690, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11000722.

Herman:2011:BER

[342] P. Edward Herman. Beyond endoscopy for the Rankin–Selberg Lfunction. Journal of Number Theory, 131(9):1691–1722, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11001041.

Murty:2011:PCR

[343] M. Ram Murty and V. Kumar Murty. A problem of Chowla revisited. Journal of Number Theory, 131(9):1723–1733, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1100103X.

Lu:2011:DPR

[344] Guangshi Lü, Jie Wu, and Wenguang Zhai. On a divisor problem related to the Epstein zeta-function, II. Journal of Number Theory, 131 (9):1734–1742, September 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001004.

Anonymous:2011:PO

[345] Anonymous. Pages 1743–2012 (October 2011). Journal of Number Theory, 131(10):??, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBj

[346] Anonymous. Editorial Board. Journal of Number Theory, 131(10): ifc, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1100134X.

Bocherer:2011:EHD

[347] Siegfried Böcherer and Bernhard Heim. Equivariant holomorphic differential operators and finite averages of values of L-functions. Journal of Number Theory, 131(10):1743–1769, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001077.

Suzuki:2011:PCF

[348] Masatoshi Suzuki. Positivity of certain functions associated with analysis on elliptic surfaces. Journal of Number Theory, 131(10):1770– 1796, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001065.

Hogadi:2011:CWV

[349] Amit Hogadi and Supriya Pisolkar. On the cohomology of Witt vectors of p-adic integers and a conjecture of Hesselholt. Journal of Number Theory, 131(10):1797–1807, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001016.

Kim:2011:ICN

[350] Heon Kim, Paul van Wamelen, and Helena A. Verrill. Infinite class of new sign ambiguities. Journal of Number Theory, 131(10):1808– 1823, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001028.

Endo:2011:RPP

[351] Akira Endô. Remark on Polický’s paper on circular units of a compositum of quadratic number fields. Journal of Number Theory, 131 (10):1824–1826, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001089.

Lev:2011:TIG

[352] Vsevolod F. Lev. Translation invariance in groups of prime order. Journal of Number Theory, 131(10):1827–1832, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11000990.

Dokchitser:2011:NMW

[353] Tim Dokchitser and Vladimir Dokchitser. A note on the Mordell– Weil rank modulo n. Journal of Number Theory, 131(10):1833–1839, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001090.

Yang:2011:DEP

[354] Shichun Yang, Alain Togbé, and Bo He. Diophantine equations with products of consecutive values of a quadratic polynomial. Journal of Number Theory, 131(10):1840–1851, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001053.

Fintzen:2011:CPC

[355] Jessica Fintzen. Cyclotomic polynomial coefficients a(n, k) with n and k in prescribed residue classes. Journal of Number Theory, 131 (10):1852–1863, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001120.

Fan:2011:EGZ

[356] Yushuang Fan, Weidong Gao, and Qinghai Zhong. On the Erdős– Ginzburg–Ziv constant of finite abelian groups of high rank. Journal of Number Theory, 131(10):1864–1874, October 2011. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001107.

Cha:2011:CBG

[357] Byungchul Cha and Bo-Hae Im. Chebyshev’s bias in Galois extensions of global function fields. Journal of Number Theory, 131(10):1875– 1886, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001119.

Scarpello:2011:HFC

[358] Giovanni Mingari Scarpello and Daniele Ritelli. π and the hypergeometric functions of complex argument. Journal of Number Theory, 131 (10):1887–1900, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001223.

Knapp:2011:EVF

[359] Michael P. Knapp. Exact values of the function Γ^{∗}(k). Journal of Number Theory, 131(10):1901–1911, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001247.

Hajdu:2011:AFP

[360] L. Hajdu and T. Kovács. Almost fifth powers in arithmetic progression. Journal of Number Theory, 131(10):1912–1923, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001272.

Lu:2011:ABC

[361] Guangshi Lü and Zhishan Yang. The average behavior of the coefficients of Dedekind zeta function over square numbers. Journal of Number Theory, 131(10):1924–1938, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11000734.

Karabulut:2011:SAZ

[362] Yunus Karabulut and Cem Yalçın Yıldırım. On some averages at the zeros of the derivatives of the Riemann zeta-function. Journal of Number Theory, 131(10):1939–1961, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001259.

Franze:2011:SLS

[363] C. S. Franze. Sifting limits for the Λ^{2}Λ^{−} sieve. Journal of Number Theory, 131(10):1962–1982, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11001260.

Schmidt:2011:DAF

[364] Asmus L. Schmidt. Diophantine approximation in the field Q(i\sqrt{2}). Journal of Number Theory, 131(10):1983–2012, October 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11000977.

Anonymous:2011:PN

[365] Anonymous. Pages 2013–2246 (November 2011). Journal of Number Theory, 131(11):??, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBk

[366] Anonymous. Editorial Board. Journal of Number Theory, 131(11):ifc, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001594.

Allouche:2011:HGP

[367] J.-P. Allouche and M. Mendès France. Hadamard grade of power series. Journal of Number Theory, 131(11):2013–2022, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001296.

Choi:2011:CHE

[368] SoYoung Choi and Chang Heon Kim. Congruences for Hecke eigenvalues in higher level cases. Journal of Number Theory, 131(11):2023–2036, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001314.

Masser:2011:RVR

[369] D. Masser. Rational values of the Riemann zeta function. Journal of Number Theory, 131(11):2037–2046, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001235.

Pribitkin:2011:OBC

[370] Wladimir de Azevedo Pribitkin. On the oscillatory behavior of certain arithmetic functions associated with automorphic forms. Journal of Number Theory, 131(11):2047–2060, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001326.

Hirschhorn:2011:NSQ

[371] Michael D. Hirschhorn. A note on Somos’ quadratic recurrence constant. Journal of Number Theory, 131(11):2061–2063, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001284.

Liedtke:2011:TSS

[372] Christian Liedtke. The p-torsion subgroup scheme of an elliptic curve. Journal of Number Theory, 131(11):2064–2077, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001399.

Fang:2011:CCM

[373] Yingjue Fang and Binyong Sun. On conjugacy classes in metaplectic groups. Journal of Number Theory, 131(11):2078–2080, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11001466.

Rodriguez:2011:RBM

[374] José Alejandro Lara Rodríguez. Relations between multizeta values in characteristic p. Journal of Number Theory, 131(11):2081–2099, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001478.

Jones:2011:PCC

[375] Lenny Jones. Polynomial Cunningham chains. Journal of Number Theory, 131(11):2100–2106, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001521. See

corrigendum [517].

Zhao:2011:SCM

[376] Yufei Zhao. Sets characterized by missing sums and differences. Journal of Number Theory, 131(11):2107–2134, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001430.

Hoshi:2011:CBS

[377] Akinari Hoshi. On correspondence between solutions of a family of cubic Thue equations and isomorphism classes of the simplest cubic fields. Journal of Number Theory, 131(11):2135–2150, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001417.

Nakamura:2011:AHU

[378] Takashi Nakamura and Łukasz Pańkowski. Applications of hybrid universality to multivariable zeta-functions. Journal of Number Theory, 131 (11):2151–2161, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001442.

Creutz:2011:PAV

[379] Brendan Creutz. Potential [Cyrillicsha] for abelian varieties. Journal of Number Theory, 131(11):2162–2174, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001533.

David:2011:BUP

[380] Agnès David. Borne uniforme pour les homothéties dans l’image de Galois associée aux courbes elliptiques. (French) [Uniform bound for the homothies in the Galois image associated with elliptic curves]. Journal of Number Theory, 131(11):2175–2191, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1100059X.

Alaca:2011:RQQ

[381] Ayşe Alaca. Representations by quaternary quadratic forms whose coefficients are 1, 4, 9 and 36. Journal of Number Theory, 131(11): 2192–2218, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001302.

Sun:2011:CRC

[382] Zhi-Wei Sun. On congruences related to central binomial coefficients. Journal of Number Theory, 131(11):2219–2238, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001211.

Bhagwat:2011:MOP

[383] Chandrasheel Bhagwat and C. S. Rajan. On a multiplicity one property for the length spectra of even dimensional compact hyperbolic spaces. Journal of Number Theory, 131(11):2239–2244, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001491.

Alkan:2011:AMS

[384] Emre Alkan. Addendum to “On the mean square average of special values of L-functions” [J. Number Theory 131 (8) (2011) 1470–1485]. Journal of Number Theory, 131(11):2245, November 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001405. See

[326].

Anonymous:2011:PD

[385] Anonymous. Pages 2247–2460 (December 2011). Journal of Number Theory, 131(12):??, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2011:EBl

[386] Anonymous. Editorial Board. Journal of Number Theory, 131(12):ifc, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001843.

Liu:2011:HPM

[387] Kui Liu and Haiyan Wang. Higher power moments of the Riesz mean error term of symmetric square L-function. Journal of Number Theory, 131(12):2247–2261, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001557.

Lau:2011:QVJ

[388] Yuk-Kam Lau and Yingnan Wang. Quantitative version of the joint distribution of eigenvalues of the Hecke operators. Journal of Number Theory, 131(12):2262–2281, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11001545.

Fuchs:2011:SAA

[389] Elena Fuchs. Strong approximation in the Apollonian group. Journal of Number Theory, 131(12):2282–2302, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001508.

Sivertsen:2011:HCH

[390] Tormod Kalberg Sivertsen and Ragnar Soleng. Hyperelliptic curves and homomorphisms to ideal class groups of quadratic number fields. Journal of Number Theory, 131(12):2303–2309, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001429.

Rosales:2011:FPN

[391] J. C. Rosales and M. B. Branco. The Frobenius problem for numerical semigroups. Journal of Number Theory, 131(12):2310–2319, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1100151X.

Lennon:2011:TFH

[392] Catherine Lennon. Trace formulas for Hecke operators, Gaussian hypergeometric functions, and the modularity of a threefold. Journal of Number Theory, 131(12):2320–2351, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001454.

Mishou:2011:JUT

[393] Hidehiko Mishou. The joint universality theorem for a pair of Hurwitz zeta functions. Journal of Number Theory, 131(12):2352–2367, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1100148X.

Sikiric:2011:IHO

[394] Mathieu Dutour Sikirić, Graham Ellis, and Achill Schürmann. On the integral homology of PSL_{4}(Z) and other arithmetic groups. Journal of Number Theory, 131(12):2368–2375, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001685.

Eum:2011:R

[395] Ick Sun Eum, Dong Hwa Shin, and Dong Sung Yoon. Representations by x_{1}^{2} + 2x_{2}^{2} + x_{3}^{2} + x_{4}^{2} + x_{1}x_{3} + x_{1}x_{4} + x_{2}x_{4}. Journal of Number Theory, 131(12):2376–2386, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001703.

Sun:2011:DNS

[396] Zhi-Wei Sun. On Delannoy numbers and Schröder numbers. Journal of Number Theory, 131(12):2387–2397, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001715.

Maze:2011:NDD

[397] Gérard Maze. Natural density distribution of Hermite normal forms of integer matrices. Journal of Number Theory, 131(12):2398–2408, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001764.

Gregor:2011:MMT

[398] Aryeh Gregor and Yu Yasufuku. Monomial maps on P^{2} and their arithmetic dynamics. Journal of Number Theory, 131(12):2409–2425, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001788.

Xia:2011:GEP

[399] Li meng Xia. A general estimation for partitions with large difference: For even case. Journal of Number Theory, 131(12):2426–2435, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002453.

Gilbreath:2011:PPG

[400] Norman Gilbreath. Processing process: the Gilbreath conjecture. Journal of Number Theory, 131(12):2436–2441, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001740.

Anselm:2011:GCF

[401] Maxwell Anselm and Steven H. Weintraub. A generalization of continued fractions. Journal of Number Theory, 131(12):2442–2460, December 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001739.

Anonymous:2012:PJa

[402] Anonymous. Pages 1–300 (January 2012). Journal of Number Theory, 132(1):??, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBa

[403] Anonymous. Editorial Board. Journal of Number Theory, 132(1): ifc, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002034.

deMathan:2012:LFB

[404] Bernard de Mathan. Linear forms at a basis of an algebraic number field. Journal of Number Theory, 132(1):1–25, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001752.

Najman:2012:TEC

[405] Filip Najman. Torsion of elliptic curves over cubic fields. Journal of Number Theory, 132(1):26–36, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1100179X.

Goswick:2012:SSO

[406] Lee M. Goswick, Emil W. Kiss, Gábor Moussong, and Nándor Simányi. Sums of squares and orthogonal integral vectors. Journal of Number Theory, 132(1):37–53, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1100182X.

Okazaki:2012:F

[407] Takeo Okazaki. L-functions of S_{3}(Γ_{2}(2, 4, 8)). Journal of Number Theory, 132(1):54–78, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001818.

Rouymi:2012:MNA

[408] D. Rouymi. Mollification et non annulation de fonctions L automorphes en niveau primaire. (French) [Mollification and not cancellation of automorphic L functions in primary level]. Journal of Number Theory, 132(1):79–93, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001727.

Hegyvari:2012:SRM

[409] Norbert Hegyvári. Some remarks on multilinear exponential sums with an application. Journal of Number Theory, 132(1):94–102, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001806.

Mollin:2012:RGT

[410] Richard A. Mollin and Anitha Srinivasan. Residuacity and genus theory of forms. Journal of Number Theory, 132(1):103–116, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001697. See corrigendum [550].

Li:2012:STA

[411] Hongze Li and Hao Pan. A Schur-type addition theorem for primes. Journal of Number Theory, 132(1):117–126, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001570.

Gekeler:2012:ZES

[412] Ernst-Ulrich Gekeler. Zeroes of Eisenstein series for principal congruence subgroups over rational function fields. Journal of Number Theory, 132(1):127–143, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001983.

Ohno:2012:CMS

[413] Yasuo Ohno, Jun ichi Okuda, and Wadim Zudilin. Cyclic q-MZSV sum. Journal of Number Theory, 132(1):144–155, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001910.

Sasaki:2012:GPB

[414] Yoshitaka Sasaki. On generalized poly-Bernoulli numbers and related L-functions. Journal of Number Theory, 132(1):156–170, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001971.

Liu:2012:ESI

[415] Kui Liu and Xiumin Ren. On exponential sums involving Fourier coefficients of cusp forms. Journal of Number Theory, 132(1):171–181, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001934.

Harper:2012:PKS

[416] Adam J. Harper. On a paper of K. Soundararajan on smooth numbers in arithmetic progressions. Journal of Number Theory, 132(1):182– 199, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001958.

Wakatsuki:2012:DFS

[417] Satoshi Wakatsuki. Dimension formulas for spaces of vector-valued Siegel cusp forms of degree two. Journal of Number Theory, 132(1):200– 253, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001922.

Shparlinski:2012:CHS

[418] Igor E. Shparlinski. On the convex hull of solutions to polynomial congruences. Journal of Number Theory, 132(1):254–257, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11001995.

Berkovich:2012:RIS

[419] Alexander Berkovich and William C. Jagy. On representation of an integer as the sum of three squares and ternary quadratic forms with the discriminants p^{2}, 16p^{2}. Journal of Number Theory, 132(1):258– 274, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002009.

Fili:2012:NER

[420] Paul Fili and Zachary Miner. Norms extremal with respect to the Mahler measure. Journal of Number Theory, 132(1):275–300, January 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1100196X.

Anonymous:2012:PF

[421] Anonymous. Pages 301–370 (February 2012). Journal of Number Theory, 132(2):??, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBb

[422] Anonymous. Editorial Board. Journal of Number Theory, 132(2):ifc, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002137.

Choie:2012:MTA

[423] YoungJu Choie and Winfried Kohnen. Mellin transforms attached to certain automorphic integrals. Journal of Number Theory, 132(2):301– 313, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002010.

Shen:2012:SIM

[424] Zhongyan Shen and Tianxin Cai. Some identities for multiple zeta values. Journal of Number Theory, 132(2):314–323, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001776.

Knopfmacher:2012:DSB

[425] Arnold Knopfmacher and Florian Luca. Digit sums of binomial sums. Journal of Number Theory, 132(2):324–331, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11001946.

Bayad:2012:AHD

[426] Abdelmejid Bayad and Abdelaziz Raouj. Arithmetic of higher dimensional Dedekind–Rademacher sums. Journal of Number Theory, 132(2):332–347, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11001909.

Smith:2012:NCC

[427] Barry R. Smith. Non-cyclic class groups and the Brumer–Stark conjecture. Journal of Number Theory, 132(2):348–370, February 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11001569.

Anonymous:2012:PMa

[428] Anonymous. Pages 371–484 (March 2012). Journal of Number Theory, 132(3):??, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBc

[429] Anonymous. Editorial Board. Journal of Number Theory, 132(3): ifc, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002459.

Bannai:2012:PDI

[430] Eiichi Bannai and Tsuyoshi Miezaki. On a property of 2-dimensional integral Euclidean lattices. Journal of Number Theory, 132(3):371– 378, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002113.

Xarles:2012:SAP

[431] Xavier Xarles. Squares in arithmetic progression over number fields. Journal of Number Theory, 132(3):379–389, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11002101.

Chen:2012:DGF

[432] Yong-Gao Chen, Ying Shi, and Jie Wu. Dynamics of Goldring’s w-function. Journal of Number Theory, 132(3):390–409, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11002186.

Bzdega:2012:SBC

[433] Bartłomiej Bzdega. Sparse binary cyclotomic polynomials. Journal of Number Theory, 132(3):410–413, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002198.

Seveso:2012:CRS

[434] Marco Adamo Seveso. Congruences and rationality of Stark–Heegner points. Journal of Number Theory, 132(3):414–447, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11002204.

Fujita:2012:MWG

[435] Yasutsugu Fujita and Tadahisa Nara. On the Mordell–Weil group of the elliptic curve y^{2} = x^{3} + n. Journal of Number Theory, 132(3):448– 466, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002241.

Roditty-Gershon:2012:SPT

[436] Edva Roditty-Gershon. Statistics for products of traces of high powers of the Frobenius class of hyperelliptic curves. Journal of Number Theory, 132(3):467–484, March 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002344.

Anonymous:2012:PAa

[437] Anonymous. Pages 485–868 (April 2012). Journal of Number Theory, 132(4):??, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBd

[438] Anonymous. Editorial Board. Journal of Number Theory, 132(4): ifc, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000157.

Cao:2012:PIA

[439] Wei Cao. A partial improvement of the Ax–Katz theorem. Journal of Number Theory, 132(4):485–494, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002277.

Rodriguez:2012:SBC

[440] José Alejandro Lara Rodríguez. On von Staudt for Bernoulli– Carlitz numbers. Journal of Number Theory, 132(4):495–501, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11002411.

Eilers:2012:SNI

[441] Søren Eilers and Ian Kiming. On some new invariants for shift equivalence for shifts of finite type. Journal of Number Theory, 132(4):502– 510, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1100223X.

Feng:2012:ZRZ

[442] Shaoji Feng. Zeros of the Riemann zeta function on the critical line. Journal of Number Theory, 132(4):511–542, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002319.

Johnson-Leung:2012:SMF

[443] Jennifer Johnson-Leung and Brooks Roberts. Siegel modular forms of degree two attached to Hilbert modular forms. Journal of Number Theory, 132(4):543–564, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002332.

Igarashi:2012:GOR

[444] Masahiro Igarashi. A generalization of Ohno’s relation for multiple zeta values. Journal of Number Theory, 132(4):565–578, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11002265.

Jeong:2012:LAF

[445] Sangtae Jeong. Locally analytic functions in a local field of positive characteristic. Journal of Number Theory, 132(4):579–589, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X11002381.

Taylor:2012:LCM

[446] Graeme Taylor. Lehmer’s conjecture for matrices over the ring of integers of some imaginary quadratic fields. Journal of Number Theory, 132(4):590–607, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002289.

Kumchev:2012:SAE

[447] Angel Kumchev and Taiyu Li. Sums of almost equal squares of primes. Journal of Number Theory, 132(4):608–636, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1100240X.

Torrey:2012:SCI

[448] Rebecca Torrey. On Serre’s conjecture over imaginary quadratic fields. Journal of Number Theory, 132(4):637–656, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11002368.

Occhipinti:2012:FEC

[449] Thomas Occhipinti. A family of elliptic curves of large rank. Journal of Number Theory, 132(4):657–665, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002393.

Munshi:2012:NSN

[450] Ritabrata Munshi. A note on simultaneous nonvanishing twists. Journal of Number Theory, 132(4):666–674, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002253.

Sun:2012:CSS

[451] Zhi-Hong Sun. Congruences for sequences similar to Euler numbers. Journal of Number Theory, 132(4):675–700, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1100237X.

Jia:2012:MVR

[452] Chaohua Jia. Mean value from representation of rational number as sum of two Egyptian fractions. Journal of Number Theory, 132(4):701– 713, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002307.

Banerjee:2012:HOL

[453] Abhishek Banerjee. Hecke operators on line bundles over modular curves. Journal of Number Theory, 132(4):714–734, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X11002320.

Liu:2012:DSF

[454] Zhixin Liu. Density of the sums of four cubes of primes. Journal of Number Theory, 132(4):735–747, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X11002290.

Kadiri:2012:EZD

[455] Habiba Kadiri and Nathan Ng. Explicit zero density theorems for Dedekind zeta functions. Journal of Number Theory, 132(4):748– 775, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X11002228.

Filaseta:2012:LPG

[456] Michael Filaseta, Travis Kidd, and Ognian Trifonov. Laguerre polynomials with Galois group A_{m} for each m. Journal of Number Theory, 132(4):776–805, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002502.

Okano:2012:CGG

[457] Keiji Okano. The commutativity of the Galois groups of the maximal unramified pro-p-extensions over the cyclotomic Z_{p}-extensions. Journal of Number Theory, 132(4):806–819, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002423.

Hiary:2012:UAF

[458] Ghaith A. Hiary and Michael O. Rubinstein. Uniform asymptotics for the full moment conjecture of the Riemann zeta function. Journal of Number Theory, 132(4):820–868, April 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002356.

Anonymous:2012:PMb

[459] Anonymous. Pages 869–1150 (May 2012). Journal of Number Theory, 132(5):??, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBe

[460] Anonymous. Editorial Board. Journal of Number Theory, 132(5): ifc, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000388.

Lau:2012:VHE

[461] Yuk-Kam Lau and Lilu Zhao. On a variance of Hecke eigenvalues in arithmetic progressions. Journal of Number Theory, 132(5):869– 887, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000108.

Corrales-Rodriganez:2012:RUQ

[462] Capi Corrales-Rodrigáñez. Rotations and units in quaternion algebras. Journal of Number Theory, 132(5):888–895, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000078.

Buium:2012:RDF

[463] Alexandru Buium and Arnab Saha. The ring of differential Fourier expansions. Journal of Number Theory, 132(5):896–937, May 2012. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X12000054.

Xiang:2012:CFE

[464] Zhengyu Xiang. A construction of the full eigenvariety of a reductive group. Journal of Number Theory, 132(5):938–952, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1200008X.

Cojocaru:2012:DGE

[465] Alina Carmen Cojocaru and Árpád Tóth. The distribution and growth of the elementary divisors of the reductions of an elliptic curve over a function field. Journal of Number Theory, 132(5):953– 965, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000029.

Buium:2012:HOD

[466] Alexandru Buium and Arnab Saha. Hecke operators on differential modular forms modp. Journal of Number Theory, 132(5):966– 997, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000042.

Reynolds:2012:PPE

[467] Jonathan Reynolds. Perfect powers in elliptic divisibility sequences. Journal of Number Theory, 132(5):998–1015, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000091.

Lu:2012:BSM

[468] Qing Lu. Bounds for the spectral mean value of central values of L-functions. Journal of Number Theory, 132(5):1016–1037, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000066.

Thompson:2012:PDE

[469] Lola Thompson. Polynomials with divisors of every degree. Journal of Number Theory, 132(5):1038–1053, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200011X.

Iyer:2012:GMS

[470] Geoffrey Iyer, Oleg Lazarev, Steven J. Miller, and Liyang Zhang. Generalized more sums than differences sets. Journal of Number Theory, 132(5):1054–1073, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000133.

Morra:2012:SRI

[471] Stefano Morra. On some representations of the Iwahori subgroup. Journal of Number Theory, 132(5):1074–1150, May 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11002435.

Anonymous:2012:PJb

[472] Anonymous. Pages 1151–1384 (June 2012). Journal of Number Theory, 132(6):??, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBf

[473] Anonymous. Editorial Board. Journal of Number Theory, 132(6): ifc, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000522.

Don:2012:PBO

[474] Henk Don. Polygons in billiard orbits. Journal of Number Theory, 132(6):1151–1163, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000236.

Zhu:2012:NGC

[475] Jun-Ming Zhu. A note on a generalized circular summation formula of theta functions. Journal of Number Theory, 132(6):1164– 1169, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000364.

Wei:2012:MFD

[476] Fu-Tsun Wei and Chia-Fu Yu. Mass formula of division algebras over global function fields. Journal of Number Theory, 132(6):1170– 1184, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000327.

Pollack:2012:ALQ

[477] Paul Pollack. The average least quadratic nonresidue modulo m and other variations on a theme of Erdős. Journal of Number Theory, 132(6):1185–1202, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000285.

Takemori:2012:ASE

[478] Sho Takemori. p-adic Siegel–Eisenstein series of degree two. Journal of Number Theory, 132(6):1203–1264, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12000248.

Languasco:2012:SMP

[479] Alessandro Languasco and Alessandro Zaccagnini. Sums of many primes. Journal of Number Theory, 132(6):1265–1283, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000352.

Liu:2012:CPA

[480] Zhixin Liu. Cubes of primes and almost prime. Journal of Number Theory, 132(6):1284–1294, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12000303.

Holmes:2012:CNT

[481] David Holmes. Computing Néron–Tate heights of points on hyperelliptic Jacobians. Journal of Number Theory, 132(6):1295–1305, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000273.

Kim:2012:NPD

[482] Byungchan Kim. On the number of partitions of n into k different parts. Journal of Number Theory, 132(6):1306–1313, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000297.

Joshi:2012:RFC

[483] Kirti Joshi. Remarks on the Fourier coefficients of modular forms. Journal of Number Theory, 132(6):1314–1336, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12000030.

Ahmadi:2012:ICE

[484] Omran Ahmadi and Robert Granger. On isogeny classes of Edwards curves over finite fields. Journal of Number Theory, 132(6):1337– 1358, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200025X.

Krishnamurthy:2012:DCF

[485] M. Krishnamurthy. Determination of cusp forms on GL(2 ) by coefficients restricted to quadratic subfields (with an appendix by Dipendra Prasad and Dinakar Ramakrishnan). Journal of Number Theory, 132(6):1359–1384, June 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000121.

Anonymous:2012:PJc

[486] Anonymous. Pages 1385–1644 (July 2012). Journal of Number Theory, 132(7):??, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBg

[487] Anonymous. Editorial Board. Journal of Number Theory, 132(7): ifc, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000753.

Feng:2012:GBZ

[488] Shaoji Feng and Xiaosheng Wu. On gaps between zeros of the Riemann zeta-function. Journal of Number Theory, 132(7):1385– 1397, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000261.

Furness:2012:DCR

[489] Amanda Furness and Adam E. Parker. On Dirichlet’s conjecture on relative class number one. Journal of Number Theory, 132(7):1398– 1403, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000443.

Luca:2012:EFC

[490] Florian Luca and Pantelimon Stănică. On the Euler function of the Catalan numbers. Journal of Number Theory, 132(7):1404–1424, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000479.

Linowitz:2012:SQA

[491] Benjamin Linowitz. Selectivity in quaternion algebras. Journal of Number Theory, 132(7):1425–1437, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12000480.

Hormann:2012:RPC

[492] Fritz Hörmann. On recursive properties of certain p-adic Whittaker functions. Journal of Number Theory, 132(7):1438–1482, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000431.

Sprung:2012:ITE

[493] Florian E. Ito Sprung. Iwasawa theory for elliptic curves at supersingular primes: a pair of main conjectures. Journal of Number Theory, 132(7):1483–1506, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000224.

Liao:2012:GNB

[494] Qunying Liao. The Gaussian normal basis and its trace basis over finite fields. Journal of Number Theory, 132(7):1507–1518, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000492.

Gugg:2012:MEC

[495] Chadwick Gugg. Modular equations for cubes of the Rogers–Ramanujan and Ramanujan–Göllnitz–Gordon functions and their associated continued fractions. Journal of Number Theory, 132(7):1519–1553, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000315.

Elsenhans:2012:QGC

[496] Andreas-Stephan Elsenhans and Jörg Jahnel. On the quasi-group of a cubic surface over a finite field. Journal of Number Theory, 132(7):1554– 1571, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000455.

Cui:2012:LAP

[497] Zhen Cui, Hongze Li, and Boqing Xue. Long arithmetic progressions in A + A + A with A a prime subset. Journal of Number Theory, 132(7):1572–1582, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000467.

Devic:2012:AOD

[498] Anna Devic and Richard Pink. Adelic openness for Drinfeld modules in special characteristic. Journal of Number Theory, 132(7):1583– 1625, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000583.

Li:2012:ATF

[499] Xiaoqing Li and Matthew P. Young. Additive twists of Fourier coefficients of symmetric-square lifts. Journal of Number Theory, 132(7): 1626–1640, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000571.

Roquette:2012:HWL

[500] Peter Roquette and David Goss. Heinrich–Wolfgang Leopoldt (22.8.1927–28.7.2011). Journal of Number Theory, 132(7):1641–1644, July 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000340.

Anonymous:2012:PAb

[501] Anonymous. Pages 1645–1836 (August 2012). Journal of Number Theory, 132(8):??, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBh

[502] Anonymous. Editorial Board. Journal of Number Theory, 132(8): ifc, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001035.

Bayad:2012:MVF

[503] Abdelmejid Bayad and Abdelaziz Raouj. Mean values of L-functions and Dedekind sums. Journal of Number Theory, 132(8):1645–1652, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000509.

DiazyDiaz:2012:CCF

[504] Francisco Diaz y Diaz and Eduardo Friedman. Colmez cones for fundamental units of totally real cubic fields. Journal of Number Theory, 132(8):1653–1663, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000844.

Qi:2012:MDS

[505] Zhi Qi. Morita’s duality for split reductive groups. Journal of Number Theory, 132(8):1664–1685, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12000716.

Pinter:2012:APR

[506] Ákos Pintér and Volker Ziegler. On arithmetic progressions in recurrences — a new characterization of the Fibonacci sequence. Journal of Number Theory, 132(8):1686–1706, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12000728.

Bradshaw:2012:HPA

[507] Robert Bradshaw and William Stein. Heegner points and the arithmetic of elliptic curves over ring class extensions. Journal of Number Theory, 132(8):1707–1719, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000856.

Thiel:2012:AGP

[508] Carsten Thiel. Adelic geometry and polarity. Journal of Number Theory, 132(8):1720–1730, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000698.

Guo:2012:PSC

[509] Victor J. W. Guo and Jiang Zeng. Proof of some conjectures of Z.-W. Sun on congruences for Apéry polynomials. Journal of Number Theory, 132(8):1731–1740, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000613.

Heath-Brown:2012:CRP

[510] D. R. Heath-Brown and Lillian B. Pierce. Counting rational points on smooth cyclic covers. Journal of Number Theory, 132(8):1741– 1757, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000674.

Hancl:2012:EIM

[511] Jaroslav Hančl, Marko Leinonen, Kalle Leppälä, and Tapani Matalaaho. Explicit irrationality measures for continued fractions. Journal of Number Theory, 132(8):1758–1769, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200087X.

Shapiro:2012:FMA

[512] Ilya Shapiro. Frobenius map and the p-adic Gamma function. Journal of Number Theory, 132(8):1770–1779, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12000947.

Saber:2012:CPM

[513] Hicham Saber and Abdellah Sebbar. On the critical points of modular forms. Journal of Number Theory, 132(8):1780–1787, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000923.

Girstmair:2012:DSV

[514] Kurt Girstmair. Dedekind sums in the vicinity of quadratic irrationals. Journal of Number Theory, 132(8):1788–1792, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000972.

Ahn:2012:CNO

[515] Jeoung-Hwan Ahn, Gérard Boutteaux, Soun-Hi Kwon, and Stéphane Louboutin. The class number one problem for some non-normal CMfields of degree 2p. Journal of Number Theory, 132(8):1793–1806, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000935.

Biro:2012:ZFI

[516] András Biró and Andrew Granville. Zeta functions for ideal classes in real quadratic fields, at s = 0. Journal of Number Theory, 132 (8):1807–1829, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000601.

Jones:2012:CPC

[517] Lenny Jones. Corrigendum to “Polynomial Cunningham chains” [J. Number Theory 131 (11) (2011) 2100–2106]. Journal of Number Theory, 132(8):1830–1836, August 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1200073X. See [375].

Anonymous:2012:PS

[518] Anonymous. Pages 1837–2074 (September 2012). Journal of Number Theory, 132(9):??, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBi

[519] Anonymous. Editorial Board. Journal of Number Theory, 132(9):ifc, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001205.

Deshouillers:2012:SAS

[520] Jean-Marc Deshouillers, Michael Drmota, and Johannes F. Morgenbesser. Subsequences of automatic sequences indexed by ⌊n^{c}⌋ and correlations. Journal of Number Theory, 132(9):1837–1866, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12000959.

Lagemann:2012:DAS

[521] Thorsten Lagemann. Distribution of Artin–Schreier extensions. Journal of Number Theory, 132(9):1867–1887, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001011.

Kihel:2012:SGQ

[522] Omar Kihel and Jason Lizotte. Small generators of quadratic fields and reduced elements. Journal of Number Theory, 132(9):1888–1895, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000996.

Sakai:2012:EQF

[523] Yuichi Sakai and Hiroyuki Tsutsumi. Extremal quasimodular forms for low-level congruence subgroups. Journal of Number Theory, 132 (9):1896–1909, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000893.

Kondo:2012:LEF

[524] Satoshi Kondo and Seidai Yasuda. Local L and epsilon factors in Hecke eigenvalues. Journal of Number Theory, 132(9):1910–1948, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200090X.

Kikuta:2012:AHE

[525] Toshiyuki Kikuta and Yoshinori Mizuno. On p-adic Hermitian Eisenstein series and p-adic Siegel cusp forms. Journal of Number Theory, 132 (9):1949–1961, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12000911.

Condon:2012:AED

[526] John D. Condon. Asymptotic expansion of the difference of two Mahler measures. Journal of Number Theory, 132(9):1962–1983, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001126.

Kondo:2012:BBT

[527] Hiroki Kondo, Shingo Saito, and Tatsushi Tanaka. The Bowman–Bradley theorem for multiple zeta-star values. Journal of Number Theory, 132 (9):1984–2002, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1200114X.

Vasiu:2012:BRF

[528] Adrian Vasiu and Thomas Zink. Boundedness results for finite flat group schemes over discrete valuation rings of mixed characteristic. Journal of Number Theory, 132(9):2003–2019, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1200100X.

Lewko:2012:VBD

[529] Allison Lewko and Mark Lewko. A variational Barban–Davenport– Halberstam Theorem. Journal of Number Theory, 132(9):2020–2045, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000649.

Suvitie:2012:SCP

[530] Eeva Suvitie. On the shifted convolution problem in mean. Journal of Number Theory, 132(9):2046–2064, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000960.

Zhang:2012:TPI

[531] Yong Zhang and Tianxin Cai. n-Tuples of positive integers with the same second elementary symmetric function value and the same product. Journal of Number Theory, 132(9):2065–2074, September 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000984.

Anonymous:2012:PO

[532] Anonymous. Pages 2075–2378 (October 2012). Journal of Number Theory, 132(10):??, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBj

[533] Anonymous. Editorial Board. Journal of Number Theory, 132(10): ifc, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001461.

Dujella:2012:PDR

[534] Andrej Dujella and Clemens Fuchs. On a problem of Diophantus for rationals. Journal of Number Theory, 132(10):2075–2083, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001187.

Hattori:2012:RCF

[535] Shin Hattori. Ramification correspondence of finite flat group schemes over equal and mixed characteristic local fields. Journal of Number Theory, 132(10):2084–2102, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12001175.

Bertolin:2012:HIE

[536] Cristiana Bertolin. Homological interpretation of extensions and biextensions of 1-motives. Journal of Number Theory, 132(10):2103–2131, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001254.

Mishiba:2012:APM

[537] Yoshinori Mishiba. On v-adic periods of t-motives. Journal of Number Theory, 132(10):2132–2165, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001357.

Eakin:2012:CPM

[538] R. T. Eakin. A combinatorial partition of Mersenne numbers arising from spectroscopy. Journal of Number Theory, 132(10):2166–2183, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001308.

Yang:2012:CVM

[539] Zifeng Yang. On the characteristic p valued measure associated to Drinfeld discriminant. Journal of Number Theory, 132(10):2184–2227, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200128X.

Balasubramanian:2012:ZQM

[540] R. Balasubramanian and Sanoli Gun. On zeros of quasi-modular forms. Journal of Number Theory, 132(10):2228–2241, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001345.

Akatsuka:2012:CEE

[541] Hirotaka Akatsuka. Conditional estimates for error terms related to the distribution of zeros of ζ′(s). Journal of Number Theory, 132 (10):2242–2257, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001333.

Barman:2012:IIA

[542] Rupam Barman and Anupam Saikia. Iwasawa λ-invariants of p-adic measures on Z_{p}^{n} and their Γ-transforms. Journal of Number Theory, 132(10):2258–2266, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001369.

Charkani:2012:GPB

[543] Mohamed E. Charkani and Abdulaziz Deajim. Generating a power basis over a Dedekind ring. Journal of Number Theory, 132(10):2267– 2276, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001278.

Hegarty:2012:CFB

[544] Peter Hegarty. On m-covering families of Beatty sequences with irrational moduli. Journal of Number Theory, 132(10):2277–2296, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001370.

Hong:2012:MGI

[545] Hoon Hong, Eunjeong Lee, Hyang-Sook Lee, and Cheol-Min Park. Maximum gap in (inverse) cyclotomic polynomial. Journal of Number Theory, 132(10):2297–2315, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001291.

Berchenko-Kogan:2012:MPS

[546] Yakov Berchenko-Kogan. Minimum product sets sizes in nonabelian groups. Journal of Number Theory, 132(10):2316–2335, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001321.

Blache:2012:VES

[547] Régis Blache. Valuation of exponential sums and the generic first slope for Artin–Schreier curves. Journal of Number Theory, 132(10):2336– 2352, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001382.

Bagis:2012:CEA

[548] N. D. Bagis and M. L. Glasser. Conjectures on the evaluation of alternative modular bases and formulas approximating 1/π. Journal of Number Theory, 132(10):2353–2370, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200131X.

Vindas:2012:CEB

[549] Jasson Vindas. Chebyshev estimates for Beurling generalized prime numbers. I. Journal of Number Theory, 132(10):2371–2376, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001400.

Mollin:2012:CRG

[550] Richard A. Mollin and Anitha Srinivasan. Corrigendum to “Residuacity and genus theory of forms” [J. Number Theory 132 (1) (2012) 103–116]. Journal of Number Theory, 132(10):2377, October 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001266. See [410].

Anonymous:2012:PN

[551] Anonymous. Pages 2379–2724 (November 2012). Journal of Number Theory, 132(11):??, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBk

[552] Anonymous. Editorial Board. Journal of Number Theory, 132(11):ifc, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001850.

Cheng:2012:RTB

[553] Chuangxun Cheng. Rank two Breuil modules: Basic structures. Journal of Number Theory, 132(11):2379–2396, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001394.

Jorge:2012:RL

[554] Grasiele C. Jorge, Agnaldo J. Ferrari, and Sueli I. R. Costa. Rotated D_{n}-lattices. Journal of Number Theory, 132(11):2397–2406, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001424.

Tsuzuki:2012:SAC

[555] Masao Tsuzuki. Spectral average of central values of automorphic Lfunctions for holomorphic cusp forms on SO_{0} (m, 2 ), I. Journal of Number Theory, 132(11):2407–2454, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001412.

Koh:2012:GEF

[556] Doowon Koh and Chun-Yen Shen. The generalized Erdős–Falconer distance problems in vector spaces over finite fields. Journal of Number Theory, 132(11):2455–2473, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12001436.

Arenas-Carmona:2012:ETO

[557] Luis Arenas-Carmona. An embedding theorem for orders in central simple algebras. Journal of Number Theory, 132(11):2474–2480, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001448.

Loper:2012:GRI

[558] K. Alan Loper and Nicholas J. Werner. Generalized rings of integervalued polynomials. Journal of Number Theory, 132(11):2481–2490, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200159X.

Ward:2012:ACN

[559] Kenneth Ward. Asymptotics of class number and genus for abelian extensions of an algebraic function field. Journal of Number Theory, 132 (11):2491–2498, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001667.

Gorodnik:2012:ANH

[560] A. Gorodnik and S. Kadyrov. Algebraic numbers, hyperbolicity, and density modulo one. Journal of Number Theory, 132(11):2499–2509, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001680.

Kritzer:2012:LDP

[561] Peter Kritzer and Friedrich Pillichshammer. Low discrepancy polynomial lattice point sets. Journal of Number Theory, 132(11):2510–2534, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001564.

Vallieres:2012:GRO

[562] Daniel Vallières. On a generalization of the rank one Rubin–Stark conjecture. Journal of Number Theory, 132(11):2535–2567, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001588.

Tarnauceanu:2012:GMI

[563] Marius Tărnăuceanu. A generalization of Menon’s identity. Journal of Number Theory, 132(11):2568–2573, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200162X.

Sha:2012:GGG

[564] Min Sha and Linsheng Yin. Galois groups and genera of a kind of quasi-cyclotomic function fields. Journal of Number Theory, 132 (11):2574–2581, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001655.

Javaheri:2012:CTC

[565] Mohammad Javaheri. A Collatz-type conjecture on the set of rational numbers. Journal of Number Theory, 132(11):2582–2594, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001552.

Kaczorowski:2012:EVS

[566] J. Kaczorowski and G. Molteni. Extremal values for the sum ∑_{r=1}^{τ} e(a2^{r}/q). Journal of Number Theory, 132(11):2595–2603, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001783.

Luca:2012:DE

[567] Florian Luca and Gökhan Soydan. On the Diophantine equation 2^{m} + nx^{2} = y^{n}. Journal of Number Theory, 132(11):2604–2609, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001606.

Siksek:2012:NMW

[568] Samir Siksek. On the number of Mordell–Weil generators for cubic surfaces. Journal of Number Theory, 132(11):2610–2629, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001710.

Zhang:2012:PN

[569] Wenpeng Zhang and Tingting Wang. Powerful numbers in (1^{k} + 1)(2^{k} + 1) ⋯ (n^{k} + 1). Journal of Number Theory, 132(11):2630–2635, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200176X.

Lim:2012:PTD

[570] Meng Fai Lim. Poitou–Tate duality over extensions of global fields. Journal of Number Theory, 132(11):2636–2672, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001576.

Sun:2012:SAP

[571] Zhi-Wei Sun. On sums of Apéry polynomials and related congruences. Journal of Number Theory, 132(11):2673–2699, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001643.

Walling:2012:HER

[572] Lynne H. Walling. Hecke eigenvalues and relations for degree 2 Siegel Eisenstein series. Journal of Number Theory, 132(11):2700–2723, November 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001692.

Anonymous:2012:PD

[573] Anonymous. Pages 2725–3108 (December 2012). Journal of Number Theory, 132(12):??, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2012:EBl

[574] Anonymous. Editorial Board. Journal of Number Theory, 132(12):ifc, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200217X.

Chen:2012:WSC

[575] Feng-Juan Chen and Yong-Gao Chen. Weighted sums of consecutive values of a polynomial. Journal of Number Theory, 132(12):2725–2735, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001631.

Lapkova:2012:ELB

[576] Kostadinka Lapkova. Effective lower bound for the class number of a certain family of real quadratic fields. Journal of Number Theory, 132 (12):2736–2747, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001801.

Arenas-Carmona:2012:MSO

[577] Luis Arenas-Carmona. Maximal selectivity for orders in fields. Journal of Number Theory, 132(12):2748–2755, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001825.

Madritsch:2012:SDF

[578] Manfred G. Madritsch. The sum-of-digits function of canonical number systems: Distribution in residue classes. Journal of Number Theory, 132 (12):2756–2772, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001746.

Miezaki:2012:FOL

[579] Tsuyoshi Miezaki. Frames in the odd Leech lattice. Journal of Number Theory, 132(12):2773–2778, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001813.

Lamzouri:2012:RCS

[580] Youness Lamzouri and Alexandru Zaharescu. Randomness of character sums modulo m. Journal of Number Theory, 132(12):2779–2792, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001758.

Andrade:2012:MVH

[581] J. C. Andrade and J. P. Keating. The mean value of L(\frac{1}{2}, χ) in the hyperelliptic ensemble. Journal of Number Theory, 132(12):2793–2816, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001679.

Xiao:2012:CIN

[582] Xiao Xiao. Computing isomorphism numbers of F -crystals using the level torsions. Journal of Number Theory, 132(12):2817–2835, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200193X.

Berrizbeitia:2012:CNW

[583] Pedro Berrizbeitia, J. G. Fernandes, Marcos J. González, Florian Luca, and V. Janitzio Mejía Huguet. On Cullen numbers which are both Riesel and Sierpiński numbers. Journal of Number Theory, 132(12): 2836–2841, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001722.

Laurincikas:2012:JUR

[584] Antanas Laurinčikas. On joint universality of the Riemann zetafunction and Hurwitz zeta-functions. Journal of Number Theory, 132 (12):2842–2853, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001771.

Kim:2012:IIF

[585] Taekyun Kim. Identities involving Frobenius–Euler polynomials arising from non-linear differential equations. Journal of Number Theory, 132 (12):2854–2865, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001904. See corrigendum [655].

Miller:2012:LLZ

[586] Steven J. Miller and Ryan Peckner. Low-lying zeros of number field L-functions. Journal of Number Theory, 132(12):2866–2891, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001928.

Adam:2012:TAL

[587] David Adam and Youssef Fares. On two affine-like dynamical systems in a local field. Journal of Number Theory, 132(12):2892–2906, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001618.

Weingartner:2012:DFI

[588] Andreas Weingartner. The distribution functions of σ(n)/n and n/φ(n), II. Journal of Number Theory, 132(12):2907–2921, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001953.

Bubboloni:2012:CSS

[589] Daniela Bubboloni, Florian Luca, and Pablo Spiga. Compositions of n satisfying some coprimality conditions. Journal of Number Theory, 132 (12):2922–2946, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002089.

Jo:2012:ERC

[590] Sihun Jo. An entry of Ramanujan on continued fraction involving the gamma function in his notebooks. Journal of Number Theory, 132 (12):2947–2954, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002077.

Liu:2012:TFI

[591] Zhi-Guo Liu. A theta function identity of degree eight and Eisenstein series identities. Journal of Number Theory, 132(12):2955–2966, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200203X.

Li:2012:GSS

[592] Yan Li and Su Hu. Gauss sums over some matrix groups. Journal of Number Theory, 132(12):2967–2976, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002065.

Kim:2012:AHT

[593] Min-Soo Kim and Su Hu. On p-adic Hurwitz-type Euler zeta functions. Journal of Number Theory, 132(12):2977–3015, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002028.

Languasco:2012:DPP

[594] Alessandro Languasco and Alessandro Zaccagnini. A Diophantine problem with a prime and three squares of primes. Journal of Number Theory, 132(12):3016–3028, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002120.

Byeon:2012:HPE

[595] Dongho Byeon. Heegner points on elliptic curves with a rational torsion point. Journal of Number Theory, 132(12):3029–3036, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002090.

Pollack:2012:PNP

[596] Paul Pollack and Vladimir Shevelev. On perfect and near-perfect numbers. Journal of Number Theory, 132(12):3037–3046, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002041.

Yang:2012:PNN

[597] Quan-Hui Yang and Yong-Gao Chen. Partitions of natural numbers with the same weighted representation functions. Journal of Number Theory, 132(12):3047–3055, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002004.

Hajdu:2012:GCT

[598] L. Hajdu and M. Szikszai. On the GCD-s of k consecutive terms of Lucas sequences. Journal of Number Theory, 132(12):3056–3069, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001734.

Miller:2012:FCA

[599] Stephen D. Miller and Siddhartha Sahi. Fourier coefficients of automorphic forms, character variety orbits, and small representations. Journal of Number Theory, 132(12):3070–3108, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001837.

Anonymous:2013:PJa

[600] Anonymous. Pages 1–368 (January 2013). Journal of Number Theory, 133(1):??, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBa

[601] Anonymous. Editorial Board. Journal of Number Theory, 133(1): ifc, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002417.

Xia:2013:PEC

[602] Li meng Xia. The parameterized-Euler-constant function γ_{α}(z). Journal of Number Theory, 133(1):1–11, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002107.

Mu:2013:MTP

[603] Quanwu Mu and Qiang Wu. The measure of totally positive algebraic integers. Journal of Number Theory, 133(1):12–19, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12001989.

Matthes:2013:RTL

[604] Roland Matthes. Regularized theta lifts and Niebur-type Poincaré series on n-dimensional hyperbolic space. Journal of Number Theory, 133(1):20–47, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002260.

Nakamura:2013:RDZ

[605] Takashi Nakamura and Koji Tasaka. Remarks on double zeta values of level 2. Journal of Number Theory, 133(1):48–54, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002272.

Mukhamedov:2013:E

[606] Farrukh Mukhamedov and Mansoor Saburov. On equation x^{q} = a over Q_{p}. Journal of Number Theory, 133(1):55–58, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002284.

Leriche:2013:CQS

[607] Amandine Leriche. Cubic, quartic and sextic Pólya fields. Journal of Number Theory, 133(1):59–71, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002144.

Alnaser:2013:BWN

[608] Ala’Jamil Alnaser. Bounds for Waring’s number modP^{⇕} in number fields. Journal of Number Theory, 133(1):72–82, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002156.

Godber:2013:ATF

[609] Daniel Godber. Additive twists of Fourier coefficients of modular forms. Journal of Number Theory, 133(1):83–104, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002326.

Meng:2013:RMH

[610] Xianmeng Meng. On RSA moduli with half of the bits prescribed. Journal of Number Theory, 133(1):105–109, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002302.

Boylan:2013:QPR

[611] Hatice Boylan and Nils-Peter Skoruppa. A quick proof of reciprocity for Hecke Gauss sums. Journal of Number Theory, 133(1):110–114, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002247.

Jeon:2013:IFE

[612] Daeyeol Jeon, Chang Heon Kim, and Yoonjin Lee. Infinite families of elliptic curves over Dihedral quartic number fields. Journal of Number Theory, 133(1):115–122, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12002119.

Virdol:2013:MCZ

[613] Cristian Virdol. The meromorphic continuation of the zeta function of a product of Hilbert and Picard modular surfaces over CM-fields. Journal of Number Theory, 133(1):123–130, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002259.

Mattarei:2013:CCB

[614] Sandro Mattarei and Roberto Tauraso. Congruences for central binomial sums and finite polylogarithms. Journal of Number Theory, 133(1):131–157, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001965.

Karanikolopoulos:2013:RCG

[615] Sotiris Karanikolopoulos and Aristides Kontogeorgis. Representation of cyclic groups in positive characteristic and Weierstrass semigroups. Journal of Number Theory, 133(1):158–175, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002132.

Godinho:2013:PAS

[616] Hemar Godinho, Michael P. Knapp, and Paulo H. A. Rodrigues. Pairs of additive sextic forms. Journal of Number Theory, 133(1):176–194, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001990.

Charles:2013:PQE

[617] Richard Charles, Tim Huber, and Andoni Mendoza. Parameterizations for quintic Eisenstein series. Journal of Number Theory, 133(1):195– 214, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002016.

Chen:2013:PSS

[618] Ching-An Chen and Ming-Guang Leu. On a proposition of Samuel and 2-stage Euclidean algorithm in global fields. Journal of Number Theory, 133(1):215–225, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002314.

ElBachraoui:2013:SLT

[619] Mohamed El Bachraoui. Sums of Liouville type over primitive pairs and quadruples and some integer representations. Journal of Number Theory, 133(1):226–241, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12002053.

Albeverio:2013:RZT

[620] Sergio Albeverio and Claudio Cacciapuoti. The Riemann zeta in terms of the dilogarithm. Journal of Number Theory, 133(1):242–277, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001977.

Letang:2013:ASI

[621] Delia Letang. Automorphic spectral identities and applications to automorphic L-functions on GL_{2} . Journal of Number Theory, 133(1):278– 317, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001795.

Reinholz:2013:FNC

[622] Lindsey Reinholz, Blair K. Spearman, and Qiduan Yang. Families of non-congruent numbers with arbitrarily many prime factors. Journal of Number Theory, 133(1):318–327, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200234X.

Reznikov:2013:TPC

[623] Andre Reznikov. Torus period coefficients on PGL(2) and Dirichlet series. Journal of Number Theory, 133(1):328–342, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002363.

Ezome:2013:EPP

[624] Tony Ezome and Reynald Lercier. Elliptic periods and primality proving. Journal of Number Theory, 133(1):343–368, January 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002296.

Anonymous:2013:PF

[625] Anonymous. Pages 369–824 (February 2013). Journal of Number Theory, 133(2):??, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBb

[626] Anonymous. Editorial Board. Journal of Number Theory, 133(2):ifc, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002788.

Burhanuddin:2013:ECL

[627] Iftikhar A. Burhanuddin and Ming-Deh A. Huang. Elliptic curves with large Shafarevich–Tate group. Journal of Number Theory, 133 (2):369–374, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002399.

Akiyama:2013:DSP

[628] Shigeki Akiyama and Vilmos Komornik. Discrete spectra and Pisot numbers. Journal of Number Theory, 133(2):375–390, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002387.

Wang:2013:CPC

[629] Venecia Wang. A counterexample to the prime conjecture of expressing numbers using just ones. Journal of Number Theory, 133(2):391– 397, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002491. See addendum [738].

Yau:2013:SPE

[630] Stephen S. T. Yau and Linda Zhao. Sharp polynomial estimate of integral points in right-angled simplices. Journal of Number Theory, 133(2):398–425, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002351.

Auil:2013:AGS

[631] F. Auil. An algorithm to generate square-free numbers and to compute the Möbius function. Journal of Number Theory, 133(2):426–436, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200248X.

Vasuki:2013:RME

[632] K. R. Vasuki and G. Sharath. On Ramanujan’s modular equations of degree 21. Journal of Number Theory, 133(2):437–445, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002569.

Adrian:2013:NRL

[633] Moshe Adrian. A new realization of the Langlands correspondence for PGL(2, F). Journal of Number Theory, 133(2):446–474, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002508.

Koo:2013:SVP

[634] Ja Kyung Koo and Dong Hwa Shin. Singular values of principal moduli. Journal of Number Theory, 133(2):475–483, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002521.

Khrennikov:2013:CMP

[635] Andrei Khrennikov and Ekaterina Yurova. Criteria of measure-preserving for p-adic dynamical systems in terms of the van der Put basis. Journal of Number Theory, 133(2):484–491, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002594.

Im:2013:PRQ

[636] Bo-Hae Im. Positive rank quadratic twists of four elliptic curves. Journal of Number Theory, 133(2):492–500, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002697.

Dummigan:2013:PMD

[637] Neil Dummigan and Srilakshmi Krishnamoorthy. Powers of 2 in modular degrees of modular abelian varieties. Journal of Number Theory, 133(2):501–522, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002703.

Shiomi:2013:OCF

[638] Daisuke Shiomi. Ordinary cyclotomic function fields. Journal of Number Theory, 133(2):523–533, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12002545.

Finch:2013:NSR

[639] Carrie Finch, Joshua Harrington, and Lenny Jones. Nonlinear Sierpiński and Riesel numbers. Journal of Number Theory, 133(2):534–544, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200265X.

Humphries:2013:DWS

[640] Peter Humphries. The distribution of weighted sums of the Liouville function and Pólya’s conjecture. Journal of Number Theory, 133(2):545–582, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002570.

Miyazaki:2013:GCR

[641] Takafumi Miyazaki. Generalizations of classical results on Jeśmanowicz’ conjecture concerning Pythagorean triples. Journal of Number Theory, 133(2):583–595, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002648.

Li:2013:RDS

[642] Zhong hua Li. Regularized double shuffle and Ohno–Zagier relations of multiple zeta values. Journal of Number Theory, 133(2):596–610, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1200251X.

Sodaigui:2013:CSE

[643] Bouchaïb Sodaïgui. Classes de Steinitz d’extensions galoisiennes à groupe de Galois de centre non trivial. (French) [Steinitz classes of Galois extensions to the Galois group of non-trivial center]. Journal of Number Theory, 133(2):611–619, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002612.

Buckingham:2013:LGF

[644] Paul R. Buckingham. Local and global fundamental classes for multiquadratic extensions. Journal of Number Theory, 133(2):620–638, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002600.

Goulden:2013:LOT

[645] Ian P. Goulden, Duc Khiem Huynh, Rishikesh, and Michael O. Rubinstein. Lower order terms for the moments of symplectic and orthogonal families of L-functions. Journal of Number Theory, 133(2):639– 674, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002557.

Fu:2013:DAS

[646] Lei Fu. Deformation of ℓ-adic sheaves with undeformed local monodromy. Journal of Number Theory, 133(2):675–691, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002533.

Mezo:2013:RFI

[647] István Mező. A q-Raabe formula and an integral of the fourth Jacobi theta function. Journal of Number Theory, 133(2):692–704, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002715.

Sander:2013:AGC

[648] J. W. Sander and T. Sander. Adding generators in cyclic groups. Journal of Number Theory, 133(2):705–718, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12002673.

Cooper:2013:DE

[649] Shaun Cooper and Heung Yeung Lam. On the Diophantine equation n^{2} = x^{2} + by^{2} + cz^{2}. Journal of Number Theory, 133(2):719–737, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002661.

Fieker:2013:CIP

[650] Claus Fieker, István Gaál, and Michael Pohst. On computing integral points of a Mordell curve over rational function fields in characteristic > 3. Journal of Number Theory, 133(2):738–750, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002582.

Hanusa:2013:NSC

[651] Christopher R. H. Hanusa and Rishi Nath. The number of selfconjugate core partitions. Journal of Number Theory, 133(2):751–768, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002636.

Pilehrood:2013:CFE

[652] Kh. Hessami Pilehrood and T. Hessami Pilehrood. On a continued fraction expansion for Euler’s constant. Journal of Number Theory, 133(2):769–786, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002624.

Ichimura:2013:ILI

[653] H. Ichimura, S. Nakajima, and H. Sumida-Takahashi. On the Iwasawa lambda invariant of an imaginary abelian field of conductor 3p^{n+1}. Journal of Number Theory, 133(2):787–801, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002685.

Pollio:2013:MPL

[654] Timothy P. Pollio and Andrei S. Rapinchuk. The multinorm principle for linearly disjoint Galois extensions. Journal of Number Theory, 133(2):802–821, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002375.

Kim:2013:CII

[655] Taekyun Kim. Corrigendum to “Identities involving Frobenius–Euler polynomials arising from non-linear differential equations” [J. Number Theory 132 (12) (2012) 2854–2865]. Journal of Number Theory, 133(2):822–824, February 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002338. See [585].

Anonymous:2013:EBc

[656] Anonymous. Editorial Board. Journal of Number Theory, 133(3): ifc, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002922.

Rosen:2013:DRH

[657] Michael Rosen. David R. Hayes: Some remarks on his life and work. Journal of Number Theory, 133(3):825–829, March 2013. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001941.

Angles:2013:PKV

[658] Bruno Anglès and Lenny Taelman. On a problem à la Kummer– Vandiver for function fields. Journal of Number Theory, 133(3):830– 841, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000650. See addendum [739].

Banaszak:2013:SSM

[659] Grzegorz Banaszak and Cristian D. Popescu. The Stickelberger splitting map and Euler systems in the K-theory of number fields. Journal of Number Theory, 133(3):842–870, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X11001892.

Bosser:2013:HTD

[660] Vincent Bosser and Federico Pellarin. Hankel-type determinants and Drinfeld quasi-modular forms. Journal of Number Theory, 133(3):871– 887, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000339.

Chang:2013:GCA

[661] Seunghwan Chang and Yoonjin Lee. Galois characters arising from Drinfeld modules. Journal of Number Theory, 133(3):888–896, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002740.

Cojocaru:2013:ACD

[662] Alina Carmen Cojocaru and Andrew Michael Shulman. An average Chebotarev Density Theorem for generic rank 2 Drinfeld modules with complex multiplication. Journal of Number Theory, 133(3):897– 914, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12001916.

Dasgupta:2013:CPF

[663] Samit Dasgupta. A conjectural product formula for Brumer–Stark units over real quadratic fields. Journal of Number Theory, 133(3):915– 925, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000704.

Dominguez:2013:QFC

[664] Carlos Dominguez, Steven J. Miller, and Siman Wong. Quadratic fields with cyclic 2-class groups. Journal of Number Theory, 133(3):926– 939, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12000881.

Gekeler:2013:ZCP

[665] Ernst-Ulrich Gekeler and Philipp Stopp. On the zeroes of certain periodic functions over valued fields of positive characteristic. Journal of Number Theory, 133(3):940–954, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12000637.

Goss:2013:SFP

[666] David Goss. On the L-series of F. Pellarin. Journal of Number Theory, 133(3):955–962, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X11002216.

Grant:2013:ALN

[667] David Grant. On an analogue of the Lutz–Nagell Theorem for hyperelliptic curves. Journal of Number Theory, 133(3):963–969, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12001163.

Gross:2013:TFR

[668] Benedict H. Gross. Trivial L-functions for the rational function field. Journal of Number Theory, 133(3):970–976, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000686.

Hoelscher:2013:ICT

[669] Jing Long Hoelscher. Infinite class towers for function fields. Journal of Number Theory, 133(3):977–982, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001138.

Lubin:2013:EAM

[670] Jonathan Lubin. Elementary analytic methods in higher ramification theory. Journal of Number Theory, 133(3):983–999, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000868.

Lutes:2013:AIV

[671] Brad A. Lutes and Matthew A. Papanikolas. Algebraic independence of values of Goss L-functions at s = 1. Journal of Number Theory, 133(3):1000–1011, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001151.

Papikian:2013:ESO

[672] Mihran Papikian. D-elliptic sheaves and odd Jacobians. Journal of Number Theory, 133(3):1012–1026, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12000625.

Eddin:2013:EUB

[673] Sumaia Saad Eddin. Explicit upper bounds for the Stieltjes constants. Journal of Number Theory, 133(3):1027–1044, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002739.

Tangedal:2013:ECG

[674] Brett A. Tangedal and Paul T. Young. Explicit computation of Gross– Stark units over real quadratic fields. Journal of Number Theory, 133(3):1045–1061, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12001709.

Zywina:2013:ECF

[675] David Zywina. Explicit class field theory for global function fields. Journal of Number Theory, 133(3):1062–1078, March 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12000662.

Anonymous:2013:PAa

[676] Anonymous. Pages 1079–1434 (April 2013). Journal of Number Theory, 133(4):??, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBd

[677] Anonymous. Editorial Board. Journal of Number Theory, 133(4): ifc, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003381.

Kazalicki:2013:CNC

[678] Matija Kazalicki. Congruent numbers and congruences between halfintegral weight modular forms. Journal of Number Theory, 133(4):1079– 1085, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003113.

Zhang:2013:MVI

[679] Deyu Zhang and Wenguang Zhai. On the mean value of the index of composition of an integral ideal (II). Journal of Number Theory, 133(4):1086–1110, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002752.

Xiong:2013:RSW

[680] Wei Xiong. A regularized Siegel–Weil formula on U(1,1) and application to theta lifting from U(1) to U(1, 1). Journal of Number Theory, 133(4):1111–1134, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12003149.

Peng:2013:ZN

[681] Yuejian Peng and K. P. S. Bhaskara Rao. On Zumkeller numbers. Journal of Number Theory, 133(4):1135–1155, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003137.

Tang:2013:NFC

[682] Hengcai Tang. A note on the Fourier coefficients of Hecke–Maass forms. Journal of Number Theory, 133(4):1156–1167, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12003010.

Dabrowski:2013:VBR

[683] Andrzej Dabrowski and Maciej Ulas. Variations on the Brocard– Ramanujan equation. Journal of Number Theory, 133(4):1168–1185, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12002880.

Louboutin:2013:SFM

[684] Stéphane R. Louboutin and Marc Munsch. The second and fourth moments of theta functions at their central point. Journal of Number Theory, 133(4):1186–1193, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12003083.

Sugiyama:2013:LCS

[685] Rin Sugiyama. Lefschetz classes of simple factors of the Jacobian variety of a Fermat curve of prime degree over finite fields. Journal of Number Theory, 133(4):1194–1206, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003162.

Ugolini:2013:GAM

[686] S. Ugolini. Graphs associated with the map X ↦ X + X^{−1} in finite fields of characteristic three and five. Journal of Number Theory, 133(4):1207–1228, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002909.

Sankaranarayanan:2013:NSD

[687] A. Sankaranarayanan, N. Saradha, and Saurabh Kumar Singh. Normality of some decimals generated by primes in a residue class. Journal of Number Theory, 133(4):1229–1250, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003174.

Giorgilli:2013:RPS

[688] A. Giorgilli and G. Molteni. Representation of a 2-power as sum of k 2-powers: a recursive formula. Journal of Number Theory, 133(4):1251– 1261, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003058.

Cwalina:2013:LBD

[689] Karol Cwalina and Tomasz Schoen. A linear bound on the dimension in Green–Ruzsa’s theorem. Journal of Number Theory, 133(4):1262– 1269, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003046.

Karpuk:2013:CWG

[690] David A. Karpuk. Cohomology of the Weil group of a p-adic field. Journal of Number Theory, 133(4):1270–1288, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003186.

McGown:2013:SSP

[691] Kevin J. McGown. On the second smallest prime non-residue. Journal of Number Theory, 133(4):1289–1299, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003034.

Choi:2013:BSW

[692] SoYoung Choi and Chang Heon Kim. Basis for the space of weakly holomorphic modular forms in higher level cases. Journal of Number Theory, 133(4):1300–1311, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003071.

Cheng:2013:MR

[693] Chuangxun Cheng. Mod p representations of SL_{2} (Q_{p}). Journal of Number Theory, 133(4):1312–1330, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003022.

Fite:2013:ARA

[694] Francesc Fité. Artin representations attached to pairs of isogenous abelian varieties. Journal of Number Theory, 133(4):1331–1345, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12002892.

Fouvry:2013:SEK

[695] Étienne Fouvry. Sum of Euler–Kronecker constants over consecutive cyclotomic fields. Journal of Number Theory, 133(4):1346–1361, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003125.

Ichikawa:2013:CBS

[696] Takashi Ichikawa. Congruences between Siegel modular forms II. Journal of Number Theory, 133(4):1362–1371, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1200306X.

Burger:2013:CDR

[697] Edward B. Burger, David C. Clyde, Cory H. Colbert, Gea Hyun Shin, and Zhaoning Wang. Canonical Diophantine representations of natural numbers with respect to quadratic “bases”. Journal of Number Theory, 133(4):1372–1388, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12002764.

Chretien:2013:MWM

[698] P. Chrétien and M. Matignon. Maximal wild monodromy in unequal characteristic. Journal of Number Theory, 133(4):1389–1408, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003150.

Cesaratto:2013:NRS

[699] Eda Cesaratto, Joachim von zur Gathen, and Guillermo Matera. The number of reducible space curves over a finite field. Journal of Number Theory, 133(4):1409–1434, April 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003101.

Anonymous:2013:PM

[700] Anonymous. Pages 1435–1808 (May 2013). Journal of Number Theory, 133(5):??, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBe

[701] Anonymous. Editorial Board. Journal of Number Theory, 133(5): ifc, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000152.

Unver:2013:DIR

[702] Sinan Ünver. Drinfel’d–Ihara relations for p-adic multi-zeta values. Journal of Number Theory, 133(5):1435–1483, May 2013. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003289.

Chen:2013:EFF

[703] Hua Chen, Jianhua Chen, Guangxing Cai, Xiaochuan Ai, and Silan Zhang. Explicit formulas for the fourth moment of mixed exponential sums. Journal of Number Theory, 133(5):1484–1491, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12003265.

Brown:2013:LSS

[704] Jim Brown and Rodney Keaton. Level stripping for Siegel modular forms with reducible Galois representations. Journal of Number Theory, 133(5):1492–1501, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12003241.

Vishe:2013:FAC

[705] Pankaj Vishe. A fast algorithm to compute L(1/2, f × χ_{q}). Journal of Number Theory, 133(5):1502–1524, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003277.

Klingler:2013:SSW

[706] L. Klingler and M. Marshall. E-sequences and the Stone–Weierstrass Theorem. Journal of Number Theory, 133(5):1525–1536, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12003253.

Weiss:2013:PGT

[707] Benjamin L. Weiss. Probabilistic Galois theory over p-adic fields. Journal of Number Theory, 133(5):1537–1563, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003356.

Sun:2013:CBP

[708] Yidong Sun, Xiaojuan Wu, and Jujuan Zhuang. Congruences on the Bell polynomials and the derangement polynomials. Journal of Number Theory, 133(5):1564–1571, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003228.

Sun:2013:CI

[709] Zhi-Hong Sun. Congruences involving \binom{2k}{k}^{2}\binom{3k}{k}. Journal of Number Theory, 133(5):1572–1595, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1200323X.

Gologlu:2013:SCK

[710] Faruk Göloğlu, Gary McGuire, and Richard Moloney. Some congruences of Kloosterman sums and their characteristic polynomials. Journal of Number Theory, 133(5):1596–1607, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12003332.

Walling:2013:FAH

[711] Lynne H. Walling. A formula for the action of Hecke operators on halfintegral weight Siegel modular forms and applications. Journal of Number Theory, 133(5):1608–1644, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003290. See [939].

Fujita:2013:GEC

[712] Yasutsugu Fujita. Generators for the elliptic curve y^{2} = x^{3} − nx of rank at least three. Journal of Number Theory, 133(5):1645– 1662, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003344.

Pellarin:2013:GCM

[713] Federico Pellarin. On the generalized Carlitz module. Journal of Number Theory, 133(5):1663–1692, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X12003319.

Schoen:2013:HMC

[714] Tomasz Schoen and Ilya D. Shkredov. Higher moments of convolutions. Journal of Number Theory, 133(5):1693–1737, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12003320.

Felix:2013:HRG

[715] Adam Tyler Felix. Higher rank generalizations of Fomenko’s conjecture. Journal of Number Theory, 133(5):1738–1751, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12003009.

Heuberger:2013:AWN

[716] Clemens Heuberger and Daniel Krenn. Analysis of width-w nonadjacent forms to imaginary quadratic bases. Journal of Number Theory, 133(5):1752–1808, May 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12003204.

Anonymous:2013:PJb

[717] Anonymous. Pages 1809–2142 (June 2013). Journal of Number Theory, 133(6):??, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBf

[718] Anonymous. Editorial Board. Journal of Number Theory, 133(6): ifc, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000449.

Graves:2013:ACN

[719] Hester Graves and M. Ram Murty. The abc conjecture and non– Wieferich primes in arithmetic progressions. Journal of Number Theory, 133(6):1809–1813, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X12003368.

Morisawa:2013:PE

[720] Takayuki Morisawa. On the ℓ-part of the Z_{p_{1}} × ⋯ × Z_{p_{s}}-extension of Q. Journal of Number Theory, 133(6):1814–1826, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12003095.

Alfes:2013:TTC

[721] Claudia Alfes and Stephan Ehlen. Twisted traces of CM values of weak Maass forms. Journal of Number Theory, 133(6):1827–1845, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003307.

Jeong:2013:CEA

[722] Sangtae Jeong. Characterization of ergodicity of T -adic maps on F_{2}[T] using digit derivatives basis. Journal of Number Theory, 133(6):1846– 1863, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000127.

El-Guindy:2013:EFD

[723] Ahmad El-Guindy and Matthew A. Papanikolas. Explicit formulas for Drinfeld modules and their periods. Journal of Number Theory, 133(6):1864–1886, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000085.

Choi:2013:CAJ

[724] Youn-Seo Choi. Cubic analogues of the Jacobian theta functions and the Jacobian elliptic functions. Journal of Number Theory, 133(6):1887– 1906, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000061.

Onodera:2013:ADR

[725] Kazuhiro Onodera. An analogue of the Dedekind–Rademacher sum and certain ray class invariants of real quadratic fields. Journal of Number Theory, 133(6):1907–1931, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000097.

Yao:2013:NRL

[726] Olivia X. M. Yao and Ernest X. W. Xia. New Ramanujan-like congruences modulo powers of 2 and 3 for overpartitions. Journal of Number Theory, 133(6):1932–1949, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000036.

Sun:2013:CCL

[727] Zhi-Hong Sun. Congruences concerning Legendre polynomials II. Journal of Number Theory, 133(6):1950–1976, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1300005X.

Agashe:2013:CNT

[728] Amod Agashe and Saikat Biswas. Constructing non-trivial elements of the Shafarevich–Tate group of an abelian variety over a number field. Journal of Number Theory, 133(6):1977–1990, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000103.

Constales:2013:FEH

[729] D. Constales, R. De Almeida, and Rolf Sören Kraußhar. The Fourier expansion of the hypermonogenic generalized trigonometric and elliptic functions. Journal of Number Theory, 133(6):1991–2004, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X12003198.

Qian:2013:PCA

[730] Guoyou Qian, Qianrong Tan, and Shaofang Hong. On the periodicity of a class of arithmetic functions associated with multiplicative functions. Journal of Number Theory, 133(6):2005–2020, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000139.

Borges:2013:CMV

[731] Herivelto Borges and Ricardo Conceição. On the characterization of minimal value set polynomials. Journal of Number Theory, 133(6):2021– 2035, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X12003216.

Arce-Nazario:2013:NSD

[732] R. Arce-Nazario, F. Castro, and R. Figueroa. On the number of solutions of ∑_{i=1}^{11} \frac{1}{x_{i}} = 1 in distinct odd natural numbers. Journal of Number Theory, 133(6):2036–2046, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000206.

Dodge:2013:RCS

[733] Joel Dodge and Cristian D. Popescu. The refined Coates–Sinnott conjecture for characteristic p global fields. Journal of Number Theory, 133(6):2047–2065, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000048.

Blondeau:2013:LNE

[734] Julien Blondeau. Lifting of nearly extraordinary Galois representations. Journal of Number Theory, 133(6):2066–2082, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X12002727.

Dujella:2013:ECT

[735] Andrej Dujella and Juan Carlos Peral. Elliptic curves and Triangles with three rational medians. Journal of Number Theory, 133(6):2083– 2091, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000231.

Katayama:2013:CNF

[736] Koji Katayama. Class number formulas, Kronecker’s limit formulas, Chowla–Selberg formulas and the generalized gamma functions. Journal of Number Theory, 133(6):2092–2120, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1300022X.

Bravo:2013:CGF

[737] Jhon J. Bravo and Florian Luca. Coincidences in generalized Fibonacci sequences. Journal of Number Theory, 133(6):2121–2137, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000073. See corrigendum [808].

Wang:2013:ACP

[738] Venecia Wang. Addendum to “A counterexample to the prime conjecture of expressing numbers using just ones” [J. Number Theory 133 (2) (2013) 391–397]. Journal of Number Theory, 133(6): 2138, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000267. See [629].

Angles:2013:SCM

[739] Bruno Anglès and Lenny Taelman. The Spiegelungssatz for the Carlitz module; Addendum to “On a problem à la Kummer–Vandiver for function fields” [J. Number Theory 133 (3) (2013) 830–841]. Journal of Number Theory, 133(6):2139–2142, June 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000292. See [658].

Anonymous:2013:PJc

[740] Anonymous. Pages 2143–2464 (July 2013). Journal of Number Theory, 133(7):??, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBg

[741] Anonymous. Editorial Board. Journal of Number Theory, 133(7): ifc, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000656.

Tang:2013:LBH

[742] Hengcai Tang. Lower bound for the higher moment of symmetric square L-functions. Journal of Number Theory, 133(7):2143– 2152, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000310.

Lascu:2013:GKT

[743] Dan Lascu. A Gauss–Kuzmin-type problem for a family of continued fraction expansions. Journal of Number Theory, 133(7):2153– 2181, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000279.

Schweiger:2013:ISS

[744] Fritz Schweiger. Invariant simplices for subtractive algorithms. Journal of Number Theory, 133(7):2182–2185, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000255.

Broderick:2013:BAS

[745] Ryan Broderick, Lior Fishman, and David Simmons. Badly approximable systems of affine forms and incompressibility on fractals. Journal of Number Theory, 133(7):2186–2205, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000243.

Wei:2013:ERT

[746] Chuanan Wei and Dianxuan Gong. Extensions of Ramanujan’s two formulas for 1/π. Journal of Number Theory, 133(7):2206–2216, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13000565.

Alkan:2013:SMF

[747] Emre Alkan and Haydar Göral. On sums over the Möbius function and discrepancy of fractions. Journal of Number Theory, 133(7):2217– 2239, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000218.

Machide:2013:GVS

[748] T. Machide. Generators for vector spaces spanned by double zeta values with even weight. Journal of Number Theory, 133(7):2240– 2246, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000309.

Petrov:2013:EDM

[749] Aleksandar Petrov. A-expansions of Drinfeld modular forms. Journal of Number Theory, 133(7):2247–2266, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000346.

Li:2013:OSE

[750] Jiyou Li. On the Odlyzko–Stanley enumeration problem and Waring’s problem over finite fields. Journal of Number Theory, 133(7):2267– 2276, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000334.

Kelmer:2013:HTF

[751] Dubi Kelmer. Hybrid trace formula for a non-uniform irreducible lattice in PSL_{2}(R)^{n}. Journal of Number Theory, 133(7):2277–2306, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000280.

Qi:2013:IRP

[752] Feng Qi. Integral representations and properties of Stirling numbers of the first kind. Journal of Number Theory, 133(7):2307– 2319, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000371.

Roulleau:2013:TCF

[753] Xavier Roulleau. On the Tate conjecture for the Fano surfaces of cubic threefolds. Journal of Number Theory, 133(7):2320–2323, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1300036X.

All:2013:AAR

[754] Timothy All. On p-adic annihilators of real ideal classes. Journal of Number Theory, 133(7):2324–2338, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000358.

Johnson:2013:HEP

[755] Matthew Leander Johnson. Hecke eigenforms as products of eigenforms. Journal of Number Theory, 133(7):2339–2362, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000024.

Chen:2013:FCM

[756] Feng-Juan Chen and Yong-Gao Chen. On the Frobenius conjecture for Markoff numbers. Journal of Number Theory, 133(7):2363– 2373, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000401.

Gyarmati:2013:RLS

[757] Katalin Gyarmati, Sergei Konyagin, and András Sárközy. On the reducibility of large sets of residues modulo p. Journal of Number Theory, 133(7):2374–2397, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000528.

Wright:2013:QRN

[758] Steve Wright. Quadratic residues and non-residues in arithmetic progression. Journal of Number Theory, 133(7):2398–2430, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000541.

Murata:2013:PSQ

[759] Manabu Murata. On primitive solutions of quadratic Diophantine equation with four variables. Journal of Number Theory, 133(7):2431– 2454, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000322.

Yamagishi:2013:NCP

[760] Masakazu Yamagishi. A note on Chebyshev polynomials, cyclotomic polynomials and twin primes. Journal of Number Theory, 133(7):2455– 2463, July 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000589.

Anonymous:2013:PAb

[761] Anonymous. Pages 2465–2842 (August 2013). Journal of Number Theory, 133(8):??, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBh

[762] Anonymous. Editorial Board. Journal of Number Theory, 133(8): ifc, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000875.

Mortici:2013:ACG

[763] Cristinel Mortici. Approximating the constants of Glaisher–Kinkelin type. Journal of Number Theory, 133(8):2465–2469, August 2013. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X13000413.

Chen:2013:P

[764] Yong-Gao Chen, Ming-Liang Gong, and Xiao-Zhi Ren. On the products (1^{ℓ} + 1)(2^{ℓ} + 1) ⋯ (n^{ℓ} + 1). Journal of Number Theory, 133 (8):2470–2474, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300053X.

Eie:2013:GED

[765] Minking Eie and Chuan-Sheng Wei. Generalizations of Euler decomposition and their applications. Journal of Number Theory, 133(8):2475– 2495, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000607.

Agarwal:2013:YLB

[766] Mahesh Agarwal and Krzysztof Klosin. Yoshida lifts and the Bloch–Kato conjecture for the convolution L-function. Journal of Number Theory, 133(8):2496–2537, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000590.

Feng:2013:LSZ

[767] ShaoJi Feng and XiaoSheng Wu. On large spacing of the zeros of the Riemann zeta-function. Journal of Number Theory, 133(8):2538– 2566, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000577.

Jeon:2013:WMP

[768] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Weak Maass– Poincaré series and weight 3/2 mock modular forms. Journal of Number Theory, 133(8):2567–2587, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000619. See

corrigendum [1471].

Nelson:2013:SAC

[769] Paul D. Nelson. Stable averages of central values of Rankin–Selberg L-functions: Some new variants. Journal of Number Theory, 133 (8):2588–2615, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000516.

An:2013:GOF

[770] Ta Thi Hoai An and Nguyen Thi Ngoc Diep. Genus one factors of curves defined by separated variable polynomials. Journal of Number Theory, 133(8):2616–2634, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000395.

Minami:2013:RV

[771] Norihiko Minami. On the random variable {1, 2, …, n}^{r} ∋ (k_{1}, k_{2}, …, k_{r}) ↦ gcd(n, k_{1}k_{2} ⋯ k_{r}) ∈ N. Journal of Number Theory, 133(8):2635–2647, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000620.

Le:2013:RDO

[772] Phong Le. Regular decomposition of ordinarity in generic exponential sums. Journal of Number Theory, 133(8):2648–2683, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13000553.

Kellner:2013:QRZ

[773] Bernd C. Kellner. On quotients of Riemann zeta values at odd and even integer arguments. Journal of Number Theory, 133(8):2684–2698, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000796.

Chen:2013:AER

[774] Chao-Ping Chen and Long Lin. Asymptotic expansions related to Glaisher–Kinkelin constant based on the Bell polynomials. Journal of Number Theory, 133(8):2699–2705, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000826.

Jung:2013:NMV

[775] Hwanyup Jung. Note on the mean value of L(\frac{1}{2}, χ) in the hyperelliptic ensemble. Journal of Number Theory, 133(8):2706–2714, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000760.

Duggan:2013:UE

[776] Mark C. Duggan and Larry J. Gerstein. Unimodular embeddings. Journal of Number Theory, 133(8):2715–2721, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1300070X.

Yang:2013:NRD

[777] Qingjie Yang. A negative result of direct sum of integral symplectic matrices. Journal of Number Theory, 133(8):2722–2728, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13000802.

DeIeso:2013:ENI

[778] Marco De Ieso. Existence de normes invariantes pour GL_{2}. (French) [Existence of invariant norms for GL_{2}]. Journal of Number Theory, 133(8):2729–2755, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000759.

Graves:2013:GRE

[779] Hester Graves. Growth results and Euclidean ideals. Journal of Number Theory, 133(8):2756–2769, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000383.

Li:2013:APP

[780] Zane Kun Li and Alexander W. Walker. Arithmetic properties of Picard– Fuchs equations and holonomic recurrences. Journal of Number Theory, 133(8):2770–2793, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000723.

Sun:2013:FTO

[781] Zhi-Wei Sun. On functions taking only prime values. Journal of Number Theory, 133(8):2794–2812, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13000747.

Cremona:2013:CAP

[782] John E. Cremona and Thotsaphon Thongjunthug. The complex AGM, periods of elliptic curves over C and complex elliptic logarithms. Journal of Number Theory, 133(8):2813–2841, August 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000735.

Anonymous:2013:PS

[783] Anonymous. Pages 2843–3206 (September 2013). Journal of Number Theory, 133(9):??, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBi

[784] Anonymous. Editorial Board. Journal of Number Theory, 133(9):ifc, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300111X.

Demeyer:2013:CSP

[785] Jeroen Demeyer and Antonella Perucca. The constant of the support problem for abelian varieties. Journal of Number Theory, 133 (9):2843–2856, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000632.

Chen:2013:HTC

[786] Yong-Gao Chen and Jin-Hui Fang. Hegyvári’s Theorem on complete sequences. Journal of Number Theory, 133(9):2857–2862, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1300084X.

Kerr:2013:DVZ

[787] Bryce Kerr and Igor E. Shparlinski. On the distribution of values and zeros of polynomial systems over arbitrary sets. Journal of Number Theory, 133(9):2863–2873, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000838.

Jeong:2013:TEA

[788] Sangtae Jeong. Toward the ergodicity of p-adic 1-Lipschitz functions represented by the van der Put series. Journal of Number Theory, 133 (9):2874–2891, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000772.

Rodriguez:2013:BDB

[789] Jose Rodriguez. Bounding the degree of Belyi polynomials. Journal of Number Theory, 133(9):2892–2900, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000711.

Zhao:2013:ECR

[790] Yu Zhao. Elliptic curves over real quadratic fields with everywhere good reduction and a non-trivial 3-division point. Journal of Number Theory, 133(9):2901–2913, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000784.

Sun:2013:CBF

[791] Zhi-Wei Sun. Connections between p = x^{2} + 3y^{2} and Franel numbers. Journal of Number Theory, 133(9):2914–2928, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000851.

Chen:2013:NPD

[792] William Y. C. Chen, Kathy Q. Ji, Hai-Tao Jin, and Erin Y. Y. Shen. On the number of partitions with designated summands. Journal of Number Theory, 133(9):2929–2938, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000814.

Vinh:2013:GEF

[793] Le Anh Vinh. On the generalized Erdős–Falconer distance problems over finite fields. Journal of Number Theory, 133(9):2939–2947, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000425.

Silverman:2013:PDD

[794] Joseph H. Silverman. Primitive divisors, dynamical Zsigmondy sets, and Vojta’s conjecture. Journal of Number Theory, 133(9):2948–2963, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000978.

Pasten:2013:PVP

[795] Hector Pasten. Powerful values of polynomials and a conjecture of Vojta. Journal of Number Theory, 133(9):2964–2998, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000930.

Hegyvari:2013:SRB

[796] Norbert Hegyvári and François Hennecart. A structure result for bricks in Heisenberg groups. Journal of Number Theory, 133(9):2999–3006, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001054.

Castillo:2013:HOS

[797] Daniel Macias Castillo. On higher order Stickelberger-type theorems. Journal of Number Theory, 133(9):3007–3032, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001005.

Kriz:2013:CCM

[798] Daniel Kriz. On a conjecture concerning the maximal cross number of unique factorization indexed sequences. Journal of Number Theory, 133 (9):3033–3056, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300098X.

Zhang:2013:NLT

[799] Chong Zhang. A note on the local theta correspondence for unitary similitude dual pairs. Journal of Number Theory, 133(9):3057–3064, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001066.

Verhoek:2013:MAV

[800] Hendrik Verhoek. Modularity of abelian varieties over Q with bad reduction in one prime only. Journal of Number Theory, 133(9):3065–3098, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001078.

Barman:2013:ECS

[801] Rupam Barman and Gautam Kalita. Elliptic curves and special values of Gaussian hypergeometric series. Journal of Number Theory, 133 (9):3099–3111, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001042.

Rault:2013:UBR

[802] Patrick X. Rault. On uniform bounds for rational points on rational curves of arbitrary degree. Journal of Number Theory, 133(9): 3112–3118, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001029.

Kaygisiz:2013:NMC

[803] Kenan Kaygısız and Adem Şahin. A new method to compute the terms of generalized order-k Fibonacci numbers. Journal of Number Theory, 133(9):3119–3126, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13000991.

Mu:2013:SRR

[804] Yan-Ping Mu. Symmetric recurrence relations and binomial transforms. Journal of Number Theory, 133(9):3127–3137, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13000954.

Bulant:2013:MGC

[805] Michal Bulant and Radan Kučera. On a modification of the group of circular units of a real abelian field. Journal of Number Theory, 133 (9):3138–3148, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001030.

Mahnkopf:2013:TIR

[806] Joachim Mahnkopf. On truncation of irreducible representations of Chevalley groups. Journal of Number Theory, 133(9):3149–3182, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300108X.

Dajani:2013:ECF

[807] Karma Dajani, Cor Kraaikamp, and Niels van der Wekken. Ergodicity of N-continued fraction expansions. Journal of Number Theory, 133 (9):3183–3204, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001091.

Bravo:2013:CCG

[808] Jhon J. Bravo and Florian Luca. Corrigendum to “Coincidences in generalized Fibonacci sequences” [J. Number Theory 133 (6) (2013) 2121–2137]. Journal of Number Theory, 133(9):3205, September 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13000929. See [737].

Anonymous:2013:PO

[809] Anonymous. Pages 3207–3564 (October 2013). Journal of Number Theory, 133(10):??, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBj

[810] Anonymous. Editorial Board. Journal of Number Theory, 133(10): ifc, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001443.

Banaszak:2013:WKD

[811] Grzegorz Banaszak. Wild kernels and divisibility in K-groups of global fields. Journal of Number Theory, 133(10):3207–3244, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001017.

Dere:2013:UPT

[812] Rahime Dere, Yilmaz Simsek, and H. M. Srivastava. A unified presentation of three families of generalized Apostol type polynomials based upon the theory of the umbral calculus and the umbral algebra. Journal of Number Theory, 133(10):3245–3263, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13000966.

Petrow:2013:TMV

[813] Ian Petrow. Transition mean values of shifted convolution sums. Journal of Number Theory, 133(10):3264–3282, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001236.

Ulas:2013:NDS

[814] Maciej Ulas. A note on Diophantine systems involving three symmetric polynomials. Journal of Number Theory, 133(10):3283–3295, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001297.

Wu:2013:ITL

[815] Chenyan Wu. Irreducibility of theta lifting for unitary groups. Journal of Number Theory, 133(10):3296–3318, October 2013. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001340.

Hall-Seelig:2013:NLB

[816] L. L. Hall-Seelig. New lower bounds for the Ihara function A(q) for small primes. Journal of Number Theory, 133(10):3319–3324, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001224.

Ozman:2013:PT

[817] Ekin Ozman. On polyquadratic twists of X_{0}(N). Journal of Number Theory, 133(10):3325–3338, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13001212.

Liu:2013:PQE

[818] Zhixin Liu. On pairs of quadratic equations in primes and powers of 2. Journal of Number Theory, 133(10):3339–3347, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001261.

Araci:2013:EKA

[819] Serkan Araci, Mehmet Acikgoz, and Erdoğan Şen. On the extended Kim’s p-adic q-deformed fermionic integrals in the p-adic integer ring. Journal of Number Theory, 133(10):3348–3361, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001273.

Bel:2013:ASN

[820] Pierre Bel. Approximation simultanée d’un nombre v-adique et de son carré par des nombres algébriques. (French) [Simultaneous approximation of a v-adic number and its square by algebraic numbers]. Journal of Number Theory, 133(10):3362–3380, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001248.

Jones:2013:PEC

[821] Nathan Jones. Pairs of elliptic curves with maximal Galois representations. Journal of Number Theory, 133(10):3381–3393, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13000942.

Wakatsuki:2013:MFD

[822] Satoshi Wakatsuki. Multiplicity formulas for discrete series representations in L^{2}(Γ \ Sp(2, R)). Journal of Number Theory, 133(10):3394– 3425, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001339.

Kariyama:2013:FDD

[823] Kazutoshi Kariyama. The formal degree of the discrete series representations of GL_{m}(D). Journal of Number Theory, 133(10):3426–3452, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001352.

Bagis:2013:RTA

[824] N. D. Bagis and M. L. Glasser. Ramanujan type 1/π approximation formulas. Journal of Number Theory, 133(10):3453–3469, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001303.

Walji:2013:SSP

[825] Nahid Walji. On the size of Satake parameters for unitary cuspidal automorphic representations for GL(4). Journal of Number Theory, 133 (10):3470–3484, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001388.

Heim:2013:SST

[826] Bernhard Heim and Atsushi Murase. Symmetries for Siegel theta functions, Borcherds lifts and automorphic Green functions. Journal of Number Theory, 133(10):3485–3499, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001285.

deVera-Piquero:2013:SCS

[827] Carlos de Vera-Piquero. The Shimura covering of a Shimura curve: automorphisms and étale subcoverings. Journal of Number Theory, 133 (10):3500–3516, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300139X.

Pacheco:2013:RPJ

[828] Amílcar Pacheco. Rational points of Jacobian varieties in pro-ℓ towers of function fields. Journal of Number Theory, 133(10):3517–3523, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300125X.

Guiraud:2013:SMB

[829] David-Alexandre Guiraud. On semisimple ℓ-modular Bernstein-blocks of a p-adic general linear group. Journal of Number Theory, 133 (10):3524–3548, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001327.

Hou:2013:CPB

[830] Xiang dong Hou. A class of permutation binomials over finite fields. Journal of Number Theory, 133(10):3549–3558, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001315.

Mei:2013:NCO

[831] Shu-Yuan Mei. A new class of ordinary integers. Journal of Number Theory, 133(10):3559–3564, October 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13001406.

Anonymous:2013:PN

[832] Anonymous. Pages 3565–3994 (November 2013). Journal of Number Theory, 133(11):??, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBk

[833] Anonymous. Editorial Board. Journal of Number Theory, 133(11):ifc, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001716.

Cosgrave:2013:SRM

[834] John B. Cosgrave and Karl Dilcher. Sums of reciprocals modulo composite integers. Journal of Number Theory, 133(11):3565–3577, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001364.

Hu:2013:UDP

[835] Su Hu and Yan Li. On a uniformly distributed phenomenon in matrix groups. Journal of Number Theory, 133(11):3578–3588, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13001418.

Kumar:2013:SCE

[836] Narasimha Kumar. On sign changes of q-exponents of generalized modular functions. Journal of Number Theory, 133(11):3589–3597, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001492.

Byott:2013:GSG

[837] Nigel P. Byott and G. Griffith Elder. Galois scaffolds and Galois module structure in extensions of characteristic p local fields of degree p^{2}. Journal of Number Theory, 133(11):3598–3610, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1300142X.

Mao:2013:IBR

[838] Renrong Mao. Inequalities between rank moments of overpartitions. Journal of Number Theory, 133(11):3611–3619, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001571.

Chen:2013:DDA

[839] Huei-Jeng Chen. Distribution of Diophantine approximation exponents for algebraic quantities in finite characteristic. Journal of Number Theory, 133(11):3620–3644, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13001546.

Ha:2013:LPP

[840] Junsoo Ha. On the least prime primitive root. Journal of Number Theory, 133(11):3645–3669, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001595.

Xu:2013:GSL

[841] Xiaodong Xu. A generalization of sets without long arithmetic progressions based on Szekeres algorithm. Journal of Number Theory, 133 (11):3670–3677, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001601.

Mao:2013:RPM

[842] Renrong Mao. Ranks of partitions modulo 10. Journal of Number Theory, 133(11):3678–3702, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001662.

Radu:2013:EAP

[843] Silviu Radu and James A. Sellers. An extensive analysis of the parity of broken 3-diamond partitions. Journal of Number Theory, 133 (11):3703–3716, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001613.

Schultz:2013:GGF

[844] Andrew Schultz and Robert Walker. A generalization of the Gaussian formula and a q-analog of Fleck’s congruence. Journal of Number Theory, 133(11):3717–3738, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001558.

Murty:2013:MFE

[845] M. Ram Murty and Hector Pasten. Modular forms and effective Diophantine approximation. Journal of Number Theory, 133(11):3739–3754, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001583.

Ionica:2013:PBA

[846] Sorina Ionica. Pairing-based algorithms for Jacobians of genus 2 curves with maximal endomorphism ring. Journal of Number Theory, 133 (11):3755–3770, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001509.

Banescu:2013:NES

[847] Magdalena Bănescu and Dumitru Popa. New extensions of some classical theorems in number theory. Journal of Number Theory, 133 (11):3771–3795, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001637.

Bourla:2013:BDJ

[848] Avraham Bourla. Bounding differences in Jager Pairs. Journal of Number Theory, 133(11):3796–3809, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001376.

Ozeki:2013:TRA

[849] Yoshiyasu Ozeki. Torsion representations arising from (φ, Ĝ)-modules. Journal of Number Theory, 133(11):3810–3861, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001625.

Zhai:2013:ABF

[850] Shuai Zhai. Average behavior of Fourier coefficients of cusp forms over sum of two squares. Journal of Number Theory, 133(11):3862–3876, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001650.

Ye:2013:ZDA

[851] Yangbo Ye and Deyu Zhang. Zero density for automorphic L-functions. Journal of Number Theory, 133(11):3877–3901, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001649.

Mortenson:2013:TIS

[852] Eric Mortenson. Threefield identities and simultaneous representations of primes by binary quadratic forms. Journal of Number Theory, 133 (11):3902–3920, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001522.

Ghitza:2013:CVV

[853] Alexandru Ghitza, Nathan C. Ryan, and David Sulon. Computations of vector-valued Siegel modular forms. Journal of Number Theory, 133 (11):3921–3940, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001510.

Fili:2013:ODS

[854] Paul Fili and Zachary Miner. Orthogonal decomposition of the space of algebraic numbers and Lehmer’s problem. Journal of Number Theory, 133(11):3941–3981, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300156X.

Kwon:2013:TDS

[855] DoYong Kwon. A two-dimensional singular function via Sturmian words in base β. Journal of Number Theory, 133(11):3982–3994, November 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13000115.

Anonymous:2013:PD

[856] Anonymous. Pages 3995–4250 (December 2013). Journal of Number Theory, 133(12):??, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2013:EBl

[857] Anonymous. Editorial Board. Journal of Number Theory, 133(12):ifc, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002060.

Young:2013:RSM

[858] Paul Thomas Young. Rational series for multiple zeta and log gamma functions. Journal of Number Theory, 133(12):3995–4009, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001686.

Seo:2013:FLE

[859] Soogil Seo. On first layers of Z_{p}-extensions. Journal of Number Theory, 133(12):4010–4023, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001698.

Henniart:2013:URS

[860] Guy Henniart and Luis Lomelí. Uniqueness of Rankin–Selberg products. Journal of Number Theory, 133(12):4024–4035, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001674.

Chen:2013:CCP

[861] Shi-Chao Chen. Congruences for t-core partition functions. Journal of Number Theory, 133(12):4036–4046, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001790.

Xia:2013:MZS

[862] Li meng Xia and Caixia Shen. Minimal zero-sum sequences of length four over cyclic group with order n = p^{α}q^{β}. Journal of Number Theory, 133(12):4047–4068, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001844.

Hajli:2013:FZA

[863] Mounir Hajli. Sur la fonction Zêta associée au Laplacien singulier Δ_{\overline{O(⇕)}_{∞}}. (French) [On the Zeta function associated with the singular Laplacien Δ_{\overline{O(⇕)}_{∞}}]. Journal of Number Theory, 133(12):4069–4139, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001832.

Grundman:2013:TTE

[864] Helen G. Grundman and Daniel P. Wisniewski. Tetranomial Thue equations. Journal of Number Theory, 133(12):4140–4174, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13001819.

Cho:2013:PPN

[865] Peter J. Cho and Henry H. Kim. Probabilistic properties of number fields. Journal of Number Theory, 133(12):4175–4187, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001856.

Frisch:2013:SGP

[866] Sophie Frisch and Daniel Krenn. Sylow p-groups of polynomial permutations on the integers modp^{n}. Journal of Number Theory, 133 (12):4188–4199, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001777.

Xu:2013:DDI

[867] Zhefeng Xu. Distribution of the difference of an integer and its m-th power modn over incomplete intervals. Journal of Number Theory, 133 (12):4200–4223, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001820.

Volkov:2013:IIV

[868] V. V. Volkov and F. V. Petrov. On the interpolation of integervalued polynomials. Journal of Number Theory, 133(12):4224–4232, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001868.

Kim:2013:ADE

[869] Min-Soo Kim and Su Hu. On p-adic Diamond–Euler Log Gamma functions. Journal of Number Theory, 133(12):4233–4250, December 2013. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13001881. See corrigendum [957].

Anonymous:2014:PJa

[870] Anonymous. Pages 1–320 (January 2014). Journal of Number Theory, 134(??):1–320, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBa

[871] Anonymous. Editorial Board. Journal of Number Theory, 134(??): ifc, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300228X.

Sanna:2014:ADS

[872] Carlo Sanna. On the asymptotic density of the support of a Dirichlet convolution. Journal of Number Theory, 134(??):1–12, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13001984.

Yasaki:2014:ICC

[873] Dan Yasaki. Integral cohomology of certain Picard modular surfaces. Journal of Number Theory, 134(??):13–28, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002023.

Zhang:2014:SCF

[874] Qizhi Zhang. On the signature calculus for finite fields of order square of prime numbers. Journal of Number Theory, 134(??):29–37, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300187X.

Lee:2014:ZDR

[875] Yoonbok Lee. Zeros of the derivatives of the Riemann zeta function on Res < 1/2. Journal of Number Theory, 134(??):38–48, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001923.

Huber:2014:TTF

[876] Tim Huber. A theory of theta functions to the quintic base. Journal of Number Theory, 134(??):49–92, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13001807.

Farhat:2014:SCN

[877] Maya Farhat and Bouchaïb Sodaïgui. On Steinitz classes of nonabelian Galois extensions and p-ary cyclic Hamming codes. Journal of Number Theory, 134(??):93–108, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13001893.

Werl:2014:BWG

[878] Milan Werl. On bases of Washington’s group of circular units of some real cyclic number fields. Journal of Number Theory, 134(??):109– 129, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002035.

Sun:2014:C

[879] Li Sun and Enlin Yang. On the GL(r) × GL(r + s) × GL(s) convolution. Journal of Number Theory, 134(??):130–141, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001996.

Ibukiyama:2014:CIM

[880] Tomoyoshi Ibukiyama, Hidenori Katsurada, Cris Poor, and David S. Yuen. Congruences to Ikeda–Miyawaki lifts and triple L-values of elliptic modular forms. Journal of Number Theory, 134(??):142–180, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300200X.

Sun:2014:ACM

[881] Zhi-Wei Sun. p-adic congruences motivated by series. Journal of Number Theory, 134(??):181–196, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13001972.

Hiraoka:2014:HAH

[882] Yoshio Hiraoka and Kaoru Okada. On Hecke algebras for Hilbert modular forms. Journal of Number Theory, 134(??):197–225, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001911.

Bergdall:2014:OMF

[883] John Bergdall. Ordinary modular forms and companion points on the eigencurve. Journal of Number Theory, 134(??):226–239, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002011.

Min:2014:ZSV

[884] Jeongwon Min. Zeros and special values of Witten zeta functions and Witten L-functions. Journal of Number Theory, 134(??):240–257, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001960.

Otake:2014:ODI

[885] Shuichi Otake. Orthogonal decompositions of integral trace forms of cyclotomic fields and their canonical forms over the ring of p-adic integers. Journal of Number Theory, 134(??):258–279, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13001935.

Trudgian:2014:IUB

[886] Timothy S. Trudgian. An improved upper bound for the argument of the Riemann zeta-function on the critical line II. Journal of Number Theory, 134(??):280–292, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002047.

Sabitova:2014:CRN

[887] Maria Sabitova. Change of root numbers of elliptic curves under extension of scalars. Journal of Number Theory, 134(??):293–319, January 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300190X.

Anonymous:2014:PF

[888] Anonymous. Pages 1–390 (February 2014). Journal of Number Theory, 135(??):1–390, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBb

[889] Anonymous. Editorial Board. Journal of Number Theory, 135(??):ifc, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1300259X.

Lin:2014:EPP

[890] Bernard L. S. Lin. Elementary proofs of parity results for broken 3-diamond partitions. Journal of Number Theory, 135(??):1–7, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002254.

Zhao:2014:APO

[891] Lilu Zhao. The additive problem with one prime and two squares of primes. Journal of Number Theory, 135(??):8–27, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13002205.

Wu:2014:AEE

[892] J. Wu. The average exponent of elliptic curves modulo p. Journal of Number Theory, 135(??):28–35, February 2014. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002217.

Chen:2014:AEB

[893] Chao-Ping Chen. Asymptotic expansions for Barnes G-function. Journal of Number Theory, 135(??):36–42, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002199.

Cox:2014:GTL

[894] David A. Cox and Trevor Hyde. The Galois theory of the lemniscate. Journal of Number Theory, 135(??):43–59, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002187.

Rudnick:2014:SFV

[895] Zeév Rudnick. Square-free values of polynomials over the rational function field. Journal of Number Theory, 135(??):60–66, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13002266.

Ohshita:2014:HFI

[896] Tatsuya Ohshita. On the higher Fitting ideals of Iwasawa modules of ideal class groups over real abelian fields. Journal of Number Theory, 135(??):67–138, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300214X.

Ito:2014:APF

[897] Hiroshi Ito. An application of a product formula for the cubic Gauss sum. Journal of Number Theory, 135(??):139–150, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002175.

Goldring:2014:NPB

[898] Wushi Goldring. A new proof of Belyi’s Theorem. Journal of Number Theory, 135(??):151–154, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002382.

Girstmair:2014:JSE

[899] Kurt Girstmair. Jacobi symbols and Euler’s number e. Journal of Number Theory, 135(??):155–166, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002357.

Guo:2014:SDP

[900] Victor J. W. Guo and C. Krattenthaler. Some divisibility properties of binomial and q-binomial coefficients. Journal of Number Theory, 135(??):167–184, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002242.

Berkovich:2014:BQF

[901] Alexander Berkovich and Frank Patane. Binary quadratic forms and the Fourier coefficients of certain weight 1 eta-quotients. Journal of Number Theory, 135(??):185–220, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002369.

Chan:2014:CNT

[902] Wai Kiu Chan and Byeong-Kweon Oh. Class numbers of ternary quadratic forms. Journal of Number Theory, 135(??):221–261, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13002230.

Larcher:2014:MLB

[903] Gerhard Larcher and Friedrich Pillichshammer. Metrical lower bounds on the discrepancy of digital Kronecker-sequences. Journal of Number Theory, 135(??):262–283, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002229.

Aribam:2014:IFS

[904] Chandrakant S. Aribam. On the μ-invariant of fine Selmer groups. Journal of Number Theory, 135(??):284–300, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002151.

Hirotsune:2014:AEK

[905] Tomoki Hirotsune. p-adic Eisenstein–Kronecker series and non-critical values of p-adic Hecke L-function of an imaginary quadratic field when the conductor is divisible by p. Journal of Number Theory, 135(??):301– 333, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002394.

Scarpello:2014:LHI

[906] Giovanni Mingari Scarpello and Daniele Ritelli. Legendre hyperelliptic integrals, π new formulae and Lauricella functions through the elliptic singular moduli. Journal of Number Theory, 135(??):334–352, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002138.

Banerjee:2014:DMF

[907] Debargha Banerjee. Differential modular forms on Shimura curves over totally real fields. Journal of Number Theory, 135(??):353–373, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002412.

Huang:2014:ANT

[908] Hsiu-Lien Huang. The average number of torsion points on elliptic curves. Journal of Number Theory, 135(??):374–389, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002163.

Schaefer:2014:CCG

[909] Edward F. Schaefer. Corrigendum to “Class groups and Selmer groups” [J. Number Theory 56 (1) (1996) 79–114]. Journal of Number Theory, 135(??):390, February 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002370. See [1].

Anonymous:2014:PMa

[910] Anonymous. Pages 1–484 (March 2014). Journal of Number Theory, 136 (??):1–484, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBc

[911] Anonymous. Editorial Board. Journal of Number Theory, 136(??): ifc, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002941.

All:2014:ACU

[912] Timothy All. On the p-adic completion of the units of a real abelian number field. Journal of Number Theory, 136(??):1–21, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002400.

Byeon:2014:ECA

[913] Dongho Byeon and Nayoung Kim. Elliptic curves with all cubic twists of the same root number. Journal of Number Theory, 136(??):22– 27, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002448.

Huynh:2014:ASE

[914] Duc Van Huynh. Artin–Schreier extensions and generalized associated orders. Journal of Number Theory, 136(??):28–45, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002540.

Bremner:2014:CDE

[915] Andrew Bremner and Maciej Ulas. On certain diophantine equations of diagonal type. Journal of Number Theory, 136(??):46–64, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002527.

Wang:2014:LSI

[916] Yingnan Wang. A large sieve inequality of Elliott–Montgomery–Vaughan type for Maass forms with applications to Linnik’s problem. Journal of Number Theory, 136(??):65–86, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002473.

Vernaeve:2014:ADI

[917] Hans Vernaeve, Jasson Vindas, and Andreas Weiermann. Asymptotic distribution of integers with certain prime factorizations. Journal of Number Theory, 136(??):87–99, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002436.

He:2014:CNI

[918] Xiaoyu He. Cross number invariants of finite abelian groups. Journal of Number Theory, 136(??):100–117, March 2014. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002497.

Ledoan:2014:ZPS

[919] Andrew Ledoan, Arindam Roy, and Alexandru Zaharescu. Zeros of partial sums of the Dedekind zeta function of a cyclotomic field. Journal of Number Theory, 136(??):118–133, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002461.

Pitale:2014:BMS

[920] Ameya Pitale and Ralf Schmidt. Bessel models for GSp(4): Siegel vectors of square-free level. Journal of Number Theory, 136(??):134– 164, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002485.

Ulmer:2014:EPL

[921] Douglas Ulmer. Explicit points on the Legendre curve. Journal of Number Theory, 136(??):165–194, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002552.

Kiss:2014:ACF

[922] Sándor Z. Kiss, Eszter Rozgonyi, and Csaba Sándor. On additive complement of a finite set. Journal of Number Theory, 136(??):195– 203, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002564.

Dick:2014:DBI

[923] Josef Dick. Discrepancy bounds for infinite-dimensional order two digital sequences over F_{2}. Journal of Number Theory, 136(??):204– 232, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002576.

Keating:2014:IIE

[924] Kevin Keating. Indices of inseparability for elementary abelian pextensions. Journal of Number Theory, 136(??):233–251, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X13002503.

Zhang:2014:DE

[925] Zhongfeng Zhang. The Diophantine equation (ax^{k} −1)(by^{k} −1) = abz^{k} − 1. Journal of Number Theory, 136(??):252–260, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002825.

Kosters:2014:AIC

[926] Michiel Kosters. Anisotropy and the integral closure. Journal of Number Theory, 136(??):261–286, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002758.

Engberg:2014:DBS

[927] Zebediah Engberg. On the distribution of balanced subgroups. Journal of Number Theory, 136(??):287–319, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002783.

Lu:2014:NQS

[928] Dawei Lu. A new quicker sequence convergent to Euler’s constant. Journal of Number Theory, 136(??):320–329, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002813.

Goss:2014:CAM

[929] David Goss. A construction of v-adic modular forms. Journal of Number Theory, 136(??):330–338, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13002795.

Pletser:2014:CFD

[930] V. Pletser. On continued fraction development of quadratic irrationals having all periodic terms but last equal and associated general solutions of the Pell equation. Journal of Number Theory, 136(??):339– 353, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002771.

Paige:2014:PEC

[931] David Paige. The projective envelope of a cuspidal representation of a finite linear group. Journal of Number Theory, 136(??):354– 374, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002837.

Kirschmer:2014:OCG

[932] Markus Kirschmer. One-class genera of maximal integral quadratic forms. Journal of Number Theory, 136(??):375–393, March 2014. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0022314X1300276X.

Almousa:2014:CPF

[933] Ayah Almousa and Melanie Matchett Wood. Counting polynomials over finite fields with given root multiplicities. Journal of Number Theory, 136(??):394–402, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002849.

Zhang:2014:SPM

[934] Wenpeng Zhang and Di Han. On the sixth power mean of the twoterm exponential sums. Journal of Number Theory, 136(??):403–413, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002916.

Lu:2014:GAR

[935] Dawei Lu. A generated approximation related to Burnside’s formula. Journal of Number Theory, 136(??):414–422, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002850.

Hutz:2014:ANS

[936] Benjamin Hutz and Lucien Szpiro. Almost Newton, sometimes Lattès. Journal of Number Theory, 136(??):423–437, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002734.

Akiyama:2014:MDM

[937] Shigeki Akiyama. Mean divisibility of multinomial coefficients. Journal of Number Theory, 136(??):438–459, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002746.

Nakagawa:2014:JSD

[938] Takahiro Nakagawa. Jacobi sum, differential equation modulo prime and logarithmic growth. Journal of Number Theory, 136(??):460– 481, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002722.

Walling:2014:CFA

[939] Lynne H. Walling. Corrigendum to “A formula for the action of Hecke operators on half-integral weight Siegel modular forms and applications” [J. Number Theory 133 (5) (2013) 1608–1644]. Journal of Number Theory, 136(??):482–483, March 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1300245X. See [711].

Anonymous:2014:PAa

[940] Anonymous. Pages 1–258 (April 2014). Journal of Number Theory, 137 (??):1–258, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBd

[941] Anonymous. Editorial Board. Journal of Number Theory, 137(??): ifc, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000080.

Shparlinski:2014:DET

[942] Igor E. Shparlinski and Daniel Sutantyo. Distribution of elliptic twin primes in isogeny and isomorphism classes. Journal of Number Theory, 137(??):1–15, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002874.

Filaseta:2014:X

[943] Michael Filaseta and Samuel Gross. 49598666989151226098104244512918. Journal of Number Theory, 137(??):16–49, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002539.

Broughan:2014:CKR

[944] Kevin Broughan, Daniel Delbourgo, and Qizhi Zhou. A conjecture of De Koninck regarding particular square values of the sum of divisors function. Journal of Number Theory, 137(??):50–66, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0022314X13002801. See comment [1346] and corrigendum [2080].

Jun:2014:EGD

[945] Byungheup Jun and Jungyun Lee. Equidistribution of generalized Dedekind sums and exponential sums. Journal of Number Theory, 137 (??):67–92, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002898.

Shapira:2014:REM

[946] Uri Shapira and Zhiren Wang. Remarks on Euclidean minima. Journal of Number Theory, 137(??):93–121, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13002710.

Zhang:2014:FUR

[947] Zhe Zhang and Qin Yue. Fundamental units of real quadratic fields of odd class number. Journal of Number Theory, 137(??):122–129, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002886.

Hajdu:2014:EVF

[948] Lajos Hajdu, Ákos Pintér, Szabolcs Tengely, and Nóra Varga. Equal values of figurate numbers. Journal of Number Theory, 137(??):130– 141, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002862.

Hattori:2014:CSB

[949] Shin Hattori. Canonical subgroups via Breuil–Kisin modules for p = 2. Journal of Number Theory, 137(??):142–159, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13003053.

Chen:2014:CE

[950] Shi-Chao Chen. On the coefficients of an eigenform. Journal of Number Theory, 137(??):160–165, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X13003065.

Pal:2014:HTM

[951] Ambrus Pál. Hodge theory and the Mordell–Weil rank of elliptic curves over extensions of function fields. Journal of Number Theory, 137(??):166–178, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13003107.

Miguel:2014:MIR

[952] C. Miguel. Menon’s identity in residually finite Dedekind domains. Journal of Number Theory, 137(??):179–185, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X13003041.

Joshi:2014:AHO

[953] Kirti Joshi and Aleksandar Petrov. On the action of Hecke operators on Drinfeld modular forms. Journal of Number Theory, 137(??):186– 200, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13003119.

Nguyen:2014:NCI

[954] Hieu D. Nguyen and Long G. Cheong. New convolution identities for hypergeometric Bernoulli polynomials. Journal of Number Theory, 137(??):201–221, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13003090.

Jedrzejak:2014:VTT

[955] Tomasz Jedrzejak and Maciej Ulas. Variations on twists of tuples of hyperelliptic curves and related results. Journal of Number Theory, 137(??):222–240, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13003120.

Peruginelli:2014:IVP

[956] Giulio Peruginelli. Integral-valued polynomials over sets of algebraic integers of bounded degree. Journal of Number Theory, 137(??):241– 255, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13003089.

Kim:2014:CAD

[957] Min-Soo Kim and Su Hu. Corrigendum to “On p-adic Diamond–Euler Log Gamma functions” [J. Number Theory 133 (12) (2013) 4233– 4250]. Journal of Number Theory, 137(??):256–257, April 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002928.

See [869].

Anonymous:2014:PMb

[958] Anonymous. Pages 1–174 (May 2014). Journal of Number Theory, 138 (??):1–174, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBe

[959] Anonymous. Editorial Board. Journal of Number Theory, 138(??): ifc, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000353.

Matringe:2014:LSL

[960] Nadir Matringe. Linear and Shalika local periods for the mirabolic group, and some consequences. Journal of Number Theory, 138(??): 1–19, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000031.

Voutier:2014:MMN

[961] Paul M. Voutier. Modules with many non-associates and norm form equations with many families of solutions. Journal of Number Theory, 138(??):20–36, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13003077.

Zhang:2014:STI

[962] Bin Zhang and Chun-Gang Ji. Sums of three integral squares in biquadratic fields. Journal of Number Theory, 138(??):37–47, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000055.

Berczes:2014:SDP

[963] Attila Bérczes, Andrej Dujella, and Lajos Hajdu. Some Diophantine properties of the sequence of S-units. Journal of Number Theory, 138(??):48–68, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13002904.

Herzog:2014:CNN

[964] Marcel Herzog, Gil Kaplan, and Arieh Lev. Critical numbers of natural intervals. Journal of Number Theory, 138(??):69–83, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000171.

Schettler:2014:AAK

[965] Jordan Schettler. An alternative approach to Kida and Ferrero’s computations of Iwasawa λ-invariants. Journal of Number Theory, 138(??): 84–96, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400016X.

Rustom:2014:GGR

[966] Nadim Rustom. Generators of graded rings of modular forms. Journal of Number Theory, 138(??):97–118, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000237.

Hedayatzadeh:2014:EPD

[967] S. Mohammad Hadi Hedayatzadeh. Exterior powers of π-divisible modules over fields. Journal of Number Theory, 138(??):119–174, May 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000183.

Anonymous:2014:PJb

[968] Anonymous. Pages 1–210 (June 2014). Journal of Number Theory, 139 (??):1–210, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBf

[969] Anonymous. Editorial Board. Journal of Number Theory, 139(??): ifc, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000584.

Schauz:2014:CPM

[970] Uwe Schauz. Classification of polynomial mappings between commutative groups. Journal of Number Theory, 139(??):1–28, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000250.

Lu:2014:SAV

[971] Guangshi Lü. Sums of absolute values of cusp form coefficients and their application. Journal of Number Theory, 139(??):29–43, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000262.

Lin:2014:SRB

[972] Bernard L. S. Lin. Some results on bipartitions with 3-core. Journal of Number Theory, 139(??):44–52, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000225.

Pal:2014:CGC

[973] Ambrus Pál. On the Chow groups of certain geometrically rational 5folds. Journal of Number Theory, 139(??):53–79, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000316.

Jeong:2014:TGA

[974] Sangtae Jeong. Two generalized abc theorems for F_{q}[t]. Journal of Number Theory, 139(??):80–90, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14000328.

Cooper:2014:EEL

[975] Shaun Cooper and Dongxi Ye. Explicit evaluations of a level 13 analogue of the Rogers–Ramanujan continued fraction. Journal of Number Theory, 139(??):91–111, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000304.

Jeong:2014:SOT

[976] Sangtae Jeong. Shift operators and two applications to F_{q}[[T]]. Journal of Number Theory, 139(??):112–137, June 2014. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000195.

Kani:2014:JIP

[977] Ernst Kani. Jacobians isomorphic to a product of two elliptic curves and ternary quadratic forms. Journal of Number Theory, 139(??):138– 174, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000213.

Milinovich:2014:MPA

[978] Micah B. Milinovich and Caroline L. Turnage-Butterbaugh. Moments of products of automorphic L-functions. Journal of Number Theory, 139(??):175–204, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000274.

Hirschhorn:2014:SSP

[979] Michael D. Hirschhorn. A short and simple proof of Ramanujan’s mod11 partition congruence. Journal of Number Theory, 139(??):205– 209, June 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000298.

Anonymous:2014:PJc

[980] Anonymous. Pages 1–426 (July 2014). Journal of Number Theory, 140 (??):1–426, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBg

[981] Anonymous. Editorial Board. Journal of Number Theory, 140(??): ifc, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400078X.

Dai:2014:IEF

[982] Li-Xia Dai and Hao Pan. Inverse Erdős–Fuchs theorem for k-fold sumsets. Journal of Number Theory, 140(??):1–12, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000286.

White:2014:BCF

[983] Paul-James White. The base change L-function for modular forms and Beyond Endoscopy. Journal of Number Theory, 140(??):13– 37, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000444.

Liu:2014:GTA

[984] Chunlei Liu and Chuanze Niu. Generic twisted T -adic exponential sums of polynomials. Journal of Number Theory, 140(??):38– 59, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000420.

Alpoge:2014:SCC

[985] Levent Alpoge. Self-conjugate core partitions and modular forms. Journal of Number Theory, 140(??):60–92, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000481.

Ghioca:2014:IPD

[986] Dragos Ghioca. Integral points for Drinfeld modules. Journal of Number Theory, 140(??):93–121, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000511.

Hiary:2014:CDC

[987] Ghaith A. Hiary. Computing Dirichlet character sums to a powerfull modulus. Journal of Number Theory, 140(??):122–146, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000201.

Alkan:2014:RSN

[988] Emre Alkan. Ramanujan sums are nearly orthogonal to powers. Journal of Number Theory, 140(??):147–168, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000456.

Wakatsuki:2014:CMD

[989] Satoshi Wakatsuki. Congruences modulo 2 for dimensions of spaces of cusp forms. Journal of Number Theory, 140(??):169–180, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000419.

Barman:2014:AGF

[990] Rupam Barman and Neelam Saikia. p-Adic gamma function and the trace of Frobenius of elliptic curves. Journal of Number Theory, 140(??):181–195, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000717.

Sun:2014:SSF

[991] Qingfeng Sun. The symmetric-square L-function on the critical line. Journal of Number Theory, 140(??):196–214, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000407.

Lu:2014:GAG

[992] Dawei Lu, Lixin Song, and Congxu Ma. A generated approximation of the gamma function related to Windschitl’s formula. Journal of Number Theory, 140(??):215–225, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000687.

Yang:2014:PIC

[993] Shichun Yang and Alain Togbé. Proof of the P -integer conjecture of Pomerance. Journal of Number Theory, 140(??):226–234, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000523.

Guo:2014:RNT

[994] Xuejun Guo, Yuzhen Peng, and Hourong Qin. On the representation numbers of ternary quadratic forms and modular forms of weight 3/2. Journal of Number Theory, 140(??):235–266, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000699.

Yao:2014:NPR

[995] Olivia X. M. Yao. New parity results for broken 11-diamond partitions. Journal of Number Theory, 140(??):267–276, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000651.

Mak:2014:DNP

[996] Kit-Ho Mak and Alexandru Zaharescu. On the distribution of the number of points on a family of curves over finite fields. Journal of Number Theory, 140(??):277–298, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400050X.

Ahmadi:2014:ESP

[997] Omran Ahmadi and Igor E. Shparlinski. Exponential sums over points of elliptic curves. Journal of Number Theory, 140(??):299– 313, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000547.

Lu:2014:NAE

[998] Dawei Lu and Xiaoguang Wang. A new asymptotic expansion and some inequalities for the gamma function. Journal of Number Theory, 140(??):314–323, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000705.

Zhao:2014:DSN

[999] Jianrong Zhao, Shaofang Hong, and Wei Zhao. Divisibility by 2 of Stirling numbers of the second kind and their differences. Journal of Number Theory, 140(??):324–348, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000432.

Watt:2014:SLS

[1000] Nigel Watt. Spectral large sieve inequalities for Hecke congruence subgroups of SL(2 , Z[i]). Journal of Number Theory, 140(??):349– 424, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000560.

Soydan:2014:NDE

[1001] Gökhan Soydan and İsmail Naci Cangül. Note on “On the Diophantine equation nx^{2} + 2^{2m} = y^{n}” [Y. Wang, T. Wang, J. Number Theory 131 (8) (2011) 1486–1491]. Journal of Number Theory, 140(??): 425–426, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000249. See [327].

Anonymous:2014:PAb

[1002] Anonymous. Pages 1–374 (August 2014). Journal of Number Theory, 141 (??):1–374, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBh

[1003] Anonymous. Editorial Board. Journal of Number Theory, 141(??): ifc, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001097.

Bzdega:2014:CFI

[1004] Bartłomiej Bzdega. On a certain family of inverse ternary cyclotomic polynomials. Journal of Number Theory, 141(??):1–12, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400064X.

Baca:2014:CSD

[1005] David Baca and Bartolomé López. Coefficients of some distinguished Drinfeld modular forms. Journal of Number Theory, 141(??):13–35, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000663.

Banerjee:2014:GGG

[1006] Pradipto Banerjee. On Galois groups of generalized Laguerre polynomials whose discriminants are squares. Journal of Number Theory, 141(??):36–58, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400047X.

Gillespie:2014:PNT

[1007] Timothy L. Gillespie and Yangbo Ye. The prime number theorem and Hypothesis H with lower-order terms. Journal of Number Theory, 141(??):59–82, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000675.

Fang:2014:ACI

[1008] Jin-Hui Fang and Yong-Gao Chen. On additive complements. III. Journal of Number Theory, 141(??):83–91, August 2014. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000730.

Jeon:2014:CIS

[1009] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Cycle integrals of a sesqui-harmonic Maass form of weight zero. Journal of Number Theory, 141(??):92–108, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14000468.

Naslund:2014:MLP

[1010] Eric Naslund. The median largest prime factor. Journal of Number Theory, 141(??):109–118, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14000729.

Xia:2014:NCM

[1011] Ernest X. W. Xia. New congruences modulo powers of 2 for broken 3-diamond partitions and 7-core partitions. Journal of Number Theory, 141(??):119–135, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000535.

Demontigny:2014:GZT

[1012] Philippe Demontigny, Thao Do, Archit Kulkarni, Steven J. Miller, David Moon, and Umang Varma. Generalizing Zeckendorf’s Theorem to f-decompositions. Journal of Number Theory, 141(??):136– 158, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000754.

Bradford:2014:SPI

[1013] Jeremy Bradford. Subrings of p-power index in endomorphism rings of simple abelian varieties over finite fields. Journal of Number Theory, 141(??):159–183, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000766.

Miyazaki:2014:GCR

[1014] Takafumi Miyazaki, Pingzhi Yuan, and Danyao Wu. Generalizations of classical results on Jeśmanowicz’ conjecture concerning Pythagorean triples II. Journal of Number Theory, 141(??):184–201, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000493.

Haensch:2014:CAU

[1015] Anna Haensch. A characterization of almost universal ternary quadratic polynomials with odd prime power conductor. Journal of Number Theory, 141(??):202–213, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000997.

Goedhart:2014:DE

[1016] Eva G. Goedhart and Helen G. Grundman. On the Diophantine equation NX^{2} + 2^{L}3^{M} = Y^{N}. Journal of Number Theory, 141(??):214– 224, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000948.

Lee:2014:ZDE

[1017] Yoonbok Lee. Zero-density estimates and fractional moments of Hecke L-functions. Journal of Number Theory, 141(??):225–241, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14000924.

Schein:2014:USR

[1018] Michael M. Schein. On the universal supersingular modp representations of GL_{2}(F). Journal of Number Theory, 141(??):242–277, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000869.

Lim:2014:CJC

[1019] Jongryul Lim. A characterization of Jacobi cusp forms of certain types. Journal of Number Theory, 141(??):278–287, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001048.

Coulangeon:2014:SDH

[1020] Renaud Coulangeon and Giovanni Lazzarini. Spherical designs and heights of Euclidean lattices. Journal of Number Theory, 141(??):288– 315, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001012.

Terai:2014:JCC

[1021] Nobuhiro Terai. On Jeśmanowicz’ conjecture concerning primitive Pythagorean triples. Journal of Number Theory, 141(??):316–323, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400095X.

Greither:2014:LFS

[1022] Cornelius Greither and Radan Kučera. Linear forms on Sinnott’s module. Journal of Number Theory, 141(??):324–342, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000894.

Fiori:2014:AVL

[1023] Andrew Fiori. Arithmetic volumes for lattices over p-adic rings. Journal of Number Theory, 141(??):343–374, August 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000973.

Anonymous:2014:PS

[1024] Anonymous. Pages 1–322 (September 2014). Journal of Number Theory, 142(??):1–322, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBi

[1025] Anonymous. Editorial Board. Journal of Number Theory, 142(??):ifc, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001358.

Keating:2014:IIR

[1026] Kevin Keating. Indices of inseparability and refined ramification breaks. Journal of Number Theory, 142(??):1–17, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000936.

Vincent:2014:TNM

[1027] Christelle Vincent. On the trace and norm maps from Γ_{0}(p) to GL_{2}(A). Journal of Number Theory, 142(??):18–43, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000985.

Lu:2014:NES

[1028] Meiying Lü. A note on Engel series expansions of Laurent series and Hausdorff dimensions. Journal of Number Theory, 142(??):44–50, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001000.

Dejou:2014:BSC

[1029] Gaelle Dejou and Xavier-François Roblot. A Brumer–Stark conjecture for non-abelian Galois extensions. Journal of Number Theory, 142(??):51–88, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001061.

Yao:2014:NCM

[1030] Olivia X. M. Yao. New congruences modulo powers of 2 and 3 for 9-regular partitions. Journal of Number Theory, 142(??):89–101, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001036.

Andrade:2014:CIM

[1031] J. C. Andrade and J. P. Keating. Conjectures for the integral moments and ratios of L-functions over function fields. Journal of Number Theory, 142(??):102–148, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400105X.

Shinder:2014:LMM

[1032] Evgeny Shinder and Masha Vlasenko. Linear Mahler measures and double L-values of modular forms. Journal of Number Theory, 142 (??):149–182, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000912.

Adolphson:2014:HIS

[1033] Alan Adolphson and Steven Sperber. Hasse invariants and modp solutions of A-hypergeometric systems. Journal of Number Theory, 142 (??):183–210, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000961.

Weingartner:2014:IDD

[1034] Andreas Weingartner. Integers with dense divisors 3. Journal of Number Theory, 142(??):211–222, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001073.

Angles:2014:FIS

[1035] Bruno Anglès and Federico Pellarin. Functional identities for L-series values in positive characteristic. Journal of Number Theory, 142(??):223– 251, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001140.

Perkins:2014:EDM

[1036] Rudolph Bronson Perkins. An exact degree for multivariate special polynomials. Journal of Number Theory, 142(??):252–263, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001152.

Wu:2014:IM

[1037] Qiang Wu and Lihong Wang. On the irrationality measure of log 3. Journal of Number Theory, 142(??):264–273, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001218.

Ugolini:2014:ICM

[1038] S. Ugolini. On the iterations of certain maps X ↦ K · (X + X^{−1}) over finite fields of odd characteristic. Journal of Number Theory, 142 (??):274–297, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001188.

Merca:2014:NUB

[1039] Mircea Merca. New upper bounds for the number of partitions into a given number of parts. Journal of Number Theory, 142(??):298–304, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001292.

Pi:2014:EDC

[1040] Qinghua Pi. On effective determination of cusp forms by L-values, level aspect. Journal of Number Theory, 142(??):305–321, September 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001267.

Anonymous:2014:PO

[1041] Anonymous. Pages 1–404 (October 2014). Journal of Number Theory, 143(??):1–404, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBj

[1042] Anonymous. Editorial Board. Journal of Number Theory, 143(??): ifc, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001644.

Hajdu:2014:DE

[1043] Lajos Hajdu and István Pink. On the Diophantine equation 1+2^{a} +x^{b} = y^{n}. Journal of Number Theory, 143(??):1–13, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001243.

Dai:2014:DOV

[1044] Haobo Dai, Chunlei Liu, and Haode Yan. On the distribution of odd values of 2^{a}-regular partition functions. Journal of Number Theory, 143(??):14–23, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400122X.

Lee:2014:EDZ

[1045] Jun Ho Lee. Evaluation of the Dedekind zeta functions at s = −1 of the simplest quartic fields. Journal of Number Theory, 143(??):24– 45, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001255.

Mahnkopf:2014:TIR

[1046] J. Mahnkopf. On truncation of irreducible representations of Chevalley groups II. Journal of Number Theory, 143(??):46–61, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001280.

DeWitt:2014:MRI

[1047] Meghan De Witt. Minimal ramification and the inverse Galois problem over the rational function field F_{p}(t). Journal of Number Theory, 143(??):62–81, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001310.

Zhu:2014:GFE

[1048] Hui June Zhu. Generic A-family of exponential sums. Journal of Number Theory, 143(??):82–101, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001279.

Bamunoba:2014:SPC

[1049] Alex Samuel Bamunoba. On some properties of Carlitz cyclotomic polynomials. Journal of Number Theory, 143(??):102–108, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001322.

Chen:2014:IGN

[1050] Yong-Gao Chen and Shu-Yuan Mei. Integers with a given number of divisors. Journal of Number Theory, 143(??):109–124, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001176.

Ghitza:2014:HES

[1051] Alexandru Ghitza and Robert Sayer. Hecke eigenvalues of Siegel modular forms of “different weights”. Journal of Number Theory, 143 (??):125–141, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001231.

Young:2014:SBP

[1052] Paul Thomas Young. Symmetries of Bernoulli polynomial series and Arakawa–Kaneko zeta functions. Journal of Number Theory, 143 (??):142–161, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001334.

Vinh:2014:MSP

[1053] Le Anh Vinh. Monochromatic sum and product in Z/mZ. Journal of Number Theory, 143(??):162–169, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000900.

Saha:2014:IRF

[1054] Arnab Saha. Invariant rational forms for correspondences of curves. Journal of Number Theory, 143(??):170–184, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001401.

Camara:2014:LCS

[1055] Alberto Cámara. Locally convex structures on higher local fields. Journal of Number Theory, 143(??):185–213, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1400119X.

Pan:2014:DSA

[1056] Hao Pan. On divisibility of sums of Apéry polynomials. Journal of Number Theory, 143(??):214–223, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001206.

Shparlinski:2014:DEF

[1057] Igor E. Shparlinski. Distribution of exponential functions modulo a prime power. Journal of Number Theory, 143(??):224–231, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001498.

Laurincikas:2014:DUT

[1058] Antanas Laurinčikas. A discrete universality theorem for the Hurwitz zeta-function. Journal of Number Theory, 143(??):232–247, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001528.

Dummigan:2014:RSC

[1059] Neil Dummigan and Daniel Fretwell. Ramanujan-style congruences of local origin. Journal of Number Theory, 143(??):248–261, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001474.

Kucuksakalli:2014:VSL

[1060] Ömer Küçüksakallı. Value sets of Lattès maps over finite fields. Journal of Number Theory, 143(??):262–278, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1400153X.

Heidaryan:2014:BPF

[1061] Bahar Heidaryan and Ali Rajaei. Biquadratic Pólya fields with only one quadratic Pólya subfield. Journal of Number Theory, 143(??):279– 285, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001425.

Zan:2014:PPB

[1062] Haixia Zan and Wei Cao. Powers of polynomials and bounds of value sets. Journal of Number Theory, 143(??):286–292, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001486.

Sun:2014:GLP

[1063] Zhi-Hong Sun. Generalized Legendre polynomials and related supercongruences. Journal of Number Theory, 143(??):293–319, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001516.

Wood:2014:ZWT

[1064] Rachael Wood and Matthew P. Young. Zeros of the weight two Eisenstein series. Journal of Number Theory, 143(??):320–333, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001462.

Mistri:2014:GSS

[1065] Raj Kumar Mistri and Ram Krishna Pandey. A generalization of sumsets of set of integers. Journal of Number Theory, 143(??):334–356, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001449.

Biswas:2014:LVH

[1066] Arunabha Biswas and Chris Monico. Limiting value of higher Mahler measure. Journal of Number Theory, 143(??):357–362, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001541.

Sargent:2014:DVL

[1067] Oliver Sargent. Density of values of linear maps on quadratic surfaces. Journal of Number Theory, 143(??):363–384, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001619.

Flammang:2014:ALP

[1068] V. Flammang. On the absolute length of polynomials having all zeros in a sector. Journal of Number Theory, 143(??):385–401, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001413.

Stirpe:2014:CLA

[1069] Claudio Stirpe. A counterexample to ‘Algebraic function fields with small class number’. Journal of Number Theory, 143(??):402–404, October 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001024.

Anonymous:2014:PN

[1070] Anonymous. Pages 1–474 (November 2014). Journal of Number Theory, 144(??):1–474, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBk

[1071] Anonymous. Editorial Board. Journal of Number Theory, 144(??):ifc, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002029.

Cho:2014:SRH

[1072] Bumkyu Cho and Ho Park. On some results of Hurwitz and Deutsch about certain quadratic forms. Journal of Number Theory, 144(??):1– 14, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001553.

Wang:2014:CCM

[1073] Liuquan Wang and Tianxin Cai. A curious congruence modulo prime powers. Journal of Number Theory, 144(??):15–24, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001437.

Keil:2014:ENS

[1074] Stefan Keil. Examples of non-simple abelian surfaces over the rationals with non-square order Tate–Shafarevich group. Journal of Number Theory, 144(??):25–69, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001577.

Andrade:2014:NCV

[1075] Julio Andrade, Alan Chang, and Steven J. Miller. Newman’s conjecture in various settings. Journal of Number Theory, 144(??):70–91, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001620.

Cilleruelo:2014:LCM

[1076] Javier Cilleruelo, Juanjo Rué, Paulius Šarka, and Ana Zumalacárregui. The least common multiple of random sets of positive integers. Journal of Number Theory, 144(??):92–104, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001504.

Cheng:2014:AEG

[1077] Jun-Xiang Cheng and Chao-Ping Chen. Asymptotic expansions of the Glaisher–Kinkelin and Choi–Srivastava constants. Journal of Number Theory, 144(??):105–110, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001565.

DHoker:2014:ZKI

[1078] Eric D’Hoker and Michael B. Green. Zhang–Kawazumi invariants and superstring amplitudes. Journal of Number Theory, 144(??):111–150, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001590.

Brumbaugh:2014:SES

[1079] J. L. Brumbaugh, Madeleine Bulkow, Patrick S. Fleming, Luis Alberto Garcia German, Stephan Ramon Garcia, Gizem Karaali, Matt Michal, Andrew P. Turner, and Hong Suh. Supercharacters, exponential sums, and the uncertainty principle. Journal of Number Theory, 144(??):151–175, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001607.

Chen:2014:PI

[1080] Yong-Gao Chen and Ming-Liang Gong. On the products (1^{ℓ} + 1)(2^{ℓ} + 1) ⋯ (n^{ℓ} + 1), II. Journal of Number Theory, 144(??):176–187, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001735.

Kim:2014:TVA

[1081] Byoung Du Kim. Two-variable p-adic L-functions of modular forms for non-ordinary primes. Journal of Number Theory, 144(??):188–218, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001747.

Lei:2014:ASP

[1082] Peng Lei, Li Guo, and Biao Ma. Applications of shuffle product to restricted decomposition formulas for multiple zeta values. Journal of Number Theory, 144(??):219–233, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001589.

Hua:2014:NMV

[1083] Chen Hua, Ai Xiaochuan, and Cai Guangxing. A note on mean value of mixed exponential sums. Journal of Number Theory, 144 (??):234–243, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001851.

Qi:2014:IRC

[1084] Feng Qi. An integral representation, complete monotonicity, and inequalities of Cauchy numbers of the second kind. Journal of Number Theory, 144(??):244–255, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001838.

Earnest:2014:SRQ

[1085] A. G. Earnest, Ji Young Kim, and N. D. Meyer. Strictly regular quaternary quadratic forms and lattices. Journal of Number Theory, 144 (??):256–266, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001772.

Kim:2014:GDS

[1086] Min-Soo Kim and Jin-Woo Son. On generalized Dedekind sums involving quasi-periodic Euler functions. Journal of Number Theory, 144 (??):267–280, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001966.

Sano:2014:GDC

[1087] Takamichi Sano. A generalization of Darmon’s conjecture for Euler systems for general p-adic representations. Journal of Number Theory, 144(??):281–324, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001814.

Dubickas:2014:FPP

[1088] Artūras Dubickas and Jonas Jankauskas. On the fractional parts of powers of Pisot numbers of length at most 4. Journal of Number Theory, 144(??):325–339, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001760.

Lu:2014:SNQ

[1089] Dawei Lu and Cristinel Mortici. Some new quicker approximations of Glaisher–Kinkelin’s and Bendersky–Adamchik’s constants. Journal of Number Theory, 144(??):340–352, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001899.

Nomura:2014:ASG

[1090] Jiro Nomura. Annihilation of Selmer groups for monomial CMextensions. Journal of Number Theory, 144(??):353–378, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001826.

Chabert:2014:MPA

[1091] Jean-Luc Chabert, Sabine Evrard, and Youssef Fares. Mahler’s polynomial approximation for odd and even p-adic functions. Journal of Number Theory, 144(??):379–407, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001954.

Lorscheid:2014:BTA

[1092] Oliver Lorscheid. Blueprints — towards absolute arithmetic? Journal of Number Theory, 144(??):408–421, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14001450.

Haessig:2014:FAF

[1093] C. Douglas Haessig and Steven Sperber. L-functions associated with families of toric exponential sums. Journal of Number Theory, 144 (??):422–473, November 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001309.

Anonymous:2014:PD

[1094] Anonymous. Pages 1–634 (December 2014). Journal of Number Theory, 145(??):1–634, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2014:EBl

[1095] Anonymous. Editorial Board. Journal of Number Theory, 145(??):ifc, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002315.

Chatterjee:2014:NVD

[1096] Tapas Chatterjee and M. Ram Murty. Non-vanishing of Dirichlet series with periodic coefficients. Journal of Number Theory, 145(??):1–21, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400184X.

Chow:2014:ATS

[1097] Sam Chow. Averaging on thin sets of diagonal forms. Journal of Number Theory, 145(??):22–50, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001887.

Tafazolian:2014:CFF

[1098] Saeed Tafazolian and Fernando Torres. On the curve y^{n} = x^{m} + x over finite fields. Journal of Number Theory, 145(??):51–66, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001930.

Movasati:2014:AFT

[1099] Hossein Movasati and Khosro M. Shokri. Automorphic forms for triangle groups: Integrality properties. Journal of Number Theory, 145 (??):67–78, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001991.

Sirolli:2014:PSM

[1100] Nicolás Sirolli. Preimages for the Shimura map on Hilbert modular forms. Journal of Number Theory, 145(??):79–98, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001802.

Bunder:2014:URI

[1101] Martin W. Bunder. Unique representations of integers using increasing sequences. Journal of Number Theory, 145(??):99–108, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001905.

Mu:2014:WGT

[1102] Quanwu Mu. A Waring–Goldbach type problem for mixed powers. Journal of Number Theory, 145(??):109–125, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14001875.

Positselski:2014:GCN

[1103] Leonid Positselski. Galois cohomology of a number field is Koszul. Journal of Number Theory, 145(??):126–152, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1400198X.

Seo:2014:FLE

[1104] Soogil Seo. On finite layers of Z_{l}-extensions and K_{2}. Journal of Number Theory, 145(??):153–180, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001796.

Quan:2014:CMA

[1105] Nguyen Ngoc Dong Quan. Carlitz module analogues of Mersenne primes, Wieferich primes, and certain prime elements in cyclotomic function fields. Journal of Number Theory, 145(??):181–193, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001863.

Laubie:2014:ARS

[1106] François Laubie. Application de réciprocité de Shafarevich et groupe de Nottingham. (French) [Shafarevich reciprocity application and Nottingham group]. Journal of Number Theory, 145(??):194–209, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002005.

Leriche:2014:AEN

[1107] Amandine Leriche. About the embedding of a number field in a Pólya field. Journal of Number Theory, 145(??):210–229, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001759.

Meher:2014:SCF

[1108] Jaban Meher and Naomi Tanabe. Sign changes of Fourier coefficients of Hilbert modular forms. Journal of Number Theory, 145(??):230–244, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001784.

Wong:2014:AOS

[1109] Siman Wong. Arithmetic of octahedral sextics. Journal of Number Theory, 145(??):245–272, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14001978.

Baran:2014:EIB

[1110] Burcu Baran. An exceptional isomorphism between modular curves of level 13. Journal of Number Theory, 145(??):273–300, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14001917.

Guo:2014:SAS

[1111] Victor J. W. Guo and Jiang Zeng. Some q-analogues of supercongruences of Rodriguez-Villegas. Journal of Number Theory, 145(??):301– 316, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002078.

Lu:2014:GAF

[1112] Dawei Lu, Jinghai Feng, and Congxu Ma. A general asymptotic formula of the gamma function based on the Burnside’s formula. Journal of Number Theory, 145(??):317–328, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002224.

Gun:2014:TGE

[1113] Sanoli Gun, Ekata Saha, and Sneh Bala Sinha. Transcendence of generalized Euler–Lehmer constants. Journal of Number Theory, 145 (??):329–339, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002157.

Entin:2014:ASF

[1114] Alexei Entin. Artin–Schreier L-functions and random unitary matrices. Journal of Number Theory, 145(??):340–351, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1400225X.

Chatterjee:2014:ZGH

[1115] T. Chatterjee and S. Gun. On the zeros of generalized Hurwitz zeta functions. Journal of Number Theory, 145(??):352–361, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002091.

Dauguet:2014:SDA

[1116] Simon Dauguet and Wadim Zudilin. On simultaneous diophantine approximations to ζ(2) and ζ(3). Journal of Number Theory, 145 (??):362–387, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002121.

Zhong:2014:GRP

[1117] Ting Zhong and Liang Tang. The growth rate of the partial quotients in a class of continued fractions with parameters. Journal of Number Theory, 145(??):388–401, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002170.

Jedrzejak:2014:TJS

[1118] Tomasz Jedrzejak. On the torsion of the Jacobians of superelliptic curves y^{q} = x^{p} + a. Journal of Number Theory, 145(??):402–425, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002194.

Stout:2014:EBH

[1119] Brian Stout and Adam Towsley. Endomorphisms of bounded height and resultant. Journal of Number Theory, 145(??):426–432, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400211X.

Ahn:2014:SEZ

[1120] Jeoung-Hwan Ahn and Soun-Hi Kwon. Some explicit zero-free regions for Hecke L-functions. Journal of Number Theory, 145(??):433–473, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002133.

Le:2014:QST

[1121] Giang Le. On the quantitative subspace theorem. Journal of Number Theory, 145(??):474–495, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002145.

Espinoza:2014:SSC

[1122] Milton Espinoza. Signed Shintani cones for number fields with one complex place. Journal of Number Theory, 145(??):496–539, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400208X.

Garaev:2014:MDA

[1123] M. Z. Garaev and S. V. Konyagin. Multiplicative decomposition of arithmetic progressions in prime fields. Journal of Number Theory, 145(??):540–553, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002169.

Adiga:2014:CFO

[1124] Chandrashekar Adiga, Zhi-Guo Liu, and A. Vanitha. On a continued fraction of order twelve and new Eisenstein series identities. Journal of Number Theory, 145(??):554–571, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002248.

Leal-Ruperto:2014:MGI

[1125] José Luis Leal-Ruperto. On the magnitude of the Gaussian integer solutions of the Legendre equation. Journal of Number Theory, 145 (??):572–578, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002297.

Rivoal:2014:HPA

[1126] T. Rivoal and J. Roques. Hadamard products of algebraic functions. Journal of Number Theory, 145(??):579–603, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002212.

Filipin:2014:EDP

[1127] Alan Filipin, Yasutsugu Fujita, and Alain Togbé. The extendibility of Diophantine pairs II: Examples. Journal of Number Theory, 145 (??):604–631, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002261.

Borisov:2014:CQI

[1128] Alexander Borisov. Corrigendum to “Quantum integers and cyclotomy” [J. Number Theory 109 (1) (2004) 120–135]. Journal of Number Theory, 145(??):632–634, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001942. See [5].

Anonymous:2015:EBa

[1129] Anonymous. Editorial Board. Journal of Number Theory, 146(??): ifc, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002662.

Cogdell:2015:SR

[1130] James Cogdell, Hervé Jacquet, Dihua Jiang, and Stephen S. Kudla. Steve Rallis (1942–2012). Journal of Number Theory, 146(??):1–3, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400239X.

Ash:2015:HGR

[1131] Avner Ash, Paul E. Gunnells, and Mark McConnell. mod2 homology for GL(4) and Galois representations. Journal of Number Theory, 146(??):4–22, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14000043.

Baruch:2015:KFA

[1132] Ehud Moshe Baruch and Orr Beit-Aharon. A kernel formula for the action of the Weyl element in the Kirillov model of SL(2 , C). Journal of Number Theory, 146(??):23–40, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1400033X.

Brubaker:2015:WFD

[1133] Ben Brubaker, Daniel Bump, and Anthony Licata. Whittaker functions and Demazure operators. Journal of Number Theory, 146(??):41– 68, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000067.

Chan:2015:LLC

[1134] Ping-Shun Chan and Wee Teck Gan. The local Langlands conjecture for GSp(4) III: Stability and twisted endoscopy. Journal of Number Theory, 146(??):69–133, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001959.

Friedberg:2015:MTF

[1135] Solomon Friedberg and David Ginzburg. Metaplectic theta functions and global integrals. Journal of Number Theory, 146(??):134–149, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14001164.

Furusawa:2015:LRN

[1136] Masaaki Furusawa and Kimball Martin. Local root numbers, Bessel models, and a conjecture of Guo and Jacquet. Journal of Number Theory, 146(??):150–170, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001765.

Garrett:2015:HDI

[1137] Paul Garrett and Bernhard Heim. Hecke duality of Ikeda lifts. Journal of Number Theory, 146(??):171–186, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000870.

Green:2015:SRS

[1138] Michael B. Green, Stephen D. Miller, and Pierre Vanhove. Small representations, string instantons, and Fourier modes of Eisenstein series. Journal of Number Theory, 146(??):187–309, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002424.

Hanzer:2015:DES

[1139] Marcela Hanzer and Goran Muić. Degenerate Eisenstein series for Sp(4 ). Journal of Number Theory, 146(??):310–342, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X13002692.

Jiang:2015:SUO

[1140] Dihua Jiang and Baiying Liu. On special unipotent orbits and Fourier coefficients for automorphic forms on symplectic groups. Journal of Number Theory, 146(??):343–389, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14000857.

Kaplan:2015:CRR

[1141] Eyal Kaplan. Complementary results on the Rankin–Selberg gamma factors of classical groups. Journal of Number Theory, 146(??):390– 447, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400002X.

Lapid:2015:CWF

[1142] Erez Lapid and Zhengyu Mao. A conjecture on Whittaker–Fourier coefficients of cusp forms. Journal of Number Theory, 146(??):448–505, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002515.

Murnaghan:2015:DDZ

[1143] Fiona Murnaghan. Distinction of depth-zero representations. Journal of Number Theory, 146(??):506–518, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14000742.

Ramakrishnan:2015:MTT

[1144] Dinakar Ramakrishnan. A mild Tchebotarev theorem for GL(n). Journal of Number Theory, 146(??):519–533, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002406.

Savin:2015:MHO

[1145] Gordan Savin and Michael Woodbury. Matching of Hecke operators for exceptional dual pair correspondences. Journal of Number Theory, 146(??):534–556, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X13001789.

Soudry:2015:LDU

[1146] David Soudry and Yaacov Tanay. On local descent for unitary groups. Journal of Number Theory, 146(??):557–626, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14000882.

Waldspurger:2015:RIO

[1147] J.-L. Waldspurger. Une remarque sur les intégrales orbitales de la fonction caractéristique d’un réseau hyperspécial de gl(n). (French) [A remark on the orbital integrals of the characteristic function of a hyperspecial network of gl(n)]. Journal of Number Theory, 146(??):627– 632, January 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13001534.

Anonymous:2015:PF

[1148] Anonymous. Pages 1–882 (February 2015). Journal of Number Theory, 147(??):1–882, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBb

[1149] Anonymous. Editorial Board. Journal of Number Theory, 147(??):ifc, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003308.

Rosales:2015:SPI

[1150] J. C. Rosales, M. B. Branco, and D. Torrão. Sets of positive integers closed under product and the number of decimal digits. Journal of Number Theory, 147(??):1–13, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002200.

Hou:2015:DTP

[1151] Xiang-Dong Hou and Stephen D. Lappano. Determination of a type of permutation binomials over finite fields. Journal of Number Theory, 147(??):14–23, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002273.

Bengoechea:2015:QPC

[1152] Paloma Bengoechea. From quadratic polynomials and continued fractions to modular forms. Journal of Number Theory, 147(??):24–43, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002182.

Do:2015:SDC

[1153] Thao Do, Archit Kulkarni, Steven J. Miller, David Moon, and Jake Wellens. Sums and differences of correlated random sets. Journal of Number Theory, 147(??):44–68, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002285.

Lu:2015:SNCa

[1154] Dawei Lu, Lixin Song, and Yang Yu. Some new continued fraction approximation of Euler’s constant. Journal of Number Theory, 147(??):69–80, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002418.

McIntosh:2015:CN

[1155] Richard J. McIntosh and Mitra Dipra. Carmichael numbers with p+1|n+ 1. Journal of Number Theory, 147(??):81–91, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002108.

Bagdasaryan:2015:ANG

[1156] Armen Bagdasaryan, Serkan Araci, Mehmet Açikgöz, and H. M. Srivastava. Analogues of Newton–Girard power-sum formulas for entire and meromorphic functions with applications to the Riemann zeta function. Journal of Number Theory, 147(??):92–102, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002455.

Ai:2015:CSS

[1157] Xiaochuan Ai, Jianhua Chen, Silan Zhang, and Hao Hu. Complete solutions of the simultaneous Pell equations x^{2} − 24y^{2} = 1 and y^{2} − pz^{2} = 1. Journal of Number Theory, 147(??):103–108, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002480.

Palmer:2015:DAC

[1158] Matthew Palmer. Diagonal approximation in completions of Q. Journal of Number Theory, 147(??):109–132, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002431.

He:2015:DPC

[1159] Bing He. On the divisibility properties of certain binomial sums. Journal of Number Theory, 147(??):133–140, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002479.

Batzaya:2015:SAP

[1160] Gantsooj Batzaya. On simultaneous approximation to powers of a real number by rational numbers. Journal of Number Theory, 147 (??):141–155, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002467.

Ratazzi:2015:CIV

[1161] Nicolas Ratazzi. Classe d’isogénie de variétés abéliennes pleinement de type GSp. (French) [Class of isogeny of fully abelian varieties of type GSp]. Journal of Number Theory, 147(??):156–171, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400256X.

Liu:2015:MVD

[1162] Huaning Liu. On the mean values of Dirichlet L-functions. Journal of Number Theory, 147(??):172–183, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002443.

Tsushima:2015:SRL

[1163] Takahiro Tsushima. On the stable reduction of the Lubin–Tate curve of level two in the equal characteristic case. Journal of Number Theory, 147(??):184–210, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002558.

Dixit:2015:SRF

[1164] Atul Dixit and Victor H. Moll. Self-reciprocal functions, powers of the Riemann zeta function and modular-type transformations. Journal of Number Theory, 147(??):211–249, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002236.

Ganguli:2015:RGR

[1165] Abhik Ganguli and Eknath Ghate. Reductions of Galois representations via the mod p Local Langlands Correspondence. Journal of Number Theory, 147(??):250–286, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002509.

Zhang:2015:PCA

[1166] Yong Zhang and Tianxin Cai. On products of consecutive arithmetic progressions. Journal of Number Theory, 147(??):287–299, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400242X.

Fang:2015:SVA

[1167] Jiangxue Fang. Special L-values of abelian t-modules. Journal of Number Theory, 147(??):300–325, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002510.

Sun:2015:SM

[1168] Zhi-Wei Sun. Supercongruences motivated by e. Journal of Number Theory, 147(??):326–341, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002522.

Jeon:2015:FEC

[1169] Daeyeol Jeon, Chang Heon Kim, and Yoonjin Lee. Families of elliptic curves with prescribed torsion subgroups over dihedral quartic fields. Journal of Number Theory, 147(??):342–363, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002534.

Bober:2015:HLZ

[1170] Jonathan Bober, J. Brian Conrey, David W. Farmer, Akio Fujii, Sally Koutsoliotas, Stefan Lemurell, Michael Rubinstein, and Hiroyuki Yoshida. The highest lowest zero of general L-functions. Journal of Number Theory, 147(??):364–373, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002625.

Han:2015:HDC

[1171] Guo-Niu Han. Hankel determinant calculus for the Thue–Morse and related sequences. Journal of Number Theory, 147(??):374–395, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002613.

Gon:2015:DST

[1172] Yasuro Gon. Differences of the Selberg trace formula and Selberg type zeta functions for Hilbert modular surfaces. Journal of Number Theory, 147(??):396–453, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002583.

Degtyarev:2015:LGP

[1173] Alex Degtyarev. Lines generate the Picard groups of certain Fermat surfaces. Journal of Number Theory, 147(??):454–477, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002595.

Childress:2015:IAL

[1174] Nancy Childress and Scott Zinzer. On λ-invariants associated to linear extensions of the product ordering on N^{d}. Journal of Number Theory, 147(??):478–489, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14001929.

Dil:2015:ESH

[1175] Ayhan Dil and Khristo N. Boyadzhiev. Euler sums of hyperharmonic numbers. Journal of Number Theory, 147(??):490–498, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002571.

Mortici:2015:GRD

[1176] Cristinel Mortici and Michael Th. Rassias. On the growth rate of divergent series. Journal of Number Theory, 147(??):499–507, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002649.

Melfi:2015:CIP

[1177] Giuseppe Melfi. On the conditional infiniteness of primitive weird numbers. Journal of Number Theory, 147(??):508–514, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002637.

Baker:2015:SBW

[1178] Simon Baker. On small bases which admit countably many expansions. Journal of Number Theory, 147(??):515–532, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002753.

Sun:2015:BMS

[1179] Chia-Liang Sun. The Brauer–Manin–Scharaschkin obstruction for subvarieties of a semi-abelian variety and its dynamical analog. Journal of Number Theory, 147(??):533–548, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002546.

Guardia:2015:HNP

[1180] Jordi Guàrdia, Jesús Montes, and Enric Nart. Higher Newton polygons and integral bases. Journal of Number Theory, 147(??):549–589, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002777.

Guo:2015:PCM

[1181] Victor J. W. Guo. Proof of a conjecture of Mircea Merca. Journal of Number Theory, 147(??):590–593, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002807.

Hu:2015:PFP

[1182] Liqun Hu and Huafeng Liu. On pairs of four prime squares and powers of two. Journal of Number Theory, 147(??):594–604, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002790.

Hu:2015:AAW

[1183] Su Hu and Min-Soo Kim. On p-adic analogue of Weil’s elliptic functions according to Eisenstein. Journal of Number Theory, 147(??):605– 619, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1400287X.

Liu:2015:SNA

[1184] Sheng-Chi Liu. Simultaneous nonvanishing of automorphic L-functions. Journal of Number Theory, 147(??):620–632, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002789.

Chida:2015:TTI

[1185] Masataka Chida, Chung Pang Mok, and Jeehoon Park. On Teitelbaum type L-invariants of Hilbert modular forms attached to definite quaternions. Journal of Number Theory, 147(??):633–665, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002856.

Lin:2015:TKS

[1186] Yi-Hsuan Lin and Fang-Ting Tu. Twisted Kloosterman sums. Journal of Number Theory, 147(??):666–690, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14002765.

Williams:2015:HRT

[1187] Kenneth S. Williams. Historical remark on a theorem of Zhang and Yue. Journal of Number Theory, 147(??):691–693, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002844.

Sun:2015:NRL

[1188] Zhi-Hong Sun. New reciprocity laws for octic residues and nonresidues. Journal of Number Theory, 147(??):694–707, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002881.

Krattenthaler:2015:GAN

[1189] C. Krattenthaler and T. W. Müller. Generalised Apéry numbers modulo 9. Journal of Number Theory, 147(??):708–720, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002893.

Yamashita:2015:SAI

[1190] Go Yamashita and Seidai Yasuda. On some applications of integral padic Hodge theory to Galois representations. Journal of Number Theory, 147(??):721–748, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002716.

Eie:2015:WSF

[1191] Minking Eie and Fu-Yao Yang. Weighted sum formulas from shuffle products of multiples of Riemann zeta values. Journal of Number Theory, 147(??):749–765, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400290X.

Bush:2015:CFT

[1192] Michael R. Bush and Daniel C. Mayer. 3-class field towers of exact length 3. Journal of Number Theory, 147(??):766–777, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002832.

VanGorder:2015:DRZ

[1193] Robert A. Van Gorder. Does the Riemann zeta function satisfy a differential equation? Journal of Number Theory, 147(??):778–788, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002868.

Xiong:2015:DZZ

[1194] Maosheng Xiong. Distribution of zeta zeroes for abelian covers of algebraic curves over a finite field. Journal of Number Theory, 147 (??):789–823, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002819.

Kwon:2015:OPF

[1195] DoYong Kwon. A one-parameter family of Dirichlet series whose coefficients are Sturmian words. Journal of Number Theory, 147(??):824– 835, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002947.

Sanna:2015:SDF

[1196] Carlo Sanna. On the sum of digits of the factorial. Journal of Number Theory, 147(??):836–841, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14002984.

Platt:2015:IEB

[1197] D. J. Platt and T. S. Trudgian. An improved explicit bound on |ζ(\frac{1}{2} +it)|. Journal of Number Theory, 147(??):842–851, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14002959.

Xia:2015:PCB

[1198] Ernest X. W. Xia. Proof of a conjecture of Baruah and Sarmah on generalized Frobenius partitions with 6 colors. Journal of Number Theory, 147(??):852–860, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002960.

Zhang:2015:HMV

[1199] Wenpeng Zhang and Di Han. A hybrid mean value of Dedekind sums and Kloosterman sums. Journal of Number Theory, 147(??):861–870, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000559.

Kim:2015:UCA

[1200] Dae San Kim and Taekyun Kim. Umbral calculus associated with Bernoulli polynomials. Journal of Number Theory, 147(??):871–882, February 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X13002709.

Anonymous:2015:PMa

[1201] Anonymous. Pages 1–592 (March 2015). Journal of Number Theory, 148 (??):1–592, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBc

[1202] Anonymous. Editorial Board. Journal of Number Theory, 148(??): ifc, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003424.

Berkovich:2015:GTS

[1203] Alexander Berkovich. On the Gauss EΥPHKA theorem and some allied inequalities. Journal of Number Theory, 148(??):1–18, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14003060.

Verma:2015:FDV

[1204] Manoj Verma. On a form of degree d in 2d + 1 variables (d ≥ 4). Journal of Number Theory, 148(??):19–32, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003084.

Lee:2015:DOG

[1205] Jun Ho Lee and Stéphane R. Louboutin. Determination of the orders generated by a cyclic cubic unit that are Galois invariant. Journal of Number Theory, 148(??):33–39, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003266.

Conley:2015:CCJ

[1206] Charles H. Conley and Rabin Dahal. Centers and characters of Jacobi group-invariant differential operator algebras. Journal of Number Theory, 148(??):40–61, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003138.

Chen:2015:RTC

[1207] William Y. C. Chen, Lisa H. Sun, Rong-Hua Wang, and Li Zhang. Ramanujan-type congruences for overpartitions modulo 5. Journal of Number Theory, 148(??):62–72, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003126.

Lengyel:2015:APS

[1208] Tamás Lengyel. On p-adic properties of the Stirling numbers of the first kind. Journal of Number Theory, 148(??):73–94, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003102.

Allouche:2015:PIP

[1209] J.-P. Allouche. Paperfolding infinite products and the gamma function. Journal of Number Theory, 148(??):95–111, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003072.

Berndt:2015:FIR

[1210] Bruce C. Berndt, Soon-Yi Kang, and Jaebum Sohn. Finite and infinite Rogers–Ramanujan continued fractions in Ramanujan’s lost notebook. Journal of Number Theory, 148(??):112–120, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1400314X.

Perucca:2015:ORA

[1211] Antonella Perucca. The order of the reductions of an algebraic integer. Journal of Number Theory, 148(??):121–136, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003175.

Gusic:2015:ISH

[1212] Ivica Gusić and Petra Tadić. Injectivity of the specialization homomorphism of elliptic curves. Journal of Number Theory, 148(??):137– 152, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003187.

Chen:2015:SDZ

[1213] Huei-Jeng Chen. On shuffle of double zeta values over F_{q}[t]. Journal of Number Theory, 148(??):153–163, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003114.

Gritsenko:2015:BPE

[1214] Valery Gritsenko, Cris Poor, and David S. Yuen. Borcherds products everywhere. Journal of Number Theory, 148(??):164–195, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002935.

Jennings-Shaffer:2015:ASC

[1215] C. Jennings-Shaffer. Another SPT crank for the number of smallest parts in overpartitions with even smallest part. Journal of Number Theory, 148(??):196–203, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400300X.

Huang:2015:ESP

[1216] Bingrong Huang and Zhiwei Wang. Exponential sums over primes in short intervals. Journal of Number Theory, 148(??):204–219, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14002996.

Jiang:2015:UES

[1217] Yujiao Jiang and Guangshi Lü. Uniform estimates for sums of coefficients of symmetric square L-function. Journal of Number Theory, 148(??):220–234, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003035.

Mistri:2015:SSD

[1218] Raj Kumar Mistri, Ram Krishna Pandey, and Om Prakash. Subsequence sums: Direct and inverse problems. Journal of Number Theory, 148(??):235–256, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003059. See corrigendum [1303].

Ji:2015:BQF

[1219] Yun-Seong Ji, Myung-Hwan Kim, and Byeong-Kweon Oh. Binary quadratic forms represented by a sum of nonzero squares. Journal of Number Theory, 148(??):257–271, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003047.

Munshi:2015:DHM

[1220] Ritabrata Munshi and Jyoti Sengupta. Determination of GL(3 ) Hecke– Maass forms from twisted central values. Journal of Number Theory, 148(??):272–287, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003254.

Lagemann:2015:DAS

[1221] Thorsten Lagemann. Distribution of Artin–Schreier–Witt extensions. Journal of Number Theory, 148(??):288–310, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003217.

Dombek:2015:DUG

[1222] Daniel Dombek, Zuzana Masáková, and Volker Ziegler. On distinct unit generated fields that are totally complex. Journal of Number Theory, 148(??):311–327, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003242.

Bloch:2015:EDS

[1223] Spencer Bloch and Pierre Vanhove. The elliptic dilogarithm for the sunset graph. Journal of Number Theory, 148(??):328–364, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14003278.

Jakubec:2015:CNR

[1224] S. Jakubec, M. Pasteka, and A. Schinzel. Class number of real Abelian fields. Journal of Number Theory, 148(??):365–371, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003229.

Hindes:2015:GUQ

[1225] Wade Hindes. Galois uniformity in quadratic dynamics over k(t). Journal of Number Theory, 148(??):372–383, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003394.

Chamberland:2015:ASP

[1226] Marc Chamberland. Averaging structure in the 3x + 1 problem. Journal of Number Theory, 148(??):384–397, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003199.

Grob:2015:STF

[1227] D. Grob and R. S. Kraußhar. A Selberg trace formula for hypercomplex analytic cusp forms. Journal of Number Theory, 148(??):398– 428, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002972.

Liu:2015:NCA

[1228] Jianya Liu and Jie Wu. The number of coefficients of automorphic L-functions for GL_{m} of same signs. Journal of Number Theory, 148(??):429–450, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003151.

Miezaki:2015:CFC

[1229] Tsuyoshi Miezaki and Matthias Waldherr. Congruences for the Fourier coefficients of the Mathieu mock theta function. Journal of Number Theory, 148(??):451–462, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14003205.

Charlton:2015:R

[1230] Steven Charlton. ζ({{2}^{m}, 1, {2}^{m}, 3}^{n}, {2}^{m})/π^{4n+2m(2n+1)} is rational. Journal of Number Theory, 148(??):463–477, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003230.

Brown:2015:SVM

[1231] Francis Brown and Oliver Schnetz. Single-valued multiple polylogarithms and a proof of the zig–zag conjecture. Journal of Number Theory, 148(??):478–506, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003023.

Coffey:2015:MTO

[1232] Mark W. Coffey and Matthew C. Lettington. Mellin transforms with only critical zeros: Legendre functions. Journal of Number Theory, 148(??):507–536, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14002601.

Blagouchine:2015:TCF

[1233] Iaroslav V. Blagouchine. A theorem for the closed-form evaluation of the first generalized Stieltjes constant at rational arguments and some related summations. Journal of Number Theory, 148(??):537– 592, March 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002820. See erratum [1291].

Anonymous:2015:PAa

[1234] Anonymous. Pages 1–434 (April 2015). Journal of Number Theory, 149 (??):1–434, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBd

[1235] Anonymous. Editorial Board. Journal of Number Theory, 149(??): ifc, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003722.

Ayad:2015:DAN

[1236] Mohamed Ayad, Abdelmejid Bayad, and Omar Kihel. Denominators of algebraic numbers in a number field. Journal of Number Theory, 149(??):1–14, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003400.

Nalli:2015:TOV

[1237] Ayse Nalli and Cagla Ozyilmaz. The third order variations on the Fibonacci universal code. Journal of Number Theory, 149(??):15– 32, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002492.

Cai:2015:NGF

[1238] Tianxin Cai, Deyi Chen, and Yong Zhang. A new generalization of Fermat’s Last Theorem. Journal of Number Theory, 149(??):33– 45, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003096.

Swanepoel:2015:GTE

[1239] Jan W. H. Swanepoel. On a generalization of a theorem by Euler. Journal of Number Theory, 149(??):46–56, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003497.

Merca:2015:NLG

[1240] Mircea Merca. A new look on the generating function for the number of divisors. Journal of Number Theory, 149(??):57–69, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003576.

Schmidt:2015:RDM

[1241] Harry Schmidt. Resultants and discriminants of multiplication polynomials for elliptic curves. Journal of Number Theory, 149(??):70– 91, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003540.

Baruah:2015:IFA

[1242] Nayandeep Deka Baruah and Kallol Nath. Infinite families of arithmetic identities and congruences for bipartitions with 3-cores. Journal of Number Theory, 149(??):92–104, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1400359X.

Takemori:2015:SES

[1243] Sho Takemori. Siegel Eisenstein series of degree n and Λ-adic Eisenstein series. Journal of Number Theory, 149(??):105–138, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003527.

Yang:2015:DSM

[1244] Qingjie Yang. Decomposability of symplectic matrices over principal ideal domain. Journal of Number Theory, 149(??):139–152, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14003643.

Petrov:2015:HSC

[1245] Aleksandar Petrov. On hyperderivatives of single-cuspidal Drinfeld modular forms with A-expansions. Journal of Number Theory, 149(??):153– 165, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003667.

Murata:2015:QDE

[1246] Manabu Murata. On quadratic Diophantine equations in four variables and orders associated with lattices II. Journal of Number Theory, 149(??):166–200, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1400362X.

Trevino:2015:LPN

[1247] Enrique Treviño. The least k-th power non-residue. Journal of Number Theory, 149(??):201–224, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14003680.

Adam:2015:CGH

[1248] Michael Adam and Gunter Malle. A class group heuristic based on the distribution of 1-eigenspaces in matrix groups. Journal of Number Theory, 149(??):225–235, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003679.

Singer:2015:PLF

[1249] Johannes Singer. Positive Limit-Fourier transform of Farey fractions. Journal of Number Theory, 149(??):236–258, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003503.

Lasjaunias:2015:HCF

[1250] A. Lasjaunias and J.-Y. Yao. Hyperquadratic continued fractions in odd characteristic with partial quotients of degree one. Journal of Number Theory, 149(??):259–284, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X14003618.

Jennings-Shaffer:2015:HOS

[1251] C. Jennings-Shaffer. Higher order SPT functions for overpartitions, overpartitions with smallest part even, and partitions with smallest part even and without repeated odd parts. Journal of Number Theory, 149(??):285–312, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003606.

Chen:2015:IAE

[1252] Chao-Ping Chen and Jing-Yun Liu. Inequalities and asymptotic expansions for the gamma function. Journal of Number Theory, 149(??):313– 326, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003011.

Cao:2015:MCF

[1253] Xiaodong Cao, Hongmin Xu, and Xu You. Multiple-correction and faster approximation. Journal of Number Theory, 149(??):327–350, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003655.

Das:2015:JFD

[1254] Soumya Das and B. Ramakrishnan. Jacobi forms and differential operators. Journal of Number Theory, 149(??):351–367, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003539.

Saez:2015:OBB

[1255] Pablo Sáez, Xavier Vidaux, and Maxim Vsemirnov. Optimal bounds for Büchi’s problem in modular arithmetic. Journal of Number Theory, 149(??):368–403, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003564.

Dixit:2015:ZCR

[1256] Atul Dixit, Nicolas Robles, Arindam Roy, and Alexandru Zaharescu. Zeros of combinations of the Riemann ξ-function on bounded vertical shifts. Journal of Number Theory, 149(??):404–434, April 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003515.

Anonymous:2015:PMb

[1257] Anonymous. Pages 1–208 (May 2015). Journal of Number Theory, 150 (??):1–208, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBe

[1258] Anonymous. Editorial Board. Journal of Number Theory, 150(??): ifc, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000062.

Chung:2015:DPR

[1259] Chan-Liang Chung, Minking Eie, Wen-Chin Liaw, and Yao Lin Ong. Decomposition of products of Riemann zeta values. Journal of Number Theory, 150(??):1–20, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003631.

ElOtmani:2015:SND

[1260] S. El Otmani, G. Rhin, and J.-M. Sac-Épée. A Salem number with degree 34 and trace −3. Journal of Number Theory, 150(??):21– 25, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003886.

Erdogan:2015:UFI

[1261] Altan Erdoğan. A universal formula for the j-invariant of the canonical lifting. Journal of Number Theory, 150(??):26–40, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003771.

Finch:2015:PPR

[1262] Carrie Finch and Lenny Jones. Perfect power Riesel numbers. Journal of Number Theory, 150(??):41–46, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003795.

Miller:2015:CNC

[1263] John C. Miller. Class numbers in cyclotomic Z_{p}-extensions. Journal of Number Theory, 150(??):47–73, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003837.

Tang:2015:ETC

[1264] Min Tang. On the Erdős–Turán conjecture. Journal of Number Theory, 150(??):74–80, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000037.

Keating:2015:IIT

[1265] Kevin Keating. Indices of inseparability in towers of field extensions. Journal of Number Theory, 150(??):81–97, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X14003849.

Coppo:2015:IBS

[1266] Marc-Antoine Coppo and Bernard Candelpergher. Inverse binomial series and values of Arakawa–Kaneko zeta functions. Journal of Number Theory, 150(??):98–119, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003825.

Troupe:2015:NPF

[1267] Lee Troupe. On the number of prime factors of values of the sumof-proper-divisors function. Journal of Number Theory, 150(??):120– 135, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003898.

Tokimoto:2015:RMR

[1268] Kazuki Tokimoto. On the reduction modulo p of representations of a quaternion division algebra over a p-adic field. Journal of Number Theory, 150(??):136–167, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003801.

Masakova:2015:DSQ

[1269] Zuzana Masáková, Kateřina Pastirčáková, and Edita Pelantová. Description of spectra of quadratic Pisot units. Journal of Number Theory, 150(??):168–190, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003862.

Bondarenko:2015:IHL

[1270] Andriy Bondarenko, Winston Heap, and Kristian Seip. An inequality of Hardy–Littlewood type for Dirichlet polynomials. Journal of Number Theory, 150(??):191–205, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003904.

Togbe:2015:CPF

[1271] Alain Togbé. Corrigendum to “A parametric family of cubic Thue equations” [J. Number Theory 107 (1) (2004) 63–79]. Journal of Number Theory, 150(??):206–207, May 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003850. See [4].

Anonymous:2015:PJa

[1272] Anonymous. Pages 1–278 (June 2015). Journal of Number Theory, 151 (??):1–278, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBf

[1273] Anonymous. Editorial Board. Journal of Number Theory, 151(??): ifc, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000463.

Endo:2015:SIQ

[1274] Akira Endô. On the Stickelberger ideal of a quadratic twist of a cyclotomic field. Journal of Number Theory, 151(??):1–6, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X14003783.

Yelton:2015:IAR

[1275] Jeffrey Yelton. Images of 2-adic representations associated to hyperelliptic Jacobians. Journal of Number Theory, 151(??):7–17, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15000025.

Hancl:2015:IEG

[1276] Jaroslav Hančl, Kalle Leppälä, Tapani Matala-aho, and Topi Törmä. On irrationality exponents of generalized continued fractions. Journal of Number Theory, 151(??):18–35, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000049.

Tan:2015:SIS

[1277] Sheng-Li Tan and Wan-Yuan Xu. On Szpiro inequality for semistable families of curves. Journal of Number Theory, 151(??):36–45, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14003874.

Saikia:2015:TTF

[1278] Nipen Saikia. Two theta-function identities for the Ramanujan–Selberg continued fraction and applications. Journal of Number Theory, 151(??): 46–53, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000335.

Ly:2015:RMA

[1279] Tony Ly. Des représentations modulo p de GL(2, D), D algèbre à division sur un corps local. (French) [Modulo p representations of GL(2, D), D division algebra on a local field]. Journal of Number Theory, 151(??): 54–106, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000268.

Zanello:2015:NOV

[1280] Fabrizio Zanello. On the number of odd values of the Klein j-function and the cubic partition function. Journal of Number Theory, 151(??): 107–115, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000359.

Jeong:2015:CEL

[1281] Sangtae Jeong. Characterization of the ergodicity of 1-Lipschitz functions on Z_{2} using the q-Mahler basis. Journal of Number Theory, 151(??):116–128, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000207.

Baruah:2015:CMD

[1282] Nayandeep Deka Baruah and Zakir Ahmed. Congruences modulo p^{2} and p^{3} for k dots bracelet partitions with k = mp^{s}. Journal of Number Theory, 151(??):129–146, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000189.

Chaubey:2015:PCF

[1283] Sneha Chaubey, Melinda Lanius, and Alexandru Zaharescu. Pair correlation of fractional parts derived from rational valued sequences. Journal of Number Theory, 151(??):147–158, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1500027X.

Arenas-Carmona:2015:SNS

[1284] Luis Arenas-Carmona and Patricio Quiroz. Spinor norm for skew-Hermitian forms over quaternion algebras. Journal of Number Theory, 151(??):159–171, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000244.

Alzer:2015:CDE

[1285] Horst Alzer and Man Kam Kwong. On the concavity of Dirichlet’s eta function and related functional inequalities. Journal of Number Theory, 151(??):172–196, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000220.

He:2015:GPF

[1286] Xiaoyu He. Geometric progression-free sequences with small gaps. Journal of Number Theory, 151(??):197–210, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000311.

Flammang:2015:MMA

[1287] V. Flammang. The Mahler measure and its areal analog for totally positive algebraic integers. Journal of Number Theory, 151(??):211– 222, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000360.

Merca:2015:CBJ

[1288] Mircea Merca. A connection between Jacobi–Stirling numbers and Bernoulli polynomials. Journal of Number Theory, 151(??):223–229, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000372.

Lanphier:2015:AHC

[1289] Dominic Lanphier and Çetin Ürtiş. Arithmeticity of holomorphic cuspforms on Hermitian symmetric domains. Journal of Number Theory, 151(??):230–262, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000281.

Bank:2015:PPV

[1290] Efrat Bank and Lior Bary-Soroker. Prime polynomial values of linear functions in short intervals. Journal of Number Theory, 151(??):263– 275, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000293.

Blagouchine:2015:ETC

[1291] Iaroslav V. Blagouchine. Erratum to “A theorem for the closedform evaluation of the first generalized Stieltjes constant at rational arguments and some related summations” [J. Number Theory 148 (2015) 537–592]. Journal of Number Theory, 151(??):276–277, June 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000165. See [1233].

Anonymous:2015:PJb

[1292] Anonymous. Pages 1–208 (July 2015). Journal of Number Theory, 152 (??):1–208, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBg

[1293] Anonymous. Editorial Board. Journal of Number Theory, 152(??): ifc, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000682.

McGinn:2015:GME

[1294] Donald McGinn. Generalized Markoff equations and Chebyshev polynomials. Journal of Number Theory, 152(??):1–20, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15000190.

Ugolini:2015:SIP

[1295] S. Ugolini. Sequences of irreducible polynomials over odd prime fields via elliptic curve endomorphisms. Journal of Number Theory, 152(??): 21–37, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000414.

Cenkci:2015:PBN

[1296] Mehmet Cenkci and Takao Komatsu. Poly-Bernoulli numbers and polynomials with a q parameter. Journal of Number Theory, 152(??):38– 54, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000177.

Farhat:2015:CRE

[1297] Maya Farhat and Bouchaïb Sodaïgui. Classes réalisables d’extensions non abéliennes de degré p^{3}. (French) [Realizable classes of non-abelian extensions of degree p^{3}]. Journal of Number Theory, 152(??):55– 89, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000232.

Chan:2015:SNA

[1298] Tsz Ho Chan. Squarefull numbers in arithmetic progression II. Journal of Number Theory, 152(??):90–104, July 2015. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000323.

Xia:2015:CSR

[1299] Ernest X. W. Xia. Congruences for some l-regular partitions modulo l. Journal of Number Theory, 152(??):105–117, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15000396.

Tretkoff:2015:TCB

[1300] Paula Tretkoff. Transcendence and CM on Borcea–Voisin towers of Calabi–Yau manifolds. Journal of Number Theory, 152(??):118– 155, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000438.

Nastasescu:2015:DEC

[1301] Maria M. Nastasescu. Determination of elliptic curves by their adjoint p-adic L-functions. Journal of Number Theory, 152(??):156– 181, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000256.

Ruiz:2015:LPF

[1302] Carlos Alexis Gómez Ruiz and Florian Luca. On the largest prime factor of the ratio of two generalized Fibonacci numbers. Journal of Number Theory, 152(??):182–203, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000426.

Mistri:2015:CSS

[1303] Raj Kumar Mistri, Ram Krishna Pandey, and Om Prakash. Corrigendum to “Subsequence sums: Direct and inverse problems” [J. Number Theory 148 (2015) 235–256]. Journal of Number Theory, 152(??): 204, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000402. See [1218].

Freeze:2015:CCN

[1304] Michael Freeze, Weidong Gao, and Alfred Geroldinger. Corrigendum to “The critical number of finite abelian groups” [J. Number Theory 129 (11) (2009) 2766–2777]. Journal of Number Theory, 152(??):205– 207, July 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000219. See [10].

Anonymous:2015:PAb

[1305] Anonymous. Pages 1–426 (August 2015). Journal of Number Theory, 153 (??):1–426, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBh

[1306] Anonymous. Editorial Board. Journal of Number Theory, 153(??): ifc, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000876.

Nunes:2015:SNA

[1307] Ramon M. Nunes. Squarefree numbers in arithmetic progressions. Journal of Number Theory, 153(??):1–36, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000384.

deReyna:2015:DNZ

[1308] J. Arias de Reyna. On the distribution (mod1) of the normalized zeros of the Riemann zeta-function. Journal of Number Theory, 153(??): 37–53, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000591.

Platt:2015:LAG

[1309] D. J. Platt and T. S. Trudgian. Linnik’s approximation to Goldbach’s conjecture, and other problems. Journal of Number Theory, 153(??):54–62, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1500061X.

Barman:2015:SIS

[1310] Rupam Barman, Neelam Saikia, and Dermot McCarthy. Summation identities and special values of hypergeometric series in the p-adic setting. Journal of Number Theory, 153(??):63–84, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15000840.

Brunotte:2015:SNC

[1311] Horst Brunotte and Jörg M. Thuswaldner. Salem numbers from a class of star-like trees. Journal of Number Theory, 153(??):85–96, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500058X.

Huynh:2015:ICL

[1312] Duc Van Huynh and Kevin Keating. k-Isomorphism classes of local field extensions. Journal of Number Theory, 153(??):97–106, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000815.

Chen:2015:ASP

[1313] Kwang-Wu Chen. Applications of stuffle product of multiple zeta values. Journal of Number Theory, 153(??):107–116, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15000566.

Bazso:2015:CPS

[1314] András Bazsó and István Mező. On the coefficients of power sums of arithmetic progressions. Journal of Number Theory, 153(??):117– 123, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000839.

Wang:2015:DCS

[1315] Guoqing Wang. Davenport constant for semigroups II. Journal of Number Theory, 153(??):124–134, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000608.

Ernvall-Hytonen:2015:RRE

[1316] Anne-Maria Ernvall-Hytönen, Jesse Jääsaari, and Esa V. Vesalainen. Resonances and Ω-results for exponential sums related to Maass forms for SL(n, Z). Journal of Number Theory, 153(??):135–157, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15000785.

Atobe:2015:PHM

[1317] Hiraku Atobe. Pullbacks of Hermitian Maass lifts. Journal of Number Theory, 153(??):158–229, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15000578.

Hajli:2015:HCS

[1318] Mounir Hajli. Hauteurs canoniques des sous-variétés toriques. Journal of Number Theory, 153(??):230–269, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000621.

Flynn:2015:DIJ

[1319] E. V. Flynn. Descent via (5, 5)-isogeny on Jacobians of genus 2 curves. Journal of Number Theory, 153(??):270–282, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15000827.

Amou:2015:APC

[1320] Masaaki Amou and Keijo Väänänen. Arithmetical properties of certain infinite products. Journal of Number Theory, 153(??):283–303, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000645.

Vasuki:2015:RME

[1321] K. R. Vasuki and R. G. Veeresha. On Ramanujan’s modular equation of degree 7. Journal of Number Theory, 153(??):304–308, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000554.

Zhao:2015:AVD

[1322] Wei Zhao, Jianrong Zhao, and Shaofang Hong. The 2-adic valuations of differences of Stirling numbers of the second kind. Journal of Number Theory, 153(??):309–320, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15000803.

Miyazaki:2015:SDE

[1323] Takafumi Miyazaki and Florian Luca. On the system of Diophantine equations (m^{2} −1)^{r} +b^{2} = c^{2} and (m^{2} −1)^{x} +b^{y} = c^{z}. Journal of Number Theory, 153(??):321–345, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15000347.

Pan:2015:DTH

[1324] Hao Pan and Yong Zhang. On the divisibility of the truncated hypergeometric function _{3}F_{2}. Journal of Number Theory, 153(??):346– 353, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000669.

Choudhry:2015:DPC

[1325] Ajai Choudhry. A diophantine problem from calculus. Journal of Number Theory, 153(??):354–363, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15000633.

Gong:2015:SCS

[1326] Ke Gong and Chaohua Jia. Shifted character sums with multiplicative coefficients. Journal of Number Theory, 153(??):364–371, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000797.

Gurevich:2015:NTA

[1327] Nadya Gurevich and Dani Szpruch. The non-tempered θ_{10} Arthur parameter and Gross–Prasad conjectures. Journal of Number Theory, 153(??):372–426, August 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003813.

Anonymous:2015:PS

[1328] Anonymous. Pages 1–380 (September 2015). Journal of Number Theory, 154(??):1–380, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBi

[1329] Anonymous. Editorial Board. Journal of Number Theory, 154(??):ifc, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001122.

Maynard:2015:BLI

[1330] James Maynard. Bounded length intervals containing two primes and an almost-prime II. Journal of Number Theory, 154(??):1–15, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000657.

Wang:2015:NCC

[1331] Liuquan Wang. A new curious congruence involving multiple harmonic sums. Journal of Number Theory, 154(??):16–31, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15000852.

Kane:2015:NAS

[1332] Daniel M. Kane. On the number of ABC solutions with restricted radical sizes. Journal of Number Theory, 154(??):32–43, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X14003163.

Laurincikas:2015:GJD

[1333] Antanas Laurinčikas. A general joint discrete universality theorem for Hurwitz zeta-functions. Journal of Number Theory, 154(??):44–62, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000967.

Heyman:2015:GCD

[1334] Randell Heyman and Igor E. Shparlinski. On the greatest common divisor of shifted sets. Journal of Number Theory, 154(??):63–73, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001018.

Goedhart:2015:DAE

[1335] Eva G. Goedhart and Helen G. Grundman. Diophantine approximation and the equation (a^{2}cx^{k}−1)(b^{2}cy^{k}−1) = (abcz^{k}−1)^{2}. Journal of Number Theory, 154(??):74–81, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001079.

Kumar:2015:FMN

[1336] K. Senthil Kumar. Fields of Mahler’s U-numbers and Schanuel’s conjecture. Journal of Number Theory, 154(??):82–100, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15001067.

Childers:2015:PCW

[1337] Kevin Childers and Darrin Doud. Proof of a conjecture of Wong concerning octahedral Galois representations of prime power conductor. Journal of Number Theory, 154(??):101–104, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001055.

Blanc:2015:OUB

[1338] Philippe Blanc. Optimal upper bound for the maximum of the kth derivative of Hardy’s function. Journal of Number Theory, 154 (??):105–117, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000980.

Benjamin:2015:CFT

[1339] Elliot Benjamin. On the 2-class field tower conjecture for imaginary quadratic number fields with 2-class group of rank 4. Journal of Number Theory, 154(??):118–143, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1500092X.

Sofo:2015:QAH

[1340] Anthony Sofo. Quadratic alternating harmonic number sums. Journal of Number Theory, 154(??):144–159, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1500102X.

Aubry:2015:CWS

[1341] Yves Aubry, Daniel J. Katz, and Philippe Langevin. Cyclotomy of Weil sums of binomials. Journal of Number Theory, 154(??):160–178, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001006.

Sarma:2015:CFA

[1342] R. Sarma, S. Kushwaha, and R. Krishnan. Continued fractions arising from F_{∞,∈}. Journal of Number Theory, 154(??):179–200, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000992.

Demangos:2015:MPW

[1343] Luca Demangos. T -modules and Pila–Wilkie estimates. Journal of Number Theory, 154(??):201–277, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001031.

Arakelian:2015:CAM

[1344] Nazar Arakelian and Gábor Korchmáros. A characterization of the Artin–Mumford curve. Journal of Number Theory, 154(??):278–291, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000955.

Greither:2015:FCG

[1345] Cornelius Greither. Families of curves with Galois action and their L-functions. Journal of Number Theory, 154(??):292–323, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15001043.

Chen:2015:CK

[1346] Yong-Gao Chen and Xin Tong. On a conjecture of de Koninck. Journal of Number Theory, 154(??):324–364, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000931. See

[944].

Mercuri:2015:CAF

[1347] Pietro Mercuri and Claudio Stirpe. Classification of algebraic function fields with class number one. Journal of Number Theory, 154 (??):365–374, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15000979.

Shen:2015:FFC

[1348] Qibin Shen and Shuhui Shi. Function fields of class number one. Journal of Number Theory, 154(??):375–379, September 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15000943.

Anonymous:2015:PO

[1349] Anonymous. Pages 1–264 (October 2015). Journal of Number Theory, 155(??):1–264, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBj

[1350] Anonymous. Editorial Board. Journal of Number Theory, 155(??): ifc, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001419.

Yang:2015:ASU

[1351] Quan-Hui Yang and Min Tang. On the addition of squares of units and nonunits modulo n. Journal of Number Theory, 155(??):1–12, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001092.

Young:2015:AAK

[1352] Paul Thomas Young. The p-adic Arakawa–Kaneko zeta functions and p-adic Lerch transcendent. Journal of Number Theory, 155(??):13– 35, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001195.

Lu:2015:SNCb

[1353] Dawei Lu and Zexi Song. Some new continued fraction estimates of the Somos’ quadratic recurrence constant. Journal of Number Theory, 155(??):36–45, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1500133X.

Eum:2015:PGC

[1354] Ick Sun Eum, Ja Kyung Koo, and Dong Hwa Shin. Primitive generators of certain class fields. Journal of Number Theory, 155(??):46– 62, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001316.

Bagis:2015:CEC

[1355] N. D. Bagis and M. L. Glasser. Conjectures on the evaluation of certain functions with algebraic properties. Journal of Number Theory, 155(??):63–84, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001225.

Rosales:2015:FPT

[1356] J. C. Rosales, M. B. Branco, and D. Torrão. The Frobenius problem for Thabit numerical semigroups. Journal of Number Theory, 155(??):85–99, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001262.

Rivertz:2015:SSA

[1357] H. J. Rivertz. Self-similar aperiodic patterns with 9-fold symmetry. Journal of Number Theory, 155(??):100–110, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001304.

Buium:2015:DMF

[1358] Alexandru Buium. Differential modular forms attached to newforms modp. Journal of Number Theory, 155(??):111–128, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15001286.

Hajdu:2015:CSC

[1359] L. Hajdu. On a conjecture of Schäffer concerning the equation 1^{k} + ⋯ + x^{k} = y^{n}. Journal of Number Theory, 155(??):129–138, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001353. See corrigendum [1627].

Zhang:2015:ZDI

[1360] Yichao Zhang. Zagier duality and integrality of Fourier coefficients for weakly holomorphic modular forms. Journal of Number Theory, 155(??):139–162, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001328.

Turkelli:2015:CCH

[1361] Seyfi Türkelli. Connected components of Hurwitz schemes and Malle’s conjecture. Journal of Number Theory, 155(??):163–201, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15001250.

Estala-Arias:2015:DCS

[1362] Samuel Estala-Arias. Distribution of cusp sections in the Hilbert modular orbifold. Journal of Number Theory, 155(??):202–225, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15000542.

Daniels:2015:IFS

[1363] Harris B. Daniels. An infinite family of Serre curves. Journal of Number Theory, 155(??):226–247, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001365.

Hu:2015:SDT

[1364] Liqun Hu and Yanjun Yao. Sums of divisors of the ternary quadratic with almost equal variables. Journal of Number Theory, 155(??):248– 263, October 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001389.

Anonymous:2015:PN

[1365] Anonymous. Pages 1–432 (November 2015). Journal of Number Theory, 156(??):1–432, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBk

[1366] Anonymous. Editorial Board. Journal of Number Theory, 156(??):ifc, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001742.

Chaves:2015:DER

[1367] Ana Paula Chaves and Diego Marques. A Diophantine equation related to the sum of powers of two consecutive generalized Fibonacci numbers. Journal of Number Theory, 156(??):1–14, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001298.

Hinkel:2015:DCV

[1368] Dustin Hinkel and Matthew P. Young. The distribution of central values of elliptic curve L-functions. Journal of Number Theory, 156 (??):15–20, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001377.

Ploner:2015:CFD

[1369] Pietro Ploner. Computation of framed deformation functors. Journal of Number Theory, 156(??):21–37, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001237.

Berend:2015:ADU

[1370] D. Berend, M. D. Boshernitzan, and G. Kolesnik. Analytically defined uniformly dense sequences. Journal of Number Theory, 156(??):38–51, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001249.

Wang:2015:SNC

[1371] Yuchao Wang. On the saturation number for cubic surfaces. Journal of Number Theory, 156(??):52–74, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001341.

Chan:2015:RIP

[1372] Wai Kiu Chan and James Ricci. The representation of integers by positive ternary quadratic polynomials. Journal of Number Theory, 156(??):75–94, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001274.

Yu:2015:NUB

[1373] Gang Yu. A new upper bound for finite additive h-bases. Journal of Number Theory, 156(??):95–104, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001535.

Dagli:2015:RFC

[1374] M. Cihat Dağlı and Mümün Can. On reciprocity formula of character Dedekind sums and the integral of products of Bernoulli polynomials. Journal of Number Theory, 156(??):105–124, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001596.

Murty:2015:ETP

[1375] M. Ram Murty and Biswajyoti Saha. On the error term in a Parseval type formula in the theory of Ramanujan expansions. Journal of Number Theory, 156(??):125–134, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001729.

Hu:2015:NRP

[1376] Shuangnian Hu, Shaofang Hong, and Wei Zhao. The number of rational points of a family of hypersurfaces over finite fields. Journal of Number Theory, 156(??):135–153, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001523.

Guo:2015:PCZ

[1377] Victor J. W. Guo and Ji-Cai Liu. Proof of a conjecture of Z.-W. Sun on the divisibility of a triple sum. Journal of Number Theory, 156 (??):154–160, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001705.

Ostafe:2015:QDM

[1378] Alina Ostafe and Min Sha. On the quantitative dynamical Mordell– Lang conjecture. Journal of Number Theory, 156(??):161–182, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0022314X15001572. See corrigendum [1628].

Yang:2015:NJC

[1379] Hai Yang and Ruiqin Fu. A note on Jeśmanowicz’ conjecture concerning primitive Pythagorean triples. Journal of Number Theory, 156 (??):183–194, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001559.

Sugiyama:2015:ABM

[1380] Shingo Sugiyama. Asymptotic behaviors of means of central values of automorphic L-functions for GL(2 ). Journal of Number Theory, 156 (??):195–246, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001493.

Haensch:2015:CAU

[1381] Anna Haensch. A characterization of almost universal ternary inhomogeneous quadratic polynomials with conductor 2. Journal of Number Theory, 156(??):247–262, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001626.

Fang:2015:LMD

[1382] Lulu Fang. Large and moderate deviation principles for alternating Engel expansions. Journal of Number Theory, 156(??):263–276, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001547.

Sairaiji:2015:CNF

[1383] Fumio Sairaiji and Takuya Yamauchi. On the class numbers of the fields of the p^{n}-torsion points of certain elliptic curves over Q. Journal of Number Theory, 156(??):277–289, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1500150X.

Patane:2015:ICT

[1384] Frank Patane. An identity connecting theta series associated with binary quadratic forms of discriminant Δ and Δ(prime)^{2}. Journal of Number Theory, 156(??):290–316, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001687.

Guo:2015:PSP

[1385] Victor Z. Guo. Piatetski–Shapiro primes in a Beatty sequence. Journal of Number Theory, 156(??):317–330, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001560.

Elmonser:2015:SPG

[1386] Hédi Elmonser. Some proprieties for the q-Gamma functions via Wielandt’s Theorem. Journal of Number Theory, 156(??):331–339, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001717.

Ernvall-Hytonen:2015:MLC

[1387] Anne-Maria Ernvall-Hytönen, Almasa Odžak, Lejla Smajlović, and Medina Sušić. On the modified Li criterion for a certain class of L-functions. Journal of Number Theory, 156(??):340–367, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001481.

Garcia-Fritz:2015:UPE

[1388] Natalia Garcia-Fritz and Hector Pasten. Uniform positive existential interpretation of the integers in rings of entire functions of positive characteristic. Journal of Number Theory, 156(??):368–393, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500164X.

Pletser:2015:GSS

[1389] Vladimir Pletser. General solutions of sums of consecutive cubed integers equal to squared integers. Journal of Number Theory, 156 (??):394–413, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001699.

Ahmadi:2015:DJC

[1390] Omran Ahmadi, Gary McGuire, and Antonio Rojas-León. Decomposing Jacobians of curves over finite fields in the absence of algebraic structure. Journal of Number Theory, 156(??):414–431, November 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001602.

Anonymous:2015:PD

[1391] Anonymous. Pages 1–546 (December 2015). Journal of Number Theory, 157(??):1–546, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2015:EBl

[1392] Anonymous. Editorial Board. Journal of Number Theory, 157(??):ifc, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002073.

Kuan:2015:EPC

[1393] Yen-Liang Kuan, Wentang Kuo, and Wei-Chen Yao. On an Erdős– Pomerance conjecture for rank one Drinfeld modules. Journal of Number Theory, 157(??):1–36, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001663.

Balasubramanian:2015:MVP

[1394] R. Balasubramanian and Sumit Giri. The mean-value of a product of shifted multiplicative functions and the average number of points of elliptic curves. Journal of Number Theory, 157(??):37–53, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001614.

Kim:2015:PCR

[1395] Sungjin Kim. Positivity of constants related to elliptic curves. Journal of Number Theory, 157(??):54–63, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001638.

Choi:2015:RPF

[1396] SoYoung Choi and Chang Heon Kim. Rational period functions in higher level cases. Journal of Number Theory, 157(??):64–78, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001675.

Coffey:2015:IRF

[1397] Mark W. Coffey. Integral representations of functions and Addisontype series for mathematical constants. Journal of Number Theory, 157(??):79–98, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001511.

Kim:2015:CNM

[1398] Bumsoo Kim. The cross number of minimal zero-sum sequences in finite abelian groups. Journal of Number Theory, 157(??):99–122, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001584.

Do:2015:SCM

[1399] Thao Do, Archit Kulkarni, Steven J. Miller, David Moon, Jake Wellens, and James Wilcox. Sets characterized by missing sums and differences in dilating polytopes. Journal of Number Theory, 157(??):123– 153, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500181X.

Chang:2015:NCF

[1400] Alan Chang, David Mehrle, Steven J. Miller, Tomer Reiter, Joseph Stahl, and Dylan Yott. Newman’s conjecture in function fields. Journal of Number Theory, 157(??):154–169, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001821.

Zhelezov:2015:SSA

[1401] Dmitrii Zhelezov. On sets with small additive doubling in product sets. Journal of Number Theory, 157(??):170–183, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001833.

Ahmed:2015:NCM

[1402] Zakir Ahmed, Nayandeep Deka Baruah, and Manosij Ghosh Dastidar. New congruences modulo 5 for the number of 2-color partitions. Journal of Number Theory, 157(??):184–198, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001845.

Cheng:2015:EEB

[1403] Chuangxun Cheng. Explicit examples for the Breuil–Mézard conjecture. Journal of Number Theory, 157(??):199–214, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001857.

Adell:2015:DEB

[1404] José A. Adell and Alberto Lekuona. Dirichlet’s eta and beta functions: Concavity and fast computation of their derivatives. Journal of Number Theory, 157(??):215–222, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001882.

Merca:2015:BPN

[1405] Mircea Merca. The bisectional pentagonal number theorem. Journal of Number Theory, 157(??):223–232, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001900.

Trudgian:2015:BND

[1406] Tim Trudgian. Bounds on the number of Diophantine quintuples. Journal of Number Theory, 157(??):233–249, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001869.

Izquierdo:2015:PLG

[1407] Diego Izquierdo. Principe local–global pour les corps de fonctions sur des corps locaux supérieurs I. (French) [Local–global principle for bodies of functions on upper local bodies I]. Journal of Number Theory, 157(??):250–270, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001870.

Monopoli:2015:GSM

[1408] Francesco Monopoli. A generalization of sumsets modulo a prime. Journal of Number Theory, 157(??):271–279, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001894.

Valverde:2015:IGR

[1409] Cesar Valverde. On irreducible generic representations distinguished by orthogonal groups. Journal of Number Theory, 157(??):280–290, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500195X.

Bauchere:2015:MHC

[1410] Hugues Bauchère. Minoration de la hauteur canonique pour les modules de Drinfeld à multiplications complexes. (French) [Canonical height reduction for Drinfeld modules with complex multiplications]. Journal of Number Theory, 157(??):291–328, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15001973.

Mossinghoff:2015:NTP

[1411] Michael J. Mossinghoff and Timothy S. Trudgian. Nonnegative trigonometric polynomials and a zero-free region for the Riemann zeta-function. Journal of Number Theory, 157(??):329–349, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001924.

Rahmani:2015:BNP

[1412] Mourad Rahmani. On p-Bernoulli numbers and polynomials. Journal of Number Theory, 157(??):350–366, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002012.

Daniels:2015:NIC

[1413] Harris B. Daniels and Álvaro Lozano-Robledo. On the number of isomorphism classes of CM elliptic curves defined over a number field. Journal of Number Theory, 157(??):367–396, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001912.

Tung:2015:MMP

[1414] Shih Ping Tung. Max–min of polynomials and exponential diophantine equations (II). Journal of Number Theory, 157(??):397–423, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002024.

Movasati:2015:QMF

[1415] Hossein Movasati. Quasi-modular forms attached to elliptic curves: Hecke operators. Journal of Number Theory, 157(??):424–441, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002036.

Choi:2015:PHM

[1416] Dohoon Choi and Subong Lim. Pairings of harmonic Maass–Jacobi forms involving special values of partial L-functions. Journal of Number Theory, 157(??):442–467, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15001948.

Han:2015:DIL

[1417] Jiyoung Han, Hyunsuk Kang, Yong-Cheol Kim, and Seonhee Lim. Distribution of integral lattice points in an ellipsoid with a diophantine center. Journal of Number Theory, 157(??):468–506, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001961.

Pach:2015:GMS

[1418] Péter Pál Pach. Generalized multiplicative Sidon sets. Journal of Number Theory, 157(??):507–529, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15002000.

Bandini:2015:CIS

[1419] Andrea Bandini, Francesc Bars, and Ignazio Longhi. Characteristic ideals and Selmer groups. Journal of Number Theory, 157(??):530–546, December 2015. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001936.

Anonymous:2016:PJa

[1420] Anonymous. Pages 1–410 (January 2016). Journal of Number Theory, 158(??):1–410, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBa

[1421] Anonymous. Editorial Board. Journal of Number Theory, 158(??): ifc, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002590.

Blackman:2016:SCM

[1422] Terrence Richard Blackman and Stefan Lemurell. Spectral correspondences for Maass waveforms on quaternion groups. Journal of Number Theory, 158(??):1–22, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001997.

Hurlimann:2016:CLC

[1423] Werner Hürlimann. Cooper and Lam’s conjecture for generalized Bell ternary quadratic forms. Journal of Number Theory, 158(??):23–32, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002334.

Chen:2016:SFD

[1424] Kwang-Wu Chen, Chan-Liang Chung, and Minking Eie. Sum formulas and duality theorems of multiple zeta values. Journal of Number Theory, 158(??):33–53, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002279.

Matera:2016:EEN

[1425] Guillermo Matera, Mariana Pérez, and Melina Privitelli. Explicit estimates for the number of rational points of singular complete intersections over a finite field. Journal of Number Theory, 158(??):54– 72, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002206.

Hu:2016:IT

[1426] Su Hu and Min-Soo Kim. The (S, {2})-Iwasawa theory. Journal of Number Theory, 158(??):73–89, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15002267.

Brown:2016:LCE

[1427] Francis Brown and Sami Omar. Li’s criterion for Epstein zeta functions, generalization of Kronecker’s limit formula and the Gauss problem. Journal of Number Theory, 158(??):90–103, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15002218.

Bremner:2016:PRD

[1428] Andrew Bremner and Maciej Ulas. Points at rational distances from the vertices of certain geometric objects. Journal of Number Theory, 158(??):104–133, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002243.

Zou:2016:PCE

[1429] Yuru Zou and Derong Kong. On a problem of countable expansions. Journal of Number Theory, 158(??):134–150, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15002309.

Chaubey:2016:PCF

[1430] Sneha Chaubey, Melinda Lanius, and Alexandru Zaharescu. Pair correlation of fractional parts derived from rational valued sequences, II. Journal of Number Theory, 158(??):151–164, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15002346.

Fares:2016:VCS

[1431] Youssef Fares and Samuel Petite. The valuative capacity of subshifts of finite type. Journal of Number Theory, 158(??):165–184, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002280.

Akbary:2016:SEN

[1432] Amir Akbary, Jeff Bleaney, and Soroosh Yazdani. On symmetries of elliptic nets and valuations of net polynomials. Journal of Number Theory, 158(??):185–216, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002188.

Axelsson:2016:GHL

[1433] Ekaterina Yurova Axelsson and Andrei Khrennikov. Generalization of Hensel’s lemma: Finding the roots of p-adic Lipschitz functions. Journal of Number Theory, 158(??):217–233, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002176.

Jang:2016:RDS

[1434] Se Wook Jang and Byeong Moon Kim. A refinement of the Dress– Scharlau theorem. Journal of Number Theory, 158(??):234–243, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002152.

Sun:2016:RTK

[1435] Chaochao Sun and Kejian Xu. On the p-rank of tame kernel of number fields. Journal of Number Theory, 158(??):244–267, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500222X.

Dupuy:2016:BBW

[1436] Taylor Dupuy and David E. Weirich. Bits of 3^{n} in binary, Wieferich primes and a conjecture of Erdős. Journal of Number Theory, 158 (??):268–280, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002139.

Adiga:2016:SNM

[1437] Chandrashekar Adiga, N. A. S. Bulkhali, D. Ranganatha, and H. M. Srivastava. Some new modular relations for the Rogers–Ramanujan type functions of order eleven with applications to partitions. Journal of Number Theory, 158(??):281–297, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002322.

Loper:2016:RMP

[1438] K. Alan Loper and Nicholas J. Werner. Resultants of minimal polynomials of maximal real cyclotomic extensions. Journal of Number Theory, 158(??):298–315, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002140.

Nakata:2016:NCP

[1439] Mitsusada Nakata. Number of class polynomials of multiple etaquotients with odd level. Journal of Number Theory, 158(??):316– 332, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002231.

Richey:2016:PIE

[1440] Joseph Richey and Noah Shutty. Polynomial identities on eigenforms. Journal of Number Theory, 158(??):333–347, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1500219X.

Khan:2016:MTR

[1441] Nadia Khan, Shin ichi Katayama, Toru Nakahara, and Tsuyoshi Uehara. Monogenity of totally real algebraic extension fields over a cyclotomic field. Journal of Number Theory, 158(??):348–355, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002310.

Andrews:2016:PIM

[1442] George E. Andrews and Stephen Hill. Partition identities with mixed mock modular forms. Journal of Number Theory, 158(??):356–364, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002358.

Blagouchine:2016:EGE

[1443] Iaroslav V. Blagouchine. Expansions of generalized Euler’s constants into the series of polynomials in π^{−2} and into the formal enveloping series with rational coefficients only. Journal of Number Theory, 158(??):365–396, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002255. See corrigendum [1817].

Bilu:2016:RPS

[1444] Yuri Bilu, Florian Luca, and Amalia Pizarro-Madariaga. Rational products of singular moduli. Journal of Number Theory, 158(??):397–410, January 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002371.

Anonymous:2016:PF

[1445] Anonymous. Pages 1–436 (February 2016). Journal of Number Theory, 159(??):1–436, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBb

[1446] Anonymous. Editorial Board. Journal of Number Theory, 159(??):ifc, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002899.

Sander:2016:SEU

[1447] J. W. Sander. Sums of exceptional units in residue class rings. Journal of Number Theory, 159(??):1–6, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002516.

Dutkay:2016:NTP

[1448] Dorin Ervin Dutkay and John Haussermann. Number theory problems from the harmonic analysis of a fractal. Journal of Number Theory, 159(??):7–26, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002425.

Dubickas:2016:PDI

[1449] Artūras Dubickas and Min Sha. Positive density of integer polynomials with some prescribed properties. Journal of Number Theory, 159(??):27–44, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002504.

Languasco:2016:SOP

[1450] Alessandro Languasco and Alessandro Zaccagnini. Sum of one prime and two squares of primes in short intervals. Journal of Number Theory, 159(??):45–58, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002437.

Vasuki:2016:RES

[1451] K. R. Vasuki and R. G. Veeresha. Ramanujan’s Eisenstein series of level 7 and 14. Journal of Number Theory, 159(??):59–75, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002528.

Shen:2016:RSA

[1452] Luming Shen and Xinqiang Li. The recurrence set arising in α-Lüroth transformation. Journal of Number Theory, 159(??):76–88, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500253X.

Qi:2016:TCF

[1453] Feng Qi and Robin J. Chapman. Two closed forms for the Bernoulli polynomials. Journal of Number Theory, 159(??):89–100, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15002541.

Gawron:2016:PIS

[1454] Maciej Gawron and Maciej Ulas. On primitive integer solutions of the Diophantine equation t^{2} = G(x, y, z) and related results. Journal of Number Theory, 159(??):101–122, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002450.

Tsuyumine:2016:STS

[1455] Shigeaki Tsuyumine. Sums of three squares under congruence condition modulo a prime. Journal of Number Theory, 159(??):123–159, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002449.

Xia:2016:APP

[1456] Ernest X. W. Xia. Arithmetic properties of partitions with designated summands. Journal of Number Theory, 159(??):160–175, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002292.

Liu:2016:DSI

[1457] H.-Q. Liu. The distribution of squarefull integers (II). Journal of Number Theory, 159(??):176–192, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15002383.

Lorscheid:2016:RTR

[1458] Oliver Lorscheid and Cecília Salgado. A remark on topologies for rational point sets. Journal of Number Theory, 159(??):193–201, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002486.

Liu:2016:DSN

[1459] H.-Q. Liu. On the distribution of squarefree numbers. Journal of Number Theory, 159(??):202–222, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15002462.

Hutchinson:2016:SHI

[1460] Kevin Hutchinson. The second homology of SL_{2} of S-integers. Journal of Number Theory, 159(??):223–272, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002553.

Wakefield:2016:GIP

[1461] Nathan Wakefield. Generalized iterations and primitive divisors. Journal of Number Theory, 159(??):273–294, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002577.

Quan:2016:SBR

[1462] Nguyen Ngoc Dong Quan. Some basic results in elementary number theory in function fields. Journal of Number Theory, 159(??):295–306, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002498.

Nair:2016:LBG

[1463] Saranya G. Nair and T. N. Shorey. Lower bounds for the greatest prime factor of product of consecutive positive integers. Journal of Number Theory, 159(??):307–328, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15002474.

Araci:2016:SCT

[1464] Serkan Araci and Mehmet Acikgoz. On the von Staudt–Clausen’s theorem related to q-Frobenius–Euler numbers. Journal of Number Theory, 159(??):329–339, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002656. See corrigendum [1743].

Alvarez:2016:CST

[1465] A. Álvarez. A calculation of L-series in terms of Jacobi sums. Journal of Number Theory, 159(??):340–354, February 2016. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002413.

Oruc:2016:NPI

[1466] A. Yavuz Oruç. On number of partitions of an integer into a fixed number of positive integers. Journal of Number Theory, 159(??):355– 369, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500236X.

Hou:2016:OFC

[1467] Fei Hou and Guangshi Lü. Oscillations of Fourier coefficients of cusp forms over primes. Journal of Number Theory, 159(??):370–383, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002401.

Bajnok:2016:MSS

[1468] Béla Bajnok and Ryan Matzke. On the minimum size of signed sumsets in elementary abelian groups. Journal of Number Theory, 159 (??):384–401, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002565.

Borges:2016:FNC

[1469] Herivelto Borges. Frobenius nonclassical components of curves with separated variables. Journal of Number Theory, 159(??):402–425, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002395.

Virdol:2016:CCA

[1470] Cristian Virdol. Cyclic components of abelian varieties ( mod ℘). Journal of Number Theory, 159(??):426–433, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002838.

Jeon:2016:CWM

[1471] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Corrigendum to “Weak Maass–Poincaré series and weight 3/2 mock modular forms” [J. Number Theory 133 (8) (2013) 2567–2587]. Journal of Number Theory, 159(??):434–435, February 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002851. See [768].

Anonymous:2016:PMa

[1472] Anonymous. Pages 1–804 (March 2016). Journal of Number Theory, 160 (??):1–804, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBc

[1473] Anonymous. Editorial Board. Journal of Number Theory, 160(??): ifc, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500325X.

Errthum:2016:DAA

[1474] Eric Errthum. A division algorithm approach to p-adic Sylvester expansions. Journal of Number Theory, 160(??):1–10, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1500284X.

Pellegrino:2016:OCM

[1475] Daniel Pellegrino. The optimal constants of the mixed (ℓ_{1}, ℓ_{2})- Littlewood inequality. Journal of Number Theory, 160(??):11–18, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002759.

Merca:2016:CSF

[1476] Mircea Merca. The cardinal sine function and the Chebyshev– Stirling numbers. Journal of Number Theory, 160(??):19–31, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15002863.

Sands:2016:TZF

[1477] Jonathan W. Sands. Twisted zeta functions of quaternion orders. Journal of Number Theory, 160(??):32–43, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002747.

Yan:2016:SES

[1478] Xiaofei Yan. On some exponential sums involving Maass forms over arithmetic progressions. Journal of Number Theory, 160(??):44–59, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002784.

Merca:2016:CIR

[1479] Mircea Merca. Combinatorial interpretations of a recent convolution for the number of divisors of a positive integer. Journal of Number Theory, 160(??):60–75, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002826.

Laishram:2016:IGL

[1480] Shanta Laishram, Saranya G. Nair, and T. N. Shorey. Irreducibility of generalized Laguerre polynomials L_{n}^{( \frac{1}{2} +u)}(x) with integer u. Journal of Number Theory, 160(??):76–107, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002723.

Sun:2016:LMW

[1481] Zhi-Wei Sun. The least modulus for which consecutive polynomial values are distinct. Journal of Number Theory, 160(??):108–116, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002693.

Adrian:2016:SRS

[1482] Moshe Adrian and Baiying Liu. Some results on simple supercuspidal representations of GL_{n}(F). Journal of Number Theory, 160(??):117– 147, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500270X.

Yu:2016:SRT

[1483] Myungjun Yu. Selmer ranks of twists of hyperelliptic curves and superelliptic curves. Journal of Number Theory, 160(??):148–185, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002772.

Franc:2016:TDI

[1484] Cameron Franc and Geoffrey Mason. Three-dimensional imprimitive representations of the modular group and their associated modular forms.

Journal of Number Theory, 160(??):186–214, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15002814.

Voutier:2016:TFI

[1485] Paul M. Voutier. Thue’s Fundamentaltheorem, II: Further refinements and examples. Journal of Number Theory, 160(??):215–236, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002760.

Burkhardt:2016:VPG

[1486] Paula Burkhardt, Alice Zhuo-Yu Chan, Gabriel Currier, Stephan Ramon Garcia, Florian Luca, and Hong Suh. Visual properties of generalized Kloosterman sums. Journal of Number Theory, 160(??):237– 253, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002875.

Yau:2016:PSU

[1487] Stephen S.-T. Yau, Beihui Yuan, and Huaiqing Zuo. On the polynomial sharp upper estimate conjecture in 7-dimensional simplex. Journal of Number Theory, 160(??):254–286, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002802.

All:2016:SEA

[1488] Timothy All. On Stickelberger elements for Q(ζ_{p^{n}+1})^{+} and p-adic L-functions. Journal of Number Theory, 160(??):287–306, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15002668.

Hou:2016:GF

[1489] Xiang dong Hou. On global P-forms. Journal of Number Theory, 160(??):307–325, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002796.

Chen:2016:FPM

[1490] Hui Chen and Tianping Zhang. On the fourth power mean of the analogous general Kloosterman sum. Journal of Number Theory, 160(??):

326–333, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002991.

Glanois:2016:MUF

[1491] Claire Glanois. Motivic unipotent fundamental groupoid of G_{m} \ μ_{N} for N = 2, 3, 4, 6, 8 and Galois descents. Journal of Number Theory, 160(??):334–384, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002711.

Dixit:2016:RTC

[1492] Atul Dixit, Arindam Roy, and Alexandru Zaharescu. Riesz-type criteria and theta transformation analogues. Journal of Number Theory, 160(??):385–408, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002735.

McKee:2016:SNT

[1493] James McKee and Pavlo Yatsyna. Salem numbers of trace −2, and a conjecture of Estes and Guralnick. Journal of Number Theory, 160(??):409–417, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003182.

Chen:2016:SIAa

[1494] Chao-Ping Chen and Wei-Wei Tong. Sharp inequalities and asymptotic expansions for the gamma function. Journal of Number Theory, 160(??):418–431, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003200.

Breuer:2016:EDM

[1495] Florian Breuer. Explicit Drinfeld moduli schemes and Abhyankar’s Generalized Iteration Conjecture. Journal of Number Theory, 160(??):432– 450, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003029.

Hu:2016:PEO

[1496] Liqun Hu and Li Yang. On pairs of equations in one prime, two prime squares and powers of 2. Journal of Number Theory, 160(??):451– 462, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003042.

Kobayashi:2016:CNP

[1497] Hirotomo Kobayashi. Class numbers of pure quintic fields. Journal of Number Theory, 160(??):463–477, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003169.

Ouyang:2016:NPF

[1498] Yi Ouyang and Jinbang Yang. Newton polygons of L functions of polynomials x^{d} + ax. Journal of Number Theory, 160(??):478–491, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003194.

Wang:2016:HCF

[1499] Victor Y. Wang. On Hilbert 2-class fields and 2-towers of imaginary quadratic number fields. Journal of Number Theory, 160(??):492– 515, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003212.

Wang:2016:IQL

[1500] Ruiqing Wang. Indecomposable quadratic lattices over global function fields. Journal of Number Theory, 160(??):516–525, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1500311X.

Hegyvari:2016:NCS

[1501] Norbert Hegyvári. Note on character sums of Hilbert cubes. Journal of Number Theory, 160(??):526–535, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003157.

Tuxanidy:2016:NFE

[1502] Aleksandr Tuxanidy and Qiang Wang. On the number of N-free elements with prescribed trace. Journal of Number Theory, 160(??):536– 565, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003078.

Uzun:2016:MHC

[1503] Mecit Kerem Uzun. Motivic homology and class field theory over p-adic fields. Journal of Number Theory, 160(??):566–585, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003030.

Im:2016:WPC

[1504] Bo-Hae Im, Daeyeol Jeon, and Chang Heon Kim. Weierstrass points on certain modular groups. Journal of Number Theory, 160(??):586– 602, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003170.

Chou:2016:TRE

[1505] Michael Chou. Torsion of rational elliptic curves over quartic Galois number fields. Journal of Number Theory, 160(??):603–628, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003121.

Mkaouar:2016:NNP

[1506] M. Mkaouar and W. Wannes. On the normal number of prime factors of φ(n) subject to certain congruence conditions. Journal of Number Theory, 160(??):629–645, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1500308X.

Moree:2016:SCD

[1507] Pieter Moree and Ana Zumalacárregui. Salajan’s conjecture on discriminating terms in an exponential sequence. Journal of Number Theory, 160(??):646–665, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003145.

Chara:2016:CKT

[1508] M. Chara and R. Toledano. On cubic Kummer type towers of Garcia, Stichtenoth and Thomas. Journal of Number Theory, 160(??):666– 678, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003017.

Palm:2016:CAF

[1509] Marc Palm. The character of GL(2) automorphic forms. Journal of Number Theory, 160(??):679–699, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15003108.

Coppola:2016:ETP

[1510] Giovanni Coppola, M. Ram Murty, and Biswajyoti Saha. On the error term in a Parseval type formula in the theory of Ramanujan expansions II. Journal of Number Theory, 160(??):700–715, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003224.

Berend:2016:JDC

[1511] D. Berend and G. Kolesnik. Joint distribution of completely q-additive functions in residue classes. Journal of Number Theory, 160(??):716– 738, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003054.

Kosters:2016:DGP

[1512] Michiel Kosters. Deterministically generating Picard groups of hyperelliptic curves over finite fields. Journal of Number Theory, 160(??): 739–758, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003066.

Tsang:2016:GMS

[1513] Cindy (Sin Yi) Tsang. On the Galois module structure of the square root of the inverse different in abelian extensions. Journal of Number Theory, 160(??):759–804, March 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15003091.

Anonymous:2016:HLW

[1514] Anonymous. In honor of the lifelong work of Wen-Ching Winnie Li. Journal of Number Theory, 161(??):??, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBd

[1515] Anonymous. Editorial Board. Journal of Number Theory, 161(??): ifc, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003364.

Hoffman:2016:P

[1516] Jerome Hoffman, Ling Long, Karl Mahlburg, Jorge Morales, Bogdan Oporowski, Robert Perlis, and Holly Swisher. Preface. Journal of Number Theory, 161(??):1–16, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15003340.

Yang:2016:SNC

[1517] Tonghai Yang and Hongbo Yin. Some non-congruence subgroups and the associated modular curves. Journal of Number Theory, 161(??): 17–48, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15001213.

Liu:2016:ACG

[1518] Tong Liu and Jiu-Kang Yu. On automorphy of certain Galois representations of GO_{4}-type. Journal of Number Theory, 161(??):49– 74, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500030X.

Tsai:2016:NOO

[1519] Pei-Yu Tsai. Newforms for odd orthogonal groups. Journal of Number Theory, 161(??):75–87, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002644.

Jiang:2016:FCA

[1520] Dihua Jiang and Chenyan Wu. On (χ, b)-factors of cuspidal automorphic representations of unitary groups I. Journal of Number Theory, 161(??):88–118, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003709.

Balakrishnan:2016:AHE

[1521] Jennifer S. Balakrishnan. On 3-adic heights on elliptic curves. Journal of Number Theory, 161(??):119–134, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002164.

Clemm:2016:ECE

[1522] Amanda Clemm and Sarah Trebat-Leder. Elliptic curves with everywhere good reduction. Journal of Number Theory, 161(??):135– 145, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500205X.

Lauter:2016:HTS

[1523] Kristin Lauter, Michael Naehrig, and Tonghai Yang. Hilbert theta series and invariants of genus 2 curves. Journal of Number Theory, 161(??):146–174, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001109.

Deines:2016:GLC

[1524] Alyson Deines, Jenny G. Fuselier, Ling Long, Holly Swisher, and Fang-Ting Tu. Generalized Legendre curves and quaternionic multiplication. Journal of Number Theory, 161(??):175–203, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001651.

Bruinier:2016:CPN

[1525] Jan Hendrik Bruinier, Ken Ono, and Andrew V. Sutherland. Class polynomials for nonholomorphic modular functions. Journal of Number Theory, 161(??):204–229, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15002127.

Duncan:2016:JDP

[1526] John F. R. Duncan and Ken Ono. The Jack Daniels problem. Journal of Number Theory, 161(??):230–239, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15002048.

Bringmann:2016:HIW

[1527] Kathrin Bringmann and Larry Rolen. Half-integral weight Eichler integrals and quantum modular forms. Journal of Number Theory, 161(??):240–254, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15001080.

Keilthy:2016:RRT

[1528] Adam Keilthy and Robert Osburn. Rogers–Ramanujan type identities for alternating knots. Journal of Number Theory, 161(??):255– 280, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500044X.

Kang:2016:RHS

[1529] Ming-Hsuan Kang. Riemann Hypothesis and strongly Ramanujan complexes from GL_{n}. Journal of Number Theory, 161(??):281–297, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1500267X.

Andrews:2016:CFN

[1530] George E. Andrews and James A. Sellers. Congruences for the Fishburn numbers. Journal of Number Theory, 161(??):298–310, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1400328X.

Lagarias:2016:FMS

[1531] Jeffrey C. Lagarias. A family of measures on symmetric groups and the field with one element. Journal of Number Theory, 161(??):311– 342, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15002681.

Kirschmer:2016:TQF

[1532] Markus Kirschmer and David Lorch. Ternary quadratic forms over number fields with small class number. Journal of Number Theory, 161(??):343–361, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003552.

Nebe:2016:AEU

[1533] Gabriele Nebe. Automorphisms of extremal unimodular lattices in dimension 72. Journal of Number Theory, 161(??):362–383, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15001390.

Papikian:2016:EIF

[1534] Mihran Papikian and Fu-Tsun Wei. On the Eisenstein ideal over function fields. Journal of Number Theory, 161(??):384–434, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15001201.

Goldfeld:2016:FMF

[1535] Dorian Goldfeld and Jyoti Sengupta. First moments of Fourier coefficients of GL(r) cusp forms. Journal of Number Theory, 161(??):435– 443, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14003588.

Murty:2016:NNV

[1536] M. Ram Murty and Naomi Tanabe. On the nature of e^{γ} and nonvanishing of derivatives of L-series at s = 1/2. Journal of Number Theory, 161(??):444–456, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X14003692.

Hoffstein:2016:MDS

[1537] Jeff Hoffstein, Thomas A. Hulse, and Andre Reznikov. Multiple Dirichlet series and shifted convolutions. Journal of Number Theory, 161(??): 457–533, April 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003005.

Anonymous:2016:PMb

[1538] Anonymous. Pages 1–614 (May 2016). Journal of Number Theory, 162 (??):1–614, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBe

[1539] Anonymous. Editorial Board. Journal of Number Theory, 162(??): ifc, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000123.

Bengoechea:2016:MLC

[1540] Paloma Bengoechea and Evgeniy Zorin. On the Mixed Littlewood Conjecture and continued fractions in quadratic fields. Journal of Number Theory, 162(??):1–10, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003492.

Fishman:2016:VDT

[1541] Lior Fishman and David Simmons. Variations on Dirichlet’s theorem. Journal of Number Theory, 162(??):11–22, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003534.

Merca:2016:NPI

[1542] Mircea Merca. A note on the partitions involving parts of k different magnitudes. Journal of Number Theory, 162(??):23–34, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003637.

Parshall:2016:SGB

[1543] Hans Parshall. Small gaps between configurations of prime polynomials. Journal of Number Theory, 162(??):35–53, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003510.

Hu:2016:RFA

[1544] Su Hu, Daeyeoul Kim, and Min-Soo Kim. On reciprocity formula of Apostol–Dedekind sum with quasi-periodic Euler functions. Journal of Number Theory, 162(??):54–67, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003649.

Erdelyi:2016:SVF

[1545] Márton Erdélyi. On the Schneider–Vigneras functor for principal series. Journal of Number Theory, 162(??):68–85, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003479.

Hess:2016:TCF

[1546] Florian Hess and Maike Massierer. Tame class field theory for global function fields. Journal of Number Theory, 162(??):86–115, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15003467.

Li:2016:WGP

[1547] Yingjie Li and Yingchun Cai. Waring–Goldbach problem: Two squares and some higher powers. Journal of Number Theory, 162(??):116– 136, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003613.

Weber:2016:AAC

[1548] Michel J. G. Weber. An arithmetical approach to the convergence problem of series of dilated functions and its connection with the Riemann zeta function. Journal of Number Theory, 162(??):137– 179, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003443.

Chen:2016:MSS

[1549] Guohua Chen. On monochromatic sums of squares of primes. Journal of Number Theory, 162(??):180–189, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003595.

Sun:2016:RSL

[1550] Zhi-Wei Sun. A result similar to Lagrange’s theorem. Journal of Number Theory, 162(??):190–211, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1500356X.

Tripathi:2016:FP

[1551] Amitabha Tripathi. On the Frobenius problem for {a, ha + d, ha + bd, ha + b^{2}d, …, ha + b^{k}d}. Journal of Number Theory, 162(??):212– 223, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003625.

Jahnel:2016:NCP

[1552] Jörg Jahnel and Damaris Schindler. On the number of certain del Pezzo surfaces of degree four violating the Hasse principle. Journal of Number Theory, 162(??):224–254, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003601.

Hulse:2016:CSD

[1553] Thomas A. Hulse, Chan Ieong Kuan, Eren Mehmet Kıral, and Li-Mei Lim. Counting square discriminants. Journal of Number Theory, 162(??):255–274, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003571.

Zemel:2016:RPM

[1554] Shaul Zemel. Regularized pairings of meromorphic modular forms and theta lifts. Journal of Number Theory, 162(??):275–311, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15003583.

Zaman:2016:EEZ

[1555] Asif Zaman. Explicit estimates for the zeros of Hecke L-functions. Journal of Number Theory, 162(??):312–375, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003455.

Latham:2016:UTS

[1556] Peter Latham. Unicity of types for supercuspidal representations of p-adic SL_{2}. Journal of Number Theory, 162(??):376–390, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15003509.

Bambozzi:2016:DGB

[1557] Federico Bambozzi and Oren Ben-Bassat. Dagger geometry as Banach algebraic geometry. Journal of Number Theory, 162(??):391– 462, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003674.

Finat:2016:DFK

[1558] Eugenio Finat. A distribution formula for Kashio’s p-adic loggamma function. Journal of Number Theory, 162(??):463–482, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X15003650.

Zhou:2016:VSF

[1559] Fan Zhou. Voronoi summation formulae on GL(n). Journal of Number Theory, 162(??):483–495, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003546.

Yao:2016:GMH

[1560] Weili Yao. A generalization of the Montgomery–Hooley theorem. Journal of Number Theory, 162(??):496–517, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003522.

Connes:2016:AAS

[1561] Alain Connes and Caterina Consani. Absolute algebra and Segal’s Γrings: au dessous de S̄pec(Z). Journal of Number Theory, 162(??): 518–551, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003662.

Hajdu:2016:GPR

[1562] L. Hajdu and N. Saradha. On generalizations of problems of Recaman and Pomerance. Journal of Number Theory, 162(??):552– 563, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003480.

Long:2016:LHC

[1563] D. D. Long and Morwen B. Thistlethwaite. Lenstra–Hurwitz cliques and the class number one problem. Journal of Number Theory, 162(??): 564–577, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003686.

Chen:2016:IAEa

[1564] Chao-Ping Chen and Long Lin. Inequalities and asymptotic expansions for the gamma function related to Mortici’s formula. Journal of Number Theory, 162(??):578–588, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003133.

Hu:2016:PNI

[1565] Yining Hu. Patterns in numbers and infinite sums and products. Journal of Number Theory, 162(??):589–600, May 2016. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X15003339.

Gao:2016:ECS

[1566] Weidong Gao, Dongchun Han, and Hanbin Zhang. The EGZ-constant and short zero-sum sequences over finite abelian groups. Journal of Number Theory, 162(??):601–613, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X15003327.

Anonymous:2016:PJb

[1567] Anonymous. Pages 1–608 (June 2016). Journal of Number Theory, 163 (??):1–608, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBf

[1568] Anonymous. Editorial Board. Journal of Number Theory, 163(??): ifc, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000561.

Bilu:2016:DRL

[1569] Yuri F. Bilu, Takao Komatsu, Florian Luca, Amalia Pizarro-Madariaga, and Pantelimon Stănică. On a divisibility relation for Lucas sequences. Journal of Number Theory, 163(??):1–18, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1600010X.

Gafni:2016:PP

[1570] Ayla Gafni. Power partitions. Journal of Number Theory, 163(??): 19–42, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000032.

Berczes:2016:EF

[1571] A. Bérczes, L. Hajdu, T. Miyazaki, and I. Pink. On the equation 1^{k} + 2^{k} + ⋯ + x^{k} = y^{n} for fixed x. Journal of Number Theory, 163(??):43–60, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1600007X.

Pankowski:2016:JUG

[1572] Łukasz Pańkowski. Joint universality and generalized strong recurrence with rational parameter. Journal of Number Theory, 163(??): 61–74, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000044.

Maier:2016:MFR

[1573] H. Maier and A. Sankaranarayanan. On multiplicative functions resembling the Möbius function. Journal of Number Theory, 163(??): 75–88, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000056.

Bottcher:2016:LZS

[1574] Albrecht Böttcher and Fuad Kittaneh. The limit of the zero set of polynomials of the Fibonacci type. Journal of Number Theory, 163(??): 89–100, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000184.

Curtin:2016:ETS

[1575] Brian Curtin and G. R. Pourgholi. An Euler totient sum inequality. Journal of Number Theory, 163(??):101–113, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000172.

Miska:2016:APS

[1576] Piotr Miska. Arithmetic properties of the sequence of derangements. Journal of Number Theory, 163(??):114–145, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000196.

Uematsu:2016:BGA

[1577] Tetsuya Uematsu. On the Brauer group of affine diagonal quadrics. Journal of Number Theory, 163(??):146–158, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000202.

Freiberg:2016:SIG

[1578] Tristan Freiberg. Short intervals with a given number of primes. Journal of Number Theory, 163(??):159–171, June 2016. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000081.

Pierce:2016:BBM

[1579] L. B. Pierce. Burgess bounds for multi-dimensional short mixed character sums. Journal of Number Theory, 163(??):172–210, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X15003236.

Camburu:2016:CCG

[1580] Oana-Maria Camburu, Emil-Alexandru Ciolan, Florian Luca, Pieter Moree, and Igor E. Shparlinski. Cyclotomic coefficients: gaps and jumps. Journal of Number Theory, 163(??):211–237, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000251.

Kaneko:2016:CCN

[1581] Hajime Kaneko and Takao Komatsu. Cauchy–Carlitz numbers. Journal of Number Theory, 163(??):238–254, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1600024X.

Xu:2016:AER

[1582] Aimin Xu. Asymptotic expansions related to the Glaisher–Kinkelin constant and its analogues. Journal of Number Theory, 163(??):255– 266, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000093.

Su:2016:KPS

[1583] Ren-He Su. The Kohnen plus space for Hilbert–Siegel modular forms. Journal of Number Theory, 163(??):267–295, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000275.

Kucera:2016:CUA

[1584] Radan Kučera and Azar Salami. Circular units of an abelian field ramified at three primes. Journal of Number Theory, 163(??):296– 315, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000287.

Gekeler:2016:GIT

[1585] Ernst-Ulrich Gekeler. The Galois image of twisted Carlitz modules. Journal of Number Theory, 163(??):316–330, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000263.

Jennings-Shaffer:2016:ORD

[1586] C. Jennings-Shaffer. Overpartition rank differences modulo 7 by Maass forms. Journal of Number Theory, 163(??):331–358, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000238.

Czarnecki:2016:RSR

[1587] Kyle Czarnecki. Resonance sums for Rankin–Selberg products of SL_{m}(Z) Maass cusp forms. Journal of Number Theory, 163(??):359– 374, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000020.

Gu:2016:DOH

[1588] Ze Gu and Xilin Tang. The doubles of one half of a numerical semigroup. Journal of Number Theory, 163(??):375–384, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000317.

Fang:2016:LTR

[1589] Lulu Fang, Min Wu, and Bing Li. Limit theorems related to betaexpansion and continued fraction expansion. Journal of Number Theory, 163(??):385–405, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000330.

Park:2016:RFM

[1590] Chol Park. Regular filtered (φ, N)-modules of dimension 3. Journal of Number Theory, 163(??):406–433, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000226.

Lu:2016:SNA

[1591] Dawei Lu, Cristinel Mortici, and Lixin Song. Some new approximations of Glaisher–Kinkelin’s and Bendersky–Adamchik’s constants by continued fraction. Journal of Number Theory, 163(??):434–448, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000391.

Nguyen:2016:GTC

[1592] Viet Kh. Nguyen. A genus two curve related to the class number one problem. Journal of Number Theory, 163(??):449–456, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000378.

Bayad:2016:MIF

[1593] Abdelmejid Bayad and Luis Navas. Möbius inversion formulas related to the Fourier expansions of two-dimensional Apostol–Bernoulli polynomials. Journal of Number Theory, 163(??):457–473, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1600038X.

Chern:2016:NCC

[1594] Shane Chern. New congruences for 2-color partitions. Journal of Number Theory, 163(??):474–481, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000512.

Zhang:2016:RGR

[1595] Zhenliang Zhang and Meiying Lü. The relative growth rate of the largest partial quotient to the sum of partial quotients in continued fraction expansions. Journal of Number Theory, 163(??):482– 492, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000329.

Wang:2016:NSS

[1596] Zhonghua Wang and Qiming Yan. A note on Schmidt’s subspace type theorem with moving hyperplanes. Journal of Number Theory, 163(??):493–509, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000366.

Marques:2016:SVM

[1597] Diego Marques. The solution of a version of a Mahler’s question. Journal of Number Theory, 163(??):510–519, June 2016. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000299.

Pioline:2016:TLR

[1598] Boris Pioline. A Theta lift representation for the Kawazumi–Zhang and Faltings invariants of genus-two Riemann surfaces. Journal of Number Theory, 163(??):520–541, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000524.

Brunault:2016:RSU

[1599] François Brunault. Regulators of Siegel units and applications. Journal of Number Theory, 163(??):542–569, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000500.

Wang:2016:UAA

[1600] Weiping Wang. Unified approaches to the approximations of the gamma function. Journal of Number Theory, 163(??):570–595, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000470.

Chen:2016:IAEb

[1601] Chao-Ping Chen. Inequalities and asymptotic expansions for the psi function and the Euler–Mascheroni constant. Journal of Number Theory, 163(??):596–607, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X15003558.

Anonymous:2016:PJc

[1602] Anonymous. Pages 1–438 (July 2016). Journal of Number Theory, 164 (??):1–438, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBg

[1603] Anonymous. Editorial Board. Journal of Number Theory, 164(??): ifc, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000925.

Ma:2016:CSI

[1604] Mi-Mi Ma and Yong-Gao Chen. On d-complete sequences of integers. Journal of Number Theory, 164(??):1–12, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000342.

Kleinbock:2016:BAN

[1605] Dmitry Kleinbock and Tue Ly. Badly approximable S-numbers and absolute Schmidt games. Journal of Number Theory, 164(??):13– 42, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000457.

Miguel:2016:MTI

[1606] C. Miguel. A Menon-type identity in residually finite Dedekind domains. Journal of Number Theory, 164(??):43–51, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000494.

Chung:2016:DCS

[1607] Fan Chung and Ron Graham. On the discrepancy of circular sequences of reals. Journal of Number Theory, 164(??):52–65, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000408.

Hu:2016:RNH

[1608] Gengran Hu, Yanbin Pan, Renzhang Liu, and Yuyun Chen. On random nonsingular Hermite Normal Form. Journal of Number Theory, 164(??):66–86, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000354.

Kulkarni:2016:LPE

[1609] Manisha Kulkarni, Vijay M. Patankar, and C. S. Rajan. Locally potentially equivalent two dimensional Galois representations and Frobenius fields of elliptic curves. Journal of Number Theory, 164(??):87– 102, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1600041X.

Gong:2016:PMK

[1610] Ke Gong, Willem Veys, and Daqing Wan. Power moments of Kloosterman sums. Journal of Number Theory, 164(??):103–126, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000536.

Sorensen:2016:LLC

[1611] Claus M. Sorensen. The local Langlands correspondence in families and Ihara’s lemma for U(n). Journal of Number Theory, 164(??):127– 165, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000433.

Kibelbek:2016:SCM

[1612] Jonas Kibelbek, Ling Long, Kevin Moss, Benjamin Sheller, and Hao Yuan. Supercongruences and complex multiplication. Journal of Number Theory, 164(??):166–178, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000445.

Zhang:2016:RTP

[1613] Yong Zhang. Right triangle and parallelogram pairs with a common area and a common perimeter. Journal of Number Theory, 164(??): 179–190, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000469.

Ye:2016:LAR

[1614] Dongxi Ye. Level 16 analogue of Ramanujan’s theories of elliptic functions to alternative bases. Journal of Number Theory, 164(??):191– 207, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000482.

Eie:2016:IAR

[1615] Minking Eie and Tung-Yang Lee. Identities among restricted sums of multiple zeta values. Journal of Number Theory, 164(??):208– 222, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000214.

Xiao:2016:DVS

[1616] Xuanxuan Xiao. Distribution of values of symmetric power L-functions at the edge of the critical strip. Journal of Number Theory, 164(??):

223–268, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000305.

Hone:2016:CFE

[1617] Andrew N. W. Hone. On the continued fraction expansion of certain Engel series. Journal of Number Theory, 164(??):269–281, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000639.

James:2016:EPE

[1618] Kevin James, Brandon Tran, Minh-Tam Trinh, Phil Wertheimer, and Dania Zantout. Extremal primes for elliptic curves. Journal of Number Theory, 164(??):282–298, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1600072X.

Yang:2016:GBC

[1619] Jongho Yang. A generalization of Beurling’s criterion for the Riemann hypothesis. Journal of Number Theory, 164(??):299–302, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000871.

Laishram:2016:IGH

[1620] Shanta Laishram and T. N. Shorey. Irreducibility of generalized Hermite– Laguerre polynomials III. Journal of Number Theory, 164(??):303– 322, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000664.

Lao:2016:ACR

[1621] Huixue Lao, Mark McKee, and Yangbo Ye. Asymptotics for cuspidal representations by functoriality from GL(2). Journal of Number Theory, 164(??):323–342, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000718.

Chen:2016:CAA

[1622] Shi-Chao Chen. Congruences and asymptotics of Andrews’ singular overpartitions. Journal of Number Theory, 164(??):343–358, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000834.

Brody:2016:RHT

[1623] Nicolas Brody and Jordan Schettler. Rational hyperbolic triangles and a quartic model of elliptic curves. Journal of Number Theory, 164(??): 359–374, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000676.

Hou:2016:ESI

[1624] Fei Hou. On exponential sums involving Fourier coefficients of cusp forms over primes. Journal of Number Theory, 164(??):375–404, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000688.

Merca:2016:FCP

[1625] Mircea Merca. Fast computation of the partition function. Journal of Number Theory, 164(??):405–416, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000809.

Chen:2016:MAA

[1626] Chao-Ping Chen. A more accurate approximation for the gamma function. Journal of Number Theory, 164(??):417–428, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000068.

Hajdu:2016:CCS

[1627] L. Hajdu. Corrigendum to “On a conjecture of Schäffer concerning the equation 1^{k} + ⋯ + x^{k} = y^{n}” [J. Number Theory 155 (2015) 129–138]. Journal of Number Theory, 164(??):429–432, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000822. See [1359].

Ostafe:2016:CQD

[1628] Alina Ostafe and Min Sha. Corrigendum to “On the quantitative dynamical Mordell–Lang conjecture” [J. Number Theory 156 (2015) 161–182]. Journal of Number Theory, 164(??):433–437, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000779. See [1378].

Anonymous:2016:PA

[1629] Anonymous. Pages 1–430 (August 2016). Journal of Number Theory, 165 (??):1–430, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBh

[1630] Anonymous. Editorial Board. Journal of Number Theory, 165(??): ifc, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300026.

Breuer:2016:DMP

[1631] Florian Breuer and Hans-Georg Rück. Drinfeld modular polynomials in higher rank II: Kronecker congruences. Journal of Number Theory, 165(??):1–14, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000640.

Banks:2016:CNS

[1632] William D. Banks and Tristan Freiberg. Carmichael numbers and the sieve. Journal of Number Theory, 165(??):15–29, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000755.

Booker:2016:EMG

[1633] Andrew R. Booker and Sean A. Irvine. The Euclid–Mullin graph. Journal of Number Theory, 165(??):30–57, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16000767.

Sivatski:2016:QPH

[1634] A. S. Sivatski. Quartic polynomials and the Hasse norm theorem modulo squares. Journal of Number Theory, 165(??):58–66, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16000858.

Tengely:2016:PDB

[1635] Sz. Tengely and M. Ulas. On products of disjoint blocks of arithmetic progressions and related equations. Journal of Number Theory, 165(??):67–83, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000810.

Xu:2016:ESI

[1636] Ce Xu, Yuhuan Yan, and Zhijuan Shi. Euler sums and integrals of polylogarithm functions. Journal of Number Theory, 165(??):84– 108, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000883.

Naika:2016:CAS

[1637] M. S. Mahadeva Naika and D. S. Gireesh. Congruences for Andrews’ singular overpartitions. Journal of Number Theory, 165(??):109–130, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000780.

Deschamps:2016:QPS

[1638] Bruno Deschamps and Ivan Suarez Atias. Quelques propriétés des sous-groupes de Gal(k((t))/k). (French) [Some properties of the subgroups of Gal(k((t))/k)]. Journal of Number Theory, 165(??):131– 168, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000548.

Ginzburg:2016:CSG

[1639] David Ginzburg. Classification of some Global Integrals related to groups of type A_{n}. Journal of Number Theory, 165(??):169–202, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1600069X.

Baluyot:2016:ZRZ

[1640] Siegfred Alan C. Baluyot. On the zeros of Riemann’s zeta-function on the critical line. Journal of Number Theory, 165(??):203–269, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000895.

Chen:2016:MF

[1641] Yong-Gao Chen, Jin-Hui Fang, Pingzhi Yuan, and Yueping Zheng. On multiplicative functions with f(p + q + n_{0}) = f(p) + f(q) + f(n_{0}). Journal of Number Theory, 165(??):270–289, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000901.

Gao:2016:NTB

[1642] Hui Gao. A note on torsion Breuil modules in the case er = p − 1. Journal of Number Theory, 165(??):290–303, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16000731.

Ge:2016:CES

[1643] Fan Ge. Conditional estimates on small distances between ordinates of zeros of ζ(s) and ζ′(s). Journal of Number Theory, 165(??):304– 313, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1600086X.

Chen:2016:SIAb

[1644] Chao-Ping Chen. Sharp inequalities and asymptotic series of a product related to the Euler–Mascheroni constant. Journal of Number Theory, 165(??):314–323, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000846.

Haan:2016:LGG

[1645] Jaeho Haan. The local Gan–Gross–Prasad conjecture for U(3) × U(2): the non-generic case. Journal of Number Theory, 165(??):324–354, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000706.

Peng:2016:SII

[1646] Wen Peng and Tianping Zhang. Some identities involving certain Hardy sum and Kloosterman sum. Journal of Number Theory, 165(??):355– 362, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16001001.

Aigner-Horev:2016:PCS

[1647] Elad Aigner-Horev and Hiêp. Hàn. Polynomial configurations in subsets of random and pseudo-random sets. Journal of Number Theory, 165(??):363–381, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000421.

Bauch:2016:CIB

[1648] Jens-Dietrich Bauch. Computation of integral bases. Journal of Number Theory, 165(??):382–407, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16000743.

Nagano:2016:HCF

[1649] A. Nagano and H. Shiga. To the Hilbert class field from the hypergeometric modular function. Journal of Number Theory, 165(??):408– 430, August 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000792.

Anonymous:2016:PS

[1650] Anonymous. Pages 1–492 (September 2016). Journal of Number Theory, 166(??):1–492, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBi

[1651] Anonymous. Editorial Board. Journal of Number Theory, 166(??):ifc, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300397.

Kane:2016:ECM

[1652] Ben Kane and Matthias Waldherr. Explicit congruences for mock modular forms. Journal of Number Theory, 166(??):1–18, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16001013.

Mornev:2016:TPU

[1653] Maxim Mornev. Tannakian properties of unit Frobenius-modules. Journal of Number Theory, 166(??):19–30, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300129.

Chen:2016:SQR

[1654] Chao-Ping Chen and Xue-Feng Han. On Somos’ quadratic recurrence constant. Journal of Number Theory, 166(??):31–40, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300257.

Sanna:2016:AVH

[1655] Carlo Sanna. On the p-adic valuation of harmonic numbers. Journal of Number Theory, 166(??):41–46, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300270.

Jakhar:2016:PDI

[1656] Anuj Jakhar, Sudesh K. Khanduja, and Neeraj Sangwan. On prime divisors of the index of an algebraic integer. Journal of Number Theory, 166(??):47–61, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300282.

Kim:2016:AROa

[1657] Sungjin Kim. Average reciprocals of the order of a modulo n. Journal of Number Theory, 166(??):62–75, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300166.

Chakraborty:2016:CRF

[1658] Debopam Chakraborty and Anupam Saikia. Congruence relations for the fundamental unit of a pure cubic field and its class number. Journal of Number Theory, 166(??):76–84, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300191.

Howe:2016:OQS

[1659] Everett W. Howe. Optimal quotients and surjections of Mordell– Weil groups. Journal of Number Theory, 166(??):85–92, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300208.

Bondarenko:2016:GTG

[1660] Andriy Bondarenko, Titus Hilberdink, and Kristian Seip. Gál-type GCD sums beyond the critical line. Journal of Number Theory, 166 (??):93–104, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300245.

Toth:2016:CTP

[1661] László Tóth. Counting r-tuples of positive integers with k-wise relatively prime components. Journal of Number Theory, 166(??):105–116, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300233.

Gun:2016:LAI

[1662] Sanoli Gun, V. Kumar Murty, and Ekata Saha. Linear and algebraic independence of generalized Euler–Briggs constants. Journal of Number Theory, 166(??):117–136, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300117.

Lazda:2016:IBC

[1663] Christopher Lazda. Incarnations of Berthelot’s conjecture. Journal of Number Theory, 166(??):137–157, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300361.

Hatley:2016:MFA

[1664] Jeffrey Hatley. Modular forms of arbitrary even weight with no exceptional primes. Journal of Number Theory, 166(??):158–165, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300336.

Ma:2016:ISM

[1665] Ding Ma. Inverse of some matrix related to double zeta values of odd weight. Journal of Number Theory, 166(??):166–180, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300142.

Sun:2016:CNQ

[1666] Liang Sun. Class numbers of quadratic Diophantine equations. Journal of Number Theory, 166(??):181–192, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300178.

Kala:2016:NII

[1667] Vítězslav Kala. Norms of indecomposable integers in real quadratic fields. Journal of Number Theory, 166(??):193–207, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300294.

Bahmanpour:2016:PNE

[1668] Kamal Bahmanpour. Prime numbers p with expression p = a^{2} ± ab ± b^{2}. Journal of Number Theory, 166(??):208–218, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300312.

Hu:2016:CDE

[1669] Hui Hu, Xin Tong, and Yueli Yu. On consecutive 0 digits in the β-expansion of 1. Journal of Number Theory, 166(??):219–234, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300269.

Kim:2016:SER

[1670] Kwang-Seob Kim. Some examples of real quadratic fields with finite nonsolvable maximal unramified extensions. Journal of Number Theory, 166(??):235–249, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300348.

Kritzinger:2016:DSG

[1671] Ralph Kritzinger and Lisa M. Kritzinger. L_{2} discrepancy of symmetrized generalized Hammersley point sets in base b. Journal of Number Theory, 166(??):250–275, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300324.

Liu:2016:GAE

[1672] Chunlei Liu, Wenxin Liu, and Chuanze Niu. Generic T -adic exponential sums in one variable. Journal of Number Theory, 166(??):276–297, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300300.

Choi:2016:ZCW

[1673] SoYoung Choi and Bo-Hae Im. On the zeros of certain weakly holomorphic modular forms for Γ_{0}^{+}(2). Journal of Number Theory, 166 (??):298–323, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300154.

Mao:2016:AFR

[1674] Renrong Mao. Asymptotic formulas for M_{2}-ranks of partitions without repeated odd parts. Journal of Number Theory, 166(??):324–343, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300105.

Gustafsson:2016:SAR

[1675] Henrik P. A. Gustafsson, Axel Kleinschmidt, and Daniel Persson. Small automorphic representations and degenerate Whittaker vectors. Journal of Number Theory, 166(??):344–399, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300099.

Li:2016:CNI

[1676] Jin-Kai Li and Yong-Gao Chen. Critical numbers of intervals. Journal of Number Theory, 166(??):400–405, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630035X.

He:2016:SCC

[1677] Bing He. Some congruences on conjectures of van Hamme. Journal of Number Theory, 166(??):406–414, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300075.

Lettl:2016:OCN

[1678] Günter Lettl and Chanwit Prabpayak. Orders in cubic number fields. Journal of Number Theory, 166(??):415–423, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300130.

Vandehey:2016:NNC

[1679] Joseph Vandehey. New normality constructions for continued fraction expansions. Journal of Number Theory, 166(??):424–451, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300087.

Chung:2016:AER

[1680] Chan-Liang Chung, Minking Eie, and Tung-Yang Lee. Another expression of the restricted sum formula of multiple zeta values. Journal of Number Theory, 166(??):452–472, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630021X.

Tolliver:2016:EBT

[1681] Jeffrey Tolliver. An equivalence between two approaches to limits of local fields. Journal of Number Theory, 166(??):473–492, September 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300221.

Anonymous:2016:PO

[1682] Anonymous. Pages 1–546 (October 2016). Journal of Number Theory, 167(??):1–546, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBj

[1683] Anonymous. Editorial Board. Journal of Number Theory, 167(??): ifc, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300725.

Faber:2016:FDC

[1684] Xander Faber and Bianca Thompson. On the field of definition of a cubic rational function and its critical points. Journal of Number Theory, 167(??):1–6, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1630018X. See corrigendum [1770].

Berkovich:2016:PFN

[1685] Alexander Berkovich and Ali Kemal Uncu. On partitions with fixed number of even-indexed and odd-indexed odd parts. Journal of Number Theory, 167(??):7–30, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300476.

Buell:2016:CPI

[1686] Duncan A. Buell and Gregory S. Call. Class pairings and isogenies on elliptic curves. Journal of Number Theory, 167(??):31–73, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300464.

Koo:2016:GCF

[1687] Ja Kyung Koo, Dong Hwa Shin, and Dong Sung Yoon. Generation of class fields by using the Weber function. Journal of Number Theory, 167(??):74–87, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300671.

Bayer-Fluckiger:2016:IMM

[1688] Eva Bayer-Fluckiger, Martino Borello, and Peter Jossen. Inhomogeneous minima of mixed signature lattices. Journal of Number Theory, 167(??):88–103, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300579.

Tran:2016:GRA

[1689] Ha Thanh Nguyen Tran. A generalization of reduced Arakelov divisors of a number field. Journal of Number Theory, 167(??):104–117, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300531.

Zaghloul:2016:QFC

[1690] Giamila Zaghloul. Quadratic forms connected with Fourier coefficients of holomorphic and Maass cusp forms. Journal of Number Theory, 167(??):118–127, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300658.

Tang:2016:SDR

[1691] Liang Tang and Ting Zhong. Some dimension relations of the Hirst sets in regular and generalized continued fractions. Journal of Number Theory, 167(??):128–140, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300646.

Childers:2016:OEG

[1692] Kevin Childers and Darrin Doud. Octahedral extensions with a given cubic subfield. Journal of Number Theory, 167(??):141–146, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300518.

Tang:2016:FCS

[1693] Hengcai Tang and Jie Wu. Fourier coefficients of symmetric power L-functions. Journal of Number Theory, 167(??):147–160, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1630052X.

Kluners:2016:NFD

[1694] Jürgen Klüners and Florin Nicolae. Are number fields determined by Artin L-functions? Journal of Number Theory, 167(??):161–168, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300701.

Fishman:2016:BMS

[1695] Lior Fishman, Vanessa Reams, and David Simmons. The Banach– Mazur–Schmidt and Banach–Mazur–McMullen games. Journal of Number Theory, 167(??):169–179, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300683.

Roldan:2016:AES

[1696] Enrico Varela Roldán. Arithmetic of Eisenstein series of level T for the function field modular group GL(2, F_{q}[T]). Journal of Number Theory, 167(??):180–201, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300622.

Ju:2016:GWT

[1697] Jangwon Ju, Inhwan Lee, and Byeong-Kweon Oh. A generalization of Watson transformation and representations of ternary quadratic forms. Journal of Number Theory, 167(??):202–231, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1630066X.

Astudillo:2016:SLB

[1698] Sergio Astudillo, Francisco Diaz y Diaz, and Eduardo Friedman. Sharp lower bounds for regulators of small-degree number fields. Journal of Number Theory, 167(??):232–258, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630049X.

Debry:2016:RAI

[1699] Christophe Debry and Antonella Perucca. Reductions of algebraic integers. Journal of Number Theory, 167(??):259–283, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300488. See addendum [2642].

Somu:2016:CD

[1700] Sai Teja Somu. On the coefficients of divisors of x^{n}−1. Journal of Number Theory, 167(??):284–293, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300373.

Laurincikas:2016:DMU

[1701] Antanas Laurinčikas. Distribution modulo 1 and universality of the Hurwitz zeta-function. Journal of Number Theory, 167(??):294–303, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300609.

Shparlinski:2016:BDM

[1702] Igor E. Shparlinski and Kam hung Yau. Bounds of double multiplicative character sums and gaps between residues of exponential functions. Journal of Number Theory, 167(??):304–316, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300695.

Bruggeman:2016:FCE

[1703] Roelof Bruggeman and Nikolaos Diamantis. Fourier coefficients of Eisenstein series formed with modular symbols and their spectral decomposition. Journal of Number Theory, 167(??):317–335, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300567.

Jun:2016:HAA

[1704] Jaiung Jun. Hyperstructures of affine algebraic group schemes. Journal of Number Theory, 167(??):336–352, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300634.

Hough:2016:ALV

[1705] Bob Hough. The angle of large values of L-functions. Journal of Number Theory, 167(??):353–393, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300555.

Balady:2016:FCC

[1706] Steve Balady. Families of cyclic cubic fields. Journal of Number Theory, 167(??):394–406, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300580.

Thas:2016:CCP

[1707] Koen Thas. The Connes–Consani plane connection. Journal of Number Theory, 167(??):407–429, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300543.

Alkan:2016:NPI

[1708] Emre Alkan and Tevekkül Mehreliyev. Number of prime ideals in short intervals. Journal of Number Theory, 167(??):430–480, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300506.

Barrett:2016:CI

[1709] Malik Barrett and Alexandru Buium. Curvature on the integers, I. Journal of Number Theory, 167(??):481–508, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300610.

Buium:2016:CII

[1710] Alexandru Buium. Curvature on the integers, II. Journal of Number Theory, 167(??):509–545, October 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300592.

Anonymous:2016:PN

[1711] Anonymous. Pages 1–510 (November 2016). Journal of Number Theory, 168(??):1–510, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBk

[1712] Anonymous. Editorial Board. Journal of Number Theory, 168(??):ifc, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301093.

Ferraguti:2016:EIN

[1713] Andrea Ferraguti and Giacomo Micheli. On the existence of infinite, non-trivial F -sets. Journal of Number Theory, 168(??):1–12, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300786.

Davis:2016:EPR

[1714] Christopher Davis and Tommy Occhipinti. Explicit points on y^{2} + xy − t^{d}y = x^{3} and related character sums. Journal of Number Theory, 168(??):13–38, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16000652.

Fei:2016:AHL

[1715] JinHua Fei. An application of the Hardy–Littlewood conjecture. Journal of Number Theory, 168(??):39–44, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300774.

Kursungoz:2016:BSI

[1716] Kağan Kurşungöz. Bressoud style identities for regular partitions and overpartitions. Journal of Number Theory, 168(??):45–63, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300798.

Lee:2016:DCC

[1717] Jun Ho Lee and Stéphane R. Louboutin. Discriminants of cyclic cubic orders. Journal of Number Theory, 168(??):64–71, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300932.

Cheon:2016:PGP

[1718] Jung Hee Cheon and Duhyeong Kim. Probability that the k-gcd of products of positive integers is B-friable. Journal of Number Theory, 168(??):72–80, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300919.

Wake:2016:FLS

[1719] Preston Wake. Full level structures revisited: Pairs of roots of unity. Journal of Number Theory, 168(??):81–100, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300890.

Vaskouski:2016:PQU

[1720] Maksim Vaskouski, Nikita Kondratyonok, and Nikolai Prochorov. Primes in quadratic unique factorization domains. Journal of Number Theory, 168(??):101–116, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301007.

Liu:2016:SID

[1721] Ji-Cai Liu. A supercongruence involving Delannoy numbers and Schröder numbers. Journal of Number Theory, 168(??):117–127, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1630097X.

Ge:2016:NIC

[1722] Fan Ge. Note on the index conjecture in zero-sum theory and its connection to a Dedekind-type sum. Journal of Number Theory, 168 (??):128–134, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300944.

Hsia:2016:RFC

[1723] Liang-Chung Hsia and Hua-Chieh Li. Ramification filtrations of certain abelian Lie extensions of local fields. Journal of Number Theory, 168(??):135–153, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300865.

Li:2016:GBC

[1724] Yan Li, Daeyeoul Kim, and Lianrong Ma. Gaussian binomial coefficients modulo cyclotomic polynomials. Journal of Number Theory, 168 (??):154–166, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1630083X.

Dummit:2016:CQC

[1725] David S. Dummit, Evan P. Dummit, and Hershy Kisilevsky. Characterizations of quadratic, cubic, and quartic residue matrices. Journal of Number Theory, 168(??):167–179, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300920.

Angles:2016:TCZ

[1726] Bruno Anglès, Tuan Ngo Dac, and Floric Tavares Ribeiro. Twisted characteristic p zeta functions. Journal of Number Theory, 168(??):180–214, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301019.

Sun:2016:STD

[1727] Qingfeng Sun and Deyu Zhang. Sums of the triple divisor function over values of a ternary quadratic form. Journal of Number Theory, 168(??):215–246, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300889.

Hopper:2016:DDP

[1728] Jackson Hopper and Paul Pollack. Digitally delicate primes. Journal of Number Theory, 168(??):247–256, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16300853.

Park:2016:ECS

[1729] Yoon Kyung Park. Evaluation of the convolution sums ∑_{ak+bl+cm=n} σ(k)σ(l)σ(m) with lcm(a, b, c) ≤ 6. Journal of Number Theory, 168(??):257–275, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301044.

Bremner:2016:DQ

[1730] Andrew Bremner and Benjamin Carrillo. On K-derived quartics. Journal of Number Theory, 168(??):276–291, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301032.

Kim:2016:RII

[1731] Namhoon Kim. Ramanujan’s integral identities of the Riemann Ξfunction and the Lerch transcendent. Journal of Number Theory, 168 (??):292–305, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300841.

Lee:2016:LAR

[1732] Yoonjin Lee and Yoon Kyung Park. The level 13 analogue of the Rogers– Ramanujan continued fraction and its modularity. Journal of Number Theory, 168(??):306–333, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300877.

Shen:2016:TGP

[1733] Quanli Shen. The ternary Goldbach problem with primes in positive density sets. Journal of Number Theory, 168(??):334–345, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300968.

Bitan:2016:HPB

[1734] Rony A. Bitan. The Hasse principle for bilinear symmetric forms over a ring of integers of a global function field. Journal of Number Theory, 168(??):346–359, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300828.

Ahlgren:2016:NCF

[1735] Scott Ahlgren and Detchat Samart. A note on cusp forms as p-adic limits. Journal of Number Theory, 168(??):360–373, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301056.

Liu:2016:SRT

[1736] Jingbo Liu and Alicia Marino. Strictly regular ternary Hermitian forms. Journal of Number Theory, 168(??):374–385, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16300907.

Lombardo:2016:EOI

[1737] Davide Lombardo. An explicit open image theorem for products of elliptic curves. Journal of Number Theory, 168(??):386–412, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300956.

Beneish:2016:APC

[1738] Lea Beneish and Claire Frechette. p-Adic properties of coefficients of certain half-integral weight modular forms. Journal of Number Theory, 168(??):413–432, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16300804.

Kucuksakalli:2016:BPM

[1739] Ömer Küçüksakallı. Bivariate polynomial mappings associated with simple complex Lie algebras. Journal of Number Theory, 168(??):433–451, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16300993.

Medvedev:2016:MGF

[1740] Alice Medvedev and Ramin Takloo-Bighash. Multiplicative groups of fields and hereditarily irreducible polynomials. Journal of Number Theory, 168(??):452–471, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16300981.

Voutier:2016:FPJ

[1741] Paul Voutier. Families of periodic Jacobi–Perron algorithms for all period lengths. Journal of Number Theory, 168(??):472–486, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16300816.

Mehta:2016:APM

[1742] Jay Mehta, Biswajyoti Saha, and G. K. Viswanadham. Analytic properties of multiple zeta functions and certain weighted variants, an elementary approach. Journal of Number Theory, 168(??):487–508, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630107X.

Araci:2016:CSC

[1743] Serkan Araci and Mehmet Acikgoz. Corrigendum to “On the von Staudt–Clausen’s theorem related to q-Frobenius–Euler numbers” [J. Number Theory 159 (2016) 329–339]. Journal of Number Theory, 168 (??):509–510, November 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301068. See [1464].

Anonymous:2016:PD

[1744] Anonymous. Pages 1–440 (December 2016). Journal of Number Theory, 169(??):1–440, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2016:EBl

[1745] Anonymous. Editorial Board. Journal of Number Theory, 169(??):ifc, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301445.

Coppo:2016:SMS

[1746] Marc-Antoine Coppo and Paul Thomas Young. On shifted Mascheroni series and hyperharmonic numbers. Journal of Number Theory, 169(??): 1–20, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301172.

Mavraki:2016:IIW

[1747] Niki Myrto Mavraki. Impossible intersections in a Weierstrass family of elliptic curves. Journal of Number Theory, 169(??):21–40, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301196.

Herrero-Miranda:2016:CTJ

[1748] Sebastián Herrero-Miranda. A converse theorem for Jacobi–Maass forms and applications. Journal of Number Theory, 169(??):41–61, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301214.

Lin:2016:AZI

[1749] Bingchen Lin and Dongwen Liu. Archimedean zeta integrals on U(n, 1). Journal of Number Theory, 169(??):62–78, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301251.

Hou:2016:CNR

[1750] Qing-Hu Hou, Hai-Tao Jin, Yan-Ping Mu, and Li Zhang. Congruences on the number of restricted m-ary partitions. Journal of Number Theory, 169(??):79–85, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301317.

Hofmann:2016:VAR

[1751] Tommy Hofmann and Yinan Zhang. Valuations of p-adic regulators of cyclic cubic fields. Journal of Number Theory, 169(??):86–102, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301329.

Bandini:2016:FGT

[1752] Andrea Bandini and Laura Paladino. Fields generated by torsion points of elliptic curves. Journal of Number Theory, 169(??):103–133, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301202.

Xu:2016:AEG

[1753] Aimin Xu, Yongcai Hu, and Peipei Tang. Asymptotic expansions for the gamma function. Journal of Number Theory, 169(??):134–143, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301366.

Steuding:2016:LPN

[1754] Jörn Steuding and Marc Technau. The least prime number in a Beatty sequence. Journal of Number Theory, 169(??):144–159, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630138X.

Basic:2016:QVE

[1755] Bojan Bašić. On quotients of values of Euler’s function on the Catalan numbers. Journal of Number Theory, 169(??):160–173, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301391.

Fan:2016:DSC

[1756] Shilei Fan and Lingmin Liao. Dynamical structures of Chebyshev polynomials on Z_{2}. Journal of Number Theory, 169(??):174–182, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301305.

Baluyot:2016:PCC

[1757] Siegfred Alan C. Baluyot. On the pair correlation conjecture and the alternative hypothesis. Journal of Number Theory, 169(??):183–226, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301238.

Jeon:2016:ZLT

[1758] Daeyeol Jeon, Soon-Yi Kang, and Chang Heon Kim. Zagier-lift type arithmetic in harmonic weak Maass forms. Journal of Number Theory, 169(??):227–249, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301287.

Luca:2016:EBN

[1759] Florian Luca and Dimbinaina Ralaivaosaona. An explicit bound for the number of partitions into roots. Journal of Number Theory, 169 (??):250–264, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301330.

Heath-Brown:2016:APT

[1760] Roger Heath-Brown and Xiannan Li. Almost prime triples and Chen’s theorem. Journal of Number Theory, 169(??):265–294, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301226.

Tang:2016:QVJ

[1761] Hengcai Tang and Yingnan Wang. Quantitative versions of the joint distributions of Hecke eigenvalues. Journal of Number Theory, 169 (??):295–314, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301275.

Zhang:2016:FPM

[1762] Wenpeng Zhang. On the fourth power mean of the general Kloosterman sums. Journal of Number Theory, 169(??):315–326, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301342.

Hilberdink:2016:AVL

[1763] Titus Hilberdink and László Tóth. On the average value of the least common multiple of k positive integers. Journal of Number Theory, 169(??):327–341, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301408.

Friedli:2016:FRF

[1764] Fabien Friedli. A functional relation for L-functions of graphs equivalent to the Riemann Hypothesis for Dirichlet L-functions. Journal of Number Theory, 169(??):342–352, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1630141X.

Kim:2016:AROb

[1765] Sungjin Kim. Average results on the order of a modulo p. Journal of Number Theory, 169(??):353–368, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301354.

Ozcag:2016:RPI

[1766] Emin Özca.ğ and İnci Ege. Remarks on polygamma and incomplete gamma type functions. Journal of Number Theory, 169(??):369–387, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301378.

Boggess:2016:SVC

[1767] Brandon Boggess. Splitting varieties for cup products with Z/3coefficients. Journal of Number Theory, 169(??):388–405, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301263.

Ma:2016:AGC

[1768] Liming Ma, Chaoping Xing, and Sze Ling Yeo. On automorphism groups of cyclotomic function fields over finite fields. Journal of Number Theory, 169(??):406–419, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301421.

Chan:2016:TSQ

[1769] Song Heng Chan, Thi Phuong Nhi Ho, and Renrong Mao. Truncated series from the quintuple product identity. Journal of Number Theory, 169(??):420–438, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301299.

Faber:2016:CFD

[1770] Xander Faber and Bianca Thompson. Corrigendum to “On the field of definition of a cubic rational function and its critical points” [J. Number Theory 167 (2016) 1–6]. Journal of Number Theory, 169 (??):439–440, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301184. See [1684].

Anonymous:2017:PJa

[1771] Anonymous. Pages 1–408 (January 2017). Journal of Number Theory, 170(??):1–408, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBa

[1772] Anonymous. Editorial Board. Journal of Number Theory, 170(??): ifc, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301779.

Omland:2017:HMP

[1773] Tron Omland. How many Pythagorean triples with a given inradius? Journal of Number Theory, 170(??):1–2, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301603.

Somu:2017:DNR

[1774] Sai Teja Somu. On the distribution of numbers related to the divisors of x^{n} − 1. Journal of Number Theory, 170(??):3–9, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1630155X.

Cohen:2017:LSF

[1775] Stephen D. Cohen and Tim Trudgian. On the least square-free primitive root modulo p. Journal of Number Theory, 170(??):10–16, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301627.

Ballantine:2017:NCN

[1776] Cristina Ballantine and Mircea Merca. New convolutions for the number of divisors. Journal of Number Theory, 170(??):17–34, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301585.

Mollahajiaghaei:2017:ASU

[1777] Mohsen Mollahajiaghaei. On the addition of squares of units modulo n. Journal of Number Theory, 170(??):35–45, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301640.

Kishi:2017:IQF

[1778] Yasuhiro Kishi and Toru Komatsu. Imaginary quadratic fields whose ideal class groups have 3-rank at least three. Journal of Number Theory, 170(??):46–54, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301706.

Merca:2017:FLR

[1779] Mircea Merca. On families of linear recurrence relations for the special values of the Riemann zeta function. Journal of Number Theory, 170(??):55–65, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301664.

Dong:2017:ATT

[1780] Xiaoqian Dong and Qiang Wu. The absolute trace of totally positive reciprocal algebraic integers. Journal of Number Theory, 170(??):66– 74, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301676.

Stadnik:2017:MFG

[1781] M. E. Stadnik. A multiquadratic field generalization of Artin’s conjecture. Journal of Number Theory, 170(??):75–102, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301639.

Guignard:2017:CAP

[1782] Quentin Guignard. Counting algebraic points of bounded height on projective spaces. Journal of Number Theory, 170(??):103–141, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301597.

Kim:2017:SDF

[1783] Sun Kim. Sums of divisors functions and Bessel function series. Journal of Number Theory, 170(??):142–184, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630169X.

Liu:2017:NME

[1784] Sheng-Chi Liu. A note on mass equidistribution of holomorphic Siegel modular forms. Journal of Number Theory, 170(??):185–190, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630172X.

Lee:2017:HHF

[1785] Jungyun Lee, Byungheup Jun, and Hi joon Chae. Higher Hickerson formula. Journal of Number Theory, 170(??):191–210, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301548.

Offen:2017:PIA

[1786] Omer Offen. On parabolic induction associated with a p-adic symmetric space. Journal of Number Theory, 170(??):211–227, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301652.

Akhilesh:2017:DTM

[1787] P. Akhilesh. Double tails of multiple zeta values. Journal of Number Theory, 170(??):228–249, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16301718.

Xia:2017:MIF

[1788] Ernest X. W. Xia. More infinite families of congruences modulo 5 for broken 2-diamond partitions. Journal of Number Theory, 170(??):250– 262, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301020.

McCarthy:2017:MRF

[1789] Dermot McCarthy. Multiplicative relations for Fourier coefficients of degree 2 Siegel eigenforms. Journal of Number Theory, 170(??):263– 281, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301731.

Ruiz:2017:MDE

[1790] Carlos Alexis Gómez Ruiz and Florian Luca. Multiplicative Diophantine equations with factors from different Lucas sequences. Journal of Number Theory, 170(??):282–301, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16301573.

Lu:2017:NGA

[1791] Dawei Lu and Peixuan Zhang. A new general asymptotic formula and inequalities involving the volume of the unit ball. Journal of Number Theory, 170(??):302–314, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16301615.

Folsom:2017:SCQ

[1792] Amanda Folsom, Caleb Ki, Yen Nhi Truong Vu, and Bowen Yang. “Strange’’ combinatorial quantum modular forms. Journal of Number Theory, 170(??):315–346, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16301561.

Yang:2017:DII

[1793] Liyang Yang. Diophantine inequalities involving a prime and an almost-prime. Journal of Number Theory, 170(??):347–367, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1630124X.

Tripathi:2017:FFN

[1794] Amitabha Tripathi. Formulae for the Frobenius number in three variables. Journal of Number Theory, 170(??):368–389, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301743.

Bodin:2017:FPT

[1795] Arnaud Bodin, Pierre Dèbes, and Salah Najib. Families of polynomials and their specializations. Journal of Number Theory, 170(??):390– 408, January 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301755.

Anonymous:2017:PF

[1796] Anonymous. Pages 1–526 (February 2017). Journal of Number Theory, 171(??):1–526, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBb

[1797] Anonymous. Editorial Board. Journal of Number Theory, 171(??):ifc, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302293.

Xia:2017:APR

[1798] Ernest X. W. Xia and Olivia X. M. Yao. Arithmetic properties for (s, t)-regular bipartition functions. Journal of Number Theory, 171(??): 1–17, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301688.

Araci:2017:VI

[1799] Serkan Araci, Ugur Duran, and Mehmet Acikgoz. (ρ, q)-Volkenborn integration. Journal of Number Theory, 171(??):18–30, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16302025.

Lin:2017:CMC

[1800] Bernard L. S. Lin. Congruences modulo 27 for cubic partition pairs. Journal of Number Theory, 171(??):31–42, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301950.

Sebe:2017:NOS

[1801] Gabriela Ileana Sebe. A near-optimal solution to the Gauss–Kuzmin– Lévy problem for θ-expansions. Journal of Number Theory, 171(??): 43–55, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630186X.

Jiang:2017:SCF

[1802] Yujiao Jiang and Guangshi Lü. Sums of coefficients of L-functions and applications. Journal of Number Theory, 171(??):56–70, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301883.

Shparlinski:2017:CHP

[1803] Igor E. Shparlinski. On the convex hull of the points on multivariate modular hyperbolas. Journal of Number Theory, 171(??):71–78, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301949.

Li:2017:SPF

[1804] Zhonghua Li and Chen Qin. Shuffle product formulas of multiple zeta values. Journal of Number Theory, 171(??):79–111, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16301962.

Hu:2017:AAD

[1805] Su Hu and Min-Soo Kim. The p-adic analytic Dedekind sums. Journal of Number Theory, 171(??):112–127, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302050.

Bibak:2017:RLC

[1806] Khodakhast Bibak, Bruce M. Kapron, Venkatesh Srinivasan, Roberto Tauraso, and László Tóth. Restricted linear congruences. Journal of Number Theory, 171(??):128–144, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302013.

vanIttersum:2017:GIV

[1807] Jan-Willem M. van Ittersum. A group-invariant version of Lehmer’s conjecture on heights. Journal of Number Theory, 171(??):145–154, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301986.

Langlois:2017:VCS

[1808] Marie-Andrée B. Langlois. The valuative capacity of the set of sums of d-th powers. Journal of Number Theory, 171(??):155–168, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16301895.

Xiao:2017:HIM

[1809] Xuanxuan Xiao. Higher integral moments of automorphic L-functions in short intervals. Journal of Number Theory, 171(??):169–197, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301871.

Heidaryan:2017:GSI

[1810] Bahar Heidaryan, Matteo Longo, and Giulio Peruginelli. Galois structure on integral valued polynomials. Journal of Number Theory, 171 (??):198–212, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301901.

Gras:2017:NGR

[1811] Georges Gras. No general Riemann–Hurwitz formula for relative pclass groups. Journal of Number Theory, 171(??):213–226, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16301998.

Matthes:2017:EDZ

[1812] Nils Matthes. Elliptic double zeta values. Journal of Number Theory, 171(??):227–251, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301937.

Sun:2017:TKS

[1813] Chaochao Sun and Kejian Xu. On tame kernels and second regulators of number fields and their subfields. Journal of Number Theory, 171(??):252–274, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301925.

Sun:2017:X

[1814] Zhi-Wei Sun. On x(ax + 1) + y(by + 1) + z(cz + 1) and x(ax + b) + y(ay + c) + z(az + d). Journal of Number Theory, 171(??):275–283, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302074.

Hurlimann:2017:UZF

[1815] Werner Hürlimann. On universal zero-free ternary quadratic form representations of primes in arithmetic progressions. Journal of Number Theory, 171(??):284–300, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16301974.

Choi:2017:SEP

[1816] Dohoon Choi, Subong Lim, Tobias Mühlenbruch, and Wissam Raji. Series expansion of the period function and representations of Hecke operators. Journal of Number Theory, 171(??):301–340, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16302037.

Anbar:2017:MIV

[1817] Nurdagül Anbar, Alp Bassa, and Peter Beelen. A modular interpretation of various cubic towers. Journal of Number Theory, 171 (??):341–357, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302086. See corrigendum [1848].

Loquias:2017:RDS

[1818] Manuel Joseph C. Loquias, Mohamed Mkaouar, Klaus Scheicher, and Jörg M. Thuswaldner. Rational digit systems over finite fields and Christol’s Theorem. Journal of Number Theory, 171(??):358–390, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302049.

deShalit:2017:SCP

[1819] Ehud de Shalit and Eyal Z. Goren. Supersingular curves on Picard modular surfaces modulo an inert prime. Journal of Number Theory, 171(??):391–421, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16301846.

Petho:2017:MDA

[1820] Attila Pethő, Michael E. Pohst, and Csanád Bertók. On multidimensional Diophantine approximation of algebraic numbers. Journal of Number Theory, 171(??):422–448, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301858.

Vatwani:2017:BGB

[1821] Akshaa Vatwani. Bounded gaps between Gaussian primes. Journal of Number Theory, 171(??):449–473, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301913.

Taleb:2017:EMC

[1822] Reza Taleb. An Equivariant Main Conjecture in Iwasawa theory without the assumption μ = 0. Journal of Number Theory, 171(??):474– 494, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302001.

Goral:2017:AAH

[1823] Haydar Göral and Doğa Can Sertbaş. Almost all hyperharmonic numbers are not integers. Journal of Number Theory, 171(??):495–526, February 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302062.

Anonymous:2017:PMa

[1824] Anonymous. Pages 1–416 (March 2017). Journal of Number Theory, 172 (??):1–416, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBc

[1825] Anonymous. Editorial Board. Journal of Number Theory, 172(??): ifc, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302724.

Borda:2017:TDG

[1826] Bence Borda. On the theorem of Davenport and generalized Dedekind sums. Journal of Number Theory, 172(??):1–20, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302244.

Ferreira:2017:NPL

[1827] E. M. Ferreira, A. K. Kohara, and J. Sesma. New properties of the Lerch’s transcendent. Journal of Number Theory, 172(??):21– 31, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302219.

Xu:2017:SCH

[1828] Zhao Xu. Sign changes of Hecke eigenvalue of primitive cusp forms. Journal of Number Theory, 172(??):32–43, March 2017. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302190.

Ikeda:2017:FES

[1829] Tamotsu Ikeda. On the functional equation of the Siegel series. Journal of Number Theory, 172(??):44–62, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302104.

Roldan:2017:RTD

[1830] Enrico Varela Roldán. Representation theory of Drinfeld modular forms of level T. Journal of Number Theory, 172(??):63–92, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16302141.

Panzer:2017:PTM

[1831] Erik Panzer. The parity theorem for multiple polylogarithms. Journal of Number Theory, 172(??):93–113, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302128.

Amou:2017:AIC

[1832] Masaaki Amou and Keijo Väänänen. On algebraic independence of a class of infinite products. Journal of Number Theory, 172(??):114– 132, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302153.

Tonini:2017:TMR

[1833] Fabio Tonini. Trace map and regularity of finite extensions of a DVR. Journal of Number Theory, 172(??):133–144, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1630227X.

Chen:2017:SIA

[1834] Chao-Ping Chen. Sharp inequalities and asymptotic series related to Somos’ quadratic recurrence constant. Journal of Number Theory, 172(??):145–159, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302189.

Harrington:2017:NPC

[1835] Joshua Harrington, Lenny Jones, and Tristan Phillips. New primitive covering numbers and their properties. Journal of Number Theory, 172(??):160–177, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302256.

Sababheh:2017:ARY

[1836] M. Sababheh and M. S. Moslehian. Advanced refinements of Young and Heinz inequalities. Journal of Number Theory, 172(??):178–199, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302177.

Kataoka:2017:CGG

[1837] Takenori Kataoka. A consequence of Greenberg’s generalized conjecture on Iwasawa invariants of Z_{p}-extensions. Journal of Number Theory, 172(??):200–233, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302165.

Fernando:2017:RDP

[1838] Neranga Fernando. Reversed Dickson polynomials of the (k + 1)- th kind over finite fields. Journal of Number Theory, 172(??):234– 255, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302268.

Lampret:2017:WSE

[1839] Vito Lampret. Wallis’ sequence estimated accurately using an alternating series. Journal of Number Theory, 172(??):256–269, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16302220.

Burrin:2017:GDS

[1840] Claire Burrin. Generalized Dedekind sums and equidistribution mod 1. Journal of Number Theory, 172(??):270–286, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302116.

Ehrenborg:2017:VWC

[1841] Richard Ehrenborg, Likith Govindaiah, Peter S. Park, and Margaret Readdy. The van der Waerden complex. Journal of Number Theory, 172(??):287–300, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302207.

Noot:2017:ARW

[1842] Rutger Noot. The p-adic representation of the Weil–Deligne group associated to an abelian variety. Journal of Number Theory, 172(??):301– 320, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630213X.

Deng:2017:DCC

[1843] Calvin Deng. Davenport constant for commutative rings. Journal of Number Theory, 172(??):321–342, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302098.

Aymone:2017:PSB

[1844] M. Aymone and V. Sidoravicius. Partial sums of biased random multiplicative functions. Journal of Number Theory, 172(??):343–382, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302347.

James:2017:EPE

[1845] Kevin James and Paul Pollack. Extremal primes for elliptic curves with complex multiplication. Journal of Number Theory, 172(??):383– 391, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302670.

Ganguli:2017:RMC

[1846] Abhik Ganguli. On the reduction modulo p of certain modular padic Galois representations. Journal of Number Theory, 172(??):392– 412, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302372.

Vatne:2017:SMD

[1847] Jon Eivind Vatne. The sequence of middle divisors is unbounded. Journal of Number Theory, 172(??):413–415, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302232.

Anbar:2017:CMI

[1848] Nurdagül Anbar, Alp Bassa, and Peter Beelen. Corrigendum to “A modular interpretation of various cubic towers” [J. Number Theory 171 (2017) 341–357]. Journal of Number Theory, 172(??):416, March 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302682. See [1817].

Anonymous:2017:PAa

[1849] Anonymous. Pages 1–632 (April 2017). Journal of Number Theory, 173 (??):1–632, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBd

[1850] Anonymous. Editorial Board. Journal of Number Theory, 173(??): ifc, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16303031.

Savala:2017:CLE

[1851] Paul Savala. Computing the Laplace eigenvalue and level of Maass cusp forms. Journal of Number Theory, 173(??):1–22, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302499.

Chaubey:2017:HZP

[1852] Sneha Chaubey. Hutchinson–Zaharescu parabolas and Ford circles. Journal of Number Theory, 173(??):23–36, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302487.

Buium:2017:AET

[1853] Alexandru Buium and Emma Previato. Arithmetic Euler top. Journal of Number Theory, 173(??):37–63, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630258X.

Garunkstis:2017:DVS

[1854] Ramūnas Garunkštis and Raivydas Šimėnas. On the distribution of the a-values of the Selberg zeta-function associated to finite volume Riemann surfaces. Journal of Number Theory, 173(??):64–86, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302700.

Ismailescu:2017:PSL

[1855] Dan Ismailescu, Lenny Jones, and Tristan Phillips. Primefree shifted Lucas sequences of the second kind. Journal of Number Theory, 173(??): 87–99, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302505.

Kong:2017:SBW

[1856] Derong Kong, Wenxia Li, and Yuru Zou. On small bases which admit points with two expansions. Journal of Number Theory, 173(??): 100–128, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302463.

Gaal:2017:IBM

[1857] István Gaál and László Remete. Integral bases and monogenity of pure fields. Journal of Number Theory, 173(??):129–146, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302438.

Kida:2017:AFC

[1858] Masanari Kida and Norihiko Namura. On Artin L-functions of certain central extensions. Journal of Number Theory, 173(??):147– 169, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302657.

Hung:2017:NVC

[1859] Pin-Chi Hung. On the non-vanishing modℓ of central L-values with anticyclotomic twists for Hilbert modular forms. Journal of Number Theory, 173(??):170–209, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302475.

Kim:2017:NRSa

[1860] Kyoungmin Kim and Byeong-Kweon Oh. The number of representations of squares by integral ternary quadratic forms (II). Journal of Number Theory, 173(??):210–229, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16302426.

Leon-Cardenal:2017:EFL

[1861] Edwin León-Cardenal. An explicit formula for the local zeta function of a Laurent polynomial. Journal of Number Theory, 173(??):230– 242, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302530.

Bautista-Ancona:2017:IQF

[1862] Victor Bautista-Ancona, Javier Diaz-Vargas, and José Alejandro Lara Rodríguez. Imaginary quadratic function fields with ideal class group of prime exponent. Journal of Number Theory, 173(??):243–253, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302554.

Sander:2017:MAD

[1863] J. W. Sander. Multiplicative atom decomposition of sets of exceptional units in residue class rings. Journal of Number Theory, 173(??):254– 271, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1630261X.

Hamahata:2017:GDS

[1864] Yoshinori Hamahata. Generalized Dedekind sums and transformation formulae in function fields. Journal of Number Theory, 173(??):272– 292, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302517.

Balkanova:2017:NET

[1865] Olga Balkanova and Dmitry Frolenkov. New error term for the fourth moment of automorphic L-functions. Journal of Number Theory, 173(??): 293–303, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302578.

Hulse:2017:SMS

[1866] Thomas A. Hulse, Chan Ieong Kuan, David Lowry-Duda, and Alexander Walker. The second moment of sums of coefficients of cusp forms. Journal of Number Theory, 173(??):304–331, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302396.

Schindler:2017:CRP

[1867] Damaris Schindler. Counting rational points on hypersurfaces and higher order expansions. Journal of Number Theory, 173(??):332– 370, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302359.

Barbero:2017:GVD

[1868] Stefano Barbero. Generalized Vandermonde determinants and characterization of divisibility sequences. Journal of Number Theory, 173(??): 371–377, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302591.

Kathiravan:2017:SNC

[1869] T. Kathiravan and S. N. Fathima. Some new congruences for Andrews’ singular overpartitions. Journal of Number Theory, 173(??):378– 393, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302608.

Hulse:2017:SIA

[1870] Thomas A. Hulse, Chan Ieong Kuan, David Lowry-Duda, and Alexander Walker. Short-interval averages of sums of Fourier coefficients of cusp forms. Journal of Number Theory, 173(??):394–415, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302384.

Kucera:2017:TT

[1871] Radan Kučera. On a theorem of Thaine. Journal of Number Theory, 173(??):416–424, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302542.

Karnataki:2017:SDA

[1872] Aditya Karnataki. Self-dual Artin representations of dimension three. Journal of Number Theory, 173(??):425–447, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302402.

Lorenzo:2017:SBC

[1873] Elisa Lorenzo, Giulio Meleleo, Piermarco Milione, and Alina Bucur. Statistics for biquadratic covers of the projective line over finite fields. Journal of Number Theory, 173(??):448–477, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302414.

Byszewski:2017:SOM

[1874] Jakub Byszewski, Mikołaj Fraçzyk, and Anna Szumowicz. Simultaneous p-orderings and minimizing volumes in number fields. Journal of Number Theory, 173(??):478–511, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1630244X.

Mishiba:2017:AIC

[1875] Yoshinori Mishiba. On algebraic independence of certain multizeta values in characteristic p. Journal of Number Theory, 173(??):512– 528, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302529.

Benjamin:2017:SRQ

[1876] Elliot Benjamin. Some real quadratic number fields with their Hilbert 2-class field having cyclic 2-class group. Journal of Number Theory, 173(??):529–546, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302645. See corrigendum [2079].

Diagana:2017:SNI

[1877] Toka Diagana and Hamadoun Maïga. Some new identities and congruences for Fubini numbers. Journal of Number Theory, 173(??):547– 569, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302669.

Newman:2017:GTE

[1878] Burton Newman. Growth of torsion of elliptic curves with full 2torsion over quadratic cyclotomic fields. Journal of Number Theory, 173(??):570–601, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302694.

Hieu:2017:DDR

[1879] Do Duy Hieu and Van Thang Pham. Distinct distances on regular varieties over finite fields. Journal of Number Theory, 173(??):602– 613, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302852.

ElBachraoui:2017:SPR

[1880] Mohamed El Bachraoui. Short proofs for q-Raabe formula and integrals for Jacobi theta functions. Journal of Number Theory, 173(??): 614–620, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302621.

He:2017:SAB

[1881] Bing He. Supercongruences arising from basic hypergeometric series. Journal of Number Theory, 173(??):621–630, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302633.

Blagouchine:2017:CEG

[1882] Iaroslav V. Blagouchine. Corrigendum to “Expansions of generalized Euler’s constants into the series of polynomials in π^{−2} and into the formal enveloping series with rational coefficients only” [J. Number Theory 158 (2016) 365–396]. Journal of Number Theory, 173(??):631– 632, April 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302992. See [1443].

Anonymous:2017:PMb

[1883] Anonymous. Pages 1–524 (May 2017). Journal of Number Theory, 174 (??):1–524, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBe

[1884] Anonymous. Editorial Board. Journal of Number Theory, 174(??): ifc, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300276.

Wu:2017:HDS

[1885] Xi Wu, Li Yan, and Ting Zhong. How the dimension of some GCF_{ϵ} sets change with proper choice of the parameter function ϵ(k). Journal of Number Theory, 174(??):1–13, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X16302955.

Zhang:2017:SGI

[1886] Xin Zhang and Chun-Gang Ji. Sums of generators of ideals in residue class ring. Journal of Number Theory, 174(??):14–25, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16303018.

Wakabayashi:2017:IAM

[1887] Noriko Wakabayashi. Interpolation of q-analogue of multiple zeta and zeta-star values. Journal of Number Theory, 174(??):26–39, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1630289X.

Xu:2017:TTS

[1888] Ce Xu and Zhonghua Li. Tornheim type series and nonlinear Euler sums. Journal of Number Theory, 174(??):40–67, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302840.

deMedeiros:2017:CCM

[1889] Davi Lopes Alves de Medeiros and Lev Birbrair. Complete classification of 3-multisets up to combinatorial equivalence. Journal of Number Theory, 174(??):68–77, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302827.

Coppola:2017:FRE

[1890] Giovanni Coppola, M. Ram Murty, and Biswajyoti Saha. Finite Ramanujan expansions and shifted convolution sums of arithmetical functions. Journal of Number Theory, 174(??):78–92, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302931.

Feaver:2017:IMF

[1891] Amy Feaver. Imaginary multiquadratic fields of class number 1. Journal of Number Theory, 174(??):93–117, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1630292X.

Zhou:2017:CPD

[1892] Haiyan Zhou, Li-Ping Wang, and Weiqiong Wang. Counting polynomials with distinct zeros in finite fields. Journal of Number Theory, 174(??):118–135, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302906.

AlTahan:2017:EHA

[1893] M. Al Tahan and B. Davvaz. On the existence of hyperrings associated to arithmetic functions. Journal of Number Theory, 174(??):136– 149, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16303006.

Kleinbock:2017:AVM

[1894] Dmitry Kleinbock, Ronggang Shi, and George Tomanov. S-adic version of Minkowski’s geometry of numbers and Mahler’s compactness criterion. Journal of Number Theory, 174(??):150–163, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X16302980.

Schmidt:2017:CNR

[1895] Claus Günther Schmidt. The critical numbers of Rankin–Selberg convolutions of cohomological representations. Journal of Number Theory, 174(??):164–188, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302864.

Gonek:2017:DPN

[1896] Steven M. Gonek and Aleksandar Ivić. On the distribution of positive and negative values of Hardy’s Z-function. Journal of Number Theory, 174(??):189–201, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302979.

Gillibert:2017:LGD

[1897] Florence Gillibert and Gabriele Ranieri. On the local–global divisibility of torsion points on elliptic curves and GL_{2}-type varieties. Journal of Number Theory, 174(??):202–220, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302943.

Balasubramanian:2017:CCM

[1898] R. Balasubramanian, Sumit Giri, and Priyamvad Srivastav. On correlations of certain multiplicative functions. Journal of Number Theory, 174(??):221–238, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302839.

Tengely:2017:CDE

[1899] Sz. Tengely and M. Ulas. On certain Diophantine equations of the form z^{2} = f(x)^{2} ± g(y)^{2}. Journal of Number Theory, 174(??):239– 257, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302967.

Nordqvist:2017:CRP

[1900] Jonas Nordqvist. Characterization of 2-ramified power series. Journal of Number Theory, 174(??):258–273, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302876. See

corrigendum [2012].

Kuhn:2017:MPR

[1901] Patrick Kühn, Nicolas Robles, and Dirk Zeindler. On a mollifier of the perturbed Riemann zeta-function. Journal of Number Theory, 174(??):274–321, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X16302566.

DelCorso:2017:WEA

[1902] Ilaria Del Corso, Roberto Dvornicich, and Maurizio Monge. On wild extensions of a p-adic field. Journal of Number Theory, 174(??):322– 342, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302888.

Bamunoba:2017:NCW

[1903] Alex Samuel Bamunoba. A note on Carlitz Wieferich primes. Journal of Number Theory, 174(??):343–357, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16303134.

Guo:2017:RVT

[1904] Victor J. W. Guo, Hao Pan, and Yong Zhang. The Rodriguez–Villegas type congruences for truncated q-hypergeometric functions. Journal of Number Theory, 174(??):358–368, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16302451.

Liu:2017:CTP

[1905] Weibin Liu and Bing Li. Chaotic and topological properties of continued fractions. Journal of Number Theory, 174(??):369–383, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X16303158.

Filaseta:2017:NPP

[1906] Michael Filaseta and Maria Markovich. Newton polygons and the Prouhet–Tarry–Escott problem. Journal of Number Theory, 174(??): 384–400, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302918.

Mao:2017:DIT

[1907] Tianyi Mao. The distribution of integers in a totally real cubic field. Journal of Number Theory, 174(??):401–418, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300136.

Zhang:2017:NFP

[1908] Wenpeng Zhang and Shimeng Shen. A note on the fourth power mean of the generalized Kloosterman sums. Journal of Number Theory, 174(??):419–426, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300185.

Lee:2017:TQM

[1909] Seok Hyeong Lee and Seungjai Lee. On the twisted quadratic moment for Dirichlet L-functions. Journal of Number Theory, 174(??):

427–435, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300082.

Alaa:2017:HRQ

[1910] Mohamed Alaa and Mohammad Sadek. High rank quadratic twists of pairs of elliptic curves. Journal of Number Theory, 174(??):436– 444, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300203.

Leonetti:2017:ULD

[1911] Paolo Leonetti and Salvatore Tringali. Upper and lower densities have the strong Darboux property. Journal of Number Theory, 174(??):445– 455, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300033.

Kisielewski:2017:CRM

[1912] Tomasz Kisielewski. Criteria for regularity of Mahler power series and Becker’s conjecture. Journal of Number Theory, 174(??):456– 486, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300021.

You:2017:AMG

[1913] Xu You. Approximation of the median of the gamma distribution. Journal of Number Theory, 174(??):487–493, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300173.

Brazelton:2017:CPR

[1914] Thomas Brazelton, Joshua Harrington, Siddarth Kannan, and Matthew Litman. On consecutive primitive n-th roots of unity modulo q. Journal of Number Theory, 174(??):494–504, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730015X.

Xu:2017:MAE

[1915] Zhihang Xu and Weiping Wang. More asymptotic expansions for the Barnes G-function. Journal of Number Theory, 174(??):505– 517, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300057.

Kohonen:2017:ILB

[1916] Jukka Kohonen. An improved lower bound for finite additive 2-bases. Journal of Number Theory, 174(??):518–524, May 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300094.

Anonymous:2017:PJb

[1917] Anonymous. Pages 1–264 (June 2017). Journal of Number Theory, 175 (??):1–264, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBf

[1918] Anonymous. Editorial Board. Journal of Number Theory, 175(??): ifc, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300586.

Kobayashi:2017:CNR

[1919] Hirotomo Kobayashi. Class number relations between pure fields and their Galois closures. Journal of Number Theory, 175(??):1– 20, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300100.

Hou:2017:HBR

[1920] Fei Hou and Meng Zhang. Hybrid bounds for Rankin–Selberg Lfunctions. Journal of Number Theory, 175(??):21–41, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300070.

Li:2017:MTIa

[1921] Yan Li and Daeyeoul Kim. A Menon-type identity with many tuples of group of units in residually finite Dedekind domains. Journal of Number Theory, 175(??):42–50, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300215.

Wang:2017:NCP

[1922] Liuquan Wang. New congruences for partitions related to mock theta functions. Journal of Number Theory, 175(??):51–65, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300161.

Wu:2017:CPH

[1923] Bing-Ling Wu and Yong-Gao Chen. On certain properties of harmonic numbers. Journal of Number Theory, 175(??):66–86, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300252.

Erdelyi:2017:DDC

[1924] Márton Erdélyi. The distribution and density of cyclic groups of the reductions of an elliptic curve over a function field. Journal of Number Theory, 175(??):87–99, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300239.

Lu:2017:SQC

[1925] Dawei Lu, Lixin Song, and Yang Yu. Some quicker continued fraction approximations and inequalities towards Euler’s constant. Journal of Number Theory, 175(??):100–116, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17300227.

Choi:2017:ODP

[1926] Youn-Seo Choi, Byungchan Kim, and Jeremy Lovejoy. Overpartitions into distinct parts without short sequences. Journal of Number Theory, 175(??):117–133, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300112.

Berrizbeitia:2017:QFR

[1927] Pedro Berrizbeitia, Robin Chapman, Florian Luca, and Alberto Mendoza. Quadratic forms representing p-th terms of Lucas sequences. Journal of Number Theory, 175(??):134–139, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300197.

Amdeberhan:2017:TTB

[1928] Tewodros Amdeberhan and Roberto Tauraso. Two triple binomial sum supercongruences. Journal of Number Theory, 175(??):140–157, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300124.

Dupuy:2017:EGB

[1929] Taylor Dupuy. Examples of geometric Bombieri–Lang–Noguchi outside Mordell–Lang: Nonrigid varieties with ample but not globally generated cotangent bundle. Journal of Number Theory, 175(??):158– 166, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16303146.

Sun:2017:RLF

[1930] Zhi-Wei Sun. Refining Lagrange’s four-square theorem. Journal of Number Theory, 175(??):167–190, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17300069.

Hong:2017:NLB

[1931] Shaofang Hong and Guoyou Qian. New lower bounds for the least common multiple of polynomial sequences. Journal of Number Theory, 175(??):191–199, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300240.

Chung:2017:DTP

[1932] Chan-Liang Chung and Yao Lin Ong. Decomposition theorems produced from differential operators. Journal of Number Theory, 175(??): 200–222, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300148.

Ghenciu:2017:HDS

[1933] A. Ghenciu, S. Munday, and M. Roy. The Hausdorff dimension spectrum of conformal graph directed Markov systems and applications to nearest integer continued fractions. Journal of Number Theory, 175(??): 223–249, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16302360.

Fili:2017:HBA

[1934] Paul Fili and Igor Pritsker. Height bounds for algebraic numbers satisfying splitting conditions. Journal of Number Theory, 175(??):250– 264, June 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300045.

Anonymous:2017:PJc

[1935] Anonymous. Pages 1–540 (July 2017). Journal of Number Theory, 176 (??):1–540, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBg

[1936] Anonymous. Editorial Board. Journal of Number Theory, 176(??): ifc, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300963.

Merca:2017:NRN

[1937] Mircea Merca. New relations for the number of partitions with distinct even parts. Journal of Number Theory, 176(??):1–12, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300483.

Balasubramanyam:2017:ADA

[1938] Baskar Balasubramanyam. p-Adic distributions attached to twisted tensor L-functions for GL_{2} over CM fields. Journal of Number Theory, 176(??):13–36, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300537.

Konig:2017:RBT

[1939] Joachim König. On the reducibility behavior of Thue polynomials. Journal of Number Theory, 176(??):37–45, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300550.

Bartoli:2017:CPP

[1940] D. Bartoli, M. Giulietti, L. Quoos, and G. Zini. Complete permutation polynomials from exceptional polynomials. Journal of Number Theory, 176(??):46–66, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300495.

Dunn:2017:DMO

[1941] Alexander Dunn. On the distribution of αp^{γ} + β modulo one. Journal of Number Theory, 176(??):67–75, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300501.

Minac:2017:CGE

[1942] Ján Mináč and Nguy˜ ên Duy Tân. Counting Galois U_{4}(F_{p})-extensions using Massey products. Journal of Number Theory, 176(??):76– 112, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730001X.

Choi:2017:LUL

[1943] Suh Hyun Choi. Local universal lifting spaces of modl Galois representations. Journal of Number Theory, 176(??):113–148, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300392.

Agoh:2017:SRR

[1944] Takashi Agoh. Shortened recurrence relations for generalized Bernoulli numbers and polynomials. Journal of Number Theory, 176(??):149– 173, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300471.

Moscariello:2017:MRD

[1945] Alessio Moscariello. Minimal relations and the Diophantine Frobenius problem in embedding dimension three. Journal of Number Theory, 176(??):174–183, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300458.

Kunisada:2017:DMA

[1946] Ryoichi Kunisada. Density measures and additive property. Journal of Number Theory, 176(??):184–203, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730046X.

Lu:2017:RIP

[1947] Xiaodong Lü and Yifei Ren. Representation of an integer as a prime plus a product of two small factors. Journal of Number Theory, 176(??): 204–210, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300410.

Zhang:2017:TQF

[1948] Deyu Zhang and Yingnan Wang. Ternary quadratic form with prime variables attached to Fourier coefficients of primitive holomorphic cusp form. Journal of Number Theory, 176(??):211–225, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300513.

Berkovich:2017:VTN

[1949] Alexander Berkovich and Ali K. Uncu. Variation on a theme of Nathan Fine. New weighted partition identities. Journal of Number Theory, 176(??):226–248, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300446.

Larsson:2017:EHL

[1950] Daniel Larsson. Equivariant hom-Lie algebras and twisted derivations on (arithmetic) schemes. Journal of Number Theory, 176(??):249– 278, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300434.

Sakugawa:2017:FEP

[1951] Kenji Sakugawa and Shin ichiro Seki. Finite and étale polylogarithms. Journal of Number Theory, 176(??):279–301, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300379.

ODorney:2017:RIC

[1952] Evan O’Dorney. On a remarkable identity in class numbers of cubic rings. Journal of Number Theory, 176(??):302–332, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300355.

Aoki:2017:IFP

[1953] Miho Aoki and Yasuhiro Kishi. An infinite family of pairs of imaginary quadratic fields with both class numbers divisible by five. Journal of Number Theory, 176(??):333–343, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300409.

Kumchev:2017:SPA

[1954] Angel Kumchev and Huafeng Liu. On sums of powers of almost equal primes. Journal of Number Theory, 176(??):344–364, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300367.

Wang:2017:ZSS

[1955] Chunlin Wang and Kevin Zhao. On zero-sum subsequences of length not exceeding a given number. Journal of Number Theory, 176(??): 365–374, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300525.

Karpenkov:2017:GCM

[1956] Oleg Karpenkov and Alexey Ustinov. Geometry and combinatoric of Minkowski–Voronoi 3-dimensional continued fractions. Journal of Number Theory, 176(??):375–419, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17300380.

Pellarin:2017:NCR

[1957] F. Pellarin. A note on certain representations in characteristic p and associated functions. Journal of Number Theory, 176(??):420– 438, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300343.

Liu:2017:GLT

[1958] Zhixin Liu. Goldbach–Linnik type problems with unequal powers of primes. Journal of Number Theory, 176(??):439–448, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300422.

Kiuchi:2017:SAGa

[1959] Isao Kiuchi. Sums of averages of gcd-sum functions. Journal of Number Theory, 176(??):449–472, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300549.

Hiranouchi:2017:HMT

[1960] Toshiro Hiranouchi. A Hermite–Minkowski type theorem of varieties over finite fields. Journal of Number Theory, 176(??):473– 499, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300562.

Folsom:2017:PMM

[1961] Amanda Folsom. Perspectives on mock modular forms. Journal of Number Theory, 176(??):500–540, July 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17300653.

Anonymous:2017:PAb

[1962] Anonymous. Pages 1–588 (August 2017). Journal of Number Theory, 177 (??):1–588, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBh

[1963] Anonymous. Editorial Board. Journal of Number Theory, 177(??): ifc, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301063.

Hindes:2017:IEO

[1964] Wade Hindes. Integrality estimates in orbits over function fields. Journal of Number Theory, 177(??):1–19, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300896.

Komatsu:2017:EAV

[1965] Takao Komatsu and Paul Thomas Young. Exact p-adic valuations of Stirling numbers of the first kind. Journal of Number Theory, 177(??): 20–27, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300884.

Luo:2017:SZS

[1966] Sammy Luo. Short zero-sum sequences over abelian p-groups of large exponent. Journal of Number Theory, 177(??):28–36, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17300860.

Guo:2017:FAS

[1967] Victor J. W. Guo and Qiang-Qiang Jiang. Factors of alternating sums of powers of q-Narayana numbers. Journal of Number Theory, 177(??):37–42, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300744.

Dubickas:2017:RAT

[1968] Artūras Dubickas. On rational approximations to two irrational numbers. Journal of Number Theory, 177(??):43–59, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300914.

Louboutin:2017:UBT

[1969] Stéphane R. Louboutin. Upper bounds on L(1, χ) taking into account ramified prime ideals. Journal of Number Theory, 177(??):60– 72, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300756.

Hu:2017:AAK

[1970] Su Hu, Daeyeoul Kim, and Min-Soo Kim. The p-adic Arakawa–Kaneko– Hamahata zeta functions and poly-Euler polynomials. Journal of Number Theory, 177(??):73–90, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300823.

Wang:2017:DME

[1971] Ruyun Wang, Binbin Wen, and Wei Cao. Degree matrices and enumeration of rational points of some hypersurfaces over finite fields. Journal of Number Theory, 177(??):91–99, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300938.

Bordelles:2017:SIR

[1972] Olivier Bordellès. Short interval results for certain prime-independent multiplicative functions. Journal of Number Theory, 177(??):100–111, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300793.

Hulse:2017:SCC

[1973] Thomas A. Hulse, Chan Ieong Kuan, David Lowry-Duda, and Alexander Walker. Sign changes of coefficients and sums of coefficients of L-functions. Journal of Number Theory, 177(??):112–135, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17300720.

Mezo:2017:DPB

[1974] István Mező and José L. Ramírez. Divisibility properties of the r-Bell numbers and polynomials. Journal of Number Theory, 177(??):136– 152, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300872.

Terakado:2017:DDC

[1975] Yasuhiro Terakado. The determinant of a double covering of the projective space of even dimension and the discriminant of the branch locus. Journal of Number Theory, 177(??):153–169, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300859.

Wu:2017:SHR

[1976] Qiang Wu and Zhuo Zhang. On the smallest houses of reciprocal algebraic integers. Journal of Number Theory, 177(??):170–180, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730077X.

Bekki:2017:PGC

[1977] Hohto Bekki. On periodicity of geodesic continued fractions. Journal of Number Theory, 177(??):181–210, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300689.

Godinho:2017:PDF

[1978] Hemar Godinho and Luciana Ventura. Pairs of diagonal forms of degree 3^{τ} · 2 and Artin’s conjecture. Journal of Number Theory, 177(??):211– 247, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17300707.

Cao:2017:RLF

[1979] Chunyun Cao and Yuanhong Chen. The run-length function of the β-expansion of the unit. Journal of Number Theory, 177(??):248– 262, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730080X.

Jurgens:2017:SCP

[1980] Michael Jürgens and Marc C. Zimmermann. Steinitz classes and partial genera of unimodular lattices over imaginary-quadratic fields. Journal of Number Theory, 177(??):263–284, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17300690.

Ahmed:2017:RNP

[1981] Suman Ahmed, Chandrakant Aribam, and Sudhanshu Shekhar. Root numbers and parity of local Iwasawa invariants. Journal of Number Theory, 177(??):285–306, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300847.

Taous:2017:PGI

[1982] Mohammed Taous and Abdelkader Zekhnini. Pólya groups of the imaginary bicyclic biquadratic number fields. Journal of Number Theory, 177(??):307–327, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300902.

Klys:2017:RPC

[1983] Jack Klys. Reflection principles for class groups. Journal of Number Theory, 177(??):328–352, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17300732.

Choi:2017:ESI

[1984] SoYoung Choi and Chang Heon Kim. Eichler–Shimura isomorphism in higher level cases and its applications. Journal of Number Theory, 177(??):353–380, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300677.

Savchev:2017:IPA

[1985] Svetoslav Savchev and Fang Chen. An inverse problem about minimal zero-sum sequences over finite cyclic groups. Journal of Number Theory, 177(??):381–427, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300665.

Vehkalahti:2017:BAD

[1986] Eeva Vehkalahti. On the binary additive divisor problem in mean. Journal of Number Theory, 177(??):428–442, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300811.

Xu:2017:MZV

[1987] Ce Xu. Multiple zeta values and Euler sums. Journal of Number Theory, 177(??):443–478, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300835.

Chen:2017:SFM

[1988] Kwang-Wu Chen, Chan-Liang Chung, and Minking Eie. Sum formulas of multiple zeta values with arguments multiples of a common positive integer. Journal of Number Theory, 177(??):479–496, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1730094X.

Kim:2017:SNI

[1989] Byeong Moon Kim and Ji Young Kim. Sums of nonvanishing integral squares in real quadratic fields. Journal of Number Theory, 177(??):497–515, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300719.

Kumar:2017:AMS

[1990] Arvind Kumar. The adjoint map of the Serre derivative and special values of shifted Dirichlet series. Journal of Number Theory, 177(??):516–527, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17300768.

Meisner:2017:DPC

[1991] Patrick Meisner. Distribution of points on cyclic curves over finite fields. Journal of Number Theory, 177(??):528–561, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300781.

Azizi:2017:SCG

[1992] A. Azizi, I. Jerrari, A. Zekhnini, and M. Talbi. On the second 2-class group Gal(K_{2}^{(2)}/K) of some imaginary quartic cyclic number field K.

Journal of Number Theory, 177(??):562–588, August 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17300926.

Anonymous:2017:PAAa

[1993] Anonymous. Pages A1–A2, 1–208 (September 2017). Journal of Number Theory, 178(??):A1–, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBi

[1994] Anonymous. Editorial Board. Journal of Number Theory, 178(??):ifc, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301427.

Stair:2017:CMD

[1995] Kristin Stair, Simone Munao, and Chris Greenwell. Condolence message on the death of Editor-in-Chief David Goss. Journal of Number Theory, 178(??):A1, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301324.

Creutz:2017:ZCD

[1996] Brendan Creutz. Zero-cycles of degree one on Skorobogatov’s bielliptic surface. Journal of Number Theory, 178(??):1–4, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1730118X.

Yegan:2017:I

[1997] Mohammad Reza Yegan. On the irrationality of π^{4} and π^{6}. Journal of Number Theory, 178(??):5–10, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301208.

Girstmair:2017:VDS

[1998] Kurt Girstmair. On the values of Dedekind sums. Journal of Number Theory, 178(??):11–18, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17301257.

Chu:2017:MQB

[1999] Wenchang Chu. Moments on quadratic binomial products. Journal of Number Theory, 178(??):19–30, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301166.

Gong:2017:SCS

[2000] K. Gong, C. Jia, and M. A. Korolev. Shifted character sums with multiplicative coefficients, II. Journal of Number Theory, 178(??):31–39, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301178.

Nguyen:2017:IQF

[2001] Viet Kh. Nguyen. Imaginary quadratic fields of class number 2 and Levy–Hendy quadratics. Journal of Number Theory, 178(??):40–46, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301269. See addendum [2139].

Wang:2017:PNC

[2002] Wen Wang. The property of a new class of symmetric functions with applications. Journal of Number Theory, 178(??):47–59, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17301130.

Jia:2017:WDP

[2003] Lirui Jia and Wenguang Zhai. A weighted divisor problem. Journal of Number Theory, 178(??):60–93, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301154.

Alkan:2017:TSS

[2004] Emre Alkan and Haydar Göral. Trigonometric series and special values of L-functions. Journal of Number Theory, 178(??):94–117, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301270.

Patankar:2017:DGR

[2005] Vijay M. Patankar and C. S. Rajan. Distinguishing Galois representations by their normalized traces. Journal of Number Theory, 178 (??):118–125, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1730121X.

Wong:2017:LRC

[2006] Peng-Jie Wong. Langlands reciprocity for certain Galois extensions. Journal of Number Theory, 178(??):126–145, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17301142.

Hirschhorn:2017:PP

[2007] Michael D. Hirschhorn. Partitions and powers of 13. Journal of Number Theory, 178(??):146–157, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301294.

Al-Saedi:2017:UPO

[2008] Ali H. Al-Saedi. Using periodicity to obtain partition congruences. Journal of Number Theory, 178(??):158–178, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17301233.

Zhao:2017:TPA

[2009] Lilu Zhao. On ternary problems in additive prime number theory. Journal of Number Theory, 178(??):179–189, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17301191.

Basson:2017:PFH

[2010] Dirk Basson. A product formula for the higher rank Drinfeld discriminant function. Journal of Number Theory, 178(??):190–200, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301221.

Haynes:2017:RVL

[2011] Alan Haynes and Henna Koivusalo. A randomized version of the Littlewood Conjecture. Journal of Number Theory, 178(??):201–207, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301300.

Nordqvist:2017:CCR

[2012] Jonas Nordqvist. Corrigendum to “Characterization of 2-ramified power series” [J. Number Theory 174 (2017) 258–273]. Journal of Number Theory, 178(??):208, September 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17301282. See [1900].

Anonymous:2017:PO

[2013] Anonymous. Pages 1–298 (October 2017). Journal of Number Theory, 179(??):1–298, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBj

[2014] Anonymous. Editorial Board. Journal of Number Theory, 179(??): ifc, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301774.

Munao:2017:ODG

[2015] Simone Munao, Ahmad El-Guindy, Ambrus Pal, and Federico Pellarin. Obituary of David Goss. Journal of Number Theory, 179(??): 1–2, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301737.

Hanson:2017:LNP

[2016] Brandon Hanson, Robert C. Vaughan, and Ruixiang Zhang. The least number with prescribed Legendre symbols and representation by binary quadratic forms of small discriminant. Journal of Number Theory, 179(??):3–16, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1730152X.

Evans:2017:SMC

[2017] Ron Evans. Some mixed character sum identities of Katz. Journal of Number Theory, 179(??):17–32, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301609.

Franze:2017:FMI

[2018] C. S. Franze. A family of multiple integrals connected with relatives of the Dickman function. Journal of Number Theory, 179(??):33–49, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301567.

Ye:2017:RIC

[2019] Dongxi Ye. Representations of integers by certain 2k-ary quadratic forms. Journal of Number Theory, 179(??):50–64, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17301634.

Hbaib:2017:CSP

[2020] M. Hbaib and Y. Laabidi. Computation of L_{⊙}(β) for some Pisot series in F_{q}((x^{−1})). Journal of Number Theory, 179(??):65–76, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301531.

Zhang:2017:MSV

[2021] Wenpeng Zhang and Ran Duan. On the mean square value of L-functions with the weight of quadratic Gauss sums. Journal of Number Theory, 179(??):77–87, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301671.

Mao:2017:PTC

[2022] Guo-Shuai Mao. Proof of two conjectural supercongruences involving Catalan–Larcombe–French numbers. Journal of Number Theory, 179(??):88–96, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301658.

Li:2017:MTIb

[2023] Yan Li and Daeyeoul Kim. Menon-type identities derived from actions of subgroups of general linear groups. Journal of Number Theory, 179(??):97–112, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301646. See corrigendum [2141].

Das:2017:JFD

[2024] Soumya Das and Ritwik Pal. Jacobi forms and differential operators: Odd weights. Journal of Number Theory, 179(??):113–125, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301695.

Kellner:2017:PCP

[2025] Bernd C. Kellner. On a product of certain primes. Journal of Number Theory, 179(??):126–141, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17301683.

Goodson:2017:CHP

[2026] Heidi Goodson. A complete hypergeometric point count formula for Dwork hypersurfaces. Journal of Number Theory, 179(??):142–171, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730166X.

Bertok:2017:DPB

[2027] Csanád Bertók, Lajos Hajdu, and Attila Pethő. On the distribution of polynomials with bounded height. Journal of Number Theory, 179(??):172–184, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301610.

Galateau:2017:HTM

[2028] Aurélien Galateau and Amílcar Pacheco. Hauteur et torsion des modules de Drinfeld de rang 2. (French) [Height and torsion of Drinfeld modules of rank 2]. Journal of Number Theory, 179(??):185–219, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301579.

Leppala:2017:RAE

[2029] Kalle Leppälä, Tapani Matala-aho, and Topi Törmä. Rational approximations of the exponential function at rational points. Journal of Number Theory, 179(??):220–239, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17301622.

Byeon:2017:STR

[2030] Dongho Byeon and Keunyoung Jeong. Sums of two rational cubes with many prime factors. Journal of Number Theory, 179(??):240– 255, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301592.

Shen:2017:SIM

[2031] Zhongyan Shen and Lirui Jia. Some identities for multiple Hurwitz zeta values. Journal of Number Theory, 179(??):256–267, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301518.

Batzaya:2017:SAP

[2032] Gantsooj Batzaya. On simultaneous approximation to powers of a real number by rational numbers II. Journal of Number Theory, 179 (??):268–298, October 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301701.

Anonymous:2017:PN

[2033] Anonymous. Pages 1–792 (November 2017). Journal of Number Theory, 180(??):1–792, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBk

[2034] Anonymous. Editorial Board. Journal of Number Theory, 180(??):ifc, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302457.

Lu:2017:NPP

[2035] Hengfei Lu. A new proof to the period problems of GL(2 ). Journal of Number Theory, 180(??):1–25, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301580.

Bitan:2017:CQF

[2036] Rony A. Bitan. On the classification of quadratic forms over an integral domain of a global function field. Journal of Number Theory, 180(??):26–44, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301555.

Gassert:2017:SSC

[2037] T. Alden Gassert and Caleb McKinley Shor. On Sylvester sums of compound sequence semigroup complements. Journal of Number Theory, 180(??):45–72, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301750.

Yang:2017:RFF

[2038] Quan-Hui Yang and Min Tang. Representation functions on finite sets with extreme symmetric differences. Journal of Number Theory, 180(??):73–85, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301749.

Jacobs:2017:ECM

[2039] Kenneth Jacobs. Equidistribution of the crucial measures in non– Archimedean dynamics. Journal of Number Theory, 180(??):86–138, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301312.

Medina:2017:PAV

[2040] Luis A. Medina, Victor H. Moll, and Eric Rowland. Periodicity in the p-adic valuation of a polynomial. Journal of Number Theory, 180 (??):139–153, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301543.

Kerr:2017:C

[2041] Bryce Kerr. On the congruence x_{1}x_{2} ≡ x_{3}x_{4} mod q. Journal of Number Theory, 180(??):154–168, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17301245.

Leitner:2017:LEM

[2042] Dominik Leitner. Linear equations over multiplicative groups in positive characteristic II. Journal of Number Theory, 180(??):169–194, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301877.

Sanna:2017:DIV

[2043] Carlo Sanna. Distribution of integral values for the ratio of two linear recurrences. Journal of Number Theory, 180(??):195–207, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302019.

Das:2017:IIT

[2044] Pradeep Das, Abhishek Juyal, and Dustin Moody. Integral isosceles triangle–parallelogram and Heron triangle–rhombus pairs with a common area and common perimeter. Journal of Number Theory, 180 (??):208–218, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301956.

Ju:2017:GCR

[2045] Jangwon Ju and Byeong-Kweon Oh. Genus-correspondences respecting spinor genus. Journal of Number Theory, 180(??):219–233, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17301865.

Wang:2017:DAO

[2046] Yuchao Wang and Weili Yao. Diophantine approximation with one prime and three squares of primes. Journal of Number Theory, 180 (??):234–250, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301993.

Huang:2017:UAG

[2047] Hau-Wen Huang and Wen-Ching Winnie Li. A unified approach to the Galois closure problem. Journal of Number Theory, 180(??):251–279, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730197X.

Molina:2017:ESI

[2048] Santiago Molina. Eichler–Shimura isomorphism and group cohomology on arithmetic groups. Journal of Number Theory, 180(??):280–296, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301932.

Takeda:2017:VLP

[2049] Wataru Takeda. Visible lattice points and the Extended Lindelöf Hypothesis. Journal of Number Theory, 180(??):297–309, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17301889.

Kiuchi:2017:SAGb

[2050] Isao Kiuchi. Sums of averages of generalized Ramanujan sums. Journal of Number Theory, 180(??):310–348, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17301853.

Walling:2017:SRF

[2051] Lynne H. Walling. Some relations on Fourier coefficients of degree 2 Siegel forms of arbitrary level. Journal of Number Theory, 180 (??):349–359, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301890.

Xu:2017:AEP

[2052] Aimin Xu and Zhongdi Cen. Asymptotic expansions for the psi function and the Euler–Mascheroni constant. Journal of Number Theory, 180(??):360–372, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302007.

Grishkov:2017:LMA

[2053] A. Grishkov and D. Logachev. Lattice map for Anderson t-motives: first approach. Journal of Number Theory, 180(??):373–402, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17301907.

Niu:2017:P

[2054] Chuanze Niu and Wenxin Liu. On the products (1^{3} + q^{3})(2^{3} + q^{3}) ⋯ (n^{3} + q^{3}). Journal of Number Theory, 180(??):403–409, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17301981.

Chen:2017:ACS

[2055] Yong-Gao Chen and Jin-Hui Fang. Additive complements of the squares. Journal of Number Theory, 180(??):410–422, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302020.

Bettin:2017:NVC

[2056] Sandro Bettin and Bruno Martin. On the non-vanishing of certain Dirichlet series. Journal of Number Theory, 180(??):423–442, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301920.

Harrington:2017:DBE

[2057] Joshua Harrington and Lenny Jones. Differences between elements of the same order in a finite field. Journal of Number Theory, 180 (??):443–459, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301944.

Zheng:2017:DHT

[2058] Zhiyong Zheng. Davenport–Hasse’s theorem for polynomial Gauss sums over finite fields. Journal of Number Theory, 180(??):460–473, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17301919.

Candelori:2017:GWW

[2059] Luca Candelori, Cameron Franc, and Gene S. Kopp. Generating weights for the Weil representation attached to an even order cyclic quadratic module. Journal of Number Theory, 180(??):474–497, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302032.

Kuznetsov:2017:DSF

[2060] Alexey Kuznetsov. On Dirichlet series and functional equations. Journal of Number Theory, 180(??):498–511, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730224X.

Yu:2017:GTF

[2061] Han Yu. On generalized trigonometric functions and series of rational functions. Journal of Number Theory, 180(??):512–532, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302172.

DaSilva:2017:RCA

[2062] Stéphane Blondeau Da Silva. On randomly chosen arrangements of q + 1 lines with different slopes in F_{q}^{2}. Journal of Number Theory, 180(??):533–543, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302159.

Sesma:2017:RHN

[2063] J. Sesma. The Roman harmonic numbers revisited. Journal of Number Theory, 180(??):544–565, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17302111.

Liu:2017:PSD

[2064] Ji-Cai Liu. Proof of some divisibility results on sums involving binomial coefficients. Journal of Number Theory, 180(??):566–572, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302081.

Varona:2017:CFE

[2065] Juan Luis Varona. The continued fraction expansion of certain Pierce series. Journal of Number Theory, 180(??):573–578, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1730207X.

Schmidt:2017:CFS

[2066] Maxie D. Schmidt. Continued fractions and q-series generating functions for the generalized sum-of-divisors functions. Journal of Number Theory, 180(??):579–605, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302251.

Miguel:2017:CFN

[2067] C. Miguel. Cesàro’s formula in number fields. Journal of Number Theory, 180(??):606–614, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302093.

Kapetanakis:2017:PCI

[2068] Giorgos Kapetanakis. Prescribing coefficients of invariant irreducible polynomials. Journal of Number Theory, 180(??):615–628, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302056.

Kim:2017:NRSb

[2069] Kyoungmin Kim and Byeong-Kweon Oh. The number of representations of squares by integral ternary quadratic forms. Journal of Number Theory, 180(??):629–642, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17302147.

Murty:2017:TPP

[2070] M. Ram Murty and Akshaa Vatwani. Twin primes and the parity problem. Journal of Number Theory, 180(??):643–659, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302135.

Lebowitz-Lockard:2017:CLC

[2071] Noah Lebowitz-Lockard and Paul Pollack. Clustering of linear combinations of multiplicative functions. Journal of Number Theory, 180 (??):660–672, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302263.

Fretwell:2017:GLE

[2072] Dan Fretwell. Generic level p Eisenstein congruences for GSp_{4}. Journal of Number Theory, 180(??):673–693, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17302068.

Ge:2017:VCF

[2073] Wenxu Ge and Feng Zhao. The values of cubic forms at prime arguments. Journal of Number Theory, 180(??):694–709, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302160.

Lapkova:2017:AND

[2074] Kostadinka Lapkova. On the average number of divisors of reducible quadratic polynomials. Journal of Number Theory, 180(??):710–729, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302044.

Sobolewski:2017:AVG

[2075] Bartosz Sobolewski. The 2-adic valuation of generalized Fibonacci sequences with an application to certain Diophantine equations. Journal of Number Theory, 180(??):730–742, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17302354.

Song:2017:DLS

[2076] Kunkun Song and Yuanyang Chang. Dimension of level sets in GCF expansion with parameters. Journal of Number Theory, 180(??):743– 755, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302299.

Somu:2017:ISP

[2077] Sai Teja Somu. On integer sequences in product sets. Journal of Number Theory, 180(??):756–768, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17302275.

Stevanovic:2017:ERZ

[2078] Milorad R. Stevanović and Predrag B. Petrović. The Euler–Riemann zeta function in some series formulae and its values at odd integer points. Journal of Number Theory, 180(??):769–786, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302184.

Benjamin:2017:CSR

[2079] Elliot Benjamin. Corrigendum to “Some real quadratic number fields with their Hilbert 2-class field having cyclic 2-class group” [J. Number Theory 173 (2017) 529–546]. Journal of Number Theory, 180 (??):787–789, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301713. See [1876].

Broughan:2017:CCK

[2080] Kevin Broughan and Daniel Delbourgo. Corrigendum: “A conjecture of De Koninck regarding particular values of the sum of divisors function” [J. Number Theory 137 (2014) 50–66]. Journal of Number Theory, 180(??):790–792, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17301968. See [944].

Anonymous:2017:PD

[2081] Anonymous. Pages 1–240 (December 2017). Journal of Number Theory, 181(??):1–240, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2017:EBl

[2082] Anonymous. Editorial Board. Journal of Number Theory, 181(??):ifc, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302585.

Katz:2017:AVW

[2083] Daniel J. Katz, Philippe Langevin, Sangman Lee, and Yakov Sapozhnikov. The p-adic valuations of Weil sums of binomials. Journal of Number Theory, 181(??):1–26, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17302226.

Moon:2017:PSS

[2084] Yong Suk Moon. Potentially semi-stable deformations of specified Hodge–Tate type and Galois type. Journal of Number Theory, 181 (??):27–50, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302202.

Troncoso:2017:BPP

[2085] Sebastian Troncoso. Bounds for preperiodic points for maps with good reduction. Journal of Number Theory, 181(??):51–72, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302421.

Liu:2017:MDP

[2086] Huaning Liu and Yuchan Qi. On multi-dimensional pseudorandom subsets. Journal of Number Theory, 181(??):73–88, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302287.

Heuberger:2017:ECI

[2087] Clemens Heuberger and Michela Mazzoli. Elliptic curves with isomorphic groups of points over finite field extensions. Journal of Number Theory, 181(??):89–98, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302561.

Ono:2017:FMZ

[2088] Masataka Ono. Finite multiple zeta values associated with 2-colored rooted trees. Journal of Number Theory, 181(??):99–116, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302214.

Srivastava:2017:PDE

[2089] H. M. Srivastava, Irem Kucukoğlu, and Yilmaz Simsek. Partial differential equations for a new family of numbers and polynomials unifying the Apostol-type numbers and the Apostol-type polynomials. Journal of Number Theory, 181(??):117–146, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730210X.

Koch:2017:SLG

[2090] Sophie Koch. On the special linear group over orders in quaternion division algebras. Journal of Number Theory, 181(??):147–163, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302196.

Makiyama:2017:AAF

[2091] Kenji Makiyama. A p-adic analytic family of the D-th Shintani lifting for a Coleman family and congruences between the central L-values. Journal of Number Theory, 181(??):164–199, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302330.

Lazar:2017:VPI

[2092] Youssef Lazar. Values of pairs involving one quadratic form and one linear form at S-integral points. Journal of Number Theory, 181 (??):200–217, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302305.

Tran:2017:FCN

[2093] Minh-Hoang Tran. A formula for the S-class number of an algebraic torus. Journal of Number Theory, 181(??):218–239, December 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302378.

Anonymous:2018:PJa

[2094] Anonymous. Pages 1–362 (January 2018). Journal of Number Theory, 182(??):1–362, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBa

[2095] Anonymous. Editorial Board. Journal of Number Theory, 182(??): ifc, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302937.

Murty:2018:ECF

[2096] M. Ram Murty and Hector Pasten. Elliptic curves, L-functions, and Hilbert’s tenth problem. Journal of Number Theory, 182(??):1–18, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302767.

Bordelles:2018:ETL

[2097] Olivier Bordellès, Florian Luca, and Igor E. Shparlinski. On the error term of a lattice counting problem. Journal of Number Theory, 182(??):19–36, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302883.

Steinerberger:2018:ESR

[2098] Stefan Steinerberger. Exponential sums and Riesz energies. Journal of Number Theory, 182(??):37–56, January 2018. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17302901.

Sakata:2018:TFJ

[2099] Hiroshi Sakata. Trace formula for Jacobi forms of odd squarefree level. Journal of Number Theory, 182(??):57–82, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302123.

Kathiravan:2018:NCC

[2100] T. Kathiravan and S. N. Fathima. New congruences for t-core partitions and Andrews’ singular overpartitions. Journal of Number Theory, 182(??):83–94, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302238.

Byott:2018:SCL

[2101] Nigel P. Byott and G. Griffith Elder. Sufficient conditions for large Galois scaffolds. Journal of Number Theory, 182(??):95–130, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302317.

Maurischat:2018:HPS

[2102] Kathrin Maurischat. On holomorphic projection for symplectic groups. Journal of Number Theory, 182(??):131–178, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302329.

Mattarei:2018:GSP

[2103] Sandro Mattarei and Roberto Tauraso. From generating series to polynomial congruences. Journal of Number Theory, 182(??):179–205, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302342.

Aistleitner:2018:PCE

[2104] Christoph Aistleitner, Thomas Lachmann, and Florian Pausinger. Pair correlations and equidistribution. Journal of Number Theory, 182 (??):206–220, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302366.

Qin:2018:KSS

[2105] Zhenzhen Qin and Tianping Zhang. Kloosterman sums over smooth numbers. Journal of Number Theory, 182(??):221–235, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1730238X.

Djankovic:2018:CRF

[2106] Goran Djanković and Rizwanur Khan. A conjecture for the regularized fourth moment of Eisenstein series. Journal of Number Theory, 182(??):236–257, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302391.

Fang:2018:SWP

[2107] Jin-Hui Fang and Yong-Gao Chen. On the shortest weakly primeadditive numbers. Journal of Number Theory, 182(??):258–270, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302408.

Kaddoura:2018:NIU

[2108] Issam Kaddoura and Bassam Mourad. On a new improved unifying closed formula for all Fibonacci-type sequences and some applications. Journal of Number Theory, 182(??):271–283, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1730241X.

Balestrieri:2018:IAE

[2109] F. Balestrieri. Iterating the algebraic étale-Brauer set. Journal of Number Theory, 182(??):284–295, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17302433.

Bridy:2018:AIZ

[2110] Andrew Bridy and Thomas J. Tucker. ABC implies a Zsigmondy principle for ramification. Journal of Number Theory, 182(??):296–310, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730255X.

Marzec:2018:NVF

[2111] Jolanta Marzec. Non-vanishing of fundamental Fourier coefficients of paramodular forms. Journal of Number Theory, 182(??):311–324, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302706.

He:2018:AGT

[2112] Bo He, Ákos Pintér, Alain Togbé, and Shichun Yang. Another generalization of a theorem of Baker and Davenport. Journal of Number Theory, 182(??):325–343, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302718.

Sun:2018:ASC

[2113] Qingfeng Sun. Averages of shifted convolution sums for GL(3) × GL(2). Journal of Number Theory, 182(??):344–362, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302731.

Anonymous:2018:PF

[2114] Anonymous. Pages 1–500 (February 2018). Journal of Number Theory, 183(??):1–500, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBb

[2115] Anonymous. Editorial Board. Journal of Number Theory, 183(??):ifc, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303682.

Lichtman:2018:EED

[2116] Jared D. Lichtman and Carl Pomerance. Explicit estimates for the distribution of numbers free of large prime factors. Journal of Number Theory, 183(??):1–23, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303402.

Andrade:2018:MVD

[2117] Julio Andrade. Mean values of derivatives of L-functions in function fields: II. Journal of Number Theory, 183(??):24–39, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303384.

Cho:2018:MLD

[2118] Peter J. Cho and Henry H. Kim. Moments of logarithmic derivatives of L-functions. Journal of Number Theory, 183(??):40–61, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303177.

Hu:2018:UBN

[2119] Yongzhong Hu and Maohua Le. An upper bound for the number of solutions of ternary purely exponential diophantine equations. Journal of Number Theory, 183(??):62–73, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730272X.

Balaz:2018:CAF

[2120] V. Baláž, J. Gogola, and T. Visnyai. I_{⌋}^{(⨿)}-convergence of arithmetical functions. Journal of Number Theory, 183(??):74–83, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302743.

Gejima:2018:EFU

[2121] Kohta Gejima. An explicit formula of the unramified Shintani functions for (GSp_{4}, GL_{2×_{GL_{1}}} GL_{2}). Journal of Number Theory, 183(??):84– 132, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302755.

Sun:2018:GCN

[2122] Liang Sun. The growth of class numbers of quadratic Diophantine equations. Journal of Number Theory, 183(??):133–145, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302779.

Sun:2018:APD

[2123] Zhi-Wei Sun. Arithmetic properties of Delannoy numbers and Schröder numbers. Journal of Number Theory, 183(??):146–171, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302792.

deJonge:2018:NEN

[2124] Jaap de Jonge and Cor Kraaikamp. Natural extensions for Nakada’s α-expansions: descending from 1 to g^{2}. Journal of Number Theory, 183(??):172–212, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302809.

Saha:2018:IZC

[2125] Ekata Saha and N. Saradha. On interlacing of the zeros of a certain family of modular forms. Journal of Number Theory, 183(??):213–232, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302810.

Sun:2018:IPH

[2126] Chia-Liang Sun and Julie Tzu-Yueh Wang. Integral points off hyperplanes in positive characteristic. Journal of Number Theory, 183 (??):233–248, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17302822.

Sheydvasser:2018:OQA

[2127] Arseniy Sheydvasser. Orders of quaternion algebras with involution. Journal of Number Theory, 183(??):249–268, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17302834.

Chim:2018:VPP

[2128] Kwok Chi Chim, István Pink, and Volker Ziegler. On a variant of Pillai’s problem II. Journal of Number Theory, 183(??):269–290, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302846.

Wang:2018:CTE

[2129] Jian Wang. On the cyclic torsion of elliptic curves over cubic number fields. Journal of Number Theory, 183(??):291–308, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302858.

Pongsriiam:2018:LAP

[2130] Prapanpong Pongsriiam. Longest arithmetic progressions in reduced residue systems. Journal of Number Theory, 183(??):309–325, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1730286X.

Berczes:2018:DE

[2131] Attila Bérczes, István Pink, Gamze Savaş, and Gökhan Soydan. On the Diophantine equation (x + 1)^{k} + (x + 2)^{k} + ⋯ + (2x)^{k} = y^{n}. Journal of Number Theory, 183(??):326–351, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17302895.

Kim:2018:RRP

[2132] Byoung Du Kim. Ranks of the rational points of abelian varieties over ramified fields, and Iwasawa theory for primes with nonordinary reduction. Journal of Number Theory, 183(??):352–387, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302913.

Jindal:2018:IGG

[2133] Ankita Jindal, Shanta Laishram, and Ritumoni Sarma. Irreducibility and Galois groups of generalized Laguerre polynomials L_{n}^{(−1−n−r)}(x). Journal of Number Theory, 183(??):388–406, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303037.

Gonzalez:2018:SSE

[2134] Oscar E. González, Chung Hang Kwan, Steven J. Miller, Roger Van Peski, and Tian An Wong. On smoothing singularities of elliptic orbital integrals on GL(n) and Beyond Endoscopy. Journal of Number Theory, 183(??):407–427, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303049.

Yaqubi:2018:SDP

[2135] Daniel Yaqubi and Madjid Mirzavaziri. Some divisibility properties of binomial coefficients. Journal of Number Theory, 183(??):428–441, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303050.

Ching:2018:LEO

[2136] Tak Wing Ching. Lagrange’s equation with one prime and three almost-primes. Journal of Number Theory, 183(??):442–465, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303074.

Kriz:2018:AEC

[2137] Igor Kriz. On the arithmetic of elliptic curves and a homotopy limit problem. Journal of Number Theory, 183(??):466–484, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303086.

Lu:2018:TCF

[2138] Guangshi Lü and Ping Xi. On triple correlations of Fourier coefficients of cusp forms. Journal of Number Theory, 183(??):485–492, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303281.

Nguyen:2018:AIQ

[2139] Viet Kh. Nguyen. Addendum to “Imaginary quadratic fields of class number 2 and Levy–Hendy quadratics” [J. Number Theory 178 (2017) 40–46]. Journal of Number Theory, 183(??):493–494, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17302871. See [2001].

Mok:2018:CAS

[2140] Chung Pang Mok. Corrigendum and addendum to “Special values of Lfunctions of elliptic curves over Q and their base change to real quadratic fields” [J. Number Theory 130 (2) (2010) 431–438]. Journal of Number Theory, 183(??):495–499, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303360. See [40].

Li:2018:CMT

[2141] Yan Li and Daeyeoul Kim. Corrigendum to “Menon-type identities derived from actions of subgroups of general linear groups” [J. Number Theory 179 (2017) 97–112]. Journal of Number Theory, 183(??): 500, February 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303414. See [2023].

Anonymous:2018:PMa

[2142] Anonymous. Pages 1–498 (March 2018). Journal of Number Theory, 184 (??):1–498, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBc

[2143] Anonymous. Editorial Board. Journal of Number Theory, 184(??): ifc, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303840.

Nathanson:2018:CEL

[2144] Melvyn B. Nathanson. Comparison estimates for linear forms in additive number theory. Journal of Number Theory, 184(??):1–26, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303062.

Coffey:2018:BIZ

[2145] Mark W. Coffey. Bernoulli identities, zeta relations, determinant expressions, Mellin transforms, and representation of the Hurwitz numbers. Journal of Number Theory, 184(??):27–67, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303098.

Im:2018:RGJ

[2146] Bo-Hae Im and Erik Wallace. Rank gain of Jacobian varieties over finite Galois extensions. Journal of Number Theory, 184(??):68–84, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303104.

Kao:2018:APP

[2147] P.-H. Kao. Almost-prime polynomials at prime arguments. Journal of Number Theory, 184(??):85–106, March 2018. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303116.

Katriel:2018:ATL

[2148] Jacob Katriel. Asymptotically trivial linear homogeneous partition inequalities. Journal of Number Theory, 184(??):107–121, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303128.

Iizuka:2018:CND

[2149] Yoshichika Iizuka. On the class number divisibility of pairs of imaginary quadratic fields. Journal of Number Theory, 184(??):122–127, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730313X.

Qu:2018:SPE

[2150] Yunyun Qu. On simultaneous Pell equations x^{2} − (a^{2} − 1)y^{2} = 1 and y^{2} − pz^{2} = 1. Journal of Number Theory, 184(??):128–132, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303141.

Su:2018:NKP

[2151] Ren-He Su. Newforms in the Kohnen plus space. Journal of Number Theory, 184(??):133–178, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303153.

Cohen:2018:AHM

[2152] Peter Cohen. An analytic heuristic for multiplicity computation for Zaremba’s Conjecture. Journal of Number Theory, 184(??):179–191, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303165.

Xi:2018:QAT

[2153] Ping Xi. A quadratic analogue of Titchmarsh divisor problem. Journal of Number Theory, 184(??):192–205, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303189.

Schweiger:2018:IMM

[2154] Fritz Schweiger. Invariant measures for Moebius maps with three branches. Journal of Number Theory, 184(??):206–215, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303190.

Lin:2018:NTP

[2155] Bernard L. S. Lin. The number of tagged parts over the partitions with designated summands. Journal of Number Theory, 184(??):216– 234, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303207.

Ji:2018:ROM

[2156] Kathy Q. Ji, Helen W. J. Zhang, and Alice X. H. Zhao. Ranks of overpartitions modulo 6 and 10. Journal of Number Theory, 184(??):235– 269, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303219.

Schleischitz:2018:CAS

[2157] Johannes Schleischitz. Cubic approximation to Sturmian continued fractions. Journal of Number Theory, 184(??):270–299, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303220.

Akatsuka:2018:ZFD

[2158] Hirotaka Akatsuka and Ade Irma Suriajaya. Zeros of the first derivative of Dirichlet L-functions. Journal of Number Theory, 184(??):300– 329, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303232.

Adzaga:2018:TWS

[2159] Nikola Adžaga, Andrej Dujella, Dijana Kreso, and Petra Tadić. Triples which are D(n)-sets for several n’s. Journal of Number Theory, 184(??): 330–341, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303244.

Logan:2018:NRM

[2160] Adam Logan. New realizations of modular forms in Calabi–Yau threefolds arising from φ^{4} theory. Journal of Number Theory, 184(??):342– 383, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303256.

Goto:2018:AIV

[2161] Akinari Goto and Taka aki Tanaka. Algebraic independence of the values of functions satisfying Mahler type functional equations under the transformation represented by a power relatively prime to the characteristic of the base field. Journal of Number Theory, 184(??):384– 410, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303268.

Chen:2018:DZL

[2162] Huei-Jeng Chen and Yen-Liang Kuan. On depth 2 zeta-like families. Journal of Number Theory, 184(??):411–427, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1730327X.

Choi:2018:HSW

[2163] Dohoon Choi and Subong Lim. Hecke structures of weakly holomorphic modular forms and their algebraic properties. Journal of Number Theory, 184(??):428–450, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303293.

Simonic:2018:LPD

[2164] Aleksander Simonič. Lehmer pairs and derivatives of Hardy’s Z-function. Journal of Number Theory, 184(??):451–460, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1730330X.

Pham:2018:STP

[2165] Thang Pham, Michael Tait, Le Anh Vinh, and Robert Won. A structure theorem for product sets in extra special groups. Journal of Number Theory, 184(??):461–472, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303311.

Chan:2018:DSP

[2166] Tsz Ho Chan, Stephen Choi, and Peter Cho-Ho Lam. Divisibility on the sequence of perfect squares minus one: the gap principle. Journal of Number Theory, 184(??):473–484, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303323.

Fu:2018:GCC

[2167] Shishuo Fu and Dazhao Tang. On a generalized crank for k-colored partitions. Journal of Number Theory, 184(??):485–497, March 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303335.

Anonymous:2018:PAa

[2168] Anonymous. Pages 1–498 (April 2018). Journal of Number Theory, 185 (??):1–498, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBd

[2169] Anonymous. Editorial Board. Journal of Number Theory, 185(??): ifc, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304341.

Shimada:2018:EEL

[2170] Ichiro Shimada. An even extremal lattice of rank 64. Journal of Number Theory, 185(??):1–15, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304201.

Coppola:2018:FRE

[2171] Giovanni Coppola and M. Ram Murty. Finite Ramanujan expansions and shifted convolution sums of arithmetical functions, II. Journal of Number Theory, 185(??):16–47, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303633.

Cheng:2018:GNS

[2172] Xiuyuan Cheng, Gal Mishne, and Stefan Steinerberger. The geometry of nodal sets and outlier detection. Journal of Number Theory, 185(??):48–64, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303621.

Ivic:2018:MZV

[2173] Aleksandar Ivić. On the multiplicities of zeros of ζ(s) and its values over short intervals. Journal of Number Theory, 185(??):65– 79, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730358X.

Akbal:2018:WGP

[2174] Yıldırım Akbal and Ahmet M. Güloğlu. On Waring–Goldbach problem with Piatetski–Shapiro primes. Journal of Number Theory, 185(??): 80–92, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303591.

Garcia:2018:PRB

[2175] Stephan Ramon Garcia, Florian Luca, and Timothy Schaaff. Primitive root biases for prime pairs I: Existence and non-totality of biases. Journal of Number Theory, 185(??):93–120, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17304158.

Panti:2018:SCF

[2176] Giovanni Panti. Slow continued fractions, transducers, and the Serret theorem. Journal of Number Theory, 185(??):121–143, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303347.

Xiang:2018:SPN

[2177] Li Xiang. The stable property of Newton slopes for general Witt towers. Journal of Number Theory, 185(??):144–159, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303359.

Wang:2018:ESS

[2178] Weiping Wang and Yanhong Lyu. Euler sums and Stirling sums. Journal of Number Theory, 185(??):160–193, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303372.

Cantarini:2018:CAN

[2179] Marco Cantarini. On the Cesàro average of the numbers that can be written as sum of a prime and two squares of primes. Journal of Number Theory, 185(??):194–217, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303426.

Cha:2018:QFT

[2180] Byungchul Cha, Emily Nguyen, and Brandon Tauber. Quadratic forms and their Berggren trees. Journal of Number Theory, 185(??):218– 256, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730344X.

Yamagami:2018:AKC

[2181] Atsushi Yamagami. On 2-adic Kaprekar constants and 2-digit Kaprekar distances. Journal of Number Theory, 185(??):257–280, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303451.

Atobe:2018:MLH

[2182] Hiraku Atobe and Hisashi Kojima. On the Miyawaki lifts of Hermitian modular forms. Journal of Number Theory, 185(??):281–318, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303463.

Jeon:2018:BMC

[2183] Daeyeol Jeon. Bielliptic modular curves X_{0}^{+}(N). Journal of Number Theory, 185(??):319–338, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303475.

Chakraborty:2018:DCN

[2184] K. Chakraborty, A. Hoque, Y. Kishi, and P. P. Pandey. Divisibility of the class numbers of imaginary quadratic fields. Journal of Number Theory, 185(??):339–348, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303487.

Ooto:2018:DEL

[2185] Tomohiro Ooto. On Diophantine exponents for Laurent series over a finite field. Journal of Number Theory, 185(??):349–378, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303499.

Quiroz:2018:SDQ

[2186] Patricio Quiroz. Selectivity in definite quaternion algebras. Journal of Number Theory, 185(??):379–395, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303505.

Laskar:2018:RWD

[2187] Abhijit Laskar. Representations of Weil–Deligne groups and Frobenius conjugacy classes. Journal of Number Theory, 185(??):396–422, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303517.

Sun:2018:ECA

[2188] Lisa Hui Sun and Mingzhi Zhang. On the enumeration and congruences for m-ary partitions. Journal of Number Theory, 185(??):423– 433, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303529.

Aygin:2018:RSQ

[2189] Zafer Selcuk Aygin. Representations by sextenary quadratic forms with coefficients 1, 2, 3 and 6 and on newforms in S_{3}(Γ_{0}(24), χ). Journal of Number Theory, 185(??):434–448, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303530.

Giberson:2018:KCA

[2190] Luke Giberson. Koblitz’s conjecture on average for elliptic curves over abelian number fields of square-free conductor. Journal of Number Theory, 185(??):449–478, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303542.

Liu:2018:SPS

[2191] Zhixin Liu. Small prime solutions of a nonlinear equation. Journal of Number Theory, 185(??):479–489, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303554.

Lu:2018:HDS

[2192] Meiying Lü and Jia Liu. Hausdorff dimensions of some exceptional sets in Engel expansions. Journal of Number Theory, 185(??):490– 498, April 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303566.

Anonymous:2018:PMb

[2193] Anonymous. Pages 1–510 (May 2018). Journal of Number Theory, 186 (??):1–510, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBe

[2194] Anonymous. Editorial Board. Journal of Number Theory, 186(??): ifc, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304420.

Garcia:2018:SAH

[2195] Stephan Ramon Garcia and Bob Lutz. A supercharacter approach to Heilbronn sums. Journal of Number Theory, 186(??):1–15, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1730402X.

Barnett:2018:QJF

[2196] Michael Barnett, Amanda Folsom, Obinna Ukogu, William J. Wesley, and Hui Xu. Quantum Jacobi forms and balanced unimodal sequences. Journal of Number Theory, 186(??):16–34, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17304146.

Gonek:2018:PCZ

[2197] Steven M. Gonek and Haseo Ki. Pair correlation of zeros of the real and imaginary parts of the Riemann zeta-function. Journal of Number Theory, 186(??):35–61, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1730416X.

Soudry:2018:UCR

[2198] David Soudry. The unramified computation of Rankin–Selberg integrals expressed in terms of Bessel models for split orthogonal groups: Part II. Journal of Number Theory, 186(??):62–102, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17304183.

Xu:2018:PDG

[2199] Wenqiang Xu. Popular differences and generalized Sidon sets. Journal of Number Theory, 186(??):103–120, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303578.

Goodson:2018:HPG

[2200] Heidi Goodson. Hypergeometric properties of genus 3 generalized Legendre curves. Journal of Number Theory, 186(??):121–136, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303608.

Mojarrad:2018:CEB

[2201] Hossein Nassajian Mojarrad and Thang Pham. Conditional expanding bounds for two-variable functions over arbitrary fields. Journal of Number Theory, 186(??):137–146, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1730361X.

Yamagata:2018:RDM

[2202] Koji Yamagata and Masakazu Yamagishi. Resultants and discriminants of the multiplication polynomials of Jacobi elliptic functions. Journal of Number Theory, 186(??):147–161, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17303645.

Takai:2018:IRC

[2203] Yuuki Takai. Indivisibility of relative class numbers of totally imaginary quadratic extensions and vanishing of these relative Iwasawa invariants. Journal of Number Theory, 186(??):162–179, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303657.

Biswas:2018:CLF

[2204] Sazzad Ali Biswas. Computation of the Lambda function for a finite Galois extension. Journal of Number Theory, 186(??):180–201, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303669.

Matala-aho:2018:EFS

[2205] Tapani Matala-aho and Wadim Zudilin. Euler’s factorial series and global relations. Journal of Number Theory, 186(??):202–210, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303815.

Hussain:2018:HMV

[2206] Mumtaz Hussain and David Simmons. The Hausdorff measure version of Gallagher’s theorem — closing the gap and beyond. Journal of Number Theory, 186(??):211–225, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303827.

Leonetti:2018:NDP

[2207] Paolo Leonetti and Salvatore Tringali. On the number of distinct prime factors of a sum of super-powers. Journal of Number Theory, 186(??):226–237, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17303943.

Hegedus:2018:EGZ

[2208] Gábor Hegedűs. The Erdős–Ginzburg–Ziv constant and progression-free subsets. Journal of Number Theory, 186(??):238–247, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303955.

Flynn:2018:ALT

[2209] E. V. Flynn. Arbitrarily large Tate–Shafarevich group on Abelian surfaces. Journal of Number Theory, 186(??):248–258, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303967.

Sharif:2018:PIH

[2210] Shahed Sharif. Period and index for higher genus curves. Journal of Number Theory, 186(??):259–268, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17303979.

Bae:2018:AVF

[2211] Sunghan Bae and Hwanyup Jung. Average values of L-functions in even characteristic. Journal of Number Theory, 186(??):269–303, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303980.

Ishii:2018:AZI

[2212] Taku Ishii. Archimedean zeta integrals for the exterior square Lfunctions on GL_{n}. Journal of Number Theory, 186(??):304–345, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17303992.

Son:2018:SST

[2213] Nguyen Thanh Son, Tran Van Tan, and Nguyen Van Thin. Schmidt’s subspace theorem for moving hypersurface targets. Journal of Number Theory, 186(??):346–369, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304006.

Megarbane:2018:COH

[2214] Thomas Mégarbané. Calcul des opérateurs de Hecke sur les classes d’isomorphisme de réseaux pairs de déterminant 2 en dimension 23 et 25. (French) [Computation of Hecke operators on the isomorphism classes of even networks of determinant 2 in dimensions 23 and 25]. Journal of Number Theory, 186(??):370–416, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17304018.

Adam:2018:PVE

[2215] David Adam. Polynômes à valeurs entières sur des ensembles invariants par rotation. (French) [Polynomials with integer values on rotationally invariant sets]. Journal of Number Theory, 186(??):417– 438, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304031.

Diaz-Vargas:2018:LEE

[2216] Javier Diaz-Vargas and Gustavo Vargas de los Santos. The Legendre equation in Euclidean imaginary quadratic number fields. Journal of Number Theory, 186(??):439–451, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17304043.

Chang:2018:ANZ

[2217] Sungkon Chang. Average number of Zeckendorf integers. Journal of Number Theory, 186(??):452–472, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17304055. See

corrigendum [2644].

Robles-Perez:2018:FNS

[2218] A. M. Robles-Pérez and J. C. Rosales. The Frobenius number for sequences of triangular and tetrahedral numbers. Journal of Number Theory, 186(??):473–492, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304067.

Ogasawara:2018:SEB

[2219] Takeshi Ogasawara. Some expressions for binary theta series by etaquotients and their applications. Journal of Number Theory, 186(??): 493–509, May 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304079.

Anonymous:2018:PJb

[2220] Anonymous. Pages 1–512 (June 2018). Journal of Number Theory, 187 (??):1–512, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBf

[2221] Anonymous. Editorial Board. Journal of Number Theory, 187(??): ifc, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300210.

Smith:2018:CMP

[2222] Barry R. Smith. Constructing minimal periods of quadratic irrationalities in Zagier’s reduction theory. Journal of Number Theory, 187(??):

1–26, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17303438.

Niu:2018:FTA

[2223] Chuanze Niu, Wenxin Liu, and Chunlei Liu. On L-functions of twisted T -adic exponential sums. Journal of Number Theory, 187(??):27– 40, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304080.

Li:2018:CS

[2224] Songsong Li and Yi Ouyang. Counting the solutions of λ_{1}x_{1}^{k_{1}} + ⋯ + λ_{t}x_{t}^{k_{t}} ≡ c mod n. Journal of Number Theory, 187(??):41– 65, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304092.

Lim:2018:FSG

[2225] Meng Fai Lim and Ramdorai Sujatha. Fine Selmer groups of congruent Galois representations. Journal of Number Theory, 187(??): 66–91, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304109.

Lee:2018:RAR

[2226] Jonathan D. Lee and Ramarathnam Venkatesan. Rigorous analysis of a randomised number field sieve. Journal of Number Theory, 187(??): 92–159, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304110.

Hajdu:2018:E

[2227] L. Hajdu, Á. Papp, and T. Szakács. On the equation A!B! = C!. Journal of Number Theory, 187(??):160–165, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X17304122.

Cheng:2018:FSM

[2228] Kaimin Cheng and Shaofang Hong. The first and second moments of reversed Dickson polynomials over finite fields. Journal of Number Theory, 187(??):166–188, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304134.

Oppenheim:2018:SPP

[2229] Amotz Oppenheim and Mark Shusterman. Squarefree polynomials with prescribed coefficients. Journal of Number Theory, 187(??):189– 197, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304171.

Tang:2018:CMP

[2230] Dazhao Tang. Congruences modulo powers of 5 for k-colored partitions. Journal of Number Theory, 187(??):198–214, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17304195.

Hu:2018:TA

[2231] Yining Hu. Transcendence of L(1, χ_{s})/Π and automata. Journal of Number Theory, 187(??):215–232, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17304213.

Chen:2018:EBT

[2232] Yong-Gao Chen, Jin-Hui Fang, and Norbert Hegyvári. Erdős–Birch type question in N^{r}. Journal of Number Theory, 187(??):233–249, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304225.

Deng:2018:NS

[2233] Naijuan Deng, Pingzhi Yuan, and Wenyu Luo. Number of solutions to ka^{x} + lb^{y} = c^{z}. Journal of Number Theory, 187(??):250– 263, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304316.

Todd:2018:CCL

[2234] George Todd. A conjectural characterization for F_{q}(t)-linear relations between multizeta values. Journal of Number Theory, 187(??):264– 287, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304328.

Hou:2018:DC

[2235] Xiang-Dong Hou. On the DLW conjectures. Journal of Number Theory, 187(??):288–295, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1730447X.

Kidwell:2018:SSG

[2236] Keenan Kidwell. On the structure of Selmer groups of p-ordinary modular forms over Z_{p}-extensions. Journal of Number Theory, 187(??):296– 331, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304481.

Nomura:2018:ISE

[2237] Jiro Nomura. Integrality of Stickelberger elements attached to unramified extensions of imaginary quadratic fields. Journal of Number Theory, 187(??):332–343, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304493.

Jell:2018:PDT

[2238] Philipp Jell and Veronika Wanner. Poincaré duality for the tropical Dolbeault cohomology of non–Archimedean Mumford curves. Journal of Number Theory, 187(??):344–371, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1730450X.

Hu:2018:DSA

[2239] Hui Hu, Yueli Yu, and Yanfen Zhao. On the digits of Schneider’s p-adic continued fractions. Journal of Number Theory, 187(??):372– 390, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304511.

AlTahan:2018:SRR

[2240] M. Al Tahan and B. Davvaz. Strongly regular relations of arithmetic functions. Journal of Number Theory, 187(??):391–402, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17304523.

Pollack:2018:LPQ

[2241] Paul Pollack. The least prime quadratic nonresidue in a prescribed residue class mod4. Journal of Number Theory, 187(??):403–414, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17304535.

Merai:2018:TLC

[2242] László Mérai and Arne Winterhof. On the N-th linear complexity of automatic sequences. Journal of Number Theory, 187(??):415– 429, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304547.

Kosters:2018:SFC

[2243] Michiel Kosters and Hui June Zhu. On slopes of L-functions of Z_{p}- covers over the projective line. Journal of Number Theory, 187(??): 430–452, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304559.

Tryba:2018:CUD

[2244] Jacek Tryba. Characterization of uniformly distributed sets and maximal density sets. Journal of Number Theory, 187(??):453–468, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X17304560.

Krumm:2018:GGF

[2245] David Krumm. Galois groups in a family of dynatomic polynomials. Journal of Number Theory, 187(??):469–511, June 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17304572.

Anonymous:2018:PJc

[2246] Anonymous. Pages 1–410 (July 2018). Journal of Number Theory, 188 (??):1–410, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBg

[2247] Anonymous. Editorial Board. Journal of Number Theory, 188(??): ifc, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300672.

Martin:2018:RDC

[2248] Kimball Martin. Refined dimensions of cusp forms, and equidistribution and bias of signs. Journal of Number Theory, 188(??):1– 17, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300453.

Hubbard:2018:IIS

[2249] David Hubbard and Lawrence C. Washington. Iwasawa invariants of some non-cyclotomic Z_{p}-extensions. Journal of Number Theory, 188(??): 18–47, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300398.

Choi:2018:BCC

[2250] SoYoung Choi and Bo-Hae Im. Bounds for the coefficients of cusp forms for Γ_{0}(3). Journal of Number Theory, 188(??):48–70, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300386.

Betina:2018:CFL

[2251] Adel Betina and Emmanuel Lecouturier. Congruence formulas for Legendre modular polynomials. Journal of Number Theory, 188(??):71– 87, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300362.

Tenenbaum:2018:NLL

[2252] Gérald Tenenbaum. Note sur les lois locales conjointes de la fonction nombre de facteurs premiers. (French) [Note on the joint local laws of the function number of prime factors]. Journal of Number Theory, 188(??):88–95, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300350.

Maier:2018:ESR

[2253] Helmut Maier and Michael Th. Rassias. Estimates of sums related to the Nyman–Beurling criterion for the Riemann Hypothesis. Journal of Number Theory, 188(??):96–120, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18300180.

Huber:2018:SL

[2254] Tim Huber, Dan Schultz, and Dongxi Ye. Series for 1/π of level 20. Journal of Number Theory, 188(??):121–136, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18300167.

Dickson:2018:PES

[2255] Martin Dickson and Michael Neururer. Products of Eisenstein series and Fourier expansions of modular forms at cusps. Journal of Number Theory, 188(??):137–164, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300192.

Bisatt:2018:FEG

[2256] Matthew Bisatt. Frobenius elements in Galois representations with SL_{n} image. Journal of Number Theory, 188(??):165–171, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300404.

Liu:2018:BFE

[2257] Kui Liu, Igor E. Shparlinski, and Tianping Zhang. Bilinear forms with exponential sums with binomials. Journal of Number Theory, 188(??): 172–185, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300179.

Lumley:2018:EBF

[2258] Allysa Lumley. Explicit bounds for L-functions on the edge of the critical strip. Journal of Number Theory, 188(??):186–209, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300313.

Gambini:2018:DAP

[2259] Alessandro Gambini, Alessandro Languasco, and Alessandro Zaccagnini. A Diophantine approximation problem with two primes and one kth power of a prime. Journal of Number Theory, 188(??):210– 228, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300325.

Blokhuis:2018:RCD

[2260] Aart Blokhuis, Xiwang Cao, Wun-Seng Chou, and Xiang-Dong Hou. On the roots of certain Dickson polynomials. Journal of Number Theory, 188(??):229–246, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300337.

Eie:2018:SFM

[2261] Minking Eie and Yao Lin Ong. Sum formulas of multiple zeta values with even arguments and polynomial weights. Journal of Number Theory, 188(??):247–262, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300349.

Agwu:2018:FDS

[2262] Anthony Agwu, Phillip Harris, Kevin James, Siddarth Kannan, and Huixi Li. Frobenius distributions in short intervals for CM elliptic curves. Journal of Number Theory, 188(??):263–280, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300374.

Nagaoka:2018:KTO

[2263] Shoyu Nagaoka and Sho Takemori. On the modp kernel of the theta operator and Eisenstein series. Journal of Number Theory, 188(??): 281–298, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300416.

Lee:2018:TRC

[2264] Yoonjin Lee and Yoon Kyung Park. Three Ramanujan continued fractions with modularity. Journal of Number Theory, 188(??):299– 323, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300428.

Namboothiri:2018:NSR

[2265] K. Vishnu Namboothiri. On the number of solutions of a restricted linear congruence. Journal of Number Theory, 188(??):324– 334, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830043X.

Sono:2018:CLB

[2266] Keiju Sono. Continuous lower bounds for the moments of Dedekind zeta-functions. Journal of Number Theory, 188(??):335–356, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18300441.

Judge:2018:DOV

[2267] Samuel D. Judge and Fabrizio Zanello. On the density of the odd values of the partition function, II: an infinite conjectural framework. Journal of Number Theory, 188(??):357–370, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18300465.

Sugiyama:2018:IRD

[2268] Yusuke Sugiyama. The integrality and reduction of Drinfeld modular forms of arbitrary rank. Journal of Number Theory, 188(??):371– 391, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300477.

Li:2018:RPF

[2269] Ya-Li Li and Yong-Gao Chen. On the r-th root partition function, II. Journal of Number Theory, 188(??):392–409, July 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300489.

Anonymous:2018:PAb

[2270] Anonymous. Pages 1–416 (August 2018). Journal of Number Theory, 189 (??):1–416, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBh

[2271] Anonymous. Editorial Board. Journal of Number Theory, 189(??): ifc, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300817.

Chung:2018:WDS

[2272] Fan Chung and Ron Graham. Well dispersed sequences in [0, 1]^{d}. Journal of Number Theory, 189(??):1–24, August 2018. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18300623.

DHoker:2018:IBM

[2273] Eric D’Hoker and Michael B. Green. Identities between modular graph forms. Journal of Number Theory, 189(??):25–80, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17304614.

Ahlgren:2018:CMM

[2274] Scott Ahlgren and Byungchan Kim. Congruences for a mock modular form on SL_{2}(Z) and the smallest parts function. Journal of Number Theory, 189(??):81–89, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X17304596.

Luca:2018:PPS

[2275] Florian Luca and Vandita Patel. On perfect powers that are sums of two Fibonacci numbers. Journal of Number Theory, 189(??):90– 96, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300520.

Swaenepoel:2018:PDF

[2276] Cathy Swaenepoel. Prescribing digits in finite fields. Journal of Number Theory, 189(??):97–114, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X17304584.

Daowsud:2018:CFR

[2277] Katthaleeya Daowsud and Thomas A. Schmidt. Continued fractions for rational torsion. Journal of Number Theory, 189(??):115–130, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17304602. See corrigendum [2648].

Choie:2018:HFA

[2278] YoungJu Choie and Winfried Kohnen. On Hecke L-functions attached to half-integral weight modular forms. Journal of Number Theory, 189(??):131–137, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300015.

Tran:2018:CRP

[2279] Manh Hung Tran. Counting rational points on smooth cubic curves. Journal of Number Theory, 189(??):138–146, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300076.

Demangos:2018:LPD

[2280] Luca Demangos. Lehmer problem and Drinfeld modules. Journal of Number Theory, 189(??):147–185, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18300088.

Gao:2018:ONT

[2281] Xia Gao. On the Ohno–Nakagawa Theorem. Journal of Number Theory, 189(??):186–210, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830009X.

Malara:2018:CDN

[2282] Grzegorz Malara, Tomasz Szemberg, and Justyna Szpond. On a conjecture of Demailly and new bounds on Waldschmidt constants in P^{N}. Journal of Number Theory, 189(??):211–219, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300106.

Giulietti:2018:SGC

[2283] M. Giulietti, M. Montanucci, L. Quoos, and G. Zini. On some Galois covers of the Suzuki and Ree curves. Journal of Number Theory, 189(??):220–254, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300118.

Sofo:2018:GOE

[2284] Anthony Sofo. General order Euler sums with multiple argument. Journal of Number Theory, 189(??):255–271, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1830012X.

Byeon:2018:ELB

[2285] Dongho Byeon and Jigu Kim. An explicit lower bound for special values of Dirichlet L-functions. Journal of Number Theory, 189(??):272– 303, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300131.

Kato:2018:ROS

[2286] Ryo Kato and Akihiko Yukie. Rational orbits of the space of pairs of exceptional Jordan algebras. Journal of Number Theory, 189(??):304– 352, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300143.

Biggs:2018:AFW

[2287] Kirsti D. Biggs. On the asymptotic formula in Waring’s problem with shifts. Journal of Number Theory, 189(??):353–379, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18300155.

Kuan:2018:HBT

[2288] Chan Ieong Kuan. Hybrid bounds on twisted L-functions associated with modular forms. Journal of Number Theory, 189(??):380– 416, August 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300519.

Anonymous:2018:PS

[2289] Anonymous. Pages 1–414 (September 2018). Journal of Number Theory, 190(??):1–414, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBi

[2290] Anonymous. Editorial Board. Journal of Number Theory, 190(??):ifc, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301124.

Forrester:2018:VDR

[2291] Peter J. Forrester and Jiyuan Zhang. Volumes and distributions for random unimodular complex and quaternion lattices. Journal of Number Theory, 190(??):1–39, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18300970.

Chan:2018:SDS

[2292] Song Heng Chan and Byungchan Kim. On some double-sum false theta series. Journal of Number Theory, 190(??):40–55, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18300611.

Defant:2018:CCC

[2293] Colin Defant. Connected components of complex divisor functions. Journal of Number Theory, 190(??):56–71, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300799.

Mao:2018:NNS

[2294] Renrong Mao. On the non-negativity of the spt-crank for partitions without repeated odd parts. Journal of Number Theory, 190(??):72–85, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300532.

Eum:2018:RSQ

[2295] Ick Sun Eum. Representations by some quinary quadratic forms of level 8. Journal of Number Theory, 190(??):86–108, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300544.

Brazelton:2018:ZNE

[2296] Thomas Brazelton and Victoria Jakicic. Zeros of newform Eisenstein series on Γ_{0}(N). Journal of Number Theory, 190(??):109–130, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300556.

Awonusika:2018:DLC

[2297] Richard Olu Awonusika. Determinants of the Laplacians on complex projective spaces P^{n}(C)(n ≥ 1). Journal of Number Theory, 190 (??):131–155, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300568.

Buium:2018:ETC

[2298] Alexandru Buium and Emma Previato. The Euler top and canonical lifts. Journal of Number Theory, 190(??):156–168, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1830057X.

Ren:2018:RTS

[2299] Yuan Ren. Rational torsion subgroups of modular Jacobian varieties. Journal of Number Theory, 190(??):169–186, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300581.

Longo:2018:EZF

[2300] Matteo Longo and Maria Rosaria Pati. Exceptional zero formulae for anticyclotomic p-adic L-functions of elliptic curves in the ramified case. Journal of Number Theory, 190(??):187–211, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300593.

He:2018:FSC

[2301] Xiaoguang He and Lilu Zhao. On the first sign change of Fourier coefficients of cusp forms. Journal of Number Theory, 190(??):212–228, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830060X.

Kim:2018:SPC

[2302] Eunmi Kim. On a sign pattern for the crank function of cubic partition pairs. Journal of Number Theory, 190(??):229–240, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18300635.

Porritt:2018:RRN

[2303] Sam Porritt. Residue races of the number of prime divisors function. Journal of Number Theory, 190(??):241–253, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300647.

Ma:2018:FDS

[2304] Liangang Ma. Full dimensional sets of reals whose sums of partial quotients increase in certain speed. Journal of Number Theory, 190 (??):254–276, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300659.

Zhou:2018:TKC

[2305] Haiyan Zhou and Zhibin Liang. Tame kernels of cubic and sextic fields. Journal of Number Theory, 190(??):277–291, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300751.

Ju:2018:USG

[2306] Jangwon Ju and Byeong-Kweon Oh. Universal sums of generalized octagonal numbers. Journal of Number Theory, 190(??):292–302, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300763.

Chen:2018:APL

[2307] Yong-Gao Chen and Cheng-Hao Lei. Arithmetic progressions in the least positive reduced residue systems. Journal of Number Theory, 190 (??):303–310, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300775.

Li:2018:DFN

[2308] Yao-Qiang Li and Bing Li. Distributions of full and non-full words in beta-expansions. Journal of Number Theory, 190(??):311–332, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300787.

Kuwata:2018:MWL

[2309] Masato Kuwata and Kazuki Utsumi. Mordell–Weil lattice of Inose’s elliptic K3 surface arising from the product of 3-isogenous elliptic curves. Journal of Number Theory, 190(??):333–351, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1830088X.

Hare:2018:PRF

[2310] Kevin G. Hare and J. C. Saunders. On (a, b) pairs in random Fibonacci sequences. Journal of Number Theory, 190(??):352–366, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300891.

Surer:2018:SCP

[2311] Paul Surer. Substitutions, coding prescriptions and integer representation. Journal of Number Theory, 190(??):367–393, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18300908.

Rump:2018:AHP

[2312] Wolfgang Rump. The L-algebra of Hurwitz primes. Journal of Number Theory, 190(??):394–413, September 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1830091X.

Anonymous:2018:PO

[2313] Anonymous. Pages 1–426 (October 2018). Journal of Number Theory, 191(??):1–426, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBj

[2314] Anonymous. Editorial Board. Journal of Number Theory, 191(??): ifc, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301598.

Bettin:2018:NIE

[2315] Sandro Bettin, Chantal David, and Christophe Delaunay. Non-isotrivial elliptic surfaces with non-zero average root number. Journal of Number Theory, 191(??):1–84, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300945.

Li:2018:VHF

[2316] Wanlin Li. Vanishing of hyperelliptic L-functions at the central point. Journal of Number Theory, 191(??):85–103, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301057.

Lichtman:2018:RSP

[2317] Jared Duker Lichtman. The reciprocal sum of primitive nondeficient numbers. Journal of Number Theory, 191(??):104–118, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301082.

Bloom:2018:SSL

[2318] Samuel Bloom. The square sieve and a Lang–Trotter question for generic abelian varieties. Journal of Number Theory, 191(??):119–157, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301215.

Banks:2018:CPB

[2319] William D. Banks and Victor Z. Guo. Consecutive primes and Beatty sequences. Journal of Number Theory, 191(??):158–174, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301227.

Nathanson:2018:EFS

[2320] Melvyn B. Nathanson. Every finite subset of an abelian group is an asymptotic approximate group. Journal of Number Theory, 191 (??):175–193, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301331.

Shparlinski:2018:ECI

[2321] Igor E. Shparlinski and Liangyi Zhao. Elliptic curves in isogeny classes. Journal of Number Theory, 191(??):194–212, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301367.

Pi:2018:CVT

[2322] Qinghua Pi. Central values of twisted base change L-functions associated to Hilbert modular forms. Journal of Number Theory, 191 (??):213–245, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300921.

Roy:2018:ZSS

[2323] Bidisha Roy and R. Thangadurai. On zero-sum subsequences in a finite abelian p-group of length not exceeding a given number. Journal of Number Theory, 191(??):246–257, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18300933.

Singh:2018:DSC

[2324] Saurabh Kumar Singh. On double shifted convolution sum of SL(2, Z) Hecke eigenforms. Journal of Number Theory, 191(??):258–272, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300957.

Guo:2018:LPS

[2325] Jingwei Guo and Weiwei Wang. Lattice points in stretched model domains of finite type in R^{d}. Journal of Number Theory, 191(??):273– 288, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300969.

Anamby:2018:BPN

[2326] Pramath Anamby and Soumya Das. Bounds for the Petersson norms of the pullbacks of Saito–Kurokawa lifts. Journal of Number Theory, 191(??):289–304, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18300982.

Sanna:2018:MLG

[2327] Carlo Sanna. The moments of the logarithm of a G.C.D. related to Lucas sequences. Journal of Number Theory, 191(??):305–315, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300994.

Isaacson:2018:CSL

[2328] Brad Isaacson. Character sums of Lee and Weintraub. Journal of Number Theory, 191(??):316–344, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301008.

Matala-aho:2018:HTE

[2329] Tapani Matala-aho and Louna Seppälä. Hermite–Thue equation: Padé approximations and Siegel’s lemma. Journal of Number Theory, 191 (??):345–383, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830101X.

Garbin:2018:EBF

[2330] Daniel Garbin. Effective bounds for Fourier coefficients of certain weakly holomorphic modular forms. Journal of Number Theory, 191 (??):384–395, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301069.

Lin:2018:AGJ

[2331] Bingchen Lin. Archimedean Godement–Jacquet zeta integrals and test functions. Journal of Number Theory, 191(??):396–426, October 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301070.

Anonymous:2018:PN

[2332] Anonymous. Pages 1–406 (November 2018). Journal of Number Theory, 192(??):1–406, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBk

[2333] Anonymous. Editorial Board. Journal of Number Theory, 192(??):ifc, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301926.

DHoker:2018:FSM

[2334] Eric D’Hoker and William Duke. Fourier series of modular graph functions. Journal of Number Theory, 192(??):1–36, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301318.

Faye:2018:FGS

[2335] Bernadette Faye and Antonin Riffaut. Fields generated by sums and products of singular moduli. Journal of Number Theory, 192(??):37– 46, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301021.

Ge:2018:VBL

[2336] Wenxu Ge, Min Zhang, and Jinjiang Li. The values of binary linear forms at prime arguments. Journal of Number Theory, 192(??):47–64, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301033.

Dubickas:2018:STT

[2337] Artūras Dubickas. On sums of two and three roots of unity. Journal of Number Theory, 192(??):65–79, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18301045.

Ahlen:2018:FCA

[2338] Olof Ahlén, Henrik P. A. Gustafsson, Axel Kleinschmidt, Baiying Liu, and Daniel Persson. Fourier coefficients attached to small automorphic representations of SL_{n}(A). Journal of Number Theory, 192(??):80–142, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301197.

Jakhar:2018:SRI

[2339] Anuj Jakhar and Neeraj Sangwan. Some results for the irreducibility of truncated binomial expansions. Journal of Number Theory, 192 (??):143–149, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301203.

Clark:2018:TBT

[2340] Pete L. Clark, Marko Milosevic, and Paul Pollack. Typically bounding torsion. Journal of Number Theory, 192(??):150–167, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301240.

Kamano:2018:WSF

[2341] Ken Kamano. Weighted sum formulas for finite multiple zeta values. Journal of Number Theory, 192(??):168–180, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301252.

Wong:2018:SQF

[2342] Ka Lun Wong. Sums of quadratic functions with two discriminants. Journal of Number Theory, 192(??):181–196, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301264.

Franc:2018:UDH

[2343] Cameron Franc, Terry Gannon, and Geoffrey Mason. On unbounded denominators and hypergeometric series. Journal of Number Theory, 192(??):197–220, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301288.

Spiegelhofer:2018:DBC

[2344] Lukas Spiegelhofer and Michael Wallner. Divisibility of binomial coefficients by powers of two. Journal of Number Theory, 192(??):221–239, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830129X.

Harrington:2018:NCE

[2345] Joshua Harrington and Lenny Jones. A new condition equivalent to the Ankeny–Artin–Chowla conjecture. Journal of Number Theory, 192 (??):240–250, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301306.

Toth:2018:MAD

[2346] László Tóth and Wenguang Zhai. On multivariable averages of divisor functions. Journal of Number Theory, 192(??):251–269, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301343.

Lee:2018:IDC

[2347] Yoonjin Lee and Jinjoo Yoo. Indivisibility of divisor class numbers of Kummer extensions over the rational function field. Journal of Number Theory, 192(??):270–292, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301355.

Gekeler:2018:DMF

[2348] Ernst-Ulrich Gekeler. On Drinfeld modular forms of higher rank III: the analogue of the k/12-formula. Journal of Number Theory, 192 (??):293–306, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301379.

Chen:2018:GER

[2349] Haibo Chen and Min Yu. A generalization of the Erdős–Rényi limit theorem and the corresponding multifractal analysis. Journal of Number Theory, 192(??):307–327, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301380.

Bertok:2018:SNT

[2350] Cs. Bertók, K. Győry, L. Hajdu, and A. Schinzel. On the smallest number of terms of vanishing sums of units in number fields. Journal of Number Theory, 192(??):328–347, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301392.

Kurokawa:2018:LFM

[2351] Nobushige Kurokawa and Hidekazu Tanaka. Limit formulas for multiple Hurwitz zeta functions. Journal of Number Theory, 192(??):348–355, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301409.

Allouche:2018:MCK

[2352] Jean-Paul Allouche and Leo Goldmakher. Mock characters and the Kronecker symbol. Journal of Number Theory, 192(??):356–372, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301410.

Li:2018:MTI

[2353] Yan Li and Daeyeoul Kim. Menon-type identities with additive characters. Journal of Number Theory, 192(??):373–385, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301422.

Walton:2018:CPP

[2354] Laura Walton. Counting periodic points on quotient varieties over F_{q}. Journal of Number Theory, 192(??):386–405, November 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301434.

Anonymous:2018:PD

[2355] Anonymous. Pages 1–424 (December 2018). Journal of Number Theory, 193(??):1–424, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2018:EBl

[2356] Anonymous. Editorial Board. Journal of Number Theory, 193(??):ifc, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302117.

Horesh:2018:PPO

[2357] Tal Horesh and Amos Nevo. Prime points in orbits: Some instances of the Bourgain–Gamburd–Sarnak conjecture. Journal of Number Theory, 193(??):1–36, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830132X.

Knapp:2018:AZD

[2358] Michael P. Knapp. 2-adic zeros of diagonal forms. Journal of Number Theory, 193(??):37–47, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301471.

Gaba:2018:GRC

[2359] Radu Gaba and Alexandru A. Popa. A generalization of Ramanujan’s congruence to modular forms of prime level. Journal of Number Theory, 193(??):48–73, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301501.

Schwagenscheidt:2018:ESW

[2360] Markus Schwagenscheidt. Eisenstein series for the Weil representation. Journal of Number Theory, 193(??):74–90, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301574.

Arenas:2018:OET

[2361] Manuel Arenas, Luis Arenas-Carmona, and Jaime Contreras. On optimal embeddings and trees. Journal of Number Theory, 193(??):91–117, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301239.

Macourt:2018:BES

[2362] Simon Macourt. Bounds on exponential sums with quadrinomials. Journal of Number Theory, 193(??):118–127, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301446.

Kumar:2018:NHI

[2363] Balesh Kumar and Manickam Murugesan. On newforms of half-integral weight and Jacobi forms. Journal of Number Theory, 193(??):128–153, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301458.

Ochiai:2018:NGC

[2364] Tsubasa Ochiai. New generalizations of congruent numbers. Journal of Number Theory, 193(??):154–170, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1830146X.

Husu:2018:BSS

[2365] Cristiano Husu. The butterfly sequence: the second difference sequence of the numbers of integer partitions with distinct parts. Journal of Number Theory, 193(??):171–188, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301483.

Griffon:2018:ABS

[2366] Richard Griffon. Analogue of the Brauer–Siegel theorem for Legendre elliptic curves. Journal of Number Theory, 193(??):189–212, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301495.

Bandini:2018:ECA

[2367] Andrea Bandini and Maria Valentino. Euler characteristic and Akashi series for Selmer groups over global function fields. Journal of Number Theory, 193(??):213–234, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301513.

Jeong:2018:BME

[2368] Sangtae Jeong. Bernoulli maps on Z_{p} in the expansions of van der Put and Mahler. Journal of Number Theory, 193(??):235–265, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301525.

Ejder:2018:TSE

[2369] Özlem Ejder. Torsion subgroups of elliptic curves over quadratic cyclotomic fields in elementary abelian 2-extensions. Journal of Number Theory, 193(??):266–301, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301537.

Kani:2018:ESC

[2370] Ernst Kani. Elliptic subcovers of a curve of genus 2 II. The refined Humbert invariant. Journal of Number Theory, 193(??):302–335, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301549.

Parenti:2018:UCF

[2371] Solly Parenti. Unitary PSL_{2} CM fields and the Colmez conjecture. Journal of Number Theory, 193(??):336–356, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301550.

Wu:2018:TGT

[2372] Jie Wu. Two generalisations of the Titchmarsh divisor problem. Journal of Number Theory, 193(??):357–372, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18301562.

Oh:2018:CPB

[2373] Byeong-Kweon Oh and Hoseog Yu. Completely p-primitive binary quadratic forms. Journal of Number Theory, 193(??):373–385, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301665.

Sharma:2018:EPP

[2374] Rajendra K. Sharma, Ambrish Awasthi, and Anju Gupta. Existence of pair of primitive elements over finite fields of characteristic 2. Journal of Number Theory, 193(??):386–394, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18301677.

McDonald:2018:TSE

[2375] Robert J. S. McDonald. Torsion subgroups of elliptic curves over function fields of genus 0. Journal of Number Theory, 193(??):395–423, December 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301689.

Anonymous:2019:PJa

[2376] Anonymous. Pages 1–426 (January 2019). Journal of Number Theory, 194(??):1–426, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBa

[2377] Anonymous. Editorial Board. Journal of Number Theory, 194(??): ifc, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302531.

Connes:2019:AWT

[2378] Alain Connes. Around Wilson’s theorem. Journal of Number Theory, 194(??):1–7, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302336.

Murty:2019:LBT

[2379] M. Ram Murty, François Séguin, and Cameron L. Stewart. A lower bound for the two-variable Artin conjecture and prime divisors of recurrence sequences. Journal of Number Theory, 194(??):8–29, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301902.

Gunns:2019:ERD

[2380] Jos Gunns and Christopher Hughes. The effect of repeated differentiation on L-functions. Journal of Number Theory, 194(??):30–43, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302270.

Baier:2019:MET

[2381] Stephan Baier and Neha Prabhu. Moments of the error term in the Sato–Tate law for elliptic curves. Journal of Number Theory, 194(??):

44–82, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302245.

Chinis:2019:CAR

[2382] Jake Chinis. Computing the average root number of an elliptic surface. Journal of Number Theory, 194(??):83–116, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302312.

Yukie:2019:DTR

[2383] Akihiko Yukie. On the density theorem related to the space of nonsplit tri-Hermitian forms I. Journal of Number Theory, 194(??):117– 169, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302348.

Trebjesanin:2019:NQ

[2384] Marija Bliznac Trebješanin and Alan Filipin. Nonexistence of D(4)- quintuples. Journal of Number Theory, 194(??):170–217, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302324.

Palojarvi:2019:EUL

[2385] Neea Palojärvi. On the explicit upper and lower bounds for the number of zeros of the Selberg class. Journal of Number Theory, 194 (??):218–250, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302257.

Tangedal:2019:TVG

[2386] Brett A. Tangedal. A two-variable generalization of the Kummer– Malmstén formula for the logarithm of the double gamma and double sine functions. Journal of Number Theory, 194(??):251–277, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302300.

Badr:2019:PMF

[2387] Eslam Badr and Francesc Bars. Plane model-fields of definition, fields of definition, and the field of moduli for smooth plane curves. Journal of Number Theory, 194(??):278–283, January 2019. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18302294.

Ren:2019:ZDE

[2388] Xiumin Ren and Wei Zhang. Zero-density estimate of L-functions for cusp forms. Journal of Number Theory, 194(??):284–296, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830221X.

Hirakawa:2019:UPT

[2389] Yoshinosuke Hirakawa and Hideki Matsumura. A unique pair of triangles. Journal of Number Theory, 194(??):297–302, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302269.

Sourmelidis:2019:MCB

[2390] Athanasios Sourmelidis. On the meromorphic continuation of Beatty zeta-functions and Sturmian Dirichlet series. Journal of Number Theory, 194(??):303–318, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302282.

Manickam:2019:EZM

[2391] M. Manickam and Karam Deo Shankhadhar. An Eichler–Zagier map for Jacobi cusp forms on H × C^{(},∞)}. Journal of Number Theory, 194(??):319–334, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302233.

Adell:2019:PGS

[2392] José A. Adell and Alberto Lekuona. A probabilistic generalization of the Stirling numbers of the second kind. Journal of Number Theory, 194(??):335–355, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302221.

Covill:2019:SGS

[2393] Elena C. Covill, Mohammad Javaheri, and Nikolai A. Krylov. On the subgroup generated by solutions of Pell’s equation. Journal of Number Theory, 194(??):356–371, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18302208.

Marques:2019:SVQ

[2394] Diego Marques and Carlos Gustavo Moreira. On a stronger version of a question proposed by K. Mahler. Journal of Number Theory, 194(??):372–380, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302038.

Caballero:2019:FIE

[2395] José Manuel Rodríguez Caballero. On a function introduced by Erdős and Nicolas. Journal of Number Theory, 194(??):381–389, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301999.

Matheus:2019:MSN

[2396] Carlos Matheus and Carlos Gustavo Moreira. Markov spectrum near Freiman’s isolated points in M \ L. Journal of Number Theory, 194 (??):390–408, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301896.

Hinman:2019:URU

[2397] J. Hinman, B. Kuca, A. Schlesinger, and A. Sheydvasser. The unreasonable rigidity of Ulam sequences. Journal of Number Theory, 194(??):409–425, January 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301884.

Anonymous:2019:PF

[2398] Anonymous. Pages 1–376 (February 2019). Journal of Number Theory, 195(??):1–376, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBb

[2399] Anonymous. Editorial Board. Journal of Number Theory, 195(??):ifc, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302646.

Doyle:2019:UFD

[2400] John R. Doyle and Joseph H. Silverman. A uniform field-of-definition/ field-of-moduli bound for dynamical systems on P^{N}. Journal of Number Theory, 195(??):1–22, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18302464.

Im:2019:RRP

[2401] Bo-Hae Im and Byoung Du Kim. Ranks of rational points of the Jacobian varieties of hyperelliptic curves. Journal of Number Theory, 195(??):23–50, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302373.

Sebe:2019:CRG

[2402] Gabriela Ileana Sebe and Dan Lascu. On convergence rate in the Gauss–Kuzmin problem for θ-expansions. Journal of Number Theory, 195(??):51–71, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301690.

Hirose:2019:DSF

[2403] Minoru Hirose, Kohei Iwaki, Nobuo Sato, and Koji Tasaka. Duality/sum formulas for iterated integrals and their application to multiple zeta values. Journal of Number Theory, 195(??):72–83, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301707.

Wang:2019:SLI

[2404] Guoqing Wang. Structure of the largest idempotent-product free sequences in semigroups. Journal of Number Theory, 195(??):84–95, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301719.

Dose:2019:DCC

[2405] Valerio Dose, Pietro Mercuri, and Claudio Stirpe. Double covers of Cartan modular curves. Journal of Number Theory, 195(??):96–114, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301720.

Booher:2019:PGD

[2406] Jeremy Booher. Producing geometric deformations of orthogonal and symplectic Galois representations. Journal of Number Theory, 195 (??):115–158, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18301732.

Marcolli:2019:FQM

[2407] Matilde Marcolli and Gonçalo Tabuada. Feynman quadrics-motive of the massive sunset graph. Journal of Number Theory, 195(??):159– 183, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301744.

Leung:2019:BSE

[2408] Chi-Wai Leung and Chi-Keung Ng. Berkovich spectra of elements in Banach rings. Journal of Number Theory, 195(??):184–216, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301756.

Pandey:2019:DPR

[2409] Manish K. Pandey, Sudhir Pujahari, and Jyoti Prakash Saha. Distinguishing pure representations by normalized traces. Journal of Number Theory, 195(??):217–225, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18301768.

Komornik:2019:BWS

[2410] Vilmos Komornik and Derong Kong. Bases in which some numbers have exactly two expansions. Journal of Number Theory, 195(??):226– 268, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830177X.

Yin:2019:MRS

[2411] Qiuyu Yin, Shaofang Hong, Liping Yang, and Min Qiu. Multiple reciprocal sums and multiple reciprocal star sums of polynomials are almost never integers. Journal of Number Theory, 195(??):269–292, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301781.

Xu:2019:FSH

[2412] Peng Xu. Freeness of spherical Hecke modules of unramified U(2, 1) in characteristic p. Journal of Number Theory, 195(??):293–311, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301793.

Hauseux:2019:FPG

[2413] Julien Hauseux, Tobias Schmidt, and Claus Sorensen. Functorial properties of generalised Steinberg representations. Journal of Number Theory, 195(??):312–329, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830180X.

Salami:2019:RPC

[2414] Sajad Salami. The rational points on certain Abelian varieties over function fields. Journal of Number Theory, 195(??):330–337, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S0022314X18301811. See corrigendum [2646].

Keating:2019:NRD

[2415] Kevin Keating. On the numerology of ramification data for power series in characteristic p. Journal of Number Theory, 195(??):338–357, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301823.

Aygin:2019:EST

[2416] Zafer Selcuk Aygin. On Eisenstein series in M_{2k}(Γ_{0}(N)) and their applications. Journal of Number Theory, 195(??):358–375, February 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301835.

Anonymous:2019:PMa

[2417] Anonymous. Pages 1–420 (March 2019). Journal of Number Theory, 196 (??):1–420, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBc

[2418] Anonymous. Editorial Board. Journal of Number Theory, 196(??): ifc, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303160.

Baier:2019:LSI

[2419] Stephan Baier and Rajneesh Kumar Singh. Large sieve inequality with power moduli for function fields. Journal of Number Theory, 196(??): 1–13, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302725. See erratum [2643].

Ye:2019:GZT

[2420] Dongxi Ye. Gross–Zagier type CM value formulas on X_{0}^{∗}(p). Journal of Number Theory, 196(??):14–34, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18302713.

Smoot:2019:FCR

[2421] Nicolas Allen Smoot. A family of congruences for Rogers–Ramanujan subpartitions. Journal of Number Theory, 196(??):35–60, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302981.

Kleine:2019:RIM

[2422] Sören Kleine. T -ranks of Iwasawa modules. Journal of Number Theory, 196(??):61–86, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302907.

Tan:2019:SBW

[2423] Bo Tan and Liling Zhang. On small bases which admit countably many expansions with multiple digits. Journal of Number Theory, 196(??):87–113, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302890.

Giulietti:2019:DGZ

[2424] Massimo Giulietti, Gábor Korchmáros, and Marco Timpanella. On the Dickson–Guralnick–Zieve curve. Journal of Number Theory, 196(??): 114–138, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302889.

Cho:2019:LDT

[2425] Peter J. Cho and Jeongho Park. n-level densities for twisted cubic Dirichlet L-functions. Journal of Number Theory, 196(??):139– 155, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302877.

Tran:2019:LFN

[2426] Long Tran. Local L-factors for GSp(4, F) via Novodvorsky’s zeta integrals. Journal of Number Theory, 196(??):156–167, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302865.

Damir:2019:WRT

[2427] Mohamed Taoufiq Damir and David Karpuk. Well-rounded twists of ideal lattices from real quadratic fields. Journal of Number Theory, 196(??):168–196, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302853.

Zaghloul:2019:NZG

[2428] Giamila Zaghloul. A note on the zeros of generalized Hurwitz zeta functions. Journal of Number Theory, 196(??):197–204, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302841.

Hayashida:2019:CPA

[2429] Daiki Hayashida. The characteristic polynomials of abelian varieties of higher dimension over finite fields. Journal of Number Theory, 196(??):205–222, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302828.

Wu:2019:TPD

[2430] Libo Wu and Xiaosheng Wu. The k-tuple prime difference champion. Journal of Number Theory, 196(??):223–243, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302816.

Hughes:2019:WBA

[2431] Adam M. Hughes. A Weil Banach algebra for multiplicative algebraic numbers. Journal of Number Theory, 196(??):244–271, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302798.

Wong:2019:CST

[2432] Peng-Jie Wong. On the Chebotarev–Sato–Tate phenomenon. Journal of Number Theory, 196(??):272–290, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18302786.

Cheng:2019:SNF

[2433] Weidong Cheng and Xuejun Guo. Some new families of noncongruent numbers. Journal of Number Theory, 196(??):291–305, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302774.

Choi:2019:APM

[2434] SoYoung Choi, Chang Heon Kim, and Kyung Seung Lee. Arithmetic properties for the minus space of weakly holomorphic modular forms. Journal of Number Theory, 196(??):306–339, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302749.

Bhanja:2019:DIT

[2435] Jagannath Bhanja and Ram Krishna Pandey. Direct and inverse theorems on signed sumsets of integers. Journal of Number Theory, 196(??):340–352, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302737.

Berczes:2019:SUR

[2436] A. Bérczes, L. Hajdu, I. Pink, and S. S. Rout. Sums of S-units in recurrence sequences. Journal of Number Theory, 196(??):353– 363, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303093.

Farmer:2019:VTF

[2437] David W. Farmer, Sally Koutsoliotas, and Stefan Lemurell. Varieties via their L-functions. Journal of Number Theory, 196(??):364– 380, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18300490.

DHoker:2019:PMR

[2438] Eric D’Hoker, Michael B. Green, and Pierre Vanhove. Proof of a modular relation between 1-, 2- and 3-loop Feynman diagrams on a torus. Journal of Number Theory, 196(??):381–419, March 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X17303396.

Anonymous:2019:PAa

[2439] Anonymous. Pages 1–410 (April 2019). Journal of Number Theory, 197 (??):1–410, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBd

[2440] Anonymous. Editorial Board. Journal of Number Theory, 197(??): ifc, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303317.

Pollack:2019:HOE

[2441] Paul Pollack. How often is Euler’s totient a perfect power? Journal of Number Theory, 197(??):1–12, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18303111.

Koo:2019:FCG

[2442] Ja Kyung Koo, Dong Hwa Shin, and Dong Sung Yoon. Form class groups for extended ring class fields. Journal of Number Theory, 197(??): 13–36, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18301847.

Rojas-Leon:2019:FMS

[2443] Antonio Rojas-León. Finite monodromy of some families of exponential sums. Journal of Number Theory, 197(??):37–48, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301859.

Mastrostefano:2019:SIC

[2444] Daniele Mastrostefano. Short intervals containing a prescribed number of primes. Journal of Number Theory, 197(??):49–61, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18301860.

Yan:2019:SLM

[2445] Dong Yan. Stable lattices in modular Galois representations and Hida deformation. Journal of Number Theory, 197(??):62–88, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18301872.

Maier:2019:MFM

[2446] Helmut Maier and Saurabh Kumar Singh. On multiplicative functions with mean-value close to −1 on primes. Journal of Number Theory, 197(??):89–105, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302361.

Kural:2019:DSE

[2447] Michael Kural. Dedekind sums with even denominators. Journal of Number Theory, 197(??):106–120, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18302385.

Kirschmer:2019:AEU

[2448] Markus Kirschmer. Automorphisms of even unimodular lattices over number fields. Journal of Number Theory, 197(??):121–134, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302397.

Virdol:2019:CCQ

[2449] Cristian Virdol. Cyclic components of quotients of abelian varieties mod p. Journal of Number Theory, 197(??):135–144, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302403.

Hockman:2019:GGI

[2450] Meira Hockman. The geometry of Gaussian integer continued fractions. Journal of Number Theory, 197(??):145–167, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302415.

Lee:2019:TMD

[2451] Seok Hyeong Lee and Seungjai Lee. The twisted moments and distribution of values of Dirichlet L-functions at 1. Journal of Number Theory, 197(??):168–184, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302427.

Bekki:2019:RHI

[2452] Hohto Bekki. The relative Hecke integral formula for an arbitrary extension of number fields. Journal of Number Theory, 197(??):185– 217, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302439.

Miska:2019:ADQ

[2453] Piotr Miska, Nadir Murru, and Carlo Sanna. On the p-adic denseness of the quotient set of a polynomial image. Journal of Number Theory, 197(??):218–227, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302452.

Sha:2019:ERS

[2454] Min Sha. Effective results on the Skolem Problem for linear recurrence sequences. Journal of Number Theory, 197(??):228–249, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18302476.

Williams:2019:RTD

[2455] Brandon Williams. Remarks on the theta decomposition of vectorvalued Jacobi forms. Journal of Number Theory, 197(??):250–267, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302488. See erratum [2548].

Jiang:2019:AAM

[2456] Yujiao Jiang and Guangshi Lü. On automorphic analogues of the Möbius randomness principle. Journal of Number Theory, 197(??):268– 296, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830249X.

Girard:2019:DZS

[2457] Benjamin Girard and Wolfgang A. Schmid. Direct zero-sum problems for certain groups of rank three. Journal of Number Theory, 197(??):297–316, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302518.

Stabler:2019:IEP

[2458] Axel Stäbler. Intermediate extensions of perverse constructible F_{p}- sheaves commute with smooth pullbacks. Journal of Number Theory, 197(??):317–340, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302695.

Tengely:2019:PVS

[2459] Szabolcs Tengely and Maciej Ulas. Power values of sums of certain products of consecutive integers and related results. Journal of Number Theory, 197(??):341–360, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302701.

Gassert:2019:FMQ

[2460] T. Alden Gassert, Hanson Smith, and Katherine E. Stange. A family of monogenic S_{4} quartic fields arising from elliptic curves. Journal of Number Theory, 197(??):361–382, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18302993.

Brevig:2019:HPR

[2461] Ole Fredrik Brevig and Winston Heap. High pseudomoments of the Riemann zeta function. Journal of Number Theory, 197(??):383– 410, April 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303007.

Anonymous:2019:PMb

[2462] Anonymous. Pages 1–396 (May 2019). Journal of Number Theory, 198 (??):1–396, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBe

[2463] Anonymous. Editorial Board. Journal of Number Theory, 198(??): ifc, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300022.

Diaconu:2019:TMR

[2464] Adrian Diaconu. On the third moment of L(\frac{1}{2}, χ_{d}) I: the rational function field case. Journal of Number Theory, 198(??):1–42, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18302932.

Khan:2019:ITP

[2465] Nadia Khan, Toru Nakahara, and Hiroshi Sekiguchi. An ideal theoretic proof on monogenity of cyclic sextic fields of prime power conductor. Journal of Number Theory, 198(??):43–51, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303019.

Zhu:2019:AFF

[2466] Yongchang Zhu. Addition formula and γ-factors for local fields. Journal of Number Theory, 198(??):52–73, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303020.

Wan:2019:LGR

[2467] Chen Wan. The local Ginzburg–Rallis model for generic representations. Journal of Number Theory, 198(??):74–123, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303032.

Klangwang:2019:ZCC

[2468] Jetjaroen Klangwang. Zeros of certain combinations of Eisenstein series of weight 2k, 3k, and k + l. Journal of Number Theory, 198(??): 124–138, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303044.

Song:2019:PMH

[2469] Ping Song, Wenguang Zhai, and Deyu Zhang. Power moments of Hecke eigenvalues for congruence group. Journal of Number Theory, 198(??):139–158, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303056.

Han:2019:EGZ

[2470] Dongchun Han and Hanbin Zhang. Erdős–Ginzburg–Ziv theorem and Noether number for C_{m}[ltimes]_{ϕ}C_{mn}. Journal of Number Theory, 198(??):159–175, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303068.

Matsumoto:2019:DFV

[2471] Kohji Matsumoto and Yumiko Umegaki. On the density function for the value-distribution of automorphic L-functions. Journal of Number Theory, 198(??):176–199, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830307X.

Kalmynin:2019:LVS

[2472] A. B. Kalmynin. Large values of short character sums. Journal of Number Theory, 198(??):200–210, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303081.

Hirano:2019:ATI

[2473] Hikaru Hirano and Masanori Morishita. Arithmetic topology in Ihara theory II: Milnor invariants, dilogarithmic Heisenberg coverings and triple power residue symbols. Journal of Number Theory, 198(??):211– 238, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1830310X.

Balog:2019:PGT

[2474] Antal Balog, András Biró, Gergely Harcos, and Péter Maga. The prime geodesic theorem in square mean. Journal of Number Theory, 198(??):239–249, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303123.

Vu:2019:HML

[2475] An Hoa Vu. Hermitian Maass lift for general level. Journal of Number Theory, 198(??):250–292, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303135.

Ma:2019:SHF

[2476] Liming Ma and Chaoping Xing. On subfields of the Hermitian function field involving the involution automorphism. Journal of Number Theory, 198(??):293–317, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303147.

Lou:2019:TCT

[2477] Miao Lou and Guangshi Lü. Triple correlations of ternary divisor functions. Journal of Number Theory, 198(??):318–345, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303275.

Farfan:2019:TAV

[2478] Victoria Cantoral Farfán. Torsion for abelian varieties of type III. Journal of Number Theory, 198(??):346–380, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303287.

Friedman:2019:FGT

[2479] Eduardo Friedman and Gabriel Ramirez-Raposo. Filling the gap in the table of smallest regulators up to degree 7. Journal of Number Theory, 198(??):381–385, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18302506.

Battistoni:2019:MDN

[2480] Francesco Battistoni. The minimum discriminant of number fields of degree 8 and signature (2, 3). Journal of Number Theory, 198(??):386– 395, May 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302750.

Anonymous:2019:PJb

[2481] Anonymous. Pages 1–494 (June 2019). Journal of Number Theory, 199 (??):1–494, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBf

[2482] Anonymous. Editorial Board. Journal of Number Theory, 199(??): ifc, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300320.

Young:2019:ECE

[2483] Matthew P. Young. Explicit calculations with Eisenstein series. Journal of Number Theory, 199(??):1–48, June 2019. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303457.

Andrade:2019:EEA

[2484] J. C. Andrade, A. Shamesaldeen, and C. Summersby. On elementary estimates of arithmetic sums for polynomial rings over finite fields. Journal of Number Theory, 199(??):49–62, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303263.

Banerjee:2019:VOS

[2485] Debika Banerjee, Ehud Moshe Baruch, and Evgeny Tenetov. A Voronoi– Oppenheim summation formula for totally real number fields. Journal of Number Theory, 199(??):63–97, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18301276.

Yu:2019:LST

[2486] Myungjun Yu. Large Shafarevich–Tate groups over quadratic number fields. Journal of Number Theory, 199(??):98–109, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300241.

Balasubramanian:2019:NSD

[2487] Kumar Balasubramanian. A note on self-dual representations of Sp(4, F). Journal of Number Theory, 199(??):110–125, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303391.

Kaltenbock:2019:BRI

[2488] Lisa Kaltenböck and Gerhard Larcher. On bounded remainder intervals and strongly non-bounded remainder intervals for sequences ({a_{n}α})_{n≥1}. Journal of Number Theory, 199(??):126–149, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303421.

Nadimpalli:2019:ESR

[2489] Santosh Nadimpalli. On extensions of supersingular representations of SL_{2} (Q_{p}). Journal of Number Theory, 199(??):150–167, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X18303433.

Chern:2019:AFC

[2490] Shane Chern. Asymptotics for the Fourier coefficients of eta-quotients. Journal of Number Theory, 199(??):168–191, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303469.

Alvarenga:2019:GHO

[2491] Roberto Alvarenga. On graphs of Hecke operators. Journal of Number Theory, 199(??):192–228, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303470.

Kalloniatis:2019:FFD

[2492] Tristan Kalloniatis. On flagged framed deformation problems of local crystalline Galois representations. Journal of Number Theory, 199(??): 229–250, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303482.

Tsaltas:2019:DCT

[2493] Konstantinos Tsaltas and Frazer Jarvis. Descending congruences of theta lifts on GSp_{4}. Journal of Number Theory, 199(??):251– 288, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303500.

Ye:2019:TFM

[2494] Dongxi Ye. On two families of modular subgroups. Journal of Number Theory, 199(??):289–309, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303512.

Takeda:2019:NCM

[2495] Wataru Takeda. A non-commutative multiple Dirichlet power product and an application. Journal of Number Theory, 199(??):310– 321, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303524.

Boca:2019:EOP

[2496] Florin P. Boca and Claire Merriman. α-Expansions with odd partial quotients. Journal of Number Theory, 199(??):322–341, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303536.

Lu:2019:ASC

[2497] Guangshi Lü and Dan Wang. Averages of shifted convolutions of general divisor sums involving Hecke eigenvalues. Journal of Number Theory, 199(??):342–351, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303548.

Chapman:2019:TLQ

[2498] Adam Chapman and Andrew Dolphin. Types of linkage of quadratic Pfister forms. Journal of Number Theory, 199(??):352–362, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X1830355X.

Li:2019:DCD

[2499] Junxian Li and George Shakan. On distinct consecutive r-differences. Journal of Number Theory, 199(??):363–376, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303561.

Garaev:2019:DES

[2500] M. Z. Garaev. Double exponential sums and congruences with intervals and exponential functions modulo a prime. Journal of Number Theory, 199(??):377–388, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X18303573.

Jiu:2019:OPC

[2501] Lin Jiu and Diane Yahui Shi. Orthogonal polynomials and connection to generalized Motzkin numbers for higher-order Euler polynomials. Journal of Number Theory, 199(??):389–402, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303597.

Kashio:2019:RBM

[2502] Tomokazu Kashio. On the ratios of Barnes’ multiple gamma functions to the p-adic analogues. Journal of Number Theory, 199(??):403– 435, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303603.

Stromberg:2019:NSM

[2503] Fredrik Strömberg. Noncongruence subgroups and Maass waveforms. Journal of Number Theory, 199(??):436–493, June 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18303585.

Anonymous:2019:PJc

[2504] Anonymous. Pages 1–486 (July 2019). Journal of Number Theory, 200 (??):1–486, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBg

[2505] Anonymous. Editorial Board. Journal of Number Theory, 200(??): ifc, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300666.

Ginzburg:2019:CGC

[2506] David Ginzburg. On certain global constructions of automorphic forms related to a small representation of F_{4}. Journal of Number Theory, 200(??):1–95, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300599.

Ivic:2019:DVA

[2507] Aleksandar P. Ivić and Maxim A. Korolev. On the distribution of values of the argument of the Riemann zeta-function. Journal of Number Theory, 200(??):96–131, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930023X.

Hirose:2019:TNS

[2508] Minoru Hirose. On the theory of normalized Shintani L-functions and its application to Hecke L-functions, I: real quadratic fields. Journal of Number Theory, 200(??):132–153, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X18303615.

Finat:2019:GRC

[2509] Eugenio Finat. On the growth of the (S, {2})-refined class number. Journal of Number Theory, 200(??):154–160, July 2019. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300137.

Gun:2019:FSS

[2510] Sanoli Gun, Balesh Kumar, and Biplab Paul. The first simultaneous sign change and non-vanishing of Hecke eigenvalues of newforms. Journal of Number Theory, 200(??):161–184, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300149.

Chen:2019:ECM

[2511] Imin Chen and Parinaz Salari Sharif. An explicit correspondence of modular curves. Journal of Number Theory, 200(??):185–204, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300150.

Li:2019:SIM

[2512] Zhonghua Li and Noriko Wakabayashi. Sum of interpolated multiple q-zeta values. Journal of Number Theory, 200(??):205–259, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300162.

Bachmann:2019:ABB

[2513] Henrik Bachmann. The algebra of bi-brackets and regularized multiple Eisenstein series. Journal of Number Theory, 200(??):260– 294, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300186.

Overkamp:2019:FDO

[2514] Otto Overkamp. Finite descent obstruction and non-abelian reciprocity. Journal of Number Theory, 200(??):295–315, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300198.

Yu:2019:MRU

[2515] Han Yu. Multi-rotations on the unit circle. Journal of Number Theory, 200(??):316–328, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300204.

Chirre:2019:EVN

[2516] Andrés Chirre. Extreme values for S_{n}(σ, t) near the critical line. Journal of Number Theory, 200(??):329–352, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300216.

Coccia:2019:HPI

[2517] Simone Coccia. The Hilbert Property for integral points of affine smooth cubic surfaces. Journal of Number Theory, 200(??):353– 379, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18303627.

Morton:2019:PFD

[2518] Patrick Morton. Product formulas for the 5-division points on the Tate normal form and the Rogers–Ramanujan continued fraction. Journal of Number Theory, 200(??):380–396, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300253.

Wu:2019:DHN

[2519] Bing-Ling Wu and Yong-Gao Chen. On the denominators of harmonic numbers, II. Journal of Number Theory, 200(??):397–406, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300290.

McMeekin:2019:APS

[2520] Christine McMeekin. On the asymptotics of a prime spin relation. Journal of Number Theory, 200(??):407–426, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300307. See

corrigendum [2647].

Grau:2019:FCN

[2521] José María Grau and Antonio M. Oller-Marcén. Fast computation of the number of solutions to x_{1}^{2} + ⋯ + x_{k}^{2} ≡ λ(modn). Journal of Number Theory, 200(??):427–440, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1830283X.

Grenie:2019:ECD

[2522] Loïc Grenié and Giuseppe Molteni. An explicit Chebotarev density theorem under GRH. Journal of Number Theory, 200(??):441– 485, July 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300174.

Anonymous:2019:PAb

[2523] Anonymous. Pages 1–474 (August 2019). Journal of Number Theory, 201 (??):1–474, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBh

[2524] Anonymous. Editorial Board. Journal of Number Theory, 201(??): ifc, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301295.

Murty:2019:PDS

[2525] M. Ram Murty and François Séguin. Prime divisors of sparse values of cyclotomic polynomials and Wieferich primes. Journal of Number Theory, 201(??):1–22, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300927.

Donnay:2019:AQS

[2526] Christopher Donnay, Stephan Ramon Garcia, and Jeremy Rouse. padic quotient sets II: Quadratic forms. Journal of Number Theory, 201(??):23–39, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930099X.

Chan:2019:SRI

[2527] Wai Kiu Chan and Lenny Fukshansky. Small representations of integers by integral quadratic forms. Journal of Number Theory, 201(??): 40–52, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300861.

Gallagher:2019:GFP

[2528] Katherine Gallagher, Lucia Li, Naomi Sweeting, Katja Vassilev, and Katharine Woo. Generating functions for power moments of elliptic curves over F_{p}. Journal of Number Theory, 201(??):53–67, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300587.

Bordignon:2019:EBE

[2529] Matteo Bordignon. Explicit bounds on exceptional zeroes of Dirichlet L-functions. Journal of Number Theory, 201(??):68–76, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300794.

Berkovich:2019:PII

[2530] Alexander Berkovich and Ali Kemal Uncu. Polynomial identities implying Capparelli’s partition theorems. Journal of Number Theory, 201(??):77–107, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301040.

Babinkostova:2019:APE

[2531] L. Babinkostova, J. C. Bahr, Y. H. Kim, E. Neyman, and G. K. Taylor. Anomalous primes and the elliptic Korselt criterion. Journal of Number Theory, 201(??):108–123, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19300897.

Roche-Newton:2019:IST

[2532] Oliver Roche-Newton and Ilya D. Shkredov. If A + A is small then AAA is superquadratic. Journal of Number Theory, 201(??):124–134, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301027.

Chen:2019:PEN

[2533] Yong-Gao Chen and Ya-Li Li. On a problem of Erdős, Nathanson and Sárközy. Journal of Number Theory, 201(??):135–147, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301015.

Li:2019:SIF

[2534] Zhonghua Li and Ende Pan. Sum of interpolated finite multiple harmonic q-series. Journal of Number Theory, 201(??):148–175, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301003.

Mundici:2019:TDA

[2535] Daniele Mundici. Triangles in diophantine approximation. Journal of Number Theory, 201(??):176–189, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300873.

Sakugawa:2019:MFP

[2536] Kenji Sakugawa. On modified finite polylogarithms. Journal of Number Theory, 201(??):190–205, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19300885.

Horinaga:2019:RGE

[2537] Shuji Horinaga. On the representations generated by Eisenstein series of weight \frac{n+3}{2}. Journal of Number Theory, 201(??):206–227, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300836.

Healey:2019:SCH

[2538] Vivian Olsiewski Healey and Wade Hindes. Stochastic canonical heights. Journal of Number Theory, 201(??):228–256, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300964.

Shiga:2019:SCM

[2539] Hironori Shiga. The Shimura canonical model for the quaternion algebra of discriminant 6 on the Ihara–Kurihara conic. Journal of Number Theory, 201(??):257–279, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300952.

Zhang:2019:FCP

[2540] Wei Zhang. Fourier coefficients at primes twisted with exponential functions. Journal of Number Theory, 201(??):280–291, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300940.

Banerjee:2019:SRT

[2541] Debargha Banerjee and Tathagata Mandal. Supercuspidal ramifications and traces of adjoint lifts. Journal of Number Theory, 201(??):292– 321, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300939.

Jiang:2019:GHR

[2542] Tiefeng Jiang and Ke Wang. A generalized Hardy–Ramanujan formula for the number of restricted integer partitions. Journal of Number Theory, 201(??):322–353, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300824.

Szikszai:2019:APL

[2543] Márton Szikszai and Volker Ziegler. On arithmetic progressions in Lucas sequences — II. Journal of Number Theory, 201(??):354–376, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300915.

Zhang:2019:TKI

[2544] Long Zhang and Kejian Xu. On the tame kernels of imaginary cyclic quartic fields with class number one. Journal of Number Theory, 201(??):377–397, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300812.

Murata:2019:PSN

[2545] Manabu Murata. Primitive solutions of nonscalar quadratic Diophantine equations. Journal of Number Theory, 201(??):398–435, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300800.

Amato:2019:PAW

[2546] Gianluca Amato, Maximilian F. Hasler, Giuseppe Melfi, and Maurizio Parton. Primitive abundant and weird numbers with many prime factors. Journal of Number Theory, 201(??):436–459, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301039.

Argaez-Garcia:2019:PPS

[2547] Alejandro Argáez-García. On perfect powers that are sums of cubes of a five term arithmetic progression. Journal of Number Theory, 201(??):460–472, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300903.

Williams:2019:ERT

[2548] Brandon Williams. Erratum to “Remarks on the theta decomposition of vector-valued Jacobi forms” [J. Number Theory 197 (2019) 250–267]. Journal of Number Theory, 201(??):473–474, August 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300976. See [2455].

Anonymous:2019:PS

[2549] Anonymous. Pages 1–458 (September 2019). Journal of Number Theory, 202(??):1–458, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBi

[2550] Anonymous. Editorial Board. Journal of Number Theory, 202(??):ifc, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301659.

Lalin:2019:TPP

[2551] Matilde Lalín and Xinchun Ma. θ-triangle and ω-parallelogram pairs with common area and common perimeter. Journal of Number Theory, 202(??):1–26, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300782.

Petsche:2019:DCS

[2552] Clayton Petsche and Emerald Stacy. A dynamical construction of small totally p-adic algebraic numbers. Journal of Number Theory, 202(??):27–36, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300630.

Ulas:2019:EIP

[2553] Maciej Ulas. Is every irreducible polynomial reducible after a polynomial substitution? Journal of Number Theory, 202(??):37–59, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300575.

Zheng:2019:DAR

[2554] Lixuan Zheng. Diophantine approximation and run-length function on β-expansions. Journal of Number Theory, 202(??):60–90, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19300563.

Bhand:2019:DSA

[2555] Ajit Bhand and Karam Deo Shankhadhar. On Dirichlet series attached to quasimodular forms. Journal of Number Theory, 202(??):91–106, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300551.

Balasubramanyam:2019:PCS

[2556] Baskar Balasubramanyam and Kaneenika Sinha. Pair correlation statistics for Sato–Tate sequences. Journal of Number Theory, 202(??):107– 140, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930054X.

Sun:2019:DFM

[2557] Haiwei Sun and Yangbo Ye. Double first moment for L(\frac{1}{2}, Sym^{2}f × g) by applying Petersson’s formula twice. Journal of Number Theory, 202 (??):141–159, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300538. See corrigendum [2649].

Jung:2019:GNP

[2558] Junehyuk Jung. On the growth of the number of primitive totally geodesic surfaces in some hyperbolic 3-manifolds. Journal of Number Theory, 202(??):160–175, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300514.

Suzuki:2019:QDR

[2559] Miyu Suzuki. Quaternion distinguished representations and unstable base change for unitary groups. Journal of Number Theory, 202 (??):176–199, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300502.

Chommaux:2019:DSR

[2560] Marion Chommaux. Distinction of the Steinberg representation and a conjecture of Prasad and Takloo-Bighash. Journal of Number Theory, 202(??):200–219, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300484.

Zhang:2019:DIP

[2561] Min Zhang and Jinjiang Li. On a Diophantine inequality over primes. Journal of Number Theory, 202(??):220–253, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300472.

Dabrowski:2019:ECR

[2562] Andrzej Dabrowski and Tomasz Jedrzejak. Elliptic curves over the rationals with good reduction outside two odd primes. Journal of Number Theory, 202(??):254–277, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300460.

Bordelles:2019:SIE

[2563] Olivier Bordellès, Lixia Dai, Randell Heyman, Hao Pan, and Igor E. Shparlinski. On a sum involving the Euler function. Journal of Number Theory, 202(??):278–297, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300459. See

corrigendum [2645].

Choi:2019:IBC

[2564] Dohoon Choi and Subong Lim. Independence between coefficients of two modular forms. Journal of Number Theory, 202(??):298–315, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300447.

Garcia:2019:SEC

[2565] Stephan Ramon Garcia and George Todd. Supercharacters, elliptic curves, and the sixth moment of Kloosterman sums. Journal of Number Theory, 202(??):316–331, September 2019. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19300435.

Aikawa:2019:BMW

[2566] Yusuke Aikawa. The bounds of the Mordell–Weil ranks in cyclotomic towers of function fields. Journal of Number Theory, 202(??):332–346, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19300423.

He:2019:CQT

[2567] Bo He, Japhet Odjoumani, and Alain Togbé. On a class of quartic Thue equations with three parameters. Journal of Number Theory, 202 (??):347–387, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300411.

Gonzalez-Aviles:2019:NMC

[2568] Cristian D. González-Avilés. The norm map and the capitulation kernel. Journal of Number Theory, 202(??):388–411, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930040X.

Bilu:2019:GCN

[2569] Yuri Bilu, Diego Marques, and Alain Togbé. Generalized Cullen numbers in linear recurrence sequences. Journal of Number Theory, 202 (??):412–425, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300289.

Driver:2019:CSN

[2570] Eric D. Driver and John W. Jones. Computing septic number fields. Journal of Number Theory, 202(??):426–429, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300988.

Springer:2019:CER

[2571] Caleb Springer. Computing the endomorphism ring of an ordinary abelian surface over a finite field. Journal of Number Theory, 202 (??):430–457, September 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300526.

Anonymous:2019:PO

[2572] Anonymous. Pages 1–428 (October 2019). Journal of Number Theory, 203(??):1–428, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBj

[2573] Anonymous. Editorial Board. Journal of Number Theory, 203(??): ifc, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301933.

Shparlinski:2019:BAV

[2574] Igor E. Shparlinski. Bounds on average values of double incomplete Kloosterman sums. Journal of Number Theory, 203(??):1–11, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301143.

Hou:2019:LAS

[2575] Fei Hou and Bin Chen. Level aspect subconvexity for twisted L-functions. Journal of Number Theory, 203(??):12–31, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301106.

Binder:2019:FRA

[2576] John Binder. Fields of rationality of automorphic representations: the case of unitary groups. Journal of Number Theory, 203(??):32–67, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730269X.

Cohen:2019:LNP

[2577] Stephen D. Cohen and Tim Trudgian. Lehmer numbers and primitive roots modulo a prime. Journal of Number Theory, 203(??):68– 79, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930109X.

Oliveira:2019:ZDP

[2578] Willian D. Oliveira. Zeros of Dirichlet polynomials via a density criterion. Journal of Number Theory, 203(??):80–94, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X18300507.

Du:2019:TAS

[2579] Tuoping Du and Tonghai Yang. Twisted arithmetic Siegel Weil formula on X_{0}(N). Journal of Number Theory, 203(??):95–117, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301751.

Babnik:2019:IKA

[2580] Maurice Babnik. Irreduzible Komponenten von 2-adischen Deformationsräumen. (German) [Irreducible components of 2-adic deformation spaces]. Journal of Number Theory, 203(??):118–138, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X18302804.

Wan:2019:CNR

[2581] Daqing Wan. Class numbers and p-ranks in Z_{p}^{d}-towers. Journal of Number Theory, 203(??):139–154, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X18303445.

Yu:2019:CVG

[2582] Peng Yu. CM values of Green functions associated to special cycles on Shimura varieties with applications to Siegel 3-fold X_{2}(2). Journal of Number Theory, 203(??):155–210, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301763.

Acharya:2019:SOB

[2583] Ratnadeep Acharya. Strong orthogonality between the Möbius function, additive characters and the coefficients of the L-functions of degree three. Journal of Number Theory, 203(??):211–229, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19300770.

Zeng:2019:MZS

[2584] Xiangneng Zeng and Guixin Deng. Minimal zero-sum sequences over [ − m, n]. Journal of Number Theory, 203(??):230–241, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301052.

Dan:2019:EVI

[2585] Ananyo Dan and Inder Kaur. Examples of varieties with index one on C_{1}-fields. Journal of Number Theory, 203(??):242–248, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301064.

Jung:2019:MGT

[2586] Ho Yun Jung, Chang Heon Kim, Soonhak Kwon, and Yeong-Wook Kwon. Modularity of Galois traces of Weber’s resolvents. Journal of Number Theory, 203(??):249–275, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19301076.

Tafazolian:2019:CMH

[2587] Saeed Tafazolian and Jaap Top. On certain maximal hyperelliptic curves related to Chebyshev polynomials. Journal of Number Theory, 203(??):276–293, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301088.

Erazo:2019:PPC

[2588] Harold S. Erazo, Carlos A. Gómez, and Florian Luca. On Pillai’s problem with X-coordinates of Pell equations and powers of 2. Journal of Number Theory, 203(??):294–309, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19301155.

Kafle:2019:CPE

[2589] Bir Kafle, Florian Luca, Amanda Montejano, László Szalay, and Alain Togbé. On the X-coordinates of Pell equations which are products of two Fibonacci numbers. Journal of Number Theory, 203(??):310– 333, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301167.

Jin:2019:SCJ

[2590] Seokho Jin and Subong Lim. On sign changes of Jacobi forms. Journal of Number Theory, 203(??):334–349, October 2019. CODEN JNUTA9.

ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301118.

Xue:2019:CVD

[2591] Hang Xue. Central values of degree six L-functions. Journal of Number Theory, 203(??):350–359, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X1930112X.

Chabert:2019:PFP

[2592] Jean-Luc Chabert. From Pólya fields to Pólya groups (I) Galois extensions. Journal of Number Theory, 203(??):360–375, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301131.

Jahnel:2019:ABC

[2593] Jörg Jahnel and Damaris Schindler. On the algebraic Brauer classes on open degree four del Pezzo surfaces. Journal of Number Theory, 203(??):376–427, October 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19300496.

Anonymous:2019:PN

[2594] Anonymous. Pages 1–682 (November 2019). Journal of Number Theory, 204(??):1–682, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBk

[2595] Anonymous. Editorial Board. Journal of Number Theory, 204(??):ifc, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302227.

Mellit:2019:EDP

[2596] Anton Mellit. Elliptic dilogarithms and parallel lines. Journal of Number Theory, 204(??):1–24, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19301246.

Brown:2019:EPI

[2597] Jim Brown, Hugh Geller, Rico Vicente, and Alexandra Walsh. Eigenform product identities for degree two Siegel modular forms. Journal of Number Theory, 204(??):25–40, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301180.

Sheydvasser:2019:QOS

[2598] Arseniy Sheydvasser. Quaternion orders and sphere packings. Journal of Number Theory, 204(??):41–98, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301192.

Chattopadhyay:2019:BFH

[2599] Jaitra Chattopadhyay and Subramani Muthukrishnan. Biquadratic fields having a non-principal Euclidean ideal class. Journal of Number Theory, 204(??):99–112, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301209.

Bostan:2019:LCM

[2600] Alin Bostan, Alexander Marynych, and Kilian Raschel. On the least common multiple of several random integers. Journal of Number Theory, 204(??):113–133, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301222.

Qureshi:2019:DMR

[2601] Claudio Qureshi and Lucas Reis. Dynamics of the a-map over residually finite Dedekind domains and applications. Journal of Number Theory, 204(??):134–154, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301234.

Hemelaer:2019:ATI

[2602] Jens Hemelaer. An arithmetic topos for integer matrices. Journal of Number Theory, 204(??):155–184, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301404.

Chatterjee:2019:SEC

[2603] Tapas Chatterjee and Suraj Singh Khurana. Shifted Euler constants and a generalization of Euler–Stieltjes constants. Journal of Number Theory, 204(??):185–210, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301416.

Lu:2019:PC

[2604] Hengfei Lu. The Prasad conjectures for U_{2}, SO_{4} and Sp_{4}. Journal of Number Theory, 204(??):211–245, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301428.

Sartori:2019:LPR

[2605] Andrea Sartori. Least primitive root and simultaneous power nonresidues. Journal of Number Theory, 204(??):246–263, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301441.

Lu:2019:APF

[2606] Yaming Lu. On almost primes of the form p + 2^{ℓ}. Journal of Number Theory, 204(??):264–278, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19301453.

Laurincikas:2019:URZ

[2607] Antanas Laurinčikas. Universality of the Riemann zeta-function in short intervals. Journal of Number Theory, 204(??):279–295, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301465.

Mocanu:2019:PES

[2608] Andreea Mocanu. Poincaré and Eisenstein series for Jacobi forms of lattice index. Journal of Number Theory, 204(??):296–333, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301477.

Martinez:2019:EPO

[2609] F. E. Brochero Martínez and Sávio Ribas. Extremal product-one free sequences in C_{q}[rtimes]_{s}C_{m}. Journal of Number Theory, 204(??):334– 353, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301489.

Steiner:2019:EVM

[2610] Raphael S. Steiner. Effective Vinogradov’s mean value theorem via efficient boxing. Journal of Number Theory, 204(??):354–404, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301507.

Choi:2019:AVC

[2611] Junhwa Choi. On the 2-adic valuations of central L-values of elliptic curves. Journal of Number Theory, 204(??):405–422, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301519.

Choi:2019:TZC

[2612] SoYoung Choi and Bo-Hae Im. The transcendence of zeros of canonical basis elements of the space of weakly holomorphic modular forms for Γ_{0}(2). Journal of Number Theory, 204(??):423–434, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301520.

Tenenbaum:2019:PPS

[2613] Gérald Tenenbaum, Jie Wu, and Ya-Li Li. Power partitions and saddle-point method. Journal of Number Theory, 204(??):435–445, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301532.

Choi:2019:RAP

[2614] Suh Hyun Choi, Chang Heon Kim, Yeong-Wook Kwon, and Kyu-Hwan Lee. Rationality and p-adic properties of reduced forms of half-integral weight. Journal of Number Theory, 204(??):446–470, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301544.

Nakaya:2019:NLF

[2615] Tomoaki Nakaya. The number of linear factors of supersingular polynomials and sporadic simple groups. Journal of Number Theory, 204 (??):471–496, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301556.

Florez:2019:ECI

[2616] Jorge Flórez, Cihan Karabulut, and Tian An Wong. Eisenstein cocycles over imaginary quadratic fields and special values of L-functions. Journal of Number Theory, 204(??):497–531, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301568.

Gazaki:2019:TDT

[2617] Evangelia Gazaki. A Tate duality theorem for local Galois symbols II; The semi-abelian case. Journal of Number Theory, 204(??):532–560, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930157X.

Hwang:2019:AGS

[2618] WonTae Hwang. Automorphism groups of simple polarized abelian varieties of odd prime dimension over finite fields. Journal of Number Theory, 204(??):561–578, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301581.

Kumar:2019:FCN

[2619] Balesh Kumar. On Fourier coefficients of Niebur Poincaré series. Journal of Number Theory, 204(??):579–598, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X1930160X.

Ma:2019:EBS

[2620] Shilin Ma, Kevin McGown, Devon Rhodes, and Mathias Wanner. Explicit bounds for small prime nonresidues. Journal of Number Theory, 204(??):599–607, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301611.

Cherubini:2019:SUT

[2621] G. Cherubini and A. Perelli. A spectral universality theorem for Maass L-functions. Journal of Number Theory, 204(??):608–623, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301623.

Atsuta:2019:ITC

[2622] Mahiro Atsuta. Iwasawa theory for class groups of CM fields with p = 2. Journal of Number Theory, 204(??):624–660, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301635.

Jorgenson:2019:HEP

[2623] Jay Jorgenson, Lejla Smajlović, and Holger Then. The Hauptmodul at elliptic points of certain arithmetic groups. Journal of Number Theory, 204(??):661–682, November 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X1930126X.

Anonymous:2019:PD

[2624] Anonymous. Pages 1–346 (December 2019). Journal of Number Theory, 205(??):1–346, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Anonymous:2019:EBl

[2625] Anonymous. Editorial Board. Journal of Number Theory, 205(??):ifc, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19302434.

Blomer:2019:SRF

[2626] Valentin Blomer, Xiaoqing Li, and Stephen D. Miller. A spectral reciprocity formula and non-vanishing for L-functions on GL(4) × GL(2). Journal of Number Theory, 205(??):1–43, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301908.

Faber:2019:API

[2627] Xander Faber, Michelle Manes, and Laura Walton. An alternate proof of idempotent relations among periodic points and quotients. Journal of Number Theory, 205(??):44–49, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301805.

Bernshteyn:2019:ILB

[2628] Anton Bernshteyn and Michael Tait. Improved lower bound for difference bases. Journal of Number Theory, 205(??):50–58, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19301817.

Hu:2019:IPR

[2629] Su Hu, Min-Soo Kim, Pieter Moree, and Min Sha. Irregular primes with respect to Genocchi numbers and Artin’s primitive root conjecture. Journal of Number Theory, 205(??):59–80, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301179.

Jarossay:2019:BNO

[2630] David Jarossay. A bound on the norm of overconvergent p-adic multiple polylogarithms. Journal of Number Theory, 205(??):81–121, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301799.

Muller:2019:NCU

[2631] Peter Müller. A note on a conjecture by Ulas on polynomial substitutions. Journal of Number Theory, 205(??):122–123, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19302112.

Gencev:2019:SWS

[2632] Marian Genčev. On some weighted sum formulas involving general multiple zeta-type series. Journal of Number Theory, 205(??):124–147, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301829.

Ding:2019:SSG

[2633] Yuchen Ding and Haiyan Zhou. Subset sums over Galois rings. Journal of Number Theory, 205(??):148–169, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X19301830.

Labesse:2019:CMC

[2634] Jean-Pierre Labesse and Joachim Schwermer. Central morphisms and cuspidal automorphic representations. Journal of Number Theory, 205 (??):170–193, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301842.

Byeon:2019:ELB

[2635] Dongho Byeon and Jigu Kim. An explicit lower bound for special values of Dirichlet L-functions II. Journal of Number Theory, 205 (??):194–209, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19301854.

Paladino:2019:DQC

[2636] Laura Paladino. Divisibility questions in commutative algebraic groups. Journal of Number Theory, 205(??):210–245, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X19301866.

Plessis:2019:MHW

[2637] Arnaud Plessis. Minoration de la hauteur de Weil dans un compositum de corps de rayon. (French) [Minimization of the Weil height in a composite of ray bodies]. Journal of Number Theory, 205(??):246– 276, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19301878.

Gao:2019:WMO

[2638] Hui Gao. Wach models and overconvergence of étale (ϕ, Γ)-modules. Journal of Number Theory, 205(??):277–299, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X1930188X.

Csige:2019:RIC

[2639] Tamás Csige. Reflexive ideals and completely faithful modules of Iwasawa algebras. Journal of Number Theory, 205(??):300–320, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1930191X.

Mello:2019:CHM

[2640] Jorge Mello. Canonical heights and monomial maps: On effective lower bounds for points with dense orbit. Journal of Number Theory, 205 (??):321–339, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X19302070.

Dujella:2019:RDS

[2641] Andrej Dujella, Matija Kazalicki, and Vinko Petričević. Rational Diophantine sextuples with square denominators. Journal of Number Theory, 205(??):340–346, December 2019. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X19302136.

Perucca:2020:ARA

[2642] Antonella Perucca, Pietro Sgobba, and Sebastiano Tronto. Addendum to: Reductions of algebraic integers [J. Number Theory 167 (2016) 259–283]. Journal of Number Theory, 209(??):391–395, April 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X19303361. See [1699].

Baier:2020:ELS

[2643] Stephan Baier and Rajneesh Kumar Singh. Erratum to “Large sieve inequality with power moduli for function fields” [J. Number Theory 196 (2019) 1–13]. Journal of Number Theory, 210(??):431– 432, May 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X19303774. See [2419].

Chang:2020:CAN

[2644] Sungkon Chang. Corrigendum to “Average number of Zeckendorf integers” [J. Number Theory 186 (2018) 452–472]. Journal of Number Theory, 217(??):443–444, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0022314X2030216X. See [2217].

Zhai:2021:CSI

[2645] Wenguang Zhai. Corrigendum to “On a sum involving the Euler function” [J. Number Theory 202 (2019) 278–297. Journal of Number Theory, 222(??):423–425, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0022314X20303267. See [2563].

Salami:2021:CRP

[2646] Sajad Salami. Corrigendum to “The rational points on certain Abelian varieties over function fields” [J. Number Theory 195 (2019) 330–337]. Journal of Number Theory, 227(??):306–307, October 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0022314X21001025. See [2414].

McMeekin:2022:CAP

[2647] Christine McMeekin. Corrigendum to: “On the asymptotics of a prime spin relation” [J. Number Theor. 200 (2019) 407–426]. Journal of Number Theory, 240(??):751–753, November 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X22000531. See

[2520].

Daowsud:2023:CCF

[2648] Katthaleeya Daowsud and Thomas A. Schmidt. Corrigendum to: “Continued fractions for rational torsion” [J. Number Theory 189 (2018) 115– 130]. Journal of Number Theory, 246(??):326–327, May 2023. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0022314X22002554.

Sun:2024:CDF

[2649] Haiwei Sun and Yangbo Ye. Corrigendum to “Double first moment for L(\frac{1}{2}, Sym^{2}f × g by applying Petersson’s formula twice” [J. Number Theory 202 (2019) 141–159]. Journal of Number Theory, 260(??): 196–197, July 2024. CODEN JNUTA9. ISSN 0022-314X (print), 1096-

1658 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X24000519. See [2557].